A Bibliography of Publications about the Fortran Programming Language: Part 2: 1981–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 2.95

Title word cross-reference

# [Hof87].

(a, a′) [SR84a, SR84a]. (a, b) [SMD84, SR84a]. (a, bγ) [SMD84, SR84a]. (a, bγ − γ) [SMD84, SR84a]. (a, γ) [SR84a]. (a, γ − γ) [SR84a]. (r_{∞}, n_{1/2}, s_{1/2}) [Hoc85]. (r_{∞}, n_{1/2}) [Tem89a]. 2 [Zak84]. 3 [Zak84]. 9 [RRC89]. ^{−1} [KW87b]. ^{2}Σ [Nai84, Nai86]. Ax = b [BS81a, BS81b, BS81c]. F [YK88]. f(x) [Gaf83b, Gaf83c, Gaf83a]. H_{p} [SSS84a, SSS84b, SSS84c]. Ki_{n}(x) [Amo83a]. ∫_{0}^{1} t^{−1} log^{n−1t} log^{p}(1 − t)dt [Köl82]. j [RRC89]. J_{n}(x) [Col84b]. L_{1} [Abd80]. Q [Gra86a]. dif^{−1} [KW87a]. r × c [MP86a, MP86c, MP86b]. SU(3) [AD84]. t [YK88].

-carotene [Wyl86]. -estimators [KW87a].

/370 [Ehi82].

1 [Ber82a, Cra83, EH81, Ess88, Fog85, Fog87, Fog88, Pet83, Res86, hTD88, Tem89a, Tem89b, WSL88]. 1-port [hTDT88]. 1-VX [Ess88]. 1.0B [Zho89]. 1/2 [SDH84]. 10 [BK88]. 10/ [Dig83, Dig85a]. 100/200 [MMM85]. 10R1 [Spe82]. 11 [AEL^{+}86, Cla89, CP84, Cli84, Dig82a, Dig82b, Dig82d, Dig82c, Dig82e, Dig82f, Dig84e, Dig88a, hHtM81, Hue83, Mic83, Sul88, Wie82]. 11/ [sT85]. 115 [CL83]. 117 [WH87]. 11R1 [Spe85]. 11th [VV86]. 128K [Mic84f]. 16 [hH82, Sho85]. 16-Bit-Rechnern [Kal85a].

164 [Tou84]. 165 [Stu81a]. 16BST [Kaw84]. 172 [OM82]. 192 [Joh83]. 1978 [AC87, Com89, RH84b]. 1980 [Bis81]. 1981-01 [IEC81]. 1982 [Jap82, Mor82]. 1985 [Ame85d]. 1985-12 [IEC85]. 1988 [Wri89]. 1988-12 [IEC88]. 1989 [Com89]. 198O [Fri84]. 198X [AC87, Com89, Ame87b]. 1a [Cha86a].

2 [AGS88, Ano87a, AS89a, AS89b, Bai87, Ber88b, Cal86, HV83, Int88c, LtW88, LCH^{+}88, SS88]. 2-CPU [Hoc85]. 20.00 [Wil87b]. 200 [BK88, MMM85, VH87]. 205 [BK88, LtW88, LW88b, Riz85, sT85, VSH83]. 25th [AW82]. 2B [KGRY81]. 2C [GKRY82, KRYG82b, KRYG82c, KRYG82a, KRYG82d]. 2E [CM81a]. 2nd [HF81, RWA84, Rod84].

3 [Gre88, MS84b, Sch84c, Sch84d, Sch84e, Sch84b]. 3-D [Kli89]. 3000 [Wie86a, Wie86b]. 3090 [AC86, ALPC88, Car88b, GSZ88, RV89]. 3090/VF [ALPC88]. 32 [Cro87a]. 32k [Kal85a]. 360 [Int83c]. 370 [ACG^{+}86, Chi85b, Chi86a, Int83c]. 370-III [Int85b]. 3A [Con81b]. 3d [RWA84, Rod84]. 3L [Cul88]. 3rd [APD86].

4 [Hei83]. 4-circle [Tho84e]. 4.1M [Dig85b]. 40 [WSL88].

5726 [Int83b]. 5726-FO1 [Int83b]. 590 [Van82]. 5th [IEE81].

6 [Hon81b, SM84, SK83]. 6/VM [Hew85]. 60 [Fog85, Fog87, Fog88, Res86]. 6000/B [Bur81b]. 646 [Cra86a]. 65-110 [Ass83b]. 654 [DM87b]. 658 [LK88]. 66 [BP81b]. 664 [Sch88a].

’77 [KK89b, LHP87, LH87, PDA83a, PDA86, Wat82a, Wat82b, AM81, AS88b, AG87c, A^{+}81, AHH89, BS88a, Bee81, Bee85b, BS86a, Bez88, Boi84, Boi87, Bor85b, Bor85a, Boy89b, BGG85a, Bru86, Bur87, Chu88, Col83, DH88a, DH88b, DKG89a, Edg89b, EE84, Ell83, Ett84a, Fen87b, Fen87a, FGGF86, Fuo86b, Fuo86c, Fuo86a, GY82, Gro83, Hah87, HRC87, HB83, HH85, Wor84a, Kat82, KL83a, KL83b, KL85a, KL85b, KF87, Lah88a, Lah88c, Lah88d, Lam81, Lam86, Lan88, Law83, Leh86, Leh83a, Mar83a, Mar83b, Mas83, Mas87, McC84a, MS88c, McK85b, Mei84, Met85a, Mey84, Mic84e, Mic84f, Mil87a, MA89, Mon82c, Mon89b, Mos88, Nan81b, Nic85b, NEM84, NL83, NL88, OK87, Pag84, Pol81, Pol83, Pru87, Rat87, Rou83, Rou86, Rul83a, Rul83b, SM88b, Sch87c, See81b, Smi85e, Sne88, Wor84b]. 77 [Weh85, Wor88, Wor89a, Wu83, Edm86, McC84c, MS88b, Nad86, RB82]. 77-beginning [YS84c, YS84d, Mic84f]. 77-Beginning/Book [Mic84f]. 77-Programmen [Mey84, NEM84]. 77/SFTRAN3 [Bee81].

8 [WSL88]. 8.25 [Edm86]. 8.95 [Cou85a]. 80 [Bel84, EML88, McN83]. 8088-based [Sho85]. ’82 [AW82, ACM82]. 820/80 [EML88]. ’84 [ACM84, Van84a]. 86 [Int85a]. 8651-1 [IEC88, ISO88]. ’87 [Dav89, ACM87, Wex87]. ’88 [Gia89, IEE88a, IEE88b, ML88]. ’89 [ACM89a, ACM89b, Gon89, Voe89]. 8x [BBB^{+}83, Smi83b, X3J89, AR87, Ber85b, Com89, Cha87a, Dol88, Edg89a, Edg89b, Ful87, Ame89b, Lah87a, Mei87, Mei88, Mei89a, Met87a, MR89, Miy87, RW85, Rei87a, Rei89a, Rei89b, Smi85a, ANS89, AKLS88, BBG^{+}82, BBG^{+}84, Bur86a, LCH^{+}88, MR87, MR88, Ros87, Smi83c, Smi87a, Smi87b, Smi88a, Smi88b, Ull85, Ull86].

9./10 [Wös82]. 9.95 [Cou85b, Nad86].

= [hC83, KS81c, sKcH81, MF84].

A-series [Bur85b]. A.G.A. [RH84b]. A.N.S.I. [Lav83]. A1 [Bre78b, Bre79]. abend [Bin85]. above [Dun88b]. abridgement [Miy87]. absorption [ZGS89]. abstract [BCF^{+}88]. Abstraction [Isn82, MMS88]. abstractions [Sus86]. Academic [Wil87b]. academy [JBT83]. Accelerated [GKRY82, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d]. Accelerator [AB89, GKKY89]. Access [BRK^{+}87a, BRK^{+}87b, BRK^{+}88, BDS88b, BDS88a, KW89, Sch88a, BK89, BDS88c]. accessible [Pay84a, Pay84b]. Accessing [Hay86]. Accompany [CM84b, APD86, BW84, BW87b, BGG85b, CM81a, DH82, DP84c, Ett83a, Gri85, HB81, JSW85b, MSR87, McC84c, MS88b, Moo82a, Nic85c, PDA83b, RZ89b, RWA84, Rod84, WAD81, Wei86a, WB85]. According [BBB^{+}83, Vu 89]. Accuracy [Don82a, Don82b, JW86]. acid [LB86]. ACM [Gon89, POP82, Van84a, Wex81, Ham85, HM90, RH84a]. ACM-SIGSAM [Gon89, ACM89a]. ACRITH [BRK^{+}87a, BRK^{+}87b, BRK^{+}88, JW86, KW89]. Across [BSP83]. activation [YHKM89]. Activities [Noh84]. ADA [Bro89d, Gra88, Rel89c, Rel89a, Wil87a, Boo81, Bro89a, Bro89b, Bro89c, DAG^{+}88, DM89a, FSO89, Lev89, Mor81a, PBB^{+}88, Rud83, Sch82a, SW83, Whi81a, Whi81b]. Adam [La 87]. adapt [Jac85b, Jac85a]. adaptation [MW84a, MW84b, TMjC81, Tew81]. adapted [Kir89]. Adapting [Gro87]. Adaptive [Ash85a, GKRY82, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Ash85b, Ash85c]. Adding [MMS88, Owe86]. Addison [Cou85a, Rid82a]. Addison-Wesley [Cou85a, Rid82a]. Addressing [DD86, Hol87a, Hol87b]. adjustments [Don83c]. administrator [YS84a, YS84c]. adolescents [BS86b]. ADS [Van84c, Van85]. Advanced [Edg89a, Edg89b, LOU86, PW86, BS88c, DSCP88, Int86c, IBM89b, Mic89b]. Adventures [JBT83]. aerial [LZ82]. aerodynamics [Dun85a]. aeromagnetic [Gra86b]. AFNOR [Ass83b]. AFOS [Mac81]. after [AS89a, AS89b]. Afterword [Bac84a]. Aftran [CRV^{+}89, Vag89]. age [Mar81]. agricultural [Don83c]. AHEAD [AEV89, dEV89]. AI [Ano85a]. Aid [BDS88b, BDS88a, CSD83, DBFK89, BDS88c, BDS89, KK89b]. Aided [Rao86a, Rao86b]. air [Wal85]. aircraft [ADP88, DPA87, EIT85, Kip82]. airfoil [DS82, kK89c]. airfoils [HL82a]. Airy [SJB83a, SJB83b]. AIRYROOT [SJB83a, SJB83b]. AIX [Int88c, Int88a, Int88c, IBM89b]. al [CS84, CM89, FK84, HRH81, HRC89, aHH83, McC85c]. al-barmajah [aHH83, McC85c]. al-Fortran [McC85c]. al-Musaid [aHH83]. Alamos [ZDS81a]. Algebra [Ber87, BD89, Dav89, DG82, Dod83, Don83a, DDHH84, DCHH87, DCHH88a, DCHH88b, DCHH88c, HK87a, HK87b, LHKK79a, LHKK79b, LN88b, LN88a, Pra89, DDDG89, DCHH85, LN87, Pra89]. Algebraic [ACM89a, GKKY89, Gia89, Gon89]. Algebraic-Numeric [GKKY89]. Algol [Baj81, Fog85, Fog87, Fog88, Osi82a, Res86, Sch82a, Smi81, Smi83a, DM85, Osi82b, SMD84]. ALGOL-60 [Fog85, Fog87, Fog88, Res86]. Algorithm [Abd80, Amo83a, Amo83b, Ash85a, AFS94, Ban78, BS81a, BS81b, BA85b, BA85c, Bre78b, Buc81a, Buc81b, Cas89a, CY89, CHPS85, Cod88, CGM84b, CGM85b, DFK83a, DM89a, Don82a, Don82b, DE84, DCHH88a, Dur80, EK87b, Gaf83b, Gaf83c, GGLM88, GKRY82, HK87a, HPB84, Hig88b, HMR85a, HMR85b, IZP81, JBJ84, Kah80, KRYG82b, KRYG82c, LHKK79a, LN88b, MP86a, MP86c, MC80, MGH81b, RH84a, Sau83b, Sau83c, Sch88a, She78, SSS84a, SSS84b, Sim76, Ste76, Stu81a, SS79, VVV89b, Wil83, Zak84, Ash85b, Ash85c, Bai87, BHK^{+}85, Dun87b, FS86a, Gen82, IA89, Kem87, Lec89, LtW88, LN87, McD89, Min88, PC89, PW84, Son83, The88, Tod85, Wit81, Bre79, Buc82, CL83, Cra86a, DM87b, DFK88, DG82, Dod83, FW82, Fut78, GL90, Ham85, HK87a, HK87b, Hig89, Hig91, HM90, LK88]. Algorithm [LN88b, LN88a, OM82, PCK84, Sch88a, Sim76, Ste79, Stu81a, Van82, WH87]. Algorithmen [San82]. Algorithms [Ana87, Cal86, DM85, Don83a, DBFK89, Gea86b, Gea86a, Ham85, HM90, JGD87, Lau86, Leh86, OM82, RH84a, Stu81a, Car89a, Car88b, vC87, FJS85, GQ88, Hil82a, Hil82b, Joh86, Jor86, MMM85, RS87, Swa84, TMS88a, TMS88b, Wee86]. Aliasing [Ana87]. Alliant [Cod86a, WSL88]. Allocation [Gol84, BCKT89, Dob85]. Almost [DFK83a, DFK83b, DFK88, DFK81]. along [EIT85]. alpha [SDH84]. Alternate [CM81a, DFK83a, DFK83b, DFK88, Wil84, DFK81, Ful87]. alternative [Tay84]. Alternatives [MB81]. Amdahl [CB82]. amelioration [Gra88]. American [Com89, DH82, DH83, DH84a]. amounts [McK83]. analitik [Osi82a, Osi82b]. analog [ZGK88]. analyse [O’R81, (??84, Tri84, (??87, Tri89]. analyses [FMH85]. Analysing [HO89, CSC^{+}86]. Analysis [ACK86a, AL81b, Ban88, BSP83, BDS88b, BDS88a, Chi86a, DAG^{+}88, Ell82c, Eve85, Fen87a, FN85, GBJ81, Har89, IA84, KKRK85, LK88, MSA86, OK87, RT85, Spa85b, Spa85a, Tsu85, AI88, ACK86b, ABC^{+}88, And89, AL81a, ALPC88, AEL^{+}86, AM89b, BZ85, Bli89, BDS88c, CK86a, CB86, CDHP86, CKT85, CK88, DS82, EH81, Ell82d, Fen87b, Fri84, HL82a, Hua82, IMS87a, IMS87d, IMS87l, IMS87m, IMS87k, IMS89a, IMS89f, IMS89g, Lib89c, IMS89m, IMS89n, JT88, Kem87, LB86, LN89, LKM88, Mal85, Mat83a, Mat83b, McD89, Mul88, MT84b, Ngu81, Pay84a, Pay84b, PS84, RS89, She89b, Shi88, Smi87a, Smi87b, Smi88b, Son83, Tan86, Tan88, Tan81a, Tan82, TPR85, Tho82b, Thu86, Wal3 , Wan86, WM85a, vM84a, vM84b]. Analytic [EW87]. Analytical [MHK86, Hue83, KWWK86]. analytically [Bar89]. Analyzer [Fed82b, Ell81b]. Analyzing [DS86b, DS87b, DS86a]. ancestral [Vu 89]. Anecdotes [Tro84]. anemometry [Owe87]. Angeles [Wex81]. angepaßtes [HS81]. angles [The88]. angular [CDW82, CDW84, RV84, RRC89, SDH84]. Ann [IEE81]. Anniversary [AW82]. Annotated [Lee84b]. Annual [POP82]. ANSI [AC87, Ano82a, BP81a, KK89b, Ame85d, Ame87b, Ano89, BP81b, CwL83, Col83, Col87a, Ehi82, Kha81, Law83, NJLB81, mT82b, Wie86a, Wie86b, Wu82a, Wu82b, Wu83]. Ansprache [Kna84]. Answer [Pag87]. ANSYS [HMB^{+}88]. antennas [Ric82b]. Antennenfeldern [Sch84f]. Anweisungen [Ano82a, Ehi82]. anwendungsorientierte [NEM84]. AOS [Dat84, Dat85a, Kir89]. AOS/VS [Dat84, Kir89]. apertures [Bai89]. APL [Ber82a, Chi85b, Chi86a, DO86, Fog85, Fog87, Fog88, Kar87a, Kar87b, Res86, Smi81, Smi83a]. APL/370 [Chi85b, Chi86a]. aplicaciones [Bor89]. APPLE [cC84b, BB82, BBuC84, DB82a, Gee86, Gre85, HF81, Kri83, hK85, Pre89, VN89b]. Application [Bin85, Int88b, IEC85, Mar84b, Ull85, Ull86, Vor89, Int81a, Int81b, Int81c, Int81d, Int82g, Int82h, Int83d, Int83e, Int83f, Int83g, IBM88, Int88d, She89b, Wyl86, ZGK88, BBB^{+}83]. Applications [AM89b, Cor88, DR86, Gea86b, Gea86a, Hig88b, Hig88a, Hig89, Hwa84, IBM86, IEE88a, Jur86, Kow85, LCMM88, Lin83, Mil87a, Moo82b, Mos88, Nor84, Ril83, See81b, Wu83, BK84, Bel84, Car89a, EP81b, Gro89, IMS87a, IMS87b, IMS87g, IMS87e, IMS87f, IMS87n, IMS89a, IMS89b, IMS89c, Lib89a, IMS89h, IMS89i, IMS89j, IMS89k, JS88, LPJ83, LP87, Mil87b, Moo82a, Moo86, Moo88a, RZ89a, RZ89b, See81a, VV86, Wu82a, Wu82b]. Applied [AM81, AM84, Boy84b, JSW85a, Lew81b, Sch81b, SB82, Sch85b, SB86, BR89a, JSW85b, Lew81a, Hue83]. appliquées [GB83, GB89]. Applying [Mar84a, Not89]. Apprendre [Del82, Del83, Del85]. Apprentice [Wor84a, Wor84b]. Approach [BCM87, BS81e, Col83, Col87a, CSD83, CM83a, CM87, Ell83, Fen87a, Fis83, Nan81b, Nic85b, Rul83a, Rul83b, Smi85e, WN87, BS81f, CK86a, CC84a, CwL83, Cou85a, CM83b, DKG89a, DKG89b, Ell82e, Fen87b, JT88, JBT83, Key81, Kha81, KS82b, KM83, Leh83b, LH81, MP81, MSR87, Nan81a, Nic85a, Nic85c, Sas83a, Sas83b, Smi85d, SR87, Mil82b]. Approaches [GL86]. approche [Str82, Str85]. approximating [Dun85a]. Approximation [HYP87, Uni84d, Dun80, DZ87, Dun87a, DZ88, Dun88a, Dun88b]. approximations [Gen82]. AQUIFER [VMS81, Cza83]. AQUISM [VMS81]. Arbitrarily [Ren84, Sav87]. Arbitrary [YF85]. Arbor [IEE81]. ARCECO [DFK83a, DFK88]. archetype [Gra88]. Architectural [SAB88]. Architecture [Par86a, Wat87, CGMW88, IBM88, Par86b]. Architectures [CT86b, CT86a, DSCP88]. archive [Kle89b, Kle89a]. ARCPRINT [Kle89b, Kle89a]. Ardent [Cod89]. Area [YF85, Boy84a, She89b]. areas [BT83]. Argonne [BH89]. Argonne/GMD [BH89]. Arguments [GF81, FCG83]. Arithmetic [BBG^{+}82, BBB^{+}83, BBG^{+}84, Bre78b, Bre78a, Bre79, BHY80, BW89, CHH81, CHH83, Cor88, IEE81, JW86, Sch81a, Ull85, Ull86, Vel82, WLO76, Cor82a, Cor83, Hym82]. Array [Bro81a, PS81, AKLS88, Bai89, Bro82a, Hsi83, RW85, The88, Vag89, CRV^{+}89]. array-fed [Hsi83]. arrays [EL81, GH87]. arrival [The88]. Art [Num88d, PFTV86, WB89, Pre88b]. Artificial [Gus84, IEE88a]. Artin [Gra86a]. ASCII [Ano82b, Ano85c, Spe82, Spe85, SDC82, Uni84d, Uni84e]. askfor [LO85b, LO85a]. Aspects [Wri89]. Assembleur [Ber82a]. assembly [Joh86, Bin85, Egg83]. Assessing [EP87a, EP89]. Assessment [Sim86, Boy84a, Sim85, Sim88b]. Assessment/F2812 [Sim86]. Assignment [BD80, Bur81a, NJLB81, PC89]. assignments [SDH84]. assistance [BD80]. assistant [Smi88d]. assisted [Le83, Lee85]. Associated [Chu88, Bod87, FDL86, PRWB89]. Association [Sof83b, Sof84, Usr82]. Astronomical [Gui89]. Asymmetric [HPB82]. Asynchronous [Wri89]. atmospheric [EIT85]. ATOMCC [Cha86b]. ATS [Seq89]. attached [Iwa84]. attraction [SJB83a, SJB83b]. Aufstellung [RAKK88]. Auftragsabwicklung [Ste87]. Aufwaertsuebersetzung [CLW81]. Aug [Rei89c]. AUGMENT [BHY80]. August [MSM84, VV86, Wri89, Rei87b, Smi83b]. Auswertung [Lep86]. automata [KS81c]. Automate [Ano85a]. Automated [Aha85a, Aha85d, Gab89, GBJ81, Sch89e, Var85, Van84c, Van85]. Automaten [KS81c]. Automatic [AK84, AK87, Arn82, BCF^{+}88, Gro89, Har89, HKP88, LG86, Maa89, MT82a, PBB^{+}88, PB86, Pol87, RW89, SR88, Tan86, Tan88, Thu86, TFI86, vvHG87a, vvHG87b, Bat85, Bli89, CGQS89, Gat85, Gin82, KS88, LCH^{+}88, Pet89, Sis85, SP87, Wan86, Wee89, ZBG88, vM84a]. Automatically [NE89]. AUTOSCAN [PRL^{+}85]. Availability [Tsu85, Joh85b]. available [BT83]. Avionics [Mar84b]. avoids [Mei89d]. awareness [Owe86]. axial [Tho81, Tho82a]. axially [Per81]. azimuthal [The88].

B [NL85a, Bur81b, Bur84a, Con81b, NL85b, PF85, Win85]. B5000 [Bur84b]. B5000/B6000/B7000 [Bur84b]. B6000 [Bur84b]. B7000 [Bur84b]. backward [Bal84]. backward-directed [Bal84]. balance [Lam89, Cod86b]. Balfour [RB82]. Ballistics [Rud83]. Band [EL81]. Banded [Sch88a]. bands [MW84a, MW84b]. barbara [SMD84]. barmajah [aHH83, McC85c]. Barrier [LO85a, LO85b]. Barron [LP85a]. basado [DH84a]. Base [Gre81, Ano81a, BNZ87, Fri84, Rin83]. Based [BCM87, Bos88, GGLM88, GL90, Miz83, PP82a, Cal86, Cre89, DH82, DH83, EGP81, Gui88, Hur82, Isn82, KS81b, Kip82, Lee85, PP82b, Roc86, Sho85, Wat82a, Wat82b, ZGK88, KS81a, RMFG85]. BASIC [RWA84, Rod84, WAD81, AB83, DDHH84, DCHH87, DCHH88b, DCHH88c, HK87a, HK87b, Kur85, DCHH85, Key81, LN87, Bro81b, CM81a, CM81b, CM84a, CM89, Ell82a, Ell82b, Fog85, Fog87, Fog88, Hua82, Res86, Rat89, Rel89c, Rel89a, Rao81a, Rao83, SS84, TS88, WG84, jYsS89, Ber82a, Bru84, Cas81, CM84b, DG82, Dod83, Don81, DCHH88a, mK84, LHKK79a, LHKK79b, LN88b, LN88a, Rao86a, Rao86b, Wol85, RWA84, Rod84]. BASIC/FORTRAN [jYsS89]. basis [BHK^{+}85, JC82, Hel85b]. Bayesian [TSU88]. Bazisnyj [Bru82]. bcc [DM89b]. BDMG [Som86]. be [Per81]. beam [MT84b, NE81, Tho84a, Tho84c, Tho84e, ZGS89]. bedload [Tat87]. beef [PMBK82a, PMBK82b]. Beginner [Phi87, Phi86]. Beginners [Gus84, Lam84, Bal85]. Beginning [Wor84a, Wor84b, YS84a, YS84b, YS84c, YS84d]. Beginning/Book [Mic84f]. Begriffe [Ano82a, Ehi82]. BEHAVIOR [VMS81]. Behavioral [GGJ^{+}89]. beings [Bro83b].

Belady [Ols83]. Benchmark [Don88a, Kaw84, Pri89, RS89]. Benchmarking [FSO89, SS88]. benchmarks [BW87a, Lan88]. Bending [GKKY89]. benefits [Wic89]. Benjamin [Cou85b]. Benjamin/Cummings [Cou85b]. Benutzung [Kal85a]. Benutzungsanleitung [Abs88]. Berechnung [Sch84f]. Bericht [CLW81, MF84]. BESM [SM84, SK83]. BESM-6 [SM84, SK83]. Bessel [Col84b]. Best [HYP87, Don83c]. Bestimmung [Sch82c]. Between [HG82a, Can81, Fun86]. beweisbare [Sto84a, Sto84b]. bi [McC85c]. biased [DZ87, DZ88, Dun88a]. Bibliography [Lee84b]. Bibliothèque [(??84, Tri84, (??87, Tri89]. Bickley [Amo83a]. Bicubic [HMR85a, HMR85b]. big [KS81b, KS81a, Wag85]. big-float [KS81b, KS81a]. Bildverarbeitung [NEM84]. Binder [Hei83]. binding [Ame85a, Gra85, Ame85b]. bindings [IEC88, ISO88, Int88q, Int88r]. biologov [BKL89]. biomass [SC83]. biomedical [BR89a]. biomolecule [FDL86]. biostratigraphic [Hel83]. birth [EML88]. Bit [Kal85a, Hay86, Rob82]. bivariate [Bau88, TSU88, Wal3 ]. blade [Dul81, Red86, Tho81, Tho82a]. blade-to-blade [Dul81]. blades [HL82a]. BLAS [Aha85c, HK87a, HK87b]. Blending [Sch83b, Sch83a]. Bloch [ZGS89]. Block [Cal86, DFK83a, DFK83b, DFK88, DFK81, Joh87b, Joh87c]. Block-oriented [Cal86]. BMDP [FE82]. BNDE [TSU88]. board [ZGK88]. boiling [Dav86]. boiling-point [Dav86]. Bonniger [BD80]. Book [Bis81, Cou85a, Cou85b, Edm86, Eve85, Mil82b, Nad86, RB82, Rit89, VTPF87, Wil87b, Tha89c, Vet85, VTP89, Mic84f]. Books [Rad81, RB82]. Boston [Rei89c, Usr82]. both [Hua82, Nai86]. boundaries [Boy84a]. boundary [BMS84, EH81, MH82]. Bounds [HL82c].

Brace [KC84b]. Branches [EBS88]. Brent [BHY80]. BRENTM [MC80]. Bridge [Chu88]. BRK45 [Cas89b]. BRKF45 [Cas89a, Hig91]. brochure [(??84, Tri89]. Brownian [Chu88]. BS2000 [Abs88]. BTREE [Win85]. Buckling [RT85]. Buffered [HPB84, WH87]. Buffering [Sor84]. builders [IA89]. Building [Mar82b, And84a]. bulk [SG88]. bulletin [Int85i]. BUMP [SD89]. Business [Hon81a, Kha81, LH81, Lin83, Mil82b, Pad85, See81b, Wu83, HB81, MP81, See81a, Wu82a, Wu82b]. Butterfly [HO89]. Byte [Mas83, Mas87, Per83a, Sho85]. Byte-Wise [Per83a, Sho85].

C [Mil82b, ADH^{+}89, AJ88, All90, Bai86, Bro84a, Cul88, Dig85e, Flo89, Fog85, Fog87, Fog88, GMW86, Goo89, Lag85, LS85, McG87a, McG87b, MT84b, Pre88b, Res86, Rel88, Rel89c, Rel89b, Rel89d, Sch89e, Tho86b, WSL88]. C-1 [WSL88]. C-to-Fortran [Flo89]. C240 [CGMW88]. C5 [MGH81b]. C6201 [Jap82]. CA [Wri89]. cables [Iwa84]. CADAC [CHH81, CHH83]. CAI [Sul2 ]. calculate [Bro85, Bru86, McC86, PB84, SP84b, SP84c, You82]. calculated [SG88]. Calculating [FW82, Ste76, Bod87, CDW82, CDW84, Dal89, MW84a, MW84b, RO85, RO86, SJB83a, SJB83b, Tho81, Tho82a, YK88]. Calculation [EBS88, Sau83b, Sau83c, Sau83a, ISJ85, Kem85, Kie86, Lam89, Nai84, Nai86, Owe87, She89b, Sho85, Tho84b, Tho84c, Uni81a]. Calculations [GKKY89, Per83a, Bar89, She89b, WP84]. Calculators [BL83]. calculo [Bor89]. Calculus [Tha89a, Tha89b, Tha89a, Tha89b]. calendar [Jan84]. CALFLT [LZ82]. calibration [Lam89]. Calif. [Wex81]. Callable [Uni85a, GC84]. Calls [TFI86]. CAMAC [IEC81]. Can [Mor81a, Ber85b].

Canada [Sof83b, Van84a]. Canadian [ZSD82b]. Canonical [DNV81, HPB82]. CANSCAN [JC82]. Capsule [Smi84]. card [Per81, Wal81]. card-imaged [Per81]. Carlo [Joh85b, NE81, VH87]. carotene [Wyl86]. CAS22 [DS82]. CAS2D [Dul81]. cascade [Dul81, HL82a]. cascades [DS82]. Case [DM87c, GGJ^{+}89, Wie82, Sch89f]. Cases [Car89b]. casual [Pag87]. catchment [War86]. catchwater [War86]. Categorical [Stu81a, Stu81b]. caused [Owe87]. CDC [Ano82a, Con83a, Con85a, Ehi82, mT82b]. Cedar [Guz87, Guz88, GPHL88, GPHL90]. Celebration [Noh84]. cell [GDK89]. CEMCORP [Dir84]. censored [CCHT89]. Center [Cha87b, KM83]. Certain [DFK83a, DFK83b, DFK88, Bow82]. certains [Tel82]. Certification [Ste79]. certifying [Sch82b]. CFT [Cra83, Cra84, Cra86b, Cra89a]. CFTP [KK89b]. chain [Law88]. chaining [hTDT88]. chaksong [hA84]. chang [hH82]. changes [AW89, Owe87, Tho84c]. channel [SDH84]. channel-spin-1 [SDH84]. channel-spin-1/2 [SDH84]. Channels [PLR85]. Chapel [BBB^{+}83]. Chapter [HL86a, HL86b, Som86, GtTB89, Hof87, HP89, Lag85, Lio85, Mul85, PF85, dR87, dZ86]. Character [Bee85b, Hon82]. Characteristics [JGD87]. characterization [GGJ^{+}89]. Charles [Rit89]. chart [jYsS89]. Charts [Coc83, Sis85]. Cheap [Wei84]. Chebychev [Fut78, Sim76]. CHEBYCODE [Ash85a, Ash85b, Ash85c]. Chebyshev [Ash85a, Ash85b, Ash85c, Dun80, Dun87a]. check [HM82]. chemical [BDS84, Hue83, KJM89, LKM88, Rap82b, Rap82c]. Chemistry [Ano88b, Jur86]. CHEMKIN [KJM89]. Cheng [sKcH81, hHtM81, Cha83, mCaLjH84, CwL83, hH82, kH84, LcY83, mM84, RkC84, fTBcL7 , mT82b, yW85, fY84]. Chi [mM84, mCaLjH84, CwL83, hHtM81, hH82, kH84, LcY83, RkC84, yW85, fY84]. ch’iao [hHtM81]. Chichester [Edm86]. Chieh [mT82b, fY84, cC84b, cT81]. children [BS86b]. Chinese [GQ88]. Chino [RL81]. Chislennyĭanaliz [BZ85]. choice [DV81]. Chonja [mK84]. Choosing [MT82a]. chuan [hC83, fTBcL7 , mT82b]. chungsim [hA84]. ciencias [Bor89]. CINDY [SR84a]. circle [Tho84e]. circular [Bai89, PM87]. City [Sof84]. Clarendon [Nad86]. clarity [Boo82]. Class [EBS88]. Classical [Sme81, CT88, SS82]. classification [PF85]. clay [Bod87]. CLAYFORM [Bod87]. Clean [CHH81]. Clinical [GBJ81, WM85a]. cliques [vC87]. Closed [Köl82]. Clunks [HG83]. Cluster [Spa85b, Spa85a, Ano87c, Red82, Eve85]. CMS [Ger83, Ber89, DW83b, Int82b, Int82c, Int82d, Int82e, Uni83c, Uni84c, Uni86b]. CMSP [JSW85b]. co [Kar87a, Kar87b]. co-operation [Kar87a, Kar87b]. coal [Lag85, Mul85, PF85]. COBOL [RWA84, Rod84, WAD81, Kur85, hA84, Bin85, Can81, CM81a, CM81b, CM84a, CM89, Fog85, Fog87, Fog88, Int82f, Int88b, IBM88, Int88d, Pay84a, Pay84b, Res86, Rat89, Rel89c, Rel89a, VD84, Ber82a, Cas81, CM84b, Don81, mK84]. CODASYL [Fri84, HT82, Jai84, GE85]. Code [Bro82b, Chi86a, HG82b, KK89a, LG86, PB86, SR84b, TvSS82, Vel82, Bat85, BSdlT87, Col89a, Col89b, CM86, DM89b, Gat85, GDK89, Gro87, Hig86, KWM88, Kee88a, KJM89, KK89b, LR89, MH82, Mye83a, Mye83b, Owe87, Sch84a, Sis85, SR88, Tan86, Tan88, Wee89, WW89, Win85, vM84a, Pem83]. Coded [Air77, vvHG87a, vvHG87b]. Codeerzeugung [Sto84a, Sto84b]. Codes [Hig88b, Hig88a, Hig89, Sme81, vdV85b, vdV85a, BT88, Bod87, RAKK88, Sim88a, SS82, RAKK88]. Codeview [Mic87e]. codification [FDL86]. Coding [Leh86, Ass82, Ass83a, Cor82b, McA86, Riz85]. coefficient [AM89a, Fra84a, Kem85, RRC89]. coefficients [AD84, ISJ85, Iwa84, KWM88, Lam89, RV84, Vu 89, Zoh84]. coherent [SDH84]. Cohort [CDHP86]. Coleman [Rit89]. collect [vM84c]. Collected [Ham85, HM90, RH84a, EIT85]. Collection [Chr84, DH84b, Hor83b, LB86]. College [KC84b]. collisions [CT88]. Collocation [HMR85a, HMR85b]. color [Gri85]. Colorado [RH84b]. Coloring [BCKT89]. COLROW [DFK83a, DFK88]. Column [DFK83a, DFK83b, DFK88, DFK81]. Combat [Kip82]. Combination [Cra86a]. combinations [Iwa84, Wal85]. Combinatorial [Lau86]. Combined [Bee81, Hur82]. Combining [Wan85]. command [Wal81]. Comments [Bow82, KTW84, RL81, Rei89a, Rei89b]. comminution [Mul85]. common [Pra89]. commutators [DFD81, DFD84]. commute [DFD81, DFD84]. compact [PW84]. Companion [Lev89, Rel88, Rat89, Rel89c, Rel89a, Rel89b, Rel89d]. Company [Cou85b]. Comparative [vdV85b, vdV85a, LCH^{+}88]. compared [Sch82a]. Comparèes [Don84c]. comparing [Fog85, Fog87, Fog88, Res86]. Comparison [Bla87, NE89, SAN^{+}81, AGS88, GG88, Joh86, KB88, Lee84a]. comparisons [Chi85b, FDL86, LR89]. compatibility [Dig85a]. compatible [Cha87b, DH82, DH83, DH84a, Int86b, KS82b, Kri83, LB86]. compiladores [DH84a]. compilation [BK88]. Compile [Hol87a, Hol87b]. Compile-Time [Hol87a, Hol87b]. Compiler [AU77, ACK86a, BGS82, Chi86a, FW83a, Mar82b, Mer88b, Mic81, PW86, Par86a, RV8 , ACM82, ACM84, Van84a, All82, ACK86b, Ano82d, Ano89, BW87a, Bro82b, Chi85b, Ell82c, Ell82d, Fed82a, FW83b, Int82a, Int82i, Int83h, Int84c, Int84d, IBM85, Int85c, Int85d, Int85e, Int86d, MI82a, MI82b, MI84, Mic84c, MZ84, Mic84b, Mic85c, Mic85b, Mic85a, Mic86, Mic87d, Mic87f, Mic87e, Mic87a, Mic87b, Mic87c, Mic88, Mic89c, Owe86, Par86b, RG85, SK86, Tou84, Wie82]. Compilers [ASU86, All84, DDH84, NE89, Dha88, Bur87, CDL88, CM81c, DH82, DH83, JH86, Pay84a, Pay84b, Uni82c, WM85b]. Compiling [AKLS88, AJ88, DO86]. complete [HW86]. Complex [Hig88b, Hig88a, Hig89, SH88]. complexation [She89b]. Complexity [RS89, Tan81b, Wit81]. component [Int82e, Lam89]. composite [Son83]. Compositional [Moo86]. compositor [HR83b, HR84a]. Compound [HKP88, SMD84, SR84a, SDH84]. compound-nuclear [SDH84]. Comprehensive [Jai84, AEL^{+}86, BR89a]. Compressed [PLR85]. compressible [EH81]. compressor [Tho81, Tho82a]. Computation [ACM89a, BL83, BRK^{+}87a, BRK^{+}87b, BRK^{+}88, BKK^{+}81, BBB^{+}83, DR86, Gia89, Gon89, JSW85a, Kow85, KW89, Pra85, Ric84, SMD84, SR84a, VV86, Vor89, Wri89, BSdlT87, CV88a, HL82a, JSW85b, Kul83, KM83, KS88, MRS84, Mar81, Mor81a, Rap82a, SC83, VC89, Zoh84]. Computational [AB83, Ano88a, Ano88b, KM89, Wan85, DGNP88a, DGNP88b, GG88, Wee86]. Computations [CV88b, GF89, Her88, Rit89]. compute [BHK^{+}85, Hsi83, Rap82b, Rap82c, Vu 89]. Computed [Don82a, Don82b]. Computer [Ame85d, AS88a, AS88b, AKLS88, AG87c, Bel89, BB82, BM81, Cou85a, DB82a, Dav89, Gea86b, Gea86a, GL81, Gro83, HRC87, HL82b, IEC88, IEE81, IEE88a, ISO88, Jur86, LP85b, Mar81, Mor82, Rad81, Rad83b, Rad83a, Rao81b, Rao82, Rao86a, Rao86b, RG85, RB83, Rou83, Sch81a, Sim86, Tea81, WP84, Ame85c, Ame85a, AG87a, AG87b, BNZ87, Bod87, BT83, Boy84a, BS83, Cla89, Col82, CDHP86, Cra83, Cza83, DSCP88, Dun85b, Eva81, Fog85, Fog87, Fog88, GF89, Gra81a, Gui87, Hew85, Hel83, HMB^{+}88, HF81, Hsi83, Int88q, IS84a, IS84b, Ame85b, Joh86, Jus88, KWM88, Kee88a, Key81, mK84, Kir85, LZ82, Le83, LD87, Lee85, MF84, McC81, McG87b, McM86, Rod87, Min88, Mul88, O’R81, Res86, Rao81a, Rao83, RO85, RO86, Ril83, Rin83, Rub83, Sel83, Sav87, Sim85, Sim88b]. computer [Int88r, Tou84, VSH83, VH87, Wie86a, Wie86b, ZDS81a, ZDS81b, ZSD82a, ZSD82b, BW87a, CDW83a, CDW83b, LP85a, McG87a]. Computerized [Stu81b]. Computerpraxis [McC85a, McC85b]. Computers [BL83, BD89, Cas81, Don81, Don84a, Don84d, Don85a, DD86, Don87b, Don88b, Don89, EBS88, Pol83, vdV86, CT88, CMM^{+}88, DDDG89, Don83b, Don84b, Don85b, Don87a, Don88c, Hud88a, Hud88b, Kir89, Kri83, Lec89, McC86, Num84d, PW84, Roc86, RS87, Sch89a, Sho85, Swa84, van86, vM84a]. Computing [AS88b, Bem84, Bow81, Chu88, DM87b, DM87a, Dyc81, Gaf83b, Gaf83c, Gaf83a, McC84a, McC84c, MS88c, MS88b, MBP^{+}85b, Moo88b, MS83, Nan81a, Nan81b, PCK84, PFTV86, Van82, VVV89b, VVV89a, ??88, AM89a, AS88a, AM89b, BS86b, BDJ^{+}89, Hue83, Iwa84, JR81, KP86, McK83, MB81, MBP^{+}85a, Pre88b, RS81, RS84, Sav87, Sun88a, TSU88, Wat82a, Wat82b, Num88d]. con [Bor89, CM89, DH84a, FK84, Zwa85]. concept [DS82, AW82]. Concepts [GJ82]. Concerning [Mei87]. Concurrent [LH88, Wee86, Ker82]. Condition [Hig88b, Hig88a, Hig89]. Conditional [EBS88]. conditions [BMS84, RMS82]. conductivities [KWM88]. Conference [Ano88a, Dav89, Ass86, IEE88a, LCMM88, Mor82, RH84b, POP82, Sof83b, Sof84, Usr82, Wri89, Wex81]. configuration [CDW82, CDW84, Dav86]. Confined [MHK86]. Confluent [GN89]. conformational [She89b]. Congress [VV86]. congruential [Gui88]. Conjugate [DM89a, RV89, vdV86, LN89, van86]. Conjugate-gradient [RV89, LN89]. Connection [Flo89, AKLS88, SS88]. CONPAR [JR89]. Conserving [PFF83]. Considerations [PS81, RG85]. Constant [CCKT86]. constants [Bod87, BDS84]. Constrained [GHM^{+}86, Sch86]. Constraints [HG82b, Zho89]. Construct [Stu81a]. Construction [Ano87e, ACM82, ACM84, Sch87a, GQ88, Stu81b, Van84a, WF85]. Constructive [Boe87]. constructs [Gal89]. contained [AI88, IA84]. containing [Jac85b, Jac85a]. contaminant [KWWK86]. contamination [PMBK82a, PMBK82b]. Contemporary [BKK^{+}81, Rey80]. content [BMS84, RMS82]. Contingency [MP86a, MP86c, MP86b]. continuation [Gra84b]. continuing [Smi83c]. continuous [CSC^{+}86, Hur82]. continuous/discrete [Hur82]. Contour [Coc83]. Contributions [Sam81, Aha85d]. Control [Hon81b, Sch83b, Sch83a, Shi88, SP82, WD81a, WD81b, CGQS89, CR84, CSD82, Fra84b, FJS85, HS83, HM82, Hig86, IEC85, Rom81]. Controlled [CHH81, CHH83]. Controlled-Precision [CHH83]. convergent [Bar89, ZGS89]. Conversion [Bun86, Hus84, SW83, Aha85a, Aha85d, Hey85, Roc86, Sis85, Tei86, Uni82a]. Convert [AK82, AK81, Cre89]. Converter [Sal84]. converting [Gro89]. Convex [GHM^{+}86, CGMW88, Mer88b, WSL88]. Convolution [AC86, Ess88]. Cookbook [Den84b, Den84a]. Cooperative [MSM84]. Core [BBF^{+}82, CY89]. Corporation [Fed81, Sof83a]. correction [Gra86b, Tho84b, Tho84d]. Corrections [Hop81, SP84b, SP84c]. Correctness [BM81]. correlating [PM87]. correlation [DFD81, DFD84, ISJ85, Kem85]. Correspondence [Hyb87, Pem83, Tan83b]. Corrigenda [DCHH88b, EP89]. Corrigendum [Dod83, Hig89]. Cost [Ric82a, BT88]. Couger [WAD81, RWA84, Rod84]. Count [Chr84]. counting [CB86, CSD82, Wei84]. counts [Dal88b]. Coupled [KBRM^{+}86, WN87, CDW82, CDW84]. coupling [Bai89]. cours [Ain89]. Course [Cal83, CM84b, Mon89b, Pre88c, Rat87, YS84a, Baj81, CM81a, CM81b, CM84a, EGP81, Gri85, HA83, HPR81, LD87, MSM84, MB81, Rat81, Rat86, RWA84, Rod84, Spe83, Tea81, WAD81, YS84c]. courses [Lee85, Ric82a]. Covariance [VVV89b, VVV89a]. covariates [CCHT89]. covering [YM85]. CP [Hon81b]. CP-6 [Hon81b]. cpu [Hei83, Hoc85, Lan88]. CPU-benchmarks [Lan88]. Crash [Mon89b, ALPC88, ALPC88]. CRAY [Hoc85, AGS88, DE84, RS85, Tem89a, vdV85b, vdV85a, Bai87, Cal86, EBS88, Fon85, Gue86, KK89b, Pet83, Sch89a, hTD88, Tem89b, Wee89, WW89, ZM86]. CRAY-1 [Tem89a, vdV85b, vdV85a, Cra83, Pet83, hTD88, Tem89b]. CRAY-2 [AGS88, Bai87, Cal86]. CRC [Per83a, Sho85]. CRC-16 [Sho85]. Creation [Col89a, Col89b, Mac81, Sis85]. criticism [BS81g]. cross [PB84, RS81, RS84, SMD84, SR84a]. cryogens [McC81]. Cryptography [Bur84c]. crystal [Joh81, Tho84e]. crystalline [SPS84, SS87a, Sav87]. crystallographic [MHS81]. Cubic [Car89b, Dur80, vM84f]. Cummings [Cou85b]. currencies [O’R81]. Current [EW87, BT83, JC82, Tat87, Tay84]. curvature [Wat82a, Wat82b]. Curve [Fut78, Sim76, Ano87a]. Curves [Pal86, SD89]. customers [Int88i, Int88j, Int89c]. customization [Int83i, Int85d, Int85f, Int88k, Int88l, IBM89a]. cutting [HM82]. CWG [Bai89]. CYBER [LW88b, Riz85, VH87, vdV85b, vdV85a, BK88, Con82a, Con83a, Con85a, Joh85a, LtW88, LN87, LN88b, LN88a, MW83, sT85, VSH83]. Cyber-205 [sT85]. CYBERPLUS [KRW88]. Cycle [CB86]. Cycle-counting [CB86]. cyclic [Gra86a]. cycliques [Gra86a]. cyclodextrins [She89b]. cyclomatic [Tan81b]. cyfrowych [Rzy84].

D [Cou85b, RB82, Lee84a, Kli89, Obl85, Tho84b, Tho84c, Wie86a, Wie86b]. D3 [Ban78, Ste79, SS79]. DAFOR [Ber87]. Dalil [McC85c]. DALL [IA89]. damit [KS81c]. Dannykh [MK86]. DAP [CHPS85, DM85, Gro87, Int81e, Sch84a]. DAP-Algol [DM85]. DAP-FORTRAN [Sch84a]. Darstellung [NEM84]. d’Artin [Gra86a]. Data [BSP83, Bla87, BP81a, CM84b, Ell82c, Gab89, GPKK82, GPKK84, Gre81, GBJ81, HK87a, HK87b, Hol87a, Hol87b, Isn82, JL81b, Pou87, Stu81a, Abe89, AI88, ACK86b, And84a, Ano81a, BK89, BNZ87, Bod87, BP81b, CCHT89, CM81a, CM81b, CM84a, Cza83, DGNP88a, DGNP88b, Ell82d, EIT85, Fri84, Fun86, GRB88, Gra84b, Gra86b, Gre88, HS86, Hig86, Hue83, IA84, Jac85b, Jac85a, JT88, JC82, JL81a, LZ82, Mal85, Mat83a, Mat83b, MH82, MHS81, MMS88, Ngu81, O’N81, Per81, RWA84, Rod84, Sal84, SM87, SDC82, Stu81b, Sus86, Tan81a, Tan82, Wal3 , WAD81, Wat82a, Wat82b, WD81a, WD81b, ZDS81a, ZDS81b, ZSD82a, ZSD82b, vMF81, vMF84, vM84d, vMT84, vM84b, vM84f, Kir89]. data-type [SM87]. databanks [FDL86]. Database [DNV81, Did86, GE85, Gol81, VLV^{+}86]. Databases [Gab89, Jai84]. date [Art81, Rin83]. datenkonverter [Fis82]. datos [CM89]. Davidons [IA89]. Davis [Dal89]. Day [AW82, Rin83]. Days [Bac84b]. DBMS [SAS86, SIR82a, SIR82b]. deadlock [Mei89d]. Debate [KTW84, KW84, McG84, RL81]. debug [Bur85b, Int82b, Int82c, Int82d, Int84d, Int85e, Int86e, Int87c, Int87g, Int88e, Int88m].

Debugger [GWM88, Ano89, Bur86b, Ell82a, Ell82b, Kie83]. Debugging [AP87, Utt89, Bin85]. DEC [Ano82a, Ehi82]. December [MSM84]. Decimal [CHH81, CHH83]. decks [BJ81a, BJ81b, BJ84a, BJ84b, Jac85b, Jac85a]. DECLIB [GC84]. decomposition [Her81]. deconvolution [Kem87]. decreasing [MT84a]. defects [Ott81]. define [BG84]. Definite [Cra86a, GL81]. Definition [HKP88, Con83c, Con84, Con85b, Con87a, Con87b, Con88a, Jia86, LB89]. Deflating [PCK84, Van82]. degenerate [MW84a, MW84b]. degree [Jes82, Rap82a, Rin83]. Demonstration [SAB88]. Denelcor [DH84b, Dun86, HLM84]. Dense [RRS88, VVV89b, VVV89a]. density [MT84a, SG88, TSU88, TFH86]. DENTS [CSC^{+}86]. Denver [RH84b]. Department [McA86]. Dependence [Ban88, CC87]. dependent [Joh85b, PMBK82a, PMBK82b]. depict [BT83]. depreciation [TMjC81, Tew81]. derivation [Fra84a, Tan86, Tan88]. Derivatives [Amo83b, KKRK85]. derive [ADP88, DPA87]. Describing [Boo81]. Description [Art81, Ame85c, Int84b, Int85b, Pyr84]. description/operations [Int84b, Int85b]. Descriptors [Hol87a, Hol87b]. Design [ADH^{+}89, AU77, BA86, DNV81, DBFK89, GMPW79, Gol82b, Goo89, HS81, Hwa84, Jai84, Law83, MSA86, PS81, Rao86a, Rao86b, RG85, Rud83, Sch83b, Sch83a, Tsu85, Boo81, Bur86b, CKT85, DS82, Ell81a, Ful87, Sel83, Sch82a, Van84c, Van85, WM85a]. Designing [Int86a, Rao81a, Rao83, Gra81a]. despiker [Eva81]. Detailed [BB82, DB82a]. detection [vC87, Whe84a, Whe84b]. determination [Gra86a, Tho84a, Tho84e, VH87, Wal85, Gra86a]. Determine [Cod88]. determining [Ack84, Bur86a, Rin83]. Developing [ASM89, DS86b, DS87b, HO89, DDDG89, DS86a, DS87a, Ear85]. Development [Cow84, Cre89, DM85, For85, Gro82, IBM86, JRS88a, JRS88b, McA86, Rod87, Mos88, PBB^{+}88, Zim86, Dha88, BK84, CKT85, Int88b, IBM88, Int88d, Jus88, Pay84a, Pay84b, RH84b, Sel83, Wie86a, Wie86b, Wyl86, Ano81a]. d’exemples [Str82, Str85]. DI-3000 [Wie86a, Wie86b]. diagnosis [Int82i, Int84c, Int87f]. Diagonal [CL83, DFK83a, DFK83b, DFK88, DFK81]. Dialects [GPHL90, Guz88, GPHL88, KB88]. diatomic [Nai84, Nai86]. Dictionary [Bro84b, Bro83a]. dies [Rao81a, Rao83]. DIET.FOR [Tho82b]. dietary [Tho82b]. dif [KW87b]. Diff [Tho84d]. difference [BMS84, GF89, SR88]. different [CSD82, Lan88, Sch89f]. Differential [Ber87, Cas89a, Cas89b, CC82, Gaf84, Hig91, LK88, MHK86, Ste79, SS79, Uni84e, Ada89, GtTB89, Pet89, SMD84, SR84a]. Differentiation [Cor88, Maa89, SP87]. differentiator [Hil82a, Hil82b]. Diffraction [JL81b, And89, JL81a, Tho84b, Tho84d, vM84f]. diffractometer [Tho84e]. diffusion [KWM88, vMF81, vMF84, vM84d]. Digital [JSW85a, Son83, Alb86, Boy84a, Gra81a, Gre88, JSW85b, Fed81]. Dimensional [PP82a, CM86, DM89b, EH81, GRB88, HS86, HK83, Kee88b, KWWK86, kK89c, PP82b, PRL^{+}85, Tho81, Tho82a, ZGS89]. dimensions [LW88b, Sun88a, VMS81]. DIN [Ano82a, Ehi82]. Dinic [GG88]. Direct [DD86, OM82]. directed [Bal84]. Directions [Sch88c]. directory [Int88i, Int88j, Int89c]. Dirichlet [Bow81]. disciplinado [DH84a]. Disciplined [DH88b, DH82, DH83]. discovery [BCF^{+}88]. discrepancy [Lec89]. Discrete [BCM87, Dur80, Fut78, Sch82d, Sim76, Dun80, DZ87, Hur82, Lee84a, MSG86, VMS81]. Discrete-Event [BCM87]. Discriminant [Stu81a, Stu81b]. discussion [Dol88, Rei84a]. disjoint [Per83b]. Disjunctive [WN87]. Disk [Pre87a, Pre88a, Sch88a]. diskette [Num85c, Num89, Pre87b, VN89a, VN89b, Mic84f]. Diskette/Self [Mic84f]. Diskette/Self-Teach [Mic84f]. dispersion [RAKK88]. Dispersionsgleichung [RAKK88]. dispersive [MT84b]. display [GRB88, HM82]. displaying [RO85, RO86]. Dissection [Eve85, Spa85b, Spa85a]. Distinguishing [Est82]. distorted [vM84e]. Distributed [Ren84, Tsu85, Per81]. Distributing [HG82a]. Distribution [RH84b, Art81]. distributions [Mal85, SDH84, YK88]. diversity [Bro85]. DL [Mor81b]. DLG [Gre88]. DLG-3 [Gre88]. DLG2ISM [Gre88]. dlia [BKL89, SM84, Sko88]. DM [Hon85]. DM-IV [Hon85]. DML [HT82]. DNA [AEL^{+}86, VLV^{+}86]. Do [FN85, Ber85b, CC87, CA86, CT86a, CT86b, LO85a, LO85b, OM82]. DO-loop [LO85a, LO85b]. DO-loops [CA86, OM82]. DOC4029 [Joh83]. DOC4029-192 [Joh83]. Docking [She89b]. document [Ame87d, Tri84]. Documentation [BA86, JR81, Dig84j, FE82, Mye83a, Mye83b, (??84, Tri89, Uni85a]. Does [Ros84]. Dokumentation [Sch84f]. domain [Dun87b, Dun85a, Apo83, Apo86]. Domains [HMR85a, HMR85b]. donnée [Gra86a]. données [(??84, Tri84, (??87, Tri89]. dose [PS82, PS83, SD89]. dose-response [SD89]. double [SMD84]. double-differential [SMD84]. down [Leh83b]. downturns [SD89]. Draft [Ame87b, Ame87c, X3J89, Kne81, Ame87e, ANS89, Ame87a, Ame87d, AR87, AW89, Com89, Dig83, Ame89b, Mei87, Mei89a]. drag [Iwa84]. drain [War86]. Drawing [Pal86]. drilling [BT83]. driver [DJM87, Moo81, Moo83]. DRIVERS [Gra85]. DSP [ZGK88]. DSUBSP [PCK84, Van82]. due [Tho84b]. d’un [LB89]. dung [Dah81]. dust [Col89a, Col89b]. duty [DJM87]. DVED [Div85]. DWBA [PB84]. dynamic [BG84, CK86a, Red86, Wat86, Wei84]. Dynamically [Cod88]. Dynamics [Sme81, GDK89, LR89, Sch89a, SS82, VSH83, Wee86]. Dynamischer [Jac82].

E. [Mil82b]. E1 [Dur80]. E2 [Fut78, Sim76]. E3 [Dur80]. E4 [MGH81b]. Early [All84, Bac84b, Bri84, Hug84, McC84b, RS85]. earthquake [RO85, RO86]. easily [MT84a]. Easy [CS83, LP85b, LP85a]. EASYTRIEVE [HG83]. EASYTRIEVE-11 [HG83]. echo [vMF81, vMF84, vM84d]. Eclipse [Kir89]. Ecological [Kre88]. écriture [AFV85]. ed [RWA84, Rod84]. eddy [JC82, JC82]. eddy-current [JC82]. edit [ZDS81a]. Edition [Bee82, Mic84f, APD86, Dig84a, Dig86d, HF81, WAD81, Wil87b, YS84d]. editor [BKK^{+}89, Dig85d, Mic87e]. EDLIN [Div85]. EDM [EIT85]. Edouard [KTW84]. education [Chi85c, MSM84]. Educational [RB82]. Edward [Rei89a, Rei89b]. Effect [PLR85, CSD82, CK88, Dav86, Nai84]. Effective [BS88b, BS88a, Met85a, Met86, Met87b, Met89b, Gui81, MRS84, Ric82a, Nad86]. effects [Don83c, Lee85, Red86]. efficiency [CMM^{+}88]. Efficient [Hou83, Kah80, Ste79, SS79, Ada89, Con82a, CT88, JT88, Mil89]. Effort [Sch89e]. ego [Osi82a, Osi82b]. EGR1.LB [Egg83]. eigenparameter [Sha89, Sha87]. eigensolution [CHPS85]. Eigenvalue [DM84]. Eigenvalues [CL83, Don82a, Don82b, FW82, Ste76, Par84]. Eighth [Ass86]. Einfuhrung [NEM84, HV83, Lam81, Sch84c]. EISPACK [DM84]. elastic [Red86, SPS84]. elasticity [SH88]. electric [Nai86]. electron [EBS88, Hue83, MT84b, NE81]. electrostatic [HM81, HM84]. ELEFUNT [Cod86b, Cod86a, Cod89]. elegant [EL81]. elektrischen [Wie85]. Elektroencephalogramm [San82].

Element [Fen87a, Rei84b, Fen87b, Rei84a, Tan86, Wan86]. Elementary [ACG^{+}86, Aya84, Lin83]. Elements [CD84, CL83, Cor81, Cor82c, Doh82, FGGF86, Fuo86b, Fuo86c, Bru86]. Elimination [DFK83a, DFK83b, DFK88, She78, DFK81, PW84]. Ellips [Ber84a]. ellipsoid [Joh81]. ellipsometer [Ber84a]. Elliptic [Car87, Car88a, Car89b, Ste79, SS79, Ada89, Lio85, dZ86]. Ellis [Cou85a]. els [Lee84a]. embedded [Rel88, Rat89, Rel89a, Rel89b, Rel89d]. Emergence [Dav84b, Hei84]. emerging [Met87a]. Empirical [VD84]. employing [Ano87a]. EMR [RV8 ]. EMS [O’R81]. end [Gul86, ZGK88]. endliche [NEM84]. energy [Kip82, MT84b]. energy-dispersive [MT84b]. Engineering [BL83, BRK^{+}87a, BRK^{+}87b, BRK^{+}88, Cha87b, Fen87b, Fen87a, KW89, Mar84a, MBP^{+}85b, Mos88, Rey80, Sch85b, SAN^{+}81, SP84a, BEE^{+}85a, Cre89, EP81b, GSZ88, JS88, MSR87, MSM84, MBP^{+}85a, Osy84, Sel83, SL82, Sch81b, SB82, SB86, SP85a, SP85b, SR87, Tan88, MS88a]. Engineering/Scientific [BRK^{+}87b, BRK^{+}88, KW89]. Engineers [AB83, BS88a, Bor85b, Bor85a, DF89, HH85, Lei87, McC84a, MS88c, Mil87a, Mil88b, NL88, Val85, Wei89a, Wor88, BS88b, Cou85b, DcF89, Ett83a, Ett83b, Ett87, McC84c, MS88b, Mil82a, NL85c, NL85d, RZ89a, RZ89b, Wor89b, Mil87b]. engines [RAKK88]. England [JR89]. Engr [Dun85b]. Enhanced [Bos88, Cor82a, Cor83, Obl85]. enhancing [BK89]. ensemble [LB89]. entire [FDL86]. Entwicklung [Fis82, Ste87]. Entwurf [HS81, Mey84]. enumeration [Red82]. envelope [And84a]. Environment [AK85, Ber88a, Sof87, Did86, Don84a, Don84d, Don85a, Don87b, Don88b, Don89, HO89, HK84, Mar82b, Pra85, ALPC88, CKT85, Dat84, Dat85a, Don83b, Don84b, Don85b, DS87a, Don87a, Don88c, Jor86, Jus88, Kim86, May89, Sto85c, Sus86, VP84]. environmental [PM87]. Environments [WN87, BS84, Col89a, Col89b]. EPEX [Bol89, BDR87, DGNP88a, DGNP88b, Sto85a, Sto85b]. EPEX/FORTRAN [BDR87, DGNP88a, DGNP88b]. equality [Zho89]. Equation [Cal86, KW87b, KW87a, LK88, Sha87, CK86a, CMM^{+}88, HK83, Sha89, Tan88]. Equations [Abd80, BA85b, BA85c, BA85a, Cas89a, Cas89b, CC82, Don84a, Don84d, Don85a, Don87b, Don88b, Don89, DR82, Gaf84, Hig91, Hop81, MHK86, MC80, Ste79, SS79, Ada89, Bow82, GtTB89, Don83b, DH84b, Don84b, Don85b, Don87a, Don88c, HL86a, HL86b, Hof87, HP89, IZP81, JBJ84, Pet89, Red86, Rei84a, RAKK88, dR87, Tod85, Uni84e]. EQUEL [Rel86, Rel88, Rel83, Rat89, Rel89c, Rel89b, Rel89d]. EQUEL/FORTRAN [Rel83]. equilibrium [NJLB81]. Equipment [Fed81, Lag85, Mul85, PF85]. Equivalence [Tan81a, Tan82]. Erkennung [San82]. Error [Whe84a, Whe84b, Bli89, JK83, Ott87]. error-handling [JK83]. Erstellung [Wie85]. Erster [CLW81]. Essays [KTW84]. Essentials [MS81, MS84a, She84, She89a, Smo89b, Smo89a, Edm86]. estilo [DH84a]. Estimate [Gaf83b, Gaf83c, Gaf83a]. Estimating [CGM84b, CGM84a, CGM85a, Hig88b, Hig88a, Hig89, CGM85b]. Estimation [Hig88b, Hig88a, Hig89, Mai81a, Mai81b, NM85, Szy87, TFH86, The88]. estimator [TSU88]. estimators [KW87b, KW87a]. estructurada [FK84, Zwa85]. estructurado [DH84a]. ETA [BK88]. ETA-10 [BK88]. Etter [Cou85b]. Euclidean [Blu78]. EUROCAL [Dav89]. European [Dav89, LCMM88, RW86]. EUSIPCO [LCMM88]. EUSIPCO-88 [LCMM88]. evaluate [Don83c]. evaluates [HM81, HM84]. evaluating [Lib87, IMS87c, IMS87j, IMS87h, IMS87i, IMS89d, IMS89e, Lib89b, IMS89l]. Evaluation [Arn82, AB89, Gaf84, GN89, Kim86, Rey80, Bli89, BDJ^{+}89, BK88, DFD81, DFD84, Gul86, KWM88, Kee88a, Kie83, Kip82, MSM84, MHS81, Mil89, Ras84, hTD88, LZ82]. evaporation [Dav86]. Event [BCM87]. events [Hel83]. everything [Col87b]. EVM [SK83, Aka88, Bri85]. evolution [Van84b]. EWALD [Tho84a]. Ewing [For85]. Exact [MP86a, MP86c, MP86b, Dal88a, SPS84]. EXAFS [AI88, IA84]. Example [Pre88a, VTPF87, Vor89, Dun85b, Vet85, VTP89, VN89a, VN89b]. Examples [Spa85b, Spa85a, CV88a, RW85, Eve85]. EXCHNG [FW82, Ste76]. EXCHQZ [PCK84, Van82]. excitations [Fra84a, MSG86]. Execution [Ana87, BDR87, CC87, Pol87, Wie82, LH88]. Executive [Cro85b, Gri82, Cro85a, IBM89b]. exemple [Dub84]. exercices [Lig82b, Lig88c]. Exercise [EA87, HWS^{+}88]. exercises [DV81]. Exhibit [Lee84c]. existing [Dha88]. Expander [SL82]. Expansion [JRS88a, JRS88b, AJ88, Gra84b]. expansions [Gro89, Rap82a]. Experience [Bri84, RS85]. Experiences [HG82a, HLM84]. Experiment [PD81, BCF^{+}88, Wie83, Wyl86]. experimental [Gat82, War86]. Experiments [LG86]. Expert [KBRM^{+}86, Miz83, CGQS89, Cre89]. Explained [MR87, MR89, MR88]. explanation [Don88a]. Explicit [LK88]. Exploitation [Rei87a]. exploiting [MR86]. exponent [Int84b]. exposures [NE81]. expression [Uni83a]. Expressions [Köl82, Vel82]. Extended [DDHH84, DCHH87, DCHH88a, DCHH88b, DCHH88c, Int84b, WN87, WLO76, BCF^{+}88, Con83b, DCHH85, HS83, HL82b, IBM88].

Extension [BRK^{+}87a, BRK^{+}87b, BRK^{+}88, KW89, Sch89d, HT82, Hym82, Kul83, SAS86]. Extensions [Col83, Gre86, Ric84, Gra86a, VLV^{+}86, Wee86]. extract [JC82, Kle89b, Kle89a, WD81a, WD81b]. Extracting [RV8 ]. extraction [Par84]. extrapolation [Gen82]. Extremal [Sau83b, Sau83c, Sau83a]. EZPLOT [GRB88]. EZVIDEO [Pat89].

F [Rit89, hC83, LN87, Num88d, Uni81b]. F-80 [hC83]. F1 [DG82, LHKK79a, NSB2 , Uni82c]. F2 [FW82, Ste76]. F2812 [Sim86]. F4 [Abd80, BS81b, She78]. F5 [Buc81b, Buc82, MC80]. F66 [Tri84]. F77 [MC82]. F77L [Lah88a]. fa [hH82, RkC84]. FABSTAT [Per81]. Fachtagung [Wös82]. Facilitate [Bro81a]. Facility [Gre81, Ano81a, Bat85, Fri84, Gat85, AC86, MS88a, RV89]. factor [Tho84b]. factoring [LtW88]. Factorizations [Buc81a, Buc81b, Buc81c, Buc82]. Factors [HPB82]. Factory [KODG^{+}87]. failure [Joh85b]. Fakturierung [Ste87]. Fall [MSM84]. farming [BDJ^{+}89]. farms [GH87]. Fast [Gui89, YF85, Bai87, Bun86, DS82, Gui88, Joh86, MT84a, Rod87, NM85, Per83b]. fatigue [CB86]. fault [HS86, RO85, RO86]. fault-plane [RO85, RO86]. fcc [DM89b]. Feasibility [RV8 , SAB88]. feasible [Zho89]. Features [Rei87a, AKLS88, JC82, RW85, VD84]. Featuring [AM81, Mei84, MO82, MO84]. February [Rei87c]. fed [Bai89, Hsi83]. Federal [RW86]. feed [Dav86]. Fehlertolerante [MF84]. Fermion [BBF^{+}82]. ferrocenes [She89b]. FEXACT [MP86a, MP86c, MP86b]. FFT [Swa84]. FFTs [Bun86]. fi [aHH83]. fibel [KW87c]. Fichiers [Lig88c]. fictitious [DS82]. fictitious-gas [DS82]. field [Gra84b, vMF81, vMF84, vM84d]. field-gradient [vMF81, vMF84, vM84d]. fields [Hay86]. fighter [Kip82]. file [Bod87, Kir89, OO86]. files [BJ81a, BJ81b, BJ84a, BJ84b, Gre88, Kle89b, Kle89a]. Filling [YF85, Pal86]. Film [Lee84c]. Filter [VVV89b, VVV89a]. Filtering [LG86]. filters [Alb86, Eva81, Gra81a]. final [Dha88]. Find [Blu78, Cra86a]. Finding [CL83]. Fine [CF85]. Fine-tuning [CF85]. FINGER [Wan86]. Finite [Fen87a, Rei84b, BMS84, Fen87b, GF89, PB84, Rei84a, Riz85, SR88, Tan86, Wan86]. finite-difference [GF89]. finite-element [Rei84a]. finite-range [PB84]. finite-volume [Riz85]. FIPS [Uni83a]. First [Cas89a, Cas89b, CM81b, CM84a, CM84b, Hig91, RWA84, Rod84, WAD81, CM81a, Fra84a, MB81]. First-Order [Cas89a, Cas89b, Hig91]. Fish [Sim86, Mar81, Sim85, Sim88b]. Fisher [MP86a, MP86c, MP86b]. Fit [Fut78, Sim76, vMF81, vMF84]. FITLIB [Ano87a]. fitting [Ano87a, BDS84, Wat82a, Wat82b]. five [Kle89b, Kle89a]. fixed [GS81]. FIXSRC [BH85]. fiziki [BZ85]. flames [Kee88b]. FLEX [Rus87, Cro87a, Cro87b]. Flex/32 [Cro87a, Cro87b]. flexible [Rus87]. float [KS81b, KS81a]. Floating [Cro87a, Fat82a, Fat82b, Kaw84, Sch81a, Cro87b, Wic89, ZGK88]. Floating-Point [Cro87a, Fat82b, Sch81a, Cro87b, Wic89, ZGK88]. flood [Uni81a]. Florida [IEE88b, ML88]. flotation [Mul85]. Flow [Ell82c, GPKK82, GPKK84, jYsS89, ACK86b, BT88, Chi88, CSD82, Ell82d, GG88, JT88, kK89c, Rod87, Shi88, Tan81a, Tan82, Wal85]. Flowchart [hA84]. flows [Dul81, Sch82b, Tho81, Tho82a]. flowsheet [Roc86]. FLOWTRAN [Roc86]. fluctuations [O’R81]. fluid [LR89, Wee86]. fluids [McC81, McC86, Rap82b, Rap82c, You82]. fluorescence [TPR85]. fly [Sho85]. FO1 [Int83b]. FORCASD [Dah81]. forces [Wal85]. FORCON [Kle89b, Kle89a]. FOREET [BA86]. FORI [Ano82a]. Forlib [Alp83]. Forlib-Plus [Alp83]. Form [AK82, AK87, WN87, AK81, Sal84, Dix85]. Formal [Ban88, Dob85]. formalized [She82]. Format [BP81a, Gra81b, Hus84, Pre89, BP81b, Gre88, Jac85b, Jac85a, OO86, Sal84, VN89b]. Formatted [Whi81a, Whi81b]. Formeln [Hof84]. Formelsammlung [EmR84]. Forms [VVV89b, VVV89a]. formula [Tan86]. formulae [Bod87, Min88]. formulas [Bar89]. FORSE [SL82]. FORSIG [Hym82]. Forth [Pou87]. FORTRAM [RO85]. FORTRAN [Abs88, BZ85, Bee82, Bee85b, BP81a, CwL83, Dig82f, Dig84j, Hel83, HS81, HT82, Int83c, Law83, MF84, MSM84, Mil87b, Num83c, Num84d, NEM84, RB82, RWA84, Rod84, Sch84c, Sch87c, SDC82, WAD81, Wu82a, Wu82b, Ame87b, X3J89, AL81a, Ano82a, Bee88, BBB^{+}83, Com89, Cod86b, Col83, Cou85a, Cou85b, Edm86, EmR84, Eve85, GN89, Hei83, Hel85b, HV83, JL81a, KL83a, KL83b, KL85a, KL85b, KK89b, Kur85, Lah88a, Lam81, McC84c, McC85a, McC85b, MS88b, Mey84, MS84b, Mon82a, Mon82b, Nad86, Pre88a, Sch84d, Sch84e, Sch84b, Weh85, Whi89, Wie85, Wu83, AFN83, Ame85a, Ame87a, AHU81, Ass82, Ass83a, Abd80, Ack84, Ada84, Dha88, AM89a, Alb86, AI88, All82, All84, All87, All90, Amo83a, Amo83b, hA84, And84a, And89, AG87b, Ano81b, Ano82c, Ano82d, Ano84]. FORTRAN [Ano85b, Ano89, Ant81, ADP88, AEV89, Apo83, Apo86, AEL^{+}86, Art81, AS89a, AS89b, Ash85a, Ash85b, Ash85c, Ash81b, Ash81a, AW82, AFV85, Buc84, Bac81, Bac84b, Bai89, Bai86, Bal84, BNZ87, BW87a, BS81a, BS81b, BGCS82, BGM83, BS81d, Bau88, Bee81, Bel89, Bem84, BS81f, Ber88a, BSdlT87, Ber87, Ber82b, BS81g, BS88c, Bin85, BBuC84, BRK^{+}87a, BRK^{+}87b, BRK^{+}88, Bod87, Boi81, BW84, BW87b, BDR87, BH89, Bor89, BA85b, BA85c, BT83, Boy84a, BS86c, Boy84b, Boy85a, Boy89a, BDS88a, BDS88c, BDS89, Bri84, Bro84a, Bro89a, Bro89b, Bro89c, Bro89d, Bro81b, Bro83a, Bro82c, Bro83b, Bru86, Buc81a, Buc81b, Buc82, BD80, Bur85a, Bur81b, Bur84b, Bur84a, Bur85b, Bur85d, Bur85c, Bur86c, Con81a, Con82b, Con82c, Con82d, Con82e, Con83b, Con83c, Con83d, Con84, Con85b, Con85d]. FORTRAN [Con85c, Con86, Con87c, Con87a, Con87b, Con88a, Cad84, Cal85, Can81, CW85, CW88, CW89, Cas89a, CL83, Cha83, Cha86a, Cha87b, cC84b, CB86, Chi85a, Chi85c, hC83, mCaLjH84, Chu88, CSC^{+}86, CHPS85, CwL83, CGM85b, CDHP86, vC87, Col89a, Col89b, Con83e, Con88b, CA86, Cor81, Cor82c, CBS81, CB82, CM81a, CM81b, CM84a, CS84, CM89, CDW83a, CDW83b, DW85, CM81c, CM83b, CM86, Cro87b, Cro85a, Cul88, Dig82b, Dig82c, Dig82e, Dig83, Dig84a, Dig84b, Dig84c, Dig84d, Dig84f, Dig84e, Dig84g, Dig84h, Dig84i, Dig85a, Dig85b, Dig85e, Dig85c, Dig85d, Dig86b, Dig86c, Dig86a, Dig86d, Dig88b, Dig88a, Dig88c, Dat81, Dat84, Dat85a, Dah81, Dal88a, Dal88b, Dal89, DGNP88a, DGNP88b, Dav82, Dav84b, DH82, DH83, Dav85, GtTB89, DFD81, DFD84, Del83, Den84a, Den82, Der82, Des89, DcF89, DKG89b, DM87b]. FORTRAN [DFK81, DFK83a, DFK83b, DFK88, DP84b, DP84c, Dig84k, DW83a, DW83b, DW84, Dir84, DV81, Dod83, Doh82, Don82a, Don82b, Don84a, Don84b, DE84, Don85b, DS87a, DCHH88b, DCHH88c, Don88c, DM89b, DPA87, Dul81, DS82, DZ87, Dun87a, DZ88, Dun88b, Dun85b, Dun85a, Dur80, DLS84b, Ear85, Edg89a, Egg83, Ehi82, EH81, EK87b, Ell81a, Ell82a, Ell82b, Ell81b, Ell82e, EIT85, EP87a, EP89, Est82, ETA88, EGP81, Ett83b, Ett87, Eva81, Fed81, Fre83, Fed82b, Fen87b, Fog85, Fog87, Fog88, FDL86, For82b, FR82, FE82, Fra84a, Fra84b, FMH85, FK81, FK84, Fri84, Gaf83b, Gaf83c, Gaf84, GRB88, GGLM88, GL90, Mer86, Mer88a, Gat82, Gee86, GSZ88, GC84, GMW86, Gla88, GH87, Gol84, Goo89, Gra88, Gra86a, Gra84b, Gra86b, Gre84, Gre81, Gri82]. FORTRAN [Gro89, GQ88, Gui81, GF81, Hew85, Hew86, HO88, HL82a, HRH81, HRC89, HK87a, HS83, Har86b, Har81, HM82, Har85, HS86, Hay86, Hea81, Hei84, HS81, HPB84, HL82b, Hig88b, Hig89, Hig91, HW86, HF81, HL86b, Hof87, HP89, Hon82, HB81, Hon81b, Hon85, HG83, Hou83, Hsi83, hH82, Hua82, kH84, HPB82, Hud88a, Hud88b, Hue83, Hug84, aHH83, Hym82, Int82b, Int82c, Int82d, Int82e, Int82f, Int82g, Int82h, Int82i, Int82j, Int83b, Int83d, Int83e, Int83f, Int83g, Int83h, Int83i, Int84b, IR84, Int84c, Int84d, Int85b, IBM85, Int85c, Int85d, Int85e, Int85f, Int85g, Int85h, Int85i, Int86a, Int86b, Int86c, Int86d, Int86e, Int86f, Int86g, Int87c, Int87d, Int87e, Int87f, Int87g, Int87h, Int87i, Int88b, Int88e, Int88f, Int88h, Int88g, Int88i, Int88j, Int88k, Int88l, Int88m, Int88n, Int88o, Int88p, IBM88, Int88d, Int89a]. FORTRAN [Int89b, IBM89a, Int89c, Int89d, Int89e, Ion84, Int81e, IEC85, IEC88, IMS82, Lib84a, Lib84b, IMS84, IMS87a, Lib87, IMS87c, IMS87b, IMS87d, IMS87g, IMS87e, IMS87f, IMS87j, IMS87h, IMS87i, IMS87l, IMS87m, IMS87k, IMS87n, IMS89a, IMS89b, IMS89c, Lib89a, IMS89d, IMS89e, IMS89f, IMS89g, Lib89c, Lib89b, IMS89h, IMS89m, IMS89i, IMS89j, IMS89k, IMS89l, IMS89n, ISO88, IS84a, IS84b, ISJ85, IZP81, Ame85b, Int85a, IA89, Iwa84, Jap82, JR81, JSW85b, Jan84, JC82, Jia86, JK82, Joh81, Joh83, Joh84, Joh87b, Joh87c, JC88, Jus88, Kah80, Kal85b, KP86, KWM88, Kee88a, Kee88b, KJM89, KD84a, KD84b, Kem85, Kem87, KWWK86, Ker82, Ket85b, Kha81, Kie86, KRYG82b, KRYG82c, Kip82, Kir89, KK89b, Kle89b, Kle89a, Kne81, KTW84, Knu84, KF87, KF88, KM89, KW89, Kre88, KS82b, KW84, Kul83, sKcH81, hK85]. FORTRAN [LZ82, Lag85, Lah86, Lam89, LB86, LD87, Lee84b, Lee84a, Lee85, Lee84c, LP85a, Ler83, LH81, Lew81a, Lew82a, Lew82b, LOU86, iL82, LcY83, Lio85, LN87, LN88b, LO85b, LKM88, MAT89a, MAT89b, Mic83, Maa89, Mac81, mM84, Mai81a, Mai81b, Mai87, MRS84, Mal85, Man82, Mar82a, MR83, Mar83c, MP81, MMM85, Mat83a, Mat83b, McA86, McC86, McC84b, McD83, McG87a, McG87b, McG84, McK83, MP86a, MD88, Rod87, MP86c, Mer81, Met86, Met87b, Met89b, MH82, Mic81, MI82b, MI84, Mic84a, Mic84c, MZ84, Mic84b, Mic84d, Mic85c, Mic85b, Mic85a, Mic85e, Mic86, Mic87d, Mic87f, Mic87e, Mic87a, Mic87b, Mic87c, Mic88, Mic89a, Mic89b, Mic89c, Mic89d, MHS81, Mil82a, MMS88, Mon83, Moo82a, Moo86, Moo88a, MGH81b, Mor84, Mor81a, Mul85, Mul88, Mye83a]. FORTRAN [Mye83b, Num83a, Num83b, Num84a, Num84c, Num84d, Num87, Num88b, Num88c, Num85c, Num88d, Num89, Nag81a, Nag81b, Nag85, NL85a, Nan81a, NM85, NJLB81, Tha89c, Nic82, Nic85a, Nic85c, Noh84, NSB2 , NL85c, NL85d, O’N81, OO86, Osi82a, Osy84, Ott87, Owe86, Owe87, Pee84b, Pee84c, Pee84d, Pee85a, Pee85b, Pee85c, Pee85d, Pee85e, Pee86, Pee89, Pag83, PDA83b, Pag87, Par86a, Par86b, PBB^{+}88, Pat89, Pay84a, Pay84b, PM87, PA84, Per81, PS82, PS83, PCK84, PB86, PMBK82a, PMBK82b, PF85, Pol82, Pra89, PFTV86, Pre87b, Pre89, PRL^{+}85, PP85, Res86, Rel83, Rel86, Rel88, Rat89, Rel89c, Rel89a, Rel89b, Rel89d, Rad83a, RRS88, Rao81b, Rao81a, Rao83, Rap82a, Rap82b, Rap82c, Rat81, Rat86, Red86, Red82, RW85, RAKK88, Rid82a, Rid82b, dR87, Rin83, Rob82, RG85, Roc86, Rod86, RL81, Rom81, Ros84].

FORTRAN [RMS82, RH84b, RB83, Roy88, Rub83, Rus87, Rzy84, Sof83a, SH88, Sas83a, Sas81, Sau83b, Sau83c, SPS84, Sav87, Sch82a, Sch82b, Sch86, Sch84a, Sch87a, Sch87b, Sch89a, Sch83b, Sch89e, Sch88c, Sch81b, SB82, SB86, See81a, Seq89, Sha87, Sha89, SK83, She89b, SS84, SSS84a, SSS84b, SJB83a, SJB83b, SD89, Sim85, Sim88b, Sin81, SP85a, SP85b, SIR82a, SIR82b, Sis85, Slo88, Smi81, Smi83a, Smi84, Smi83b, Smi85d, Smi85b, Smi85c, Smo89a, Sne88, Spe82, Spe83, SP84b, SP84c, SG88, SR86, Sto85a, Sto85c, Sto85b, Str85, Stu81b, Sul2 , Sul88, Sun84, Sun85, Sun86a, Sun86b, Sun88b, Sym85, Sym86, Sym88, cT81, cTcT84, TU81, TSU88, Tan83a, Tan85a, Tan85b, TPR85, Tat87, TS88, Tel82, TMS88a, TMS88b, Tew81, TBM85, Tha89a, Tha89b, The88, Tho81, Tho82a]. FORTRAN [Tho82b, Tho84a, Tho84b, Tho84c, Tho84d, Tho84e, mT82b, Tod85, Ton82, TPS^{+}88, (??84, Tro84, TR84, sT85, Uni86a, Uni82a, Uni82b, Uni82c, Uni81b, Uni83b, Uni85a, Uni84a, Uni84b, Uni81a, Uni83a, Uni85c, Uni83c, Uni86b, Ull84, Div85, Van82, VLV^{+}86, Van84b, VVV89b, Van84c, Van85, Var85, VMS81, VTP89, VN89a, VN89b, VL81, VH87, Vu 89, WH87, WG84, Wal3 , Wal81, Wan84, yW85, War86, WB89, Wat86, Wat82a, Wat82b, Web88, Wee86, Wei86a, Wei86b, WD81a, WD81b, Whi81a, Whi81b, WM85a, WB85, Wie83, Wie86a, Wie86b, Win85, Wu82a, Wu82b, WF85, YK88, jYsS89, YS84a, YS84b, YS84c, YS84d, fY84, dZ86, Zim86, ZDS81a, ZDS81b, ZSD82a, ZSD82b, ZGS89, Zwa81, Zwa85, dB82b, dEV89, vvHG87a, vvHG87b, van86, SL82, HB84, Ass84, Ass83b, AC87, Adv86]. Fortran [Abe89, AR87, Ada89, AW89, Adm84, Adm85, AM81, Aha85a, Aha85d, Ain89, Air77, AD84, AS88a, AS88b, AKLS88, Alc82, Alc83a, Alc83b, AK81, AK82, ACK86a, ACK86b, AK87, AP87, APD86, And84b, AL81b, AG87a, AG87c, Ano81a, Ano82a, Ano82b, Ano83, Ano85c, Ano87a, Ano87b, Ano87c, Ano87e, Ano88c, A^{+}81, AHH89, Aya84, Baj81, Bal85, Ban78, Bar84, BS88b, BS88a, BK84, BS81c, BG82, BSP83, Bau82, BL83, BR89a, BS86a, BA86, BS81e, Ber82a, Ber85a, Ber89, Ber84a, Ber84b, Ber85b, Ber88b, Bez88, Bez89, BMS84, BB82, Bli89, Blu78, BKK^{+}81, BBG^{+}82, BBG^{+}84, BS86b, Boi84, Boi85, Boi87, BDJ^{+}89, BKL89, Bor85b, Bor85a, Bow82, BA85a, Boy85b, Boy89b, BGG85a, BGG85b, BR89b, BK88, Bre78b, Bre78a, Bre79, BDS88b, Bri85, Bro81a, Bro82a, BS83]. Fortran [Bro82b, Bro84b, BH85, Bro85, Bru84, Bru82, Buc81c, Bur81a, BJ81a, BJ81b, BJ84a, BJ84b, Bur86a, Bur86b, Bur87, BP81b, Com89, Con81b, Con82a, Con83a, Con85a, Cal83, Cal89b, Cal89c, Cal89a, CF85, Cas89b, Cas81, CS83, CCN^{+}79, CD84, CRV^{+}89, CDW82, CDW84, Cha87a, Chi85b, Chi81, Chi86b, Chi88, CC84a, CK86b, CR84, CCHT89, Cla86a, Cla86b, Cla86c, Cla89, Cod86a, Cod89, Col83, Col87a, Col82, CGM84b, Col84b, CGM84a, CGM85a, CSD82, Cor82a, Cor83, Cor82b, CSD83, CM84b, CT86a, CT86b, CM83a, CM87, Cra83, Cra84, Cra86b, Cra89a, Cra89b, Cre89, Cro87a, Cro85b, Cza83, DB82a, Dig82a, Dig82d, Dat85b, Dav81a, Dav84a, Dav81b, DH84a, Dav86, DH88a, DH88b, Del82, Del85, Del88, Den84b, DF89, DKG89a, DM87a, DP81, DP84a, Dil85, DM89a, Dob85, DG82, Doe88, Dol88]. Fortran [DJM87, Don81, DS86a, Don83b, Don84c, DDHH84, Don84d, DCHH85, Don85a, DD86, DS86b, DCHH87, Don87a, Don87b, DS87b, Don88b, Don89, DM87c, Dre81, Dub84, DR82, Dun80, Dun87b, Dun88a, Dur83a, Dur83b, DLS84a, Edg89b, EL81, EE84, EK87a, Ell82c, Ell82d, Ell83, Els82, Enc87, Ett83a, Ett84a, EP81a, EP81b, EP87b, Fed82a, FS86b, FW83b, FW83a, Fen87a, Fis83, FW82, Flo89, For82a, For85, FSO89, Fre81, FKSS81, FGH81, Ful87, Fun86, FGGF86, Fuo86b, Fuo86c, Fuo86a, Fut78, Gra85, Gaf83a, GY82, Gic88, GMPW79, GHM^{+}86, Gil86, Got84, GB83, GB89, Gra84a, Gra81a, Gre88, Gre86, Gre85, GDK89, Gri85, GKRY82, Gro83, Gru88, Gue86, Gui87, Gul86, Guz87, Guz88, GPHL88, GPHL90, Hah87, HG82a, HR83a, HRC87, HA83, HK87b, Har86a, HWS^{+}88, Her81, HM81, HM84]. Fortran [HL82c, Hig88a, Hig86, Hil81, HO89, HL86a, HB83, Hon81a, HK84, Hop81, Hos88, HK83, HPR81, HR83b, HR83c, HR84a, HR84b, HH85, Hur82, Int81a, Int81b, Int81c, Int81d, Int82a, Int83a, Int87a, Int87b, Int88a, Int88c, IBM87, IBM89b, Wor84a, IA84, Ame89b, Int84a, Isn82, Jac85b, Jac85a, JS88, Jam86a, Jam86b, Jan88, Jes82, Joh85a, Joh85b, Joh86, Joh87a, JL81b, JC89, KW87b, KW87a, Kal85a, KK89a, Kan88, KB88, Kat82, Kaw84, Ket82, Ket84, Ket85a, Key81, Kie83, mK84, Kim86, KGRY81, KRYG82a, KRYG82d, Kir85, KKK89, Kna84, KS82a, KC84a, KC84b, Kri83, KW87c, Kum86, La 87, Lah87a, Lah87b, Lah88b, Lah88c, Lah88d, Lam86, Lam84, Lan88, Lar81, Lau86, Lav83, LHKK79a, LHKK79b, Le83, LG86, LP85b, LN89, Leh86, Leh83a, Leh83b, Lei87, LHP87, LH87]. Fortran [LB89, LW88a, LW89, Lev89, Lew81b, Lig82a, Lig82b, Lig84, Lig85a, Lig85b, Lig88c, Lig88a, Lig88b, Lin83, LPJ83, LP87, LS87, LS88, LN88a, Lud81, LCH^{+}88, LW88b, LO85a, Met89a, MS81, MS84a, Mar83a, Mar83b, MW83, Mar81, Mar83d, MW84a, Mas83, Mas87, MW84b, May89, McC81, MSR87, McC84a, McC85c, MS88c, McD85, McK85b, McM86, McN83, MO82, Mei84, MO84, Mei87, Mei88, Mei89a, Mei89c, MP86b, Mer88b, Met82, Met85a, Met85c, Met85b, MR87, Met87a, MR88, MR89, Met89c, MI82a, Mic84e, Mic84f, Mid84, Mil82b, Mil87a, Mil88a, Mil88b, Miy87, MA89, Mon89a, Mon82c, Mon89b, Moo85a, Moo81, Moo83, MM81, Moo82b, Moo85b, MC80, MGH81a, MSG86, Mos88, MT84b, Num84b, Num88a, Num86, Nai84, Nai86, NL85b, Nan81b, NSV1 , Nic85b, NE89, Nor83]. Fortran [NL83, NL88, OM82, O’R81, OK87, Oli81, Oni85, Osi82b, Pad85, PDA83a, Pag84, PDA86, Pag88, Pal86, PB84, Pee84a, PD81, PA83, Pet83, Pet87, Pol81, Pol83, Pol87, PS84, Pre87a, Pre88c, PP82a, PP82b, Pru87, Pyr84, Rad81, Rad83b, Rai84, Rao82, RV84, Rao86a, Rao86b, Rao87, RRC89, Ras84, Rat87, RO86, RZ89a, RZ89b, Rei84a, Rei84b, Rei87a, Rei89a, Rei89b, Rey80, Ric84, Ric82b, Rid82c, Rob83, RV8 , Ros87, Rou83, Rou86, RS81, RS84, Rul83a, Rul83b, RkC84, Rya86, SAS86, SIG84, Sel83, Sal84, SM84, Sas83b, Sau83a, SK86, Sch82d, Sch89c, Sch89b, SM88a, SM88b, Sch89d, Sch79, Sch85b, SB83, See81b, Ser85, Ser89, SMD84, SR84a, SDH84, She84, She89a, She82, She78, SSS84c, Sim88a, Sim76, Sim86, SC83, SP84a, SW83]. Fortran [Sme81, SS82, Smi83c, Smi85e, Smi85a, Smi87a, Smi87b, Smi88a, Smi88d, Smi88b, Smi88c, Smo89b, Sol89, Som86, Son83, Spa85b, Spa85a, Spe85, SS87b, SR87, SR84b, SP87, Ste79, Ste76, Str82, Stu81a, Sun88a, Sus86, SS79, Szy87, Tan86, Tan81a, Tan82, hTD88, TFH86, Tei86, Ter87, TMjC81, Tho86a, TW87, Tou84, Tri84, (??87, Tri89, TFI86, TWI88, Tuc86, Tur86, Uni87, Uni85b, Uni84c, Uni84d, Uni84e, Uni88, Ull85, Ull86, Vag89, Val85, VVV89a, Vet85, VTPF87, VPH82, VSH83, vdV85b, vdV86, Wor84b, Wag84, Wag85, Wal85, Wat85, Web85a, Web85b, Wee89, WW89, Weh85, Wei89a, Wei89b, WS84, Whe84a, Whe84b, Wid88, Wil87a, WP84, Wis81, Wol85, Woo89, Wor88, Wor89a, Wor89b, Wu83, WLO76, Wyl86, Ame87e, ANS89, You82, hYsA82, Zho89, Zoh84, dB84]. Fortran [vdV85a, vMF81, vM84a, vM84c, vMF84, vM84d, vMT84, vM84b, vM84e, vM84f, FCG83, SFKS81, Wil87b, Bis81].

FORTRAN-10 [Dig85a]. FORTRAN-10/ [Dig85a]. FORTRAN-77 [Hay86, Enc87]. FORTRAN-86 [Int85a, Int84a]. FORTRAN-C [Sch89e]. FORTRAN-Coded [vvHG87a, vvHG87b]. Fortran-fibel [KW87c]. Fortran-IV [SDH84]. FORTRAN-Lexikon [Ehi82, Ano82a]. FORTRAN-like [Whi81a, Whi81b, HL82c]. FORTRAN-orientiertes [Dah81]. Fortran-preprozessor [Els82]. Fortran-Programmen [Wie85, Wid88]. Fortran-Programmpaket [Kna84]. Fortran-Programs [Bur81a]. Fortran-Routinen [Wis81]. FORTRAN-SC [BRK^{+}87a, BRK^{+}88, KW89]. Fortran-to-Pcode [CCN^{+}79]. FORTRAN/ [DLS84b, Hew86]. Fortran/2 [Ano87b, Int87b, IBM87]. Fortran/77 [DLS84a]. Fortran/ANSI [KK89b]. FORTRAN/WATFOR/WATFIV [BS81d]. FORTRAN77 [Mic85d]. FORTRANe [BZ85]. FORTRANHybrid [RAKK88]. Fortrans [Aha85b]. FORTRANSIT [Hem86]. forum [Ass84, For82b, SIG84]. Foulkes [Dal89]. Foulkes-Davis [Dal89]. Fourier [NEM84, AC86, AI88, Bai87, IA84, Joh86, Mil89, PS84, TFH86]. Fourier- [NEM84]. Fourth [Cas89c, LCMM88]. FPFIT [RO85, RO86]. FPPAGE [RO85, RO86]. FPPLOT [RO85, RO86]. FPS [Tou84]. FPS-164 [Tou84]. fraction [Bar89]. FRAME [MT84b]. France [Ass86, LCMM88]. Franconian [Rub83]. Fraser [Dun87b]. Free [BP81a, Gra81b, BP81b, DS82, Jac85b, Jac85a, OO86]. Free-Format [BP81a, Gra81b, BP81b, OO86]. French [Gra86a]. frequencies [Iwa84]. frequency [Alb86]. Friedman [Dav82]. front [Gul86, ZGK88]. FSQP [Zho89]. FTIDY [BS86c]. FTN [Ano82a, Ehi82]. FTN5 [Ano82a]. Fu [hC83, fTBcL7 , mT82b]. full [Lah88b, Lah88c, Not89, Dix85]. fully [Col89a, Col89b]. Function [Amo83b, DM87b, DM87a, HKP88, Stu81a, Woo89, Bar84, Bau88, Bru86, Col84b, Ols83, Stu81b, Whe84a, Whe84b]. functional [Ame85c]. functionality [SM87]. Functions [ACG^{+}86, Amo83a, GN89, Maa89, DFD81, DFD84, Gro89, Hil82a, Hil82b, IMS87a, Lib87, IMS87c, IMS87j, IMS87h, IMS87i, IMS89d, IMS89e, Lib89b, IMS89l, MT84a, Num88d, SP87, Wat82a, Wat82b]. Fundamental [Gro83, SR87, MSR87, Ril83]. Fundamentals [BEE^{+}85a, BGG85a, BGG86, Hor83a, Nic82, Nic85a, Nic85b, Nic85c, Pru87, BGG85b]. FURI [Bar84]. Further [Tem89a, Tem89b]. Future [AE87a, AE88, EA87, Aha85b, Met85c, AE87b]. FX [All87, Cod86a, WSL88]. FX/8 [WSL88, Cod86a]. FX/FORTRAN [All87, Cod86a].

G [BD80]. G1 [Mar81]. GALCYCL [Gra86a]. Gamma [DM87b, DM87a]. Garbage [Chr84]. gas [BT83, BDS84, DS82, KWM88, Kee88a, LKM88, Rom81]. gas-phase [KWM88, Kee88a]. GASP [Hur82]. gauge [Lam89]. Gauss [Sch88a]. Gaussian [PW84, She78]. GDR [Dav89]. GENCOL [HMR85a]. GENCOV [CCHT89]. GENCRAY [Wee89, WW89]. General [Coc83, HMR85a, Kir89, KW89, See81b, Ano82d, Bai86, Eva81, Int82e, Int83h, Int84d, Int85e, KJM89, Mai81a, Mai81b, See81a, Wal85, Zho89, ZGK88]. general-purpose [KJM89, ZGK88]. Generalized [Ban78, KW87b, KW87a, Gui88, Web85a, Web85b, Zoh84]. generate [CDW82, CDW84]. generated [BK88]. generates [CCHT89, MH82]. Generating [Dix85, TWI88, FF84, Lec89, Zho89]. Generation [Chi86a, Gab89, HG82b, LG86, PB86, Vel82, vvHG87a, Bat85, BT83, BDS89, Cas89c, CGQS89, EML88, Gat85, Gro89, IS84a, IS84b, Red86, SR88, Tan86, Tan88, Wan86, vvHG87b, VMS81]. generations [Vu 89]. Generator [Gui89, Haa87, MR83, Sch79, Gui88, HR83c, HR84b, Mil89, Mye83a, Mye83b, Smi85b, Wal81, Wee89, WW89]. Generatoren [Jan88]. generators [Ras84]. GENTRAN [Bat85, Gat85]. geochemical [BNZ87]. geodetic [Ack84]. geographic [BT83]. geological [Cre89, SDC82]. GEOMAN [SDC82]. geomechanics [SPS84, SS87a, Sav87]. geometrical [FS86a]. geometries [And84a]. geometry [Owe87]. geothermal [Rap82b, Rap82c]. geothermometers [Rap82b, Rap82c]. German [HS81, RAKK88]. Germany [RW86]. gestion [TR84]. get [Wic89]. GetData [Mai87]. Getting [Dav81b, Wag85]. GF [PW84]. given [BT83, Boy84a, Gra86a, Hil82a, Hil82b]. giving [Tho84e]. GKS [IEC88, ISO88, Ame85c, Ame85a, Ame85d, BB86, Cha86a, Int88q, Ame85b, Int88r, Thu86, VLV^{+}86]. Global [Ell82c, Ell82d, GMW86, Ack84, JT88]. GMD [BH89]. Go [Dav84b, Bal84]. Goddard [Bro89a, Bro89b, Bro89c, Bro89d]. Good [Ano88c, Smi88c, Dun85b]. Govorite [Rai84]. GPSS [HS81, HV83, MS84b, Sch84c, Sch84d, Sch84e, Sch84b, Ano87e, HS81, Hel85b, Jan88, Lee84a, Sch82d, Sch87a, Sch87b, Sch88b, Ber84b]. GPSS-FORTRAN [HS81, Sch84c, HV83, MS84b, Sch84d, Sch84e, Sch84b, HS81, Sch87a, Sch87b, Ano87e, Jan88, Sch82d, Ber84b]. GPSS-Fortran-Basis [Hel85b]. GR [Ger83]. GR-Software [Ger83]. Gradient [DM89a, LN89, RV89, vMF81, vMF84, vM84d, Obl85]. Gradient-enhanced [Obl85]. Gradients [vdV86, van86]. grading [Lee85]. Grafik [MF84]. grains [Art81, Col89a, Col89b]. Grand [Hor83b]. granichnykh [Sko88]. Graph [Mar82b, YHKM89]. Graphic [HG82a, GRB88, Int82f, VLV^{+}86]. Graphical [Ame85d, ISO88, JL81b, Nag81a, Ame85a, Int88q, Ame85b, JL81a, Num81, Num85a, Nag85, Int88r, Ame85c, IEC88]. graphically [HM82]. Graphics [Ame85d, AG87c, BB83, Egg83, IEC88, Rag86, Ame85c, Ame85a, Adv86, AG87a, AG87b, BR89a, BDS89, Har86b, Int86c, ISO88, Int88q, Ame85b, Kli89, Mac81, Pat89, Int88r, Wie86a, Wie86b]. Graphische [Ger83]. gravimetric [Rap82a]. gravitational [SJB83a, SJB83b]. gravity [For85, JR81, SPS84, Web85a, Web85b]. Grenoble [LCMM88]. GRESS [Obl85]. Grit [Lew81b, Lew81a]. grossen [Kal85a]. ground [KWWK86]. Group [Sof83b, Sof84, Usr82, Gra86a]. groupe [Gra86a]. growth [Mar81]. GSS [Gra85]. GSS-DRIVERS [Gra85]. GSYLV [KW87b, KW87a]. GSYLV- [KW87b]. Guide [Bre81, Dig84a, Den82, GHM^{+}86, IBM86, Pag84, Phi87, Pre88c, Rid82c, Sym85, Sym86, Sym88, AD84, Apo83, AFV85, Buc84, Ber85a, BW84, BW87b, Bur85a, Bur85b, Con81a, Con82a, Con82c, Con82e, Cal85, CB86, CDW83a, CDW83b, DW85, Dig82b, Dig82e, Dig84e, Dig84h, Dig84i, Dig85e, Dig85c, Dig85d, Dig86d, DW83b, Dir84, Ett83a, Fre83, Fed82a, Fog85, Fog87, Fog88, FE82, GC84, Gil86, Gue86, HO88, Hon81b, Int81a, Int81b, Int82c, Int82g, Int83d, Int85h, Int86e, Int86g, Int87c, Int87e, Int87f, Int87g, Int87i, Int88b, Int88c, Int88e, Int88h, Int88g, Int88m, Int88o, IBM88, Int88d, Int89b, Int89e, IS84a, IS84b, ISJ85, Int84a, Int85a, Joh83, Kle89b, Kle89a, Lew82b, Mar81, Mar82a, MSR87, Mer81, Mic84c, Mic85c, Mic87d, Mic87f, Num84c, Num88b, Obl85, Pag83, Pag88, Phi86, Res86, Rel83, Rel86, Rel88, Rat89]. guide [Rel89c, Rel89a, Rel89b, Rel89d, Rid82a, Rid82b, Sun85, Sun86a, Sun86b, Sun88b, Tan83a, Tan85a, TMjC81, Tew81, Uni82a, Uni83c, Uni84c, Uni86b, WB85, YS84a, YS84c, Zho89]. guide/release [Dig82b, Dig84e, Dig85c]. Guidebook [LW88a, LW89]. Guidelines [BBB^{+}83, Cor82b, Ear85, McA86]. gun [MH82].

H [Eve85, RB82]. HADY [EH81]. HADY-1 [EH81]. haesol [jYsS89]. Hajimete [TS88]. hand [KK89b]. hand-parallelizing [KK89b]. Handbook [CV88b, Fog85, Fog87, Fog88, Res86, Rit89, Sel83, Bin85, Guz87, Num83c, Num84d]. handling [JK83, Mal85, Num88a, Ott87]. Handprint [PB86]. hands [CC84a]. hands-on [CC84a]. Handtc [SP84b, SP84c]. Harcourt [KC84b]. Hard [BBF^{+}82]. Hard-Core [BBF^{+}82]. Hardware [Kna84]. Hardware-Testschnittstelle [Kna84]. harmonic [Rap82a]. harmonics [dB82b, dB84]. Harray [YHKM89]. Harris [Har81]. Hart [Dun87b]. Hartley [Bun86]. Harvard [Par86a, Par86b]. Harwell [Num88a]. HCPRVR [Gat82]. Heap [Kah80]. heaters [WP84]. HECLIB [Cha87b]. Heinemann [RB82]. helicopter [Red86]. help [Tha89c, RAKK88]. heng [hC83]. HENTRAN [Gui81]. HEP [BDJ^{+}89, Den82, DH84b, Dun86, HLM84, Kow85]. here [JH86]. HERMCOL [HMR85b]. Hermite [HMR85a, HMR85b]. Hessenberg [FW82, Ste76, VVV89b, VVV89a]. Hessian [CGM85b, CGM85a]. Hessians [vvHG87a, vvHG87b]. Heuristic [Bos88, Lau86]. heuristics [BCKT89]. Hewlett [Pol83]. Hidden [Sha87, Wit81, Col82, Sha89]. Hidden-line [Wit81]. Hierarchical [Bla87, Gal89]. hierarchical-memory [Gal89]. High [AG87a, AG87b, BD89, Fat82a, Fat82b, Hus84, JW86, Sam81, Adv86, Ano89, Bai87, DDDG89, Ell82a, Ell82b, FF84, Jes82, KRW88, Mul83, Rap82a, RW89, SDH84, Wit81, AG87c]. High-Accuracy [JW86]. High-Level [Fat82a, Fat82b, Adv86, Ell82a, Ell82b, FF84, Mul83, RW89]. high-lying [SDH84]. High-Performance [BD89, Bai87, DDDG89]. High-resolution [AG87a, AG87b, AG87c]. High-Speed [Hus84]. highly [PW84]. Hilfe [RAKK88, RS82]. Hill [BBB^{+}83]. histograms [vM84c]. histories [CB86]. History [AW82, McC84b, RL81, Sam81, WD81a, WD81b, Wex81, Bac81, Hig86, Mai87]. Hitachi [EML88]. Höhere [Hah81]. homogeneity [TBM85]. homogeneous [LKM88]. Homogenization [Sch83b, Sch83a]. Homogenization/Blending [Sch83b, Sch83a]. homotopy [Tod85]. HOOPS [Kli89]. horizontal [The88]. Horn [Min88]. HOST [SIR82a, SIR82b]. Houston [Mor82]. HP [Hew85]. HPFORTRAN [Sul88]. HQR3 [FW82, Ste76]. hsi [cC84b, hHtM81, cT81]. HSPF [AHU81, Don83c, Gri82, JK82, Uni84a, Uni84b]. HSSR [ZDS81b]. hsu [kH84, LcY83, yW85]. hsueh [Cha83]. hua [mT82b, fY84]. human [Bro83b]. Humans [DP81, DP84a, PDA83a, PDA86, APD86, DP84b, DP84c, PDA83b]. Hungarian [Sch84a]. Hutty [Bis81]. hybrid [Gra88, RAKK88]. hydraulic [Gic88]. hydrocarbon [Hig86]. hydrogeochemical [ZDS81a, ZSD82a, ZSD82b]. hydrographs [Uni81a]. Hydrologic [BK84, Cha87b]. Hydrological [AHU81, Uni84a, Uni84b, Gri82, JK82, War86]. hydroxy [She89b]. hypercube [IS84a, IS84b]. hyperfine [Nai84, Nai86]. Hypergeometric [GN89]. HYPOINVERSE [Kle89a]. Hypothetical [Jia86].

I/O [AS89a, AS89b, Hus84, Joh87b, Joh87c]. IAzyk [Osi82a]. iazyka [BKL89]. iazykami [Osi82a]. iazyke [SM84]. IBM [Ano82a, Noh84, AC86, ACG^{+}86, ALPC88, Ano87a, Ano87b, BW87a, CW85, CW88, CW89, CSZ86, Car88b, CB82, CDW83a, CDW83b, DW85, DW83b, Ehi82, Fuo86a, GSZ88, Int82e, Int83b, Int83c, Int86b, Int87b, Int88c, Wor84a, Joh86, Kir85, KM83, LB86, Lee84c, LOU86, LS88, Mar81, MS88a, McG87a, McG87b, Num84d, Pat89, Pee84a, RRS88, RV89, RMFG85, RG85, RB83, Rou83, Sof83a, Sam81, Ser89, Tei86, TPS^{+}88, Div85, Wor84b]. IBM-PC [LB86, Div85]. IBM-PC-based [RMFG85]. IBM/ [Ehi82]. IBM/Amdahl [CB82]. ICL [Gro87]. ICL-DAP [Gro87]. ICON [Dir84]. Ideas [Bee81]. identical [BS84]. identification [Uni86a]. identify [MSG86, Bur86a]. identifying [SD89]. IEEE [Fat82a, Fat82b]. IFIP [Wri89]. II [Fed81, Tea81, Bac81, CM81b, Hei84, Joh81, Kri83, MW84b, Num84b, Nai86, OO86, She89b, Wal3 ]. III [Ano88a, Bac81, Int85b, PRL^{+}85, SMD84, She89b]. ILLOD [NL85b, NL85a]. ILLOD- [NL85a, NL85b]. illustrated [CT88]. Illustrating [Alc82, Alc83a, Alc83b]. illustrations [Joh81]. illustrée [Str82, Str85]. im [Ger83, Hah81, San82]. IMACS [VV86]. image [Pra89, Pra89]. imaged [Per81]. Impact [Gre84, CKT85]. implement [Gra86b]. Implementation [ACK86a, Ash85a, Ban78, Col84a, DCHH87, DCHH88a, DBFK89, HO89, Kah80, MT82a, Pou87, Sha82, Sus86, Tha89a, Tha89b, ACK86b, Art81, Ash85b, Ash85c, BH89, Bur86b, Cha86a, Dav82, Gat82, Jia86, Ker82, Kie83, Lee85, MD88, Min88, Mye83a, Mye83b, NSB2 , PMBK82a, PMBK82b, RMFG85, Ril83, TFH86, Tod85, Wat86]. Implementations [vdV86, KTW84, Knu84, van86]. implemented [Joh86, MT84a, MMM85, PW84]. Implementierung [Mey84, Ste87, Sch84f].

Implementing [MR83, Gra81a].

Implications [Fat82a, Fat82b]. impress [TW87]. Impresslib [TW87]. Improve [CT86b, CT86a]. Improved [HL82c, PFF83, VLV^{+}86]. Improvement [Pie85, Bro82b]. Improving [Don82a, Don82b, Hon82]. IMSL [Ano84, Lib84a, Lib84b, Lib87]. in- [Mul83]. In-line [Wil87a]. inbreeding [Vu 89]. includes [Dig85b]. Including [BL83, Edg89b, Bal85, Edg89a]. Incomplete [DM87b, DM87a]. Independent [Bee85b, Ric86, SAB88, FF84, KJM89]. Index [EBS88, Dal89, Int82j]. indicators [BS86b]. indices [Bro85]. Indirect [DD86]. Industrial [IEC85, Kne81]. inelastic [CT88]. inequality [Zho89]. infiltration [BMS84, RMS82]. inflow [Red86]. Informatik [MF84]. Information [Ame85d, Com89, Cas81, Don81, IEC88, ISO88, Ame89b, KW89, MSA86, VVV89b, VVV89a, Ass83b, Ame85c, Ame85a, AC87, Ano82d, Hig86, Int82e, Int83h, Int84d, Int85e, Ame85b, Sch82b]. Informationsverarbeitung [MF84, Mau84]. informatique [Ano87d]. infrared [MW84a, MW84b]. ingenieria [Bor89]. INGRES [Rel86, Rel88, Rat89, Rel89c, Rel89a, Rel89b, Rel89d]. INGRES/ [Rel88, Rat89, Rel89a, Rel89b, Rel89d]. INGRES/EQUEL [Rat89, Rel89c, Rel89b, Rel89d]. INGRES/EQUEL/C [Rel88]. INGRES/EQUEL/FORTRAN [Rel86]. INGRES/QUEL [Rel89d]. INGRES/SQL [Rel89d]. Iniciacion [Dav85]. Initial [Cas89a, Cas89b, EP87a, EP89, Hig91, BMS84, CR84]. initial-value [CR84]. initialization [Per83b]. Initiation [Dub84]. inline [AJ88]. inner [SP84b, SP84c]. inner-zone [SP84b, SP84c]. INPROG [Smi85b]. Input [BP81a, Gab89, Gra81b, BP81b, Mei89b, OO86, Smi85b, Tho86b, Wil87a]. input/output [OO86, Wil87a]. Insecurities [CA86]. insertion [Zak84]. Instabilitäten [RAKK88]. instabilities [RAKK88]. installation [BS83, Dig82b, Dig84e, Dig85c, Int83i, Int85d, Int85f, Int88k, Int88l, IBM89a]. instantaneous [Joh85b]. Institutionalization [Ada84]. Instruction [Mil82b, Joh84, Le83, Lee85, MP81, Wei84, Wie82]. instructional [Sul2 ]. Instructions [BB82, DB82a, Ano85b]. Instructor [APD86, BS81f, BW84, BW87b, BGG85b, DH82, DP84c, Ett83a, HB81, Mar83b, MSR87, McC84c, MS88b, Moo82a, Nic85c, PDA83b, RZ89b, Wei86a, WB85, CM83b]. Instrumentation [Bli89]. insurance [McK83]. INTCOL [HMR85b]. Integer [BW89]. Integral [BA85b, BA85c, BA85a, Sha87, Sha89, Bar89, Bau88, dB82b, dB84]. integralnykh [Sko88]. Integrals [AB89, Car87, Car88a, Car89b, EBS88, Gen82, KP86]. integrated [ALPC88, Tho84b, Tho84c, Tho84d, Wat82a, Wat82b]. Integrating [Gol81, Col89a, Col89b]. Integration [SSS84a, SSS84b, SSS84c]. INTEL [HK87a, BH89, Ess88]. INTEL8087 [HK87b]. Intelligence [Gus84, IEE88a]. intensities [Tho84d]. intensity [Tho84b, Tho84c]. Interaction [PLR85, Dal88b, Nai86]. Interactions [EW87, vM84c]. Interactive [AL81b, CC84a, GM83, McC81, McC86, Sis85, You82, Ack84, AL81a, ADP88, BKK^{+}89, Gre88, Har86b, Int82b, Int82c, Int82d, Int84d, Int85e, Int86e, Int87c, Int87g, Int88e, Int88m, IBM89b, Kir85, Nor83, PS82, PS83, Smi88d, SDC82]. Interbattery [HPB82]. Interchange [AK84]. Interchangeability [RH84b]. Interface [And84b, BHY80, Did86, Dix85, GE85, Fun86, Hue83, Nor83, SIR82a, SIR82b, TW87]. interfaces [Int88i, Int88j, Int89c]. Interfacing [All90, Egg83]. interfering [Wal85]. Interim [MSG86, Tan85b].

Intermediate

[Pem83, SW83, TvSS82, BG84]. internal [Car88b, Cra83, RS87, DJM87]. International [ACM89a, Ano88a, Ano88b, Cou85a, Gia89, Gon89, Ass86, IEE88a]. interplanetary [Col89a, Col89b]. Interpolation [Dur80, Ren84, Uni84d]. Interpolatory [EK87b, EK87a]. interpret [vM84d, vM84f]. Interpretation [Boe87, For85, Uni83a]. Interpreter [OK87, Chi85b, Rom81, Zim86]. Interpreters [ACM87, Wex87]. Interpretive [ACM87, Wex87]. Interprocedural [ACK86a, CCKT86, CK88, Har89, ACK86b, CKT85]. Interprozedurale [Wid88]. Interrupted [Wil83]. Interval [Moo88b, Dun87a, DZ88]. interview [Tay86]. Introduccion [CS84, CM89, HRH81, HRC89, FK84]. introducing [SL82]. Introduction [AHH89, Ban88, Cas81, DF89, Don81, Dyc81, DLS84a, GY82, Gra84a, Gus84, HR83a, HRC87, HYP87, Hud88a, Hud88b, Key81, KC84a, KC84b, Lam86, Law83, McD83, McD85, Moo82b, Moo82a, Moo88a, Rou86, SM88a, SM88b, Spe83, Sto85a, Bol89, DcF89, DLS84b, Oli81, Tuc86]. Introductory [Der82, LD87, Num84c, Num88b, Ric82a]. invariant [MSG86]. Inverse [DM87b, DM87a, HS86]. Inversion [GGLM88, GL90, HS86, Web85a, Web85b]. investigation [RAKK88, VD84]. inviscid [Tho81, Tho82a]. Invitation [McN83]. inzynierskiej [Rzy84]. ion [PFF83]. ion-Ure [PFF83]. IPMIXD [Ano87c]. iPSC [AS89a, AS89b, BH89, Ess88]. iPSC/ [BH89]. iPSC/1 [Ess88]. iPSC/1-VX [Ess88]. iPSC/2 [AS89a, AS89b]. IQPACK [EK87b, EK87a]. Isaac [La 87]. isarithmic [Cla89]. Ising [CM86, DM89b]. ISO [Rei89c, IEC85, IEC88, Wic89]. ISO-Pascal [Wic89]. ISO/IEC [IEC85, IEC88]. ISO/WG5 [Rei89c]. isocomp [JR81]. isolated [HL82a, SPS84, VMS81]. isolating [Dal88b]. isolation [JBJ84]. isostatic [JR81, SJB83a, SJB83b]. ISSAC [ACM89a, Gia89, Gon89]. issues [DSCP88]. iterates [Zho89]. Iteration [CC87]. Iteration-level [CC87]. Iterative [ET86, GKRY82, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Gin82, GQ88]. ITPACK [GKRY82, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d]. IV [Ano82a, VMS81, AL81a, AL81b, AEL^{+}86, Aya84, CK86b, hC83, CwL83, Col83, Col82, DM87c, Dre81, Ehi82, FGH81, Got84, GB83, Gra86a, Gra81a, HRH81, HR83a, Hei84, Hel83, HF81, Hon85, Hur82, aHH83, Int82f, Int83b, Int83c, Int86b, Iwa84, JL81a, JL81b, Key81, Kha81, Kip82, LCMM88, Lav83, MRS84, Man82, Mar81, McC81, Rod87, O’N81, PMBK82a, PMBK82b, Rod86, RMS82, SR84a, SDH84, TMjC81, Tew81, Uni81a, Wu82a, Wu82b, Wu83, hYsA82, Zwa81, Zwa85]. IV-GALCYCL [Gra86a]. IV-PLUS [Ano82a, Ehi82]. IV.0 [BGCS82].

J [KM83]. Jackson [Rei89c]. Jacobian [CGM84b, CGM84a]. Jacobians [vvHG87a, vvHG87b]. JAKEF [Hil82a, Hil82b]. Jazyk [Osi82b]. jazykami [Osi82b]. jen [hK85]. jezyku [Rzy84]. JIS [Mor84]. John [Edm86, Rei87b, Rei87d, Rei87c, Rei89c]. Joint [Usr82]. Journal [BR89b, For89, Ano81a]. Jovanovich [KC84b]. JOVIAL [Sch82a]. July [Gon89, Sof83b, Usr82]. June [Dav89, Mor82, Sof84, Van84a, Wex81, Wex87].

KAP [LCH^{+}88]. kappa [AM89a]. Katzakidis [BD80]. KERMIT [Col84a]. kernal [Ame85a, Ame85b]. Kernel [Ame85c, Ame85d, Bro84a, HWS^{+}88, IEC88, ISO88, HMB^{+}88, Int88q, Int88r]. KERNELS [VMS81, McM86, MSG86]. Key [Bur84c]. Keyboard [DF89, DcF89].

Keyword [Gra81b, Tho86b]. Keywords [Ham85, HM90, RH84a]. kihon [Ton82]. kill [Aha85b]. Kind [Car87, Car88a, Sch89f]. kinetics [BDS84, KJM89, LKM88]. kinship [Vu 89]. kipon [BBuC84]. kipop [hA84]. Kit [Sym85, Sym86, Sym88, Dig84j]. km [SJB83a, SJB83b]. know [Bro81b, Col87b]. Knowledge [DK84, Cre89]. knowledge-based [Cre89]. Ko [hK85, Cha83]. Konfidenzintervalle [Sch82c]. kou [mT82b, fY84]. kraevykh [Sko88]. Kryptographie [MF84]. Kung [Cha83]. kurs [HPR81]. kyesan [mK84].

L [Hos88]. L-moments [Hos88]. label [ZSD82a]. Laboratories [Hue83]. Laboratory [Noh84]. LABPLT [ZSD82a]. Lader [Hei83]. Lagrange [Gen82]. Lake [Sof84]. laminar [Kee88b]. laminate [Son83]. Langage [AFN83, Dav84a, Ass83b, Lig82a, Lig84, Lig85a, Lig88a]. Langages [Ber82a]. Langley [HL82b]. Language [ADH^{+}89, BB82, Com89, DB82a, Egg83, Fat82a, Fat82b, GJ82, Jap82, Lei87, LS85, MT82a, Ric86, SAB88, Ame87a, AC87, Ame87c, Aha85c, All87, Apo86, BGM83, BS81g, BG84, Bur84a, Con83c, Con84, Con85b, Con87a, Con87b, Con88a, CRV^{+}89, Dig82d, Dig82c, Dig83, Dig84g, Dig85d, Dig86b, Dig88a, FF84, Gol81, Hur82, Int82h, Int83c, Int83e, Int85g, Int86f, Int87b, Int87d, Int87h, Int88f, Int88n, Int89a, Int89d, Int81e, IEC88, ISO88, Int88q, Ame89b, Joh86, Lah88b, Lah88c, Le83, Lee85, Mic83, Mer81, Mer85, Mic87d, Roy88, SAS86, Smi83c, Sne88, Sol89, Int88r, Tei86, Tha82, Vag89, Wag84, Wan84, Wee86, Ame87e]. language-sensitive [Dig85d]. Languages [Bro84a, GPKK82, GPKK84, Hor83a, Hor83b, ML87, PS81, Pra84, Rao82, POP82, Sam81, SAN^{+}81, Fog85, Fog87, Fog88, Mul83, PZA86, Res86, Rao81b, Wex81]. Laplace [GGLM88, GL90, HK83]. Laplasa [Sko88]. Large [AEV89, Bla87, Coc83, GKRY82, HWS^{+}88, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Mar84b, Rei84b, Rey80, Sch88a, dEV89, BS83, GL81, GDK89, Gui88, Hon85, LN89, Rei84a]. Large-Scale [HWS^{+}88, LN89]. laser [Owe87]. Latin [IS84a, IS84b]. Lattice [HL82b, HM81, HM84]. lattices [DM89b]. layers [EH81]. Layout [MF84]. LDEC [MSM84]. learned [Bro89a, Bro89b, Bro89c, Bro89d]. Learning [CSD83, Gee86, Gre85, Int86c, Bel84, Can81]. Least [GHM^{+}86, TU81]. Least-squares [GHM^{+}86]. lecture [Joh84]. lectures [Pet87]. Left [Vel82]. Leipzig [Dav89]. length [CK86a, HS86]. lenguaje [CS84, FK84]. Lessons [Bro89a, Bro89b, Bro89c, Bro89d]. Let [Aha85b]. letter [CF85, Eve84, Knu84]. Level [Fat82a, Fat82b, Mar81, Sam81, Adv86, Bur85b, Bur85d, Bur85c, Bur86c, Cha86a, CC87, Ell82a, Ell82b, FF84, Gra84b, Hon81b, Mul83, RW89, Spe82, Spe85]. level-surface [Gra84b]. levels [SDH84]. Levin [Gro89]. Lexically [BGS82]. lexicographic [Wil83]. Lexikon [Ano82a, Ehi82]. LFP [CBS81, CB82]. li [Rai84]. LIB99 [Con86, Con87c]. libraries [BS83, CM81c, IMS87a, IMS89a]. Library [Ano87a, BD89, Egg83, GMPW79, HL86b, IMS89a, IMS89m, MAT89a, MAT89b, Mor81b, Mor90, Mor93, Num83a, Nag81a, Nag85, Phi87, Ser85, Ser89, Woo89, WLO76, Adv86, Ano82d, Ano84, BJ81a, BJ81b, BJ84a, BJ84b, Cha87b, Cra89b, Dig85e, GtTB89, DDDG89, Fra84b, GC84, HL86a, Hof87, HP89, HP88, Int81c, Int81d, Int82i, Int83f, Int83h, Int84c, Int84d, IBM85, Int85c, Int85d, Int85e, Int85g, Int86d, Int86f, Int87d, Int87h, Int88f, Int88n, Int89a, Int89d, IMS82, Lib84a, Lib84b, IMS84, IMS87g, IMS87e, IMS87j, IMS87h, IMS87i, IMS87l, IMS87m, Lib89a, Lib89c, Lib89b, Jac85b, Jac85a, Lee85, Lio85, Num83b, Num83c, Num84a, Num84d, Num87, Num88a, Num88b, Num88c, Ott87, Pee84b, Pee84c, Pee84d, Pee85d, Pee85e, Pee86, Pee89, Pay84a, Pay84b, Phi86, dR87, Som86, TW87, dZ86, IMS87a, IMS87c, IMS87b, IMS87d, IMS87f, IMS87k, IMS87n]. LIBRARY [IMS89b, IMS89c, IMS89d, IMS89e, IMS89f, IMS89g, IMS89h, IMS89i, IMS89j, IMS89k, IMS89l, IMS89n, IMS89a, MS88a, Pee84a, Whi89]. library-based [Lee85]. life [Ols83]. life-time [Ols83]. lifecycle [Not89]. lifting [Wal85]. Like [HL82c, Whi81b, Whi81a]. likelihood [IA89, Mai81a, Mai81b]. Lincoln [CBS81, CB82]. Line [PB86, BR89a, Col82, HS86, The88, Wil87a, Wit81]. line-length [HS86]. Linear [Abd80, BD89, Cal86, Cra86a, DFK83a, DFK83b, DFK88, DG82, Dod83, Don83a, DS84, DDHH84, Don84a, Don84d, Don85a, DCHH87, Don87b, DCHH88a, DCHH88b, DCHH88c, Don88b, Don89, DR82, ET86, GS81, GHM^{+}86, Gre86, GKRY82, HK87a, HK87b, Hop81, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, LHKK79a, LHKK79b, LN88b, LN88a, Mar82b, Rei84b, Sme81, Wol85, Ada89, ADP88, BT88, CK88, CMM^{+}88, DDDG89, DFK81, Don83b, DH84b, Don84b, DCHH85, Don85b, Don87a, Don88c, DPA87, Fra84a, HL86b, Hof87, HP89, LN87, Mel88, Min88, MSG86, Num88a, O’N81, PM87, RRS88, Rei84a, dR87, SS82, Tod85, Web85a, Web85b, Zho89, ADP88, DPA87]. Linear-time [GS81, Min88]. lines [EIT85]. linguagem [Cad84]. linguaggio [SS87b]. link [GDK89]. LINPACK [DS84, Don88a]. liquids [VH87]. Lisp [LH88, Boy84b, KS81b, KS81a, NSB2 , Ols83, Sch82a, BGS82]. LISP-based [KS81a, KS81b]. List [RAKK88]. listing [War86, ZGS89]. Livermore [Hug84, McM86, hTD88]. Load [Dav84b, CB86, Rod87]. Load-and-Go [Dav84b]. Loan [Rit89]. Local [Cal86, PLR85, GS81]. Local-memory-based [Cal86]. locality [MRS84]. location [DJM87, Whe84a, Whe84b, ZSD82a, ZSD82b]. locations [Boy84a, GS81]. log [IA89]. logarithmic [O’N81]. Logic [Gol82b, Gus84, Lew81b, WN87, Lew81a]. logiciel [LB89]. London [Bis81, RB82]. Loop [AK84, Gal89, LO85b, LO85a]. looping [Tay84]. Loops [CT86b, FN85, HL82c, SP82, CC87, CA86, CT86a, hTD88, OM82]. Loosely [WN87]. Lösung [RAKK88]. low [Lec89]. LSE [Ber82a]. LSOLVE [HMB^{+}88]. LSQUMNS [dR87]. LSSOL [GHM^{+}86]. LTUR [Min88]. Lucas [KTW84]. LUCY [MD88]. lughat [McC85c]. lying [SDH84].

M [Cou85a, Cou85b, Nad86, DH84a, Tho84b, Tho84c, Tho84d, Uni81b]. M-D-test [Tho84b, Tho84c]. M-Diff [Tho84d]. M77 [DH82, DH83, Uni83b]. MA [Rei89c]. MA27 [DR82]. MacFortran [Ano89]. Mach [Bol89]. Mach/EPEX [Bol89]. MACHAR [Cod88]. Machine [Bee85b, Cod88, SS88, AP87, FF84, HM82, IBM88, Int88d, Sor84, AKLS88]. Machine-Independent [Bee85b]. Machines [AE87a, GPKK82, GPKK84, AE88, FSO89, Lan88, Rao81a, Rao83, AE87b]. Macintosh [Num88d, Pre87a, Pre88a, Pre89, VN89b]. Macmillan [Bis81]. Macro [JRS88a, JRS88b]. macros [BH89]. MACSYMA [RAKK88, SR86]. MACSYMA-FORTRAN [RAKK88]. MACSYMA-FORTRANHybrid-Codes [RAKK88]. Made [CS83]. Madrid [KTW84]. maenyuol [BBuC84]. magnetic [BJ81a, BJ84a, Web85a, Web85b, vM84b, Nai86]. magnetization [Gra86b]. Magneto [MHK86]. Magneto-Plasma [MHK86]. Magnets [GKKY89]. main [Gic88]. mainframe [Roc86]. maintain [BJ81a, BJ81b, BJ84a, BJ84b, Jac85b, Jac85a]. Maintenance [JK82, Kah80, BK84, BS83, Sel83]. Making [Flo89, SDH84]. man [Sas81]. Management [Bla87, DK84, MSA86, Don83c, OO86, SDC82, VLV^{+}86]. Manager [Dix85]. Manchester [JR89]. mandy [SMD84]. manipulating [Kle89b, Kle89a]. Manipulation [Bee85b, RT85, Hon82]. manipulator [Kir89]. manipulators [SL82]. manoeuvrability [Kip82]. Manteuffel [Ash85a, Ash85b, Ash85c]. Manual [Sof87, DFD81, DFD84, Dir81, Lib84a, Mai87, Mar83b, Num83a, NL85b, AHU81, All87, APD86, Ano82b, Ano83, Ano84, Ano85c, Ano87a, Ano87c, ADP88, BGM83, BS81f, BMS84, BGG85b, Bur81b, Bur84b, Bur84a, Bur85d, Bur85c, Bur86c, Con81b, Con82b, Con83a, Con83b, Con83d, Con85a, Cha87b, Con83e, CBS81, CB82, CM83b, Cra83, Cra84, Cra86b, Cra89a, Cra89b, Dig82c, Dig83, Dig84b, Dig85a, Dig85e, Dig86b, Dig86c, Dig88a, Dig88c, Dat81, Dat84, Dat85a, DH82, GtTB89, DP84c, DW83a, DW84, DJM87, DPA87, Dun85b, Enc87, ETA88, Fed82b, GRB88, Gic88, Gri85, Hew85, Har81, Har85, HL86a, HL86b, Hof87, HP89, HB81, Hon85, Hua82, Int83b, Int84b, Int85b, Int88a, IMS82, Lib84b, IMS84, Lib87, IMS87c, IMS87b, IMS87d, IMS87g, IMS87j, IMS87l, IMS87m, IMS87n, IMS89c, IMS89e, IMS89g, IMS89h]. manual [IMS89m, IMS89i, IMS89j, IMS89k, IMS89l, IMS89n, Iwa84, JSW85b, Ket85b, La 87, Lag85, Lah88d, Lio85, Met89a, Mic83, Mai81a, Mai81b, McC84c, MS88b, Mic84d, Mic85e, Mil87b, Moo82a, MSG86, Mul85, Num83b, Num84a, Num84b, Num84c, Num87, Num88a, Num88c, Nag81b, NL85a, Nic85c, NL85d, PDA83b, PF85, Rel89b, Rel89d, RZ89b, dR87, RMS82, SAS86, Som86, Sun87, Tan85b, Tho81, Tho82a, Uni81b, Uni83b, Uni84a, Uni84b, Uni81a, Uni84d, Uni84e, Uni88, VMS81, VL81, Wei86a, dZ86]. Manuel [Ano87d, TR84]. Many [Maa89, Bur86a]. many-to-one [Bur86a]. maple [PM87, Gro89]. Maps [GM83, BT83, Cla89, ZSD82a, ZSD82b]. March [RW86]. Marching [Ban78]. Mark [Num83a, Num88b, Num81, Num84b, Num85a, Num88c]. Martin [Mil82b]. Marwick [RB82]. März [Wös82]. Maschinen [Sto84a, Sto84b]. Maschinen-unabhaengige [Sto84a, Sto84b]. Mass [Con81b, Iwa84]. masses [Iwa84]. master [Int82j]. masters [Nic85c]. mastery [EGP81]. mastery-based [EGP81]. maszyn [Rzy84]. Matching [Bur81a, HS81, BD80, GS81]. matematicheskomu [Aka88]. Materialy [Aka88]. Math [IMS89a, Lib89a, IMS87a, IMS87b, IMS87g, IMS87e, IMS87f, IMS87n, IMS89b, IMS89c, IMS89h, IMS89i, IMS89j, IMS89k]. MATH/ [IMS87g]. Math/Library [IMS89a, Lib89a, IMS87a, IMS87b, IMS87f, IMS87n, IMS89b, IMS89c, IMS89h, IMS89i, IMS89j, IMS89k, IMS87e]. MATH77 [MAT89a, MAT89b]. Mathematical [Air77, ACG^{+}88, Cow84, DR86, MAT89a, MAT89b, Woo89, IMS82, IMS87a, IMS87b, IMS87g, IMS87e, IMS87f, IMS87n, IMS89a, IMS89b, IMS89c, IMS89h, IMS89i, IMS89j, IMS89k, Num84d, Pet88]. Mathematics [Mor81b, Mor90, Mor93, Ano84, Cor81, Cor82c, IMS84, IMS87a, IMS89a, Pet88, VV86]. Mathematik [EmR84]. Matrices [CGM84b, CGM84a, CGM85a, Cra86a, Sch88a, CHPS85, CGM85b, Fra84a, Mel88, Par84]. Matrix [CL83, CV88b, Don82a, Don82b, DM84, DBFK89, FW82, Hig88b, Hig88a, Hig89, Rit89, Ste76, Bru86, Kem87, Num88a]. maximization [IA89]. Maximum [LS88, CMM^{+}88, GG88, Mai81a, Mai81b]. May [IEE81, Rei89c, RH84b, Com89, La 87]. McFadden [RWA84, Rod84, WAD81]. McGill [Pie85]. means [RH84b]. measured [Tho84d]. Measurements [BS84, Hoc85, Tem89a, Tem89b]. mechanical [AM89b, CT88, VC89]. mechanics [CV88a]. median [Eva81]. medium [vM84a]. medium-scale [vM84a]. medizinische [MF84]. Meeting [Rei89c, Rei87b, Rei87d, Rei87c, Rei87e, Noh84]. Meetings [AW82, BBB^{+}83, Noh84]. MegaFrame [Con83e]. Meiko [BDJ^{+}89]. membrane [PS82]. Memory [Ana87, BDS88b, BDS88a, AP87, BDS88c, Cal86, Dob85, GGJ^{+}89, Gal89, Jor86, Ols83]. Message [BCM87]. Message-Based [BCM87]. Messung [RS82]. Metallen [RS82]. MetaWINDOW [Met89a]. Metcalf [Nad86, Wil87b]. meter [Tat87]. Method [Coc83, GGLM88, GL90, HYP87, KW87b, KW87a, MHK86, Gra86b, Gui88, HM81, HM84, Hos88, Kem87, kK89c, MT84a, Rod87, PJ84, Sch84a, TU81, ZGS89]. méthodes [FS86b, GB83, GB89]. Methodology [Isn82, DSCP88, Kre88]. Methods [Ano88b, AHH89, BZ85, Bor85a, Bur81a, DM87c, EP87a, EP89, GKRY82, JSW85a, KMN89, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Moo88b, Sha82, Sme81, BD80, Car89a, CB86, Des89, GG88, Her88, HP88, JSW85b, Ngu81, Ric82a, SS82]. metodom [Sko88]. Metody [BZ85, HPR81]. Metric [BSP83]. metrics [CSD82]. MGD1V [dZ86]. MGD5V [dZ86]. MGZEB [Lio85]. Michael [Tay86, Wil87b]. Michigan [IEE81]. micro [Col87b, Hey85, McC86]. Microcomputer [Oni85, PP82a, YK88, PP82b]. Microcomputer-Based [PP82a, PP82b]. microcomputers [LB86, Rod86]. Microeconomics [Aya84]. microelectronic [Mat83a, Mat83b]. Microfortran [Dir81]. microNet [CDW83b]. microprobe [Hue83]. microprocessor [Rom81]. Micros [Mar83d, Mar83c]. microscopic [vM84c]. Microsoft [All90, Chi81, Chi86b, McG87a, McG87b, Sul88, Woo89]. microstrip [Ric82b]. Microtasking [CSZ86]. MicroVMS [Dig84b, Dig84c, Dig85b]. MIDAS [MSA86]. mikroprogrammierten [Wis81]. Mikros [MF84]. MIMD [Jor86, Kow85, RS85, Sol89, Sor84, ZBG88]. MIMD/SIMD [Sol89, ZBG88]. MINCLC [FMH85]. Mine [Nor84]. mineral [Boy84a, FMH85, Uni86a]. Minerals [Sch83b, Sch83a, Bod87, Kie86]. mini [Hey85, Num84c, Nag81b]. MinIdent [Uni86a]. Minimal [YM85]. minimization [LN89, YM85]. minimum [BT88, Wat82a, Wat82b]. minisupercomputers [WSL88]. Minnesota [VL81, Wex87]. MINTDF [KP86]. MIPROPS [McC86]. Miscellaneous [Ano81a, Lag85]. mixed [Mic87d, Wic89]. mixed-language [Mic87d]. Mixing [McG87a, McG87b]. MM2 [She89b]. MMLE3 [Mai81a, Mai81b]. MNF [VL81]. mo [Lee84a]. MODCOMP [CBS81]. mode [Iwa84]. Model [Ano87e, DK84, DCHH87, DCHH88a, Hol87a, Hol87b, Jai84, Sch87a, Art81, BNZ87, CM86, DGNP88a, DGNP88b, DM89b, Gic88, IA89, KWWK86, PMBK82a, PMBK82b, War86]. modeling [DJM87, Gla88, Gre88, Kee88b, Pet88, SL82]. Modellaufbau [Sch85a]. Modellbil [Dah81]. Modellbildung [Sch82e, Sch84d, Sch84e]. modelling [FS86a, Tat87, VP84]. Models [ET86, ADP88, BDS84, DPA87, IS84a, IS84b, Lag85, Mul85, Not89, PF85, Rin83, Tan81b, vMF81, vMF84]. Modern [Wan85, Wor88, Wor89a, Wor89b]. Modified [BBB^{+}83, DFK83a, DFK83b, DFK88, Par86a, DFK81, Par86b]. modifier [Jia86]. modulad [LB89, (??87, (??84, Tri84, Tri89]. modulation [Ber84a]. Module [BBB^{+}83, Ull85, Ull86]. modules [McG87a, McG87b]. moduli [Gui88]. Modulnykh [MK86]. moisture [SG88]. molecular [Bar89, GDK89, Sch89a, She89b, VSH83]. molecule [Nai84, Nai86]. moment [Fra84a]. Momentenmethode [Sch84f]. moments [Hos88]. momentum [CDW82, CDW84, RV84, RRC89]. Monitor [Hei83]. Monitors [LO85a, LO85b]. monochromatic [Wyl86]. monolayer [She89b]. Monte [Joh85b, NE81, VH87]. MONTEC [PS83]. Montréal [Van84a]. morphometric [Wal3 ]. Most [DE84]. mostly [JH86]. Mothra [Sof87]. Motion [Chu88]. movement [Col89a, Col89b, KWWK86]. MP [Bre78b, Bre79, Bre81, Hoc85, RS85, hTD88, Tem89a, Tem89b, ZM86]. MP/24 [AGS88]. MPD [RAKK88, RAKK88]. MPD-engines [RAKK88]. MPD-Triebwerken [RAKK88]. MPQS [LtW88]. MPQS-factoring [LtW88]. MS [JC88, JC89, Mic84c, MZ84, Mic85c, Mic85b, Mic86, Mic87c, Tha89c, Div85]. MS-DOS [JC88, JC89, Mic84c, MZ84, Mic85c, Mic85b, Mic87c]. MS-FORTRAN [Tha89c]. MTS [CW85, CW88, CW89]. mu [ZGK88]. mudpack [Ada89]. Multi [Aha85c, Dav86, Gui88]. multi-effect [Dav86]. Multi-language [Aha85c]. multi-moduli [Gui88]. multicomponent [Kee88a]. Multicomputer [Cro87a, Cro87b]. Multicriterion [Osy84]. Multigrid [Ada89]. multiphoton [BSdlT87]. Multiple [Bre78b, Bre78a, Bre79, BHY80, Haa87, DGNP88a, DGNP88b, DV81, Gen82, JT88, Tho84b, Tho84d]. Multiple-Precision [Bre78b, Bre78a, Bre79]. Multiprecise [BW89]. multiprocessing [ABC^{+}88]. Multiprocessor [Dun86, HS81, KRW88, GGJ^{+}89, Gal89, GSZ88, PJ84, RRS88]. Multiprocessors [CSZ86, GL86, HKP88]. Multiprozessorsystem [Hel85b, HS81]. Multitasking [Bro84a]. multivariate [Mul88]. München [Wös82]. MUNIT [MSG86]. Musaid [aHH83]. music [Moo86, Ion84]. Muster [San82]. Mustererkennung [MF84]. Mutant [ADH^{+}89]. Mutation [DAG^{+}88, OK87, Tan81a, Tan82]. mutual [Bai89]. MV [Kir89]. MVS [Bin85, Int88k, Int88l, IBM89a].

N [Uni81b]. n$-$HLF$/ [Tem89b]. nach [Sch84f]. Nag [Phi87, HP88, Num81, Num83c, Num84b, Num84d, Num85a, Num88b, Phi86, Whi89, Num83b, Num84a, Num84c, Num87, Num88c]. Name [GM83]. Naming [Boo82]. nao [Cha83, hK85, iL82, fS82]. NASA [Bro89a, Bro89b, Bro89c, Bro89d, HL82b]. NASA-Langley [HL82b]. NASA/ [Bro89a, Bro89b, Bro89c, Bro89d]. Nassi [Sis85]. Nassi-Schneiderman [Sis85]. NATFREQ [Iwa84]. National [Com89, DH82, DH83, DH84a, Mor82, LZ82]. natural [Iwa84]. Naturwissenschaft [McC85a, McC85b]. Navier [Gro87]. Navy [Gro82]. NBS*AIDS80 [MHS81]. NBSGSC [TPR85]. NCAR [CM81c]. NCC [AW82, Smi84]. NEC [LtW88, Wat87]. need [Col87b]. NESS [KBRM^{+}86]. Nested [FN85, OM82, SP82, Sto85c, Ber88b, Uni83a]. Network [Bla87, ZM86, BT88, GG88, Gol81, Sim88a]. network-structured [Gol81]. Networks [NL85b, Tsu85, NL85a]. Netzen [Hof84]. Neues [MF84]. neutron [CSC^{+}86]. Nevada [ACM89b, Ano83, Ket82]. Newton [TU81, DM89a, La 87, Rod87]. NF [Ass83b]. nine [Sun88a]. Ninth [POP82]. NLPQL [Sch86]. NMFECC [Fon85]. NMR [vMT84, vM84e]. No [Cou85b, KW84, RB82, Wil87b, Cou85a, Edm86, Int83b, PB84, TS88, Uni83a]. no-recoil [PB84]. noise [Fra84a]. NOLLSI [TU81]. nombreux [GB83, GB89]. non [Adm84, Adm85, Wyl86]. non-scientific [Adm85]. non-scientists [Adm84].

Nonalgebraic [SL82]. Nonlinear [KKRK85, LG86, MHK86, MC80, ZM86, Dun80, DZ87, Dun87a, Gil86, IZP81, Num88a, Sch86, TU81, Tod85]. nonparametric [TSU88]. nonrelativistic [VC89]. nonrotating [Dul81]. Nonstiff [Cas89a, Cas89b, Hig91]. NOR [NL85a, NL85b, NL85a]. NOR-B [NL85a, NL85b]. Norm [Blu78, Hig88b, Hig88a, Hig89, PFF83]. Norm-Conserving [PFF83]. Normal [WN87, Bau88, Sun88a]. normalisation [Tri84]. Norme [Ass83b]. normes [LB89]. norms [CT88]. Norway [VV86]. NOS [Con83c, Con84, Con85b, Con85d, Con85c, Con86, Con87c, Con87a, Con87b, Con88a]. NOS/VE [Con83c, Con84, Con85b, Con85d, Con85c, Con86, Con87c, Con87a, Con87b, Con88a]. Note [WH87]. Notes [KW89, Div85, Dig82b, Dig84e, Dig85b, Dig85c, Smi81, Smi83a]. Nov [Rei89c]. November [ACM89b, IEE88b, ML88, Rei87d, Rei87e]. NPSOL [Gil86]. NS32000 [Cod86b]. NSPIV [She78]. NSWC [Mor81b, Mor90, Mor93]. NSWC/DL [Mor81b]. nuclear [SMD84, SR84a, SDH84]. nuclei [Nai86]. nucleic [LB86]. nucleon [SDH84]. null [BHK^{+}85]. Number [Gui89, Haa87, Pie85, Sch79, SP82, GS81, GDK89, GQ88, Ras84, Vu 89]. number-theoretic [GQ88]. Numeric [BB83, GKKY89, HK87a, HK87b]. Numerical [AHH89, BZ85, Boe87, BKK^{+}81, Bor85a, DM87c, GGLM88, GL90, HL82a, HP88, Hus84, JSW85a, KKRK85, KMN89, LG86, MC80, Num85b, Num85c, Num86, Num88d, Num89, Ngu81, PFF83, PFTV86, Pre87b, Pre87a, Pre88b, Pre88a, Pre89, SSS84a, SSS84b, Vet85, VTPF87, VTP89, VN89a, VN89b, VV86, Wan85, AM89b, Des89, Gro89, HM82, JSW85b, McM86, Mor81a, Pet83, Red86, Smi83c, Wan86, Gra86a]. numerico [Bor89]. numérique [Gra86a]. numériques [FS86b, GB83, GB89]. Numerischen [EmR84]. Numvec [Som86, GtTB89, HL86a, HL86b, Hof87, HP89, Lio85, dR87, dZ86]. nutritional [BS86b]. NYU [Gre86]. nyumon [Mor84, TS88].

O [AS89a, AS89b, Hus84, Joh87b, Joh87c]. Oak [ZDS81b]. obespecheniiu [Aka88]. object [Gol81, RW89, Pou87]. object-oriented [Gol81, RW89, Pou87]. objective [LN89]. Obrabotki [MK86]. observations [Mul83]. observed [Tho84b]. obtain [MRS84]. Occurring [AB89]. October [Ass86]. ODESSA [LK88]. OEHPC [GtTB89]. Oil [ET86, Kre86a, BT83, Kre86b]. O’Leary [Tay86]. olefins [She89b]. Olson [WF85]. OLYMPUS [HR83b, HR83c, CR84, HR84a, HR84b]. On-Line [PB86]. on-resonance [SDH84]. on-screen [Tha89c]. One [Hig88b, Hig88a, Hig89, Bur86a, Dun88b, Kee88b]. one-dimensional [Kee88b]. One-Norm [Hig88b, Hig88a, Hig89]. one-sided [Dun88b]. only [GS81]. Ontario [Sof83b]. operands [Est82]. operating [Ant81, Con83a, Con85a, Int88c, Mic84c, Mic85c, Mic85b, Mic87c, Pay84a, Pay84b]. operation [Kar87a, Kar87b]. Operations [Cro87a, Cro87b, Int84b, Int85b]. Operators [ADH^{+}89, Bru86, CDW82, CDW84, CSD82]. Opinion [GPKK82, GPKK84]. Optimal [Gaf83b, Gaf83c, Gaf83a, NL85b, Ull85, Ull86, NL85a]. Optimalnykh [MK86]. Optimierung [Jac82]. Optimising [Sch89c, Uni82c]. Optimization [DDH84, GMPW79, LtW88, Maa89, Met82, Met85b, Mon89a, MGH81b, MGH81a, Pem83, TvSS82, CKT85, Osy84, Sim88a, Zho89, Wil87b]. Optimizations [PW86]. Optimized [TWI88, Sch89a]. Optimizer [GMW86, JT88]. Optimizing [BGS82, JH86, Sch89b, GF89, Mic87d, Mic87f, Mic87e, Mic87a, Mic87b, Mic87c, Mic88, Mic89c, RG85, WM85b]. option [LD87]. Optional [GF81, FCG83, Gre88]. options [ZGK88]. Orbit [AB89]. Order [Cas89a, Cas89b, Hig91, Bru86, Col84b, Wil83]. Ordering [FW82, Ste76]. Ordinary [Cas89a, Cas89b, CC82, Gaf84, Hig91, LK88]. Ordinateurs [Don84c]. ore [Lag85, Mul85]. oreand [PF85]. OREDIT [ZDS81b]. Oregon [Gon89]. ORGT [Tho84e]. orientation [Tho84e]. oriented [Cal86, Gol81, RW89, Sav87, Tea81, Pou87]. orientiertes [Dah81]. Orlando [IEE88b, ML88]. ORSIM2 [HPB82]. orthant [Sun88a]. Orthogonal [HPB82]. OS/VS [Int82f]. oscillating [Wal85]. Oscillatory [Gaf84]. Oslo [VV86]. osnoveFortrana [HPR81]. osnovy [BKL89]. Other [GPHL90, Rao82, Bod87, Guz88, GPHL88, Rao81b]. outil [TR84]. Outline [KC84b]. Output [TWI88, Whi81a, Whi81b, HPB84, LCH^{+}88, Mei89b, Moo81, Moo83, Mul83, OO86, WH87, Wil87a]. outputs [Sim88b]. over-relaxation [PJ84]. Overall [PLR85]. Overdetermined [Abd80]. overlap [AM89a]. Overview [Rud83, ABC^{+}88, Cha87a, Tei86]. Oxford [Nad86, Pet87]. oyo [Ton82].

p [Ano82c, BG82, BGCS82, Nad86]. p-system [Ano82c, BG82, BGCS82]. P.D.Q. [Boy85a, Boy85b, Boy89b]. paced [EGP81]. pack [McC81, Chu88]. Package [Bre78b, Bre78a, Bre79, BHY80, BW89, Chu88, DM84, DS84, GGLM88, GL90, GHM^{+}86, GKRY82, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Maa89, Mar82b, Mar84b, PP82a, Rei84b, WLO76, AI88, AEL^{+}86, CR84, Dun85a, Gil86, Gre88, Har86b, Int82f, IA84, JK83, KWM88, Kee88a, KJM89, LOU86, PP82b, PRL^{+}85, PP85, Rei84a, Sul2 , Pie85]. Packages [Arn82, DFK83a, DFK83b, DFK88, EP87a, EP89, Rag86, DFK81, Num88a, ZGK88]. Packard [Pol83]. packing [She89b]. PAGE [AEV89, dEV89]. PAGE-AHEAD [AEV89, dEV89]. pages [Cou85a, Cou85b, Edm86, RB82, Wil87b]. Painters [Kre86a, Kre86b]. Paketvermittlung [Hof84]. Paketzerlegung [Hof84]. PAM [ALPC88]. PAM-CRASH [ALPC88]. paper [Joh84]. Papers [Smi84, Hor83b]. parabolic [Som86]. paraboloid [Hsi83]. Parallel [AK82, AK85, Ana87, Ano88a, ASM89, AE87b, AE87a, Ber88a, Car89a, Cul88, DM85, DS86b, DS87b, ET86, DBFK89, Gre86, GWM88, GL86, GPHL90, HL82c, HO89, Hus84, JGD87, KS81c, Kow85, Mel88, Per87, Pra85, Ric84, RV8 , SAB88, Utt89, Wri89, ??88, Dha88, AK81, AP87, AE88, BKK^{+}89, BH89, BDS89, CC87, DS86a, DH84b, DS87a, FSO89, FJS85, GSZ88, Guz88, GPHL88, HMB^{+}88, Int86a, Jor86, KB88, May89, McD89, PC89, PW84, PRWB89, Pol87, Roy88, RW89, Sne88, Sto85c, TPS^{+}88, WG84, Wee86, YHKM89, HS81]. Parallele [KS81c]. Paralleles [Abs88]. Parallelism [FN85, Kar86, Kar87c, Rei87a, BK89, Ber88b, BCF^{+}88, Jes82, Sto85c]. Parallelization [Har89, AJ88, BDR87, ZBG88]. Parallelizations [TFI86]. parallelized [PJ84]. parallelizing [KK89b, Smi88d]. Parallelstruktur [HS81]. parameter [Don83c, Mai81a, Mai81b, Szy87]. Parameters [Cod88, Moo85a]. parametric [Fra84a, KP86]. ParaScope [BKK^{+}89]. Parent [Jia86]. parentheses [Uni83a]. PARFOR [Abs88, Ber88a]. Paris [Ass86, Rei89c]. parity [SDH84]. parler [Ain89]. Parser [DDH84]. Parsing [HT82]. Part [Bur84c, Chi85a, EA87, IEC88, ISO88, She89b, Lag85, Mul85, PF85, SPS84, Sav87, Goo89]. Partial [Mei89b, She78, Ste79, SS79, Ada89, Bar89, GtTB89, ISJ85, Pet89]. Partial-record [Mei89b]. particle [GH87]. particles [GDK89]. parts [Smi85a]. Pascal [Kur85, Mil87b, Wic89, Fog85, Fog87, Fog88, JBT83, Ker82, Pay84a, Pay84b, PA84, Res86, Rel89c, Smi81, Smi83a, Ber82a, Cas81, Cul88, Don81, Fre81, GMW86, Ler83, Mil87a, PD81, PA83, Sch82a, Sun84, Ter87, WS84]. pass [JT88]. Passing [SP87]. PAT [ASM89, Smi88d]. path [Uni88]. paths [HM82]. Pattern [Ass86, Hsi83]. Patterns [BDS88b, BDS88a, BDS88c, ZGS89, vM84e]. Paul [Wex87]. PC [RMFG85, Ano87a, CW85, CW88, Cla86b, Cla86c, DW85, Fuo86a, HRC89, HK87a, HK87b, Wor84a, LB86, Pee84a, Ser89, Div85, Wor84b, WB89, ZGK88, Cla86a]. PC-BLAS [HK87a, HK87b]. PC-Portable [Cla86b, Cla86c, Cla86a]. PC50 [Num83c, Num84d]. PCFORT [CCN^{+}79]. Pcode [CCN^{+}79]. PCs [Lah88b, Lah88c]. PD77230 [ZGK88]. PDE [Pet89]. PDEs [Lio85, Som86, dZ86]. PDFIND [Cra86a]. PDP [AEL^{+}86, Cla89, Hue83, Mic83]. PDP-11 [AEL^{+}86, Cla89, Hue83, Mic83]. PDQ [Boy89a]. Peak [LS87, Tho84c]. Peephole [Pem83, TvSS82]. percentage [YK88]. Performance [AGS88, Arn82, BD89, CT86b, Cro87a, Cro87b, DM89a, Don83b, Don84a, Don84d, Don84b, Don85a, Don85b, DD86, Don87a, Don87b, Don88b, Don88c, Don89, Gaf84, Gal89, GE85, LS88, MC82, vdV85b, vdV86, vdV85a, Ano89, Bai87, BW87a, Bli89, Bow82, Chi85b, CT88, CT86a, DDDG89, DSCP88, KRW88, Kip82, Lee85, LS87, LR89, MMM85, McM86, Sor84, hTD88, WSL88, Wat87, van86]. Performances [Don84c]. period [Gui88]. Perkins [Mil82b]. permutations [Wil83]. Personal [BW87a, CSD83, CDW83a, CDW83b, HRC87, McG87a, RG85, RB83, Rou83, BNZ87, Gui87, IBM89b, Jus88, Kir85, Lah87b, Lah88b, Lah88c, Lah88d, McG87b, Mul88, Num84d]. perspective [Ott81]. Perspectives [Sch88b]. perturbation [kK89c]. petroleum [Hig86]. PFC [AK81, AK82]. phase [Alb86, BDS84, KWM88, Kee88a, LKM88]. phases [KD84a, KD84b]. Photo [KTW84]. photoelectron [Abe89]. photometry [Moo85a]. photomicrography [PRL^{+}85]. photon [BSdlT87]. physcs [GH87]. Physics [Vor89, BZ85, KM89]. physiological [PM87]. piecewise [Tod85]. piecewise-linear [Tod85]. Pioneer [AW82]. Pipeline [HG82b, CT88, MR86, Owe86]. pipelined [Sor84]. Pisces [Pra85]. PITMAN [Dal88a]. Pivoting [She78]. PL [Ber82a, Bin85, Bro81b, Bro83a, Ell82a, Ell82b, Fog85, Fog87, Fog88, Int82f, Kur85, Res86, Rel89c, Rel89a, Sch82a, Bro84b]. Pl-I [Bro84b]. PL/1 [Ber82a, Fog85, Fog87, Fog88, Res86]. PL/I [Bin85, Bro81b, Bro83a, Ell82a, Ell82b, Int82f, Kur85, Rel89c, Rel89a, Sch82a]. PLA [YM85]. Placement [GM83]. PLALGO [Tod85]. PLAN [CY89]. PLAN-I [CY89]. planar [Bai89, Per81]. planar-structural [Per81]. Plane [Ren84, RO85, RO86]. planes [Sav87]. planewise [HM81, HM84]. Planning [Nor84]. planting [Art81]. Plasma [MHK86]. Plasticating [Rao86a, Rao86b]. PLATO [Joh84, MSM84]. Plattsum [HM81, HM84]. PLEX [SAS86]. Plod [ACG^{+}88]. plot [Col82, Cza83, Joh81, LZ82, O’N81, ZSD82a, ZSD82b]. PLOT79 [BR89a]. Plotting [AL81b, AL81a, Boy84a, Nag81a, Nag85]. PLTSYM [ZSD82b]. PLUS [Ano82a, Kri83, Ehi82, Alp83]. Pocket [Pag84, Rid82b, Rid82c, Dig85d, Pag83, Rid82a, Tan83a, Tan85a]. podstawy [Rzy84]. Poincaré [Ril83]. Point [Cro87a, Fat82a, Fat82b, Kaw84, Sch81a, Ack84, Cro87b, Dav86, PJ84, Tho84a, Wic89, ZGK88]. Pointer [SM87, Mei88]. Points [Ren84, Tho84c, YK88]. Poisson [HK83]. polarization [Ber84a]. polarization-modulation [Ber84a]. Polish [BS81g]. Polycyclic [hTDT88]. Polygon [CY89]. polyhedra [Ril83]. polymer [Chi88, Rao81a, Rao83]. Polynomial [Sau83b, Sau83c, Sau83a]. Polynomials [HMR85a, HMR85b]. poor [Sas81]. populations [Mar81, PS82, PS83]. port [hTDT88]. Portability [Air77, HWS^{+}88, Lar81]. Portable [Amo83a, Amo83b, Bee82, Blu78, DDH84, FW83a, MR83, May89, OO86, Oni85, PP82a, Sch79, WLO76, Ada89, Alc82, BR89a, BH89, Bur86b, DS87a, FW83b, Gui88, Hil82a, Hil82b, Kli89, LB89, Pay84a, Pay84b, PP82b, Smi85c, VP84, WW89, Cla86b, Cla86c, Cla86a]. portables [AFV85]. Portland [Gon89]. positioning [Ack84]. Positive [Cra86a, GL81]. Possibilities [Sch89d]. possible [Tho84a, Tho84c]. post [DJM87]. post-of-duty [DJM87]. Postbuckling [PLR85, RT85]. potential [Dul81, Est82, Gra84b, Sav87]. potential-field [Gra84b]. pour [Ano85b, AFV85, LB89, (??84, Tri84, (??87, Tri89]. Powder [JL81b, JL81a, vM84e, vM84f]. power [Ash81b, Ash81a, Jan84, NM85, Tho84c]. Powerful [CY89, Kli89]. Practical [AHH89, Rul83a, Rul83b, Key81, McD89]. practice [HP88]. practices [Don83c]. Pratique [Lig82b, Lig85b, Lig88b, Ano87d, VPH82]. Praxis [BB86]. pre [BK88, Sus86]. pre-compilation [BK88]. pre-processor [Sus86]. preceding [YHKM89]. PRECI [DNV81]. Precision [Bre78b, Bre78a, Bre79, BHY80, CHH81, CHH83, WLO76, Wic89]. Precompilation [Bro81a]. Precompiler [Ber87, WLO76, GF89]. Preconditioned [vdV86, van86]. predicting [LKM88]. Prediction [Tan81b, BMS84, Gal89, KWWK86, RMS82]. PREFACE [Ber88b]. PREFACE-2 [Ber88b]. prefine [KK89b]. preliminary [Dig83]. premature [Not89]. premixed [Kee88b]. Preprocessor [CBS81, CB82, KK89a, BH85, Ell81b, Gui81, Roy88, Sto85a, Sto85b, Wal81]. Preprocessors [TWI88, Joh87a, LCH^{+}88]. preprozessor [Els82]. Prescribed [CL83]. Presence [HG82b, TFI86]. presented [RH84b]. Press [Nad86, Wil87b]. pressure [Chi88, Cza83]. PRETTY [Bee82]. Prettyprinter [Bee81, Bee82, Bee88]. Prettyprinting [Ash81b, Ash81a]. Preview [Edg89b, Edg89a]. Price [Cou85b, RB82, Wil87b, Cou85a, Edm86]. primary [McK83, Tho84c]. Prime [Haa87, Her81]. primer [Bro82c, Dig84c, Dig85b, HW86, Man82, TMS88a, TMS88b]. primers [Dig85b]. Primitives [Bee85b]. Principle [FDL86, Wat82a, Wat82b]. Principles [AU77, ASU86, Lew81b, Mar84a, POP82, Lew81a]. printing [Jan84]. Prior [Bem84]. Probabilities [Chu88, Sun88a]. probe [Owe87]. Problem [BL83, BS81e, Boi84, Boi87, BM81, Col87a, Dil85, DLS84a, Edg89b, Ett84a, Ett84b, FKSS81, FK81, FK82, Hah87, HB83, HB84, Hon81a, Lib84a, JW86, KF87, KF88, Kre86a, Lew81a, Nan81b, NL83, SFKS81, BS81f, BW84, BW87b, DLS84b, Edg89a, Ett85, HB81, IMS82, Lib84b, Lib87, KJM89, Num84d, Nan81a, PC89, Smi85d, Uni88, YM85, Lew81b, Smi85e]. problem-independent [KJM89]. Problem-Solving [JW86, Nan81b, Lib84a, Lew81a, IMS82, Lib84b, Lib87, Nan81a, Smi85d, Lew81b, Smi85e]. problemas [FK84]. problematyce [Rzy84]. Probleme [MF84]. problémes [GB83, GB89, Tel82]. Problems [Bur81a, Cas89a, Cas89b, DM84, Hig91, Adm85, BZ85, Ber85b, BT88, BD80, Cas89c, CGQS89, Dun85b, GSZ88, MR86, Num88a, Pet83, Sch86, VC89, Zho89]. Problemsolving [Kre86b]. Procedure [Pal86, TFI86, YF85, AM89a, FF84, Riz85, WP84]. Proceedings [ACM89a, ACM89b, Gia89, Gon89, IEE81, IEE88a, IEE88b, KM83, LCMM88, RW86, ACM82, ACM84, ACM87, Sof83b, Usr82, Van84a, Wex81, Wex87, Dav89, Ass86, ML88, Sof84, Smi84, Wri89]. procesamiento [CM89]. process [Cza83, Lag85, Mul85, Not89, PF85, Roc86, ZDS81b]. processed [Dun85b]. Processes [Chu88, GWM88, BZ85, BSdlT87, IEC85, KS81c]. Processing [BB83, Bro81a, Cas81, CM84b, Don81, HG83, LCMM88, LS85, PS81, Sch83b, Sch83a, Tsu85, WN87, Chi88, CM81a, CM81b, CM84a, Int86a, IEC88, ISO88, PRWB89, PF85, Pra89, Rao81a, Rao83, RMFG85, Rob82, RWA84, Rod84, Sch89a, VSH83, VH87, WAD81, Wic89, YHKM89]. Processor [CY89, HK87a, HK87b, HL82c, MR86, Red86, Sus86, Wal81]. Processors [Ano88a, DBFK89, HG82a, Wri89]. Produced [KK89a]. product [Ano82d, Ano87a, CDW82, CDW84, Int81a, Int81b, Int81c, Int82g, IBM88, Int88d]. Production [Coc83, HL82b, Cla89, Hig86]. products [Int82e]. Professional [IR84, Pee85d, Pee85e, Pee86, RG85, Ser85, HA83, LOU86, Pag88]. Professor [Tay86]. PROFGEN [FF84]. profile [Slo88]. profilers [FF84]. profiles [BMS84, RMS82, Web85a, Web85b]. Program [AK82, AL81b, AW82, BS81a, BS81b, BS81c, Bee82, Blu78, Bos88, Boy84b, Chi86a, GMPW79, Gol82b, HL82b, HPB82, JL81b, NL85b, Sau83b, Sau83c, Sau83a, Spa85a, AHU81, Abe89, Ack84, Dha88, AM89a, AI88, AK81, And89, AL81a, Ano82d, Ant81, ADP88, Bai89, BK84, BS84, BMS84, Bod87, BDR87, BT83, Boy84a, BS86c, Bro85, BJ81a, BJ81b, BJ84a, BJ84b, BDS84, CCHT89, CSC^{+}86, CDHP86, vC87, Dal88a, Dal88b, Dal89, DGNP88a, DGNP88b, DFD81, DFD84, DPA87, Dul81, DS82, Dun80, DZ87, Dun87a, Dun87b, DZ88, Dun88a, Dun88b, EH81, For85, FMH85, GRB88, Gla88, Gra86a, Gra84b, Gra86b, Gra81a, Gre88, Gri82, Gru88, HL82a, HM82, HS86, Hel83, Her81, HM81, HM84, Hig86, Hon81b, HK83, Hsi83, Int81a, Int81b, Int81c, Int82e, Int82g, Int83b, Int84b, Int85b]. program [IS84a, IS84b, ISJ85, IA84, Iwa84, JR81, Jac85b, Jac85a, Jam86a, Jam86b, Jan84, JC82, JK82, Joh81, Joh85b, JL81a, Kee88b, Kem85, Kem87, Kie86, LZ82, Lam89, LN89, LKM88, Mai81a, Mai81b, Mai87, Mar81, MW84a, MW84b, Mat83a, Mat83b, McA86, McG87a, McG87b, MSM84, McK83, Rod87, MH82, MHS81, Mil89, Mul88, Mye83a, Mye83b, MT84b, Nai84, Nai86, NL85a, NM85, NJLB81, O’N81, PB84, Per81, PS82, PS83, PS84, RRC89, Rap82a, Rap82b, Rap82c, Ras84, Red82, Ric82b, Rin83, RMS82, RH84b, RS81, RS84, Rub83, Rus87, SPS84, Sav87, Sch89a, SR84a, SDH84, She89b, SJB83a, SJB83b, SD89, Slo88, Smi85b, SP84b, SP84c, SG88, SDC82, Szy87, TPR85, Tat87, Tei86, TBM85, Tho81, Tho82a, Tho82b, Tho84a, Tho84b, Tho84c, Tho84d]. program [Tho84e, Uni86a, Uni84a, Uni84b, Uni81a, Van84c, Van85, Wal85, Web85a, Web85b, WD81a, WD81b, WM85a, Wie86a, Wie86b, WP84, WF85, Wyl86, You82, ZDS81b, Zoh84, dB82b, dB84, vMF81, vM84c, vMF84, vM84d, vMT84, vM84b, vM84e, vM84f]. program-package [AI88, IA84]. program-six [Hon81b]. Programacion [Ber82b, Bor89, Mer86, Mer88a, FK84, Zwa85]. Programando [Zwa85]. Programmation [AFN83, LPJ83, LP87, Ass83b, Ain89, Ber82a, VPH82]. programme [Gra86a, O’R81, VH87, Els82]. Programmed [Mil82b, Baj81, MP81]. Programmen [Wie85, Mey84, NEM84, Wid88, EmR84]. Programmentwicklung [Wös82]. Programmer [Buc84, Dav81a, Del88, Fre83, Lev89, Mai81a, Mai81b, Pre88c, Sch89e, Ano83, Dig84b, Dig84c, Dig85b, Dig85e, Del82, Del83, Del85, DJM87, FE82, Guz87, Hon81b, Hon85, Ket85b, Spe82, Sun85, Sun86a, Sun86b, Sun88b]. Programmering [EE84]. Programmers [Lew81b, PA83, WS84, BS88c, Bin85, Con85d, Cha87b, La 87, Lew81a, Num84d, Pag88, PA84]. Programmes [Don84c, AFV85, Chi85c, TR84]. Programmiersprache [Lam81]. Programmiersprachen [Hah81, Wös82]. Programmierstil [Kur85]. Programmierung [Sch87c, Weh85, KL83a, KL83b, KL85a, KL85b]. Programming [AM81, ADH^{+}89, AK85, Ano81b, AE87b, AE87a, A^{+}81, BS81d, Bla87, Bro84a, Com89, Cal83, Cal89b, Cal89c, Cal89a, CK86b, CP84, Cli84, Col83, Dig84d, DKG89a, DKG89b, Did86, DSCP88, Ell83, Ett84b, EP81a, EP87b, Fis83, For82a, FGGF86, Fuo86b, Fuo86c, Fuo86a, GJ82, GY82, GHM^{+}86, Got84, GL86, Gro83, Gus84, Hil81, HB83, HK84, Hor83a, Hor83b, Isn82, Jap82, Jes82, Kan88, Kar86, Kar87c, KS82a, Kum86, LP85b, Leh86, LS88, MS81, MS84a, ML87, Mas83, Mas87, McK85b, McK85a, Mei84, Nic85b, Per87, Pol81, Pol83, Pra84, Pru87, Rad81, Rad83b, Rao82, Rao87, RW86, RB83, Rou83, SAS86, POP82, Sam81, SM88b, SAN^{+}81, Sin81, Smi81, Smi83a, SAB88, Tea81, WB89, Wol85, Wu83, ZM86, Zwa81, Ame87a, AC87, Ame87c, AE88, BKK^{+}89, Bel84, Bel89]. programming [BS81g, BH89, Con82d, Con82e, CRV^{+}89, CwL83, Cor81, Cor82c, Cul88, Dig84a, Dig86d, Des89, DV81, Ell82e, Ett85, EP81b, FR82, FKSS81, FK81, FK82, Gil86, Gol81, Hea81, HB84, HPR81, Int81a, Int81b, Int81c, Int81d, Int82g, Int82h, Int83d, Int83e, Int83f, Int83g, Int85h, Int86b, Int86g, Int87e, Int87i, Int88h, Int88g, Int88i, Int88j, Int88o, Int89b, Int89c, Int89e, Ion84, IMS82, Lib84a, Lib84b, Ame89b, Joh85a, Key81, Kha81, Kim86, KF87, KF88, KS82b, Lee85, LP85a, Ler83, May89, MSR87, MSM84, MO82, MO84, Mic87d, NSV1 , Nic82, Nic85a, Nic85c, Pol82, Rad83a, Rao81b, Ric82a, Rod86, Ros87, Sel83, Sas83a, Sas83b, Sch86, SM88a, Sch88c, SFKS81, SR87, Tha82, Tur86, Vag89, Wag84, Wee86, Wex81, Wu82a, Wu82b, Ame87e, jYsS89, EA87, Cou85a, RB82]. Programmirovanie [BKL89, SM84]. programmirovnija [HPR81]. Programmpaket [Kna84]. Programmpakets [Fis82]. Programms [Mey84, Sch84f]. Programmsystem [Dah81]. Programmsystems [Ste87]. Programmtransformation [Sto84a, Sto84b]. programmy [BZ85]. programowania [Rzy84]. Programowanie [Rzy84]. Programs [AK87, Ana87, AEV89, ASM89, Aya84, BA86, Boe87, BDS88b, BDS88a, Bro81a, DAG^{+}88, DS86b, DCHH87, DS87b, DCHH88a, Eve85, HO89, How82, Hus84, Kah80, KW89, LH88, Mil88b, Oni85, Rao86a, Rao86b, Sch89b, SB83, Sim86, Spa85b, TFI86, dEV89, AD84, BZ85, Bel84, Bli89, BS86c, BDS88c, BDS89, Bro82a, BD80, CDW82, CDW84, Chi88, CR84, Cla89, Cor82b, Cza83, DS86a, DS87a, Ear85, EIT85, Est82, Eva81, FDL86, FSO89, Gin82, Gro89, GQ88, Gui87, Her81, Int86a, KD84a, KD84b, Kie83, Kip82, Kle89b, Kle89a, Kre88, LB86, Mal85, McC81, McC86, McD89, Mil82a, Moo85a, Osy84, Pol87, Rao81a, Rao83, RV84, RO85, RO86, RW89, Sch82b, Sch83b, Sch89c, SMD84, Sim85, Sim88b, Smi85c, Var85, VSH83, Wal3 , Wan86, Wat82a, Wat82b, Whe84a]. programs [Whe84b, YK88, ZDS81a, VMS81, Bur81a, Har89]. progress [Met89c]. progressif [Ain89]. project [DDDG89, MSM84, Tea81]. project-oriented [Tea81]. Projects [BSP83, Bla87]. PROLOG [Miz83, Fun86]. promptuary [KTW84]. Propagation [CCKT86, LW88b]. properties [BSdlT87, Kee88a, KD84a, KD84b, Kie86, McC81, McC86, VH87, You82]. proportional [AM89a]. proportions [TBM85]. Proposal [BBG^{+}82, BBB^{+}83, DDHH84, Aha85c, DCHH85, Wic89, Bee85b]. proposals [Mei88]. Proposed [Ame87b, Fat82a, Fat82b, Sch89d, Ame87a, Ame87c, Ame87e]. PROPOV [AM89a]. PROPOV-K [AM89a]. protein [LB86]. Proton [GKKY89]. Prototype [Did86, Gre86, Sch82b]. PROTRAN [Ric86]. protsessov. [BZ85]. provision [Dav86]. Provisional [GM83]. Prozesse [KS81c]. prozessors [Fis82]. PRP [AL81a, AL81b]. Pruefung [Wie85]. PS [Ano87a, CW89, Int88c, IBM89b]. PS/ [CW89, IBM89b]. PS/2 [Ano87a, Int88c]. pseudo [Ack84, LN89]. pseudo-range [Ack84]. pseudo-stress [LN89]. Pseudocode [PB86]. Pseudopotentials [PFF83]. pseudorandom [Ras84]. Pseudozufallszahlen [Jan88]. Pseudozufallszahlen-Generatoren [Jan88]. Psi [Amo83b]. PSOD [PS82]. PSR [Gla88]. Psychology [Lew81b, Lew81a]. PTRAN [ABC^{+}88]. Public [Bur84c, Dun87b, Smi88a]. publication [For82b]. Publications [McC84b]. Publishing [Cou85b]. pulsed [vM84d]. Pulsmethode [RS82]. PUMA [Fra84b]. puroguramingu [Har86a, TS88]. puroguramu [SS84]. Purpose [Coc83, KJM89, ZGK88].

QNAP2 [VP84]. QNX [Dir84]. QR [Buc81a, Buc81b, Buc81c, Buc82]. QRUP [Buc81a, Buc81b, Buc81c, Buc82]. Quadratic [GHM^{+}86, PC89]. Quadratures [EK87b, EK87a]. quadrupole [Nai86, vM84e]. quadrupole-distorted [vM84e]. Qualitätsprüfung [Jan88]. quantitative [MT84b, TPR85]. quantities [Rap82a]. quantization [MD88]. Quantum [Ano88b, CV88a, VC89]. quarterly [For82b]. Quasi [MHK86, Gui88, TU81]. Quasi-Analytical [MHK86]. quasi-Newton [TU81]. quasi-random [Gui88]. QUEL [Rel89d]. Quelques [Ano85b]. Queries [KTW84, RL81, WN87]. queueing [VP84]. quick [Con85c, FDL86, IMS87f, IMS87i, IMS87k, IMS89b, IMS89d, IMS89f].

R [Cou85a]. r$-$INF$/ [Tem89b]. Racah [AD84]. radiation [Hsi83, Jam86a, Jam86b, PS82, PS83]. radiometric [LZ82]. radionuclide [PMBK82a, PMBK82b]. RAGBEEF [PMBK82a, PMBK82b]. rainfall [RMS82]. Raman [MW84a, MW84b, Wyl86]. ramification [Gra86a]. Random [Gui89, Haa87, Pie85, Sch79, Sch88a, CB86, Gui88, IS84a, IS84b]. Random-Access [Sch88a]. randomization [Dal88a]. randomly [CCHT89]. range [Ack84, Int84b, McM86, PB84]. ranking [Hel83, Kem85]. Raphson [Rod87]. RASC [Hel83]. RASP [Ott87]. rate [BDS84, PS83]. rates [BMS84, Joh85b]. RATFOR [Eve84, BB83, Col84a, Mar84b, Nor84, Gro82]. ratfor-T [Gro82]. rational [Dun87b, Dun88a, Dun88b, Gro89]. Ratios [DM87b, DM87a]. Ray [JL81b, JL81a, TPR85, vM84f, Abe89, MT84b, Tho84e]. RCJOIN [Dal88b]. RDARL4 [Hue83]. Rdbm [Fis82]. Rdbm-prozessors [Fis82]. reaction [RS81, RS84]. reactions [SMD84, SR84a, SDH84]. reactors [Gla88]. read [AS89a, AS89b, Gre88, ZDS81a]. readability [Dun85b]. Real [Boe87, CL83, Cra86a, Cro85b, FW82, Gla83, Hig88b, Hig88a, Hig89, Kri83, MBP^{+}85a, MBP^{+}85b, Ste76, CHPS85, Cro85a, IEC85, Kne81]. Real-Time [Cro85b, Gla83, MBP^{+}85b, MBP^{+}85a, Cro85a, IEC85, Kne81]. réalisation [LB89]. Realities [PRWB89]. realization [HS83, Sas81]. REAME [Hua82]. REAMES [Hua82]. rearranger [Mye83a, Mye83b]. recalculating [FMH85]. Rechnerarithmetik [Ull84]. Rechnergestuetzte [Jac82]. Rechnern [Kal85a]. Rechnernetzes [Hof84]. Rechnerprogramm [RS82]. Recipes [PFTV86, Pre87a, Pre88a, VTPF87, Pre87b, Pre88b, Pre89, Vet85, VTP89]. Recognition [Ass86, PB86]. recoil [PB84]. Recollections [Hem86]. recombinant [VLV^{+}86]. reconnaissance [ZSD82a, ZSD82b]. Reconstruction [PP82a, PP82b, PRL^{+}85]. Record [POP82, Mei89b]. records [Tat87]. Rectangular [HMR85b, Ric82b]. Recursive [Sas81, WN87, Gra81a]. redefinition [Sul88]. Redesigning [Don83a]. REDUCE [Bat85, Gat85, KK89a, Sha87, Sha89]. reduced [CC87]. reduction [EIT85, Slo88, vMT84]. Reference [Chr84, Dir81, MM81, NL85b, NL85a, Ano83, Ano87a, Apo86, BG82, BGM83, Bur85a, Bur81b, Bur84b, Bur85d, Bur85c, Bur86c, Con81b, Con82b, Con83a, Con83b, Con83d, Con85c, Con85a, Cra83, Cra84, Cra86b, Cra89a, Cra89b, Dig82c, Dig85e, Dig86b, Dig88a, Dat81, Dat85b, DW83a, DW84, ETA88, Hew85, Hew86, HO88, Har81, Har85, Hon85, Int81c, Int81d, Int82c, Int82d, Int82h, Int83b, Int83e, Int83f, Int83g, IBM85, Int85g, Int86e, Int86f, Int87b, Int87c, Int87d, Int87g, Int87h, Int88a, Int88e, Int88f, Int88m, Int88n, Int88p, Int89a, Int89d, IMS87f, IMS87i, IMS87k, IMS89b, IMS89d, IMS89f, Joh83, Kem87, Ket85b, Lah88d, Lew82b, Met89a, Mic83, Mar82a, Mic84d, Mic85e, Mic87d, Mic89d, Moo85b, Rel89b, Rel89d, Spe82, Spe85, Sto85b, Tan85b, Uni81b, Uni83b, Uni84e, Uni88, VL81, Wan84]. References [Ham85, HM90, RH84a]. Refined [KKK89]. reflection [Tho84c]. reflector [Hsi83]. reformatting [Abe89]. refraction [Owe87]. regeneration [Rob83]. region [MS83]. Register [Gol84, BCKT89]. règles [LB89]. regression [ISJ85, Rus87]. regular [kK89c]. Reid [Rei87b, Rei87d, Rei87c, Rei89a, Rei89b, Rei89c]. Reihungen [Kal85a]. related [Dav82, KS81c]. Relational [GE85, IBM86]. relations [CC87]. relative [Bur85b, Bur85d, Bur85c, Bur86c, Tho84c]. relax [BT88]. RELAX3D [HK83]. relaxation [PJ84, vMT84]. Release [AHU81, Int88i, Int88j, Int88k, Int88l, Int88m, Int88n, Int88o, Int88p, IBM89a, Sof83a, Uni84a, Uni84b, Bur85b, Bur85d, Bur85c, Bur86c, Dig82b, Dig84e, Dig85b, Dig85c, Int87g, Int87h, Int87i, Int88b, Int89c, Int89d, Int89e, Pyr84, MAT89a, MAT89b]. Reliability [Moo88b, Ack84, Ber85b]. Remark [AFS94, Bre79, Buc82, DFK88, DG82, Dod83, FW82, Fut78, GL90, Ham85, Hig91, HM90, PCK84]. Remez [Dun87b]. Reno [ACM89b]. renography [Kem87]. Reorganization [HG82b]. repair [Joh85b]. replace [Mor81a]. Reply [Tan83b]. Report [RV8 , Dha88, CGQS89, Fed81, Fri84, MSM84, MSG86, Sof83a, Smi83b]. Reports [Rei89c, Rei87b, Rei87d, Rei87c]. Representation [DK84, JL81b, SW83, EL81, JL81a, Per83b]. Representations [DR86]. representing [And84a]. Republic [RW86]. Repulsion [EBS88]. requirements [Sch82a]. Research [GBJ81, Hue83, KM83, McA86, BR89a, Joh84]. Reservoir [ET86, SP84a, SP85a, SP85b]. Reshenie [Sko88]. resheniya [BZ85]. residual [JR81]. resistance [NM85]. resolucion [FK84]. résolus [GB83, GB89, Lig82b, Lig88c]. resolution [AG87a, AG87b, Min88, AG87c]. resonance [SDH84]. Resource [LZ82]. resources [Boy84a]. response [SD89]. responses [PS82, PS83]. Restructuring [Bro82a, Pol87, LH88]. result [KS88]. Results [Cod86b, Cod86a, Cod89, SH88, CDL88, Car88b, RS87]. Retargeting [Dha88, Hey85]. Retire [KW84, McG84]. retrieval [Jac85a, BJ81b, BJ84b]. retrieve [Jac85a, WD81a, WD81b]. Retrospect [AW82, Noh84]. RETURNs [Wil84]. Revenue [DJM87]. reversed [Law88]. reversion [Law88]. Review [All84, Bis81, Cou85a, Cou85b, Edm86, Eve85, Mil82b, Nad86, RB82, Rit89, Smi87a, Wil87b, All82, BEE^{+}85a, Smi87b, Smi88a, Smi88b, Whi89]. Reviews [Smi84]. Revised [Moo83, Sym88, Ame87c, BNZ87, Wil87b, Ame87e]. Revision [AC87, Com89, AR87, Wag84]. revolution [FS86a]. rhythm [Rub83]. Richtlinien [Sch87c]. Ridge [ZDS81b]. ridges [SPS84]. Right [Vel82, Tha82]. Right-to-Left [Vel82]. rigid [She89b]. ring [And84a]. rise [Dav86]. RM/Fortran [Rya86]. RNFREE [Gra81b]. Robotertechnologie [MF84]. Rock [AL81b, AL81a]. rocks [SPS84, SS87a, Sav87]. Roger [Bis81]. Role [Moo88b, Pet88]. Root [JBJ84, VVV89b, VVV89a, SJB83a, SJB83b]. Rosenbrock [Sha82]. rotatable [The88]. Rotation [HPB82, Bru86]. rotational [Nai84, Nai86, Tho81, Tho82a]. rotationally [CT88]. rotor [HL82a, Red86]. rotors [Red86]. roundoff [Bli89]. Routine [Cra86a, Som86, Col82, GtTB89, Lio85, OO86, dR87, dZ86, vM84b]. Routinen [Wis81]. Routines [Buc81a, Buc81b, Buc81c, Buc82, Dix85, Dur80, GF81, JW86, KW87b, KW87a, Ano82b, Ano85c, DH84b, FCG83, GC84, Hos88, Hou83, Kri83, Nag85, Pat89, Uni84d, Uni84e]. Row [DFK83a, DFK83b, DFK88, DFK81]. rows [Tho81, Tho82a]. rowwise [PJ84]. RPG [WAD81, CM81a, CM81b, CM84a, CM84b, CM89, RWA84, Rod84]. RPG/2E [CM81a]. RSA [Bur84c]. RT [CP84, Cli84, hHtM81, Int88c]. RT-11 [CP84, Cli84, hHtM81]. RTE [Hew85]. RTE-6 [Hew85]. RTE-6/VM [Hew85]. RTE-A [Hew85]. Rule [Bos88, Jia86, Law88]. Rule-Based [Bos88]. rules [CSD82]. run [Col87b, Dig85e, VH87]. run-time [Dig85e]. running [vC87, Eva81]. runtime [Pay84a, Pay84b]. Russian [BZ85]. RVT [SPS84].

s [Osi82a, Osi82b, BGG86, DH82, DH83, DH84a, EML88, Mer85]. S-820 [EML88]. S-820/80 [EML88]. S8 [Com89, AC87]. Saarbrucken [RW86]. SAKI [Web85a, Web85b]. Salt [Sof84]. Sample [KW89, Sim88b, ZSD82a]. samples [Dav82, IS84a, IS84b]. sampling [MT84a]. Sans [SM87]. Santa [Noh84]. SAS [Kar87a, Kar87b]. Savage [Kaw84]. Saves [Sch89e]. Savez [Ain89]. Savez-vous [Ain89]. saving [Zak84]. Scalar [ACG^{+}86, vM84c]. Scale [HWS^{+}88, LN89, O’N81, vM84a]. scaling [Hel83, PP85]. scattering [Tho84a, Tho84e]. scene [Par84]. scenes [Wit81]. SCHEDULE [DS86a, DS86b, HO89, DS87b]. schedules [TMjC81, Tew81]. Scheduling [LO85a, LO85b, hTDT88]. Schematic [PB86]. Scheme [BP81a, Har89, BP81b, YHKM89, Shi88]. Schneiderman [Sis85]. Schur [KW87b, KW87a]. Science [Ano88a, BM81, Cou85a, IEE88a, Lei87, Sch85b, Baj81, JS88, LD87, MSR87, Ott81, Sch81b, SB82, SB86, SR87, CSD82, MF84]. Sciences [Leh86]. Scientific [Ano87a, AE87b, AE87a, BRK^{+}87a, BRK^{+}87b, BRK^{+}88, BBB^{+}83, DR86, EA87, How82, JRS88a, JRS88b, KW89, KS88, Lin83, MS88a, Moo88b, Num88d, Pee84a, Pra85, PFTV86, Ser85, Ser89, Tur86, VV86, Wic89, Adm85, AE88, BR89a, Kul83, KM83, Pee84b, Pee84c, Pee84d, Pee85a, Pee85b, Pee85c, Pee85d, Pee85e, Pee86, Pee89, Pre88b, SL82, Tou84]. Scientific-Computation [BBB^{+}83]. Scientists [BS88a, HH85, McC84a, MS88c, Mil87a, Mil88b, NL88, Wei89a, Wor88, Adm84, BS88b, Cor81, Ett83a, Ett83b, Ett87, McC84c, MS88b, Mil82a, Mil87b, NL85c, NL85d, RZ89a, RZ89b, Wor89b, Cou85b]. Scoped [BGS82]. Screen [Dix85, Kie83, Tha89c, Dix85]. Screws [Rao86a, Rao86b]. SCS [WSL88]. SCS-40 [WSL88]. SDD [AGS88]. Second [Bee82, Car87, GPKK82, GPKK84, Pre88c, AI88, EML88, Fra84a, HPR81, WAD81]. section [SR84a, SDC82]. Sections [PP82a, PB84, PP82b, PRL^{+}85, RS81, RS84, SMD84]. secure [Sch82b]. Sed [KODG^{+}87]. sediment [ZSD82a, ZSD82b]. sedimentary [SPS84, SS87a, Sav87]. Seismic [HWS^{+}88]. Selby [WF85]. Selby-Olson [WF85]. SELECT [Kle89a]. Selected [Sim85, Sim86, Sim88b, McC81, WD81a, WD81b]. Selecting [Tha82, Fog85, Fog87, Fog88, Res86]. selection [Num88a]. Self [LO85a, LO85b, MHK86, AI88, Der82, EGP81, IA84, Spe83, YS84d]. Self-Confined [MHK86]. self-contained [AI88, IA84]. self-paced [EGP81]. Self-Scheduling [LO85a, LO85b]. self-study [Spe83]. self-teach [YS84d, Mic84f]. Semantic [Sch89d]. semantics [BG84]. semantikerhaltende [Sto84a, Sto84b]. semester [MSM84]. semi [She89b, ZBG88]. semi-automatic [ZBG88]. semi-rigid [She89b]. semiautomated [PRL^{+}85]. semiclassical [MD88]. SENKIN [LKM88]. sensitive [Dig85d]. Sensitivity [LK88, LKM88]. Sep [Rei89c]. Separable [Ste79, SS79]. September [LCMM88]. Sequence [EBS88, AEL^{+}86]. sequences [Gen82, Lec89]. Sequent [Cod86b]. sequential [KK89b, WM85a]. Serial [PP82a, PP82b, PRL^{+}85]. Series [CC82, Cou85a, Kir89, KC84b, Law88, Bar89, Bur85b, Bur85d, Bur85c, Bur86c, Gra84b, Gro89, Mil89, She89b, TFH86]. Service [And84b, Con82a, DJM87]. services [Int83g]. Set [Bee85b, Buc81a, Buc81b, Buc81c, Buc82, DDHH84, DCHH87, DCHH88a, DCHH88b, DCHH88c, DR82, CC87, Col87b, DCHH85, Sun87, Wie82]. Sets [DR82, Hop81, Rei84b, Per83b, Rei84a]. seven [HA83, Gra84a]. seventy [HA83, Gra84a]. Seventy-Seven [Gra84a]. Several [CV88a, Bro85, Dav82, MSG86]. SFTRAN3 [Bee88, Bee81]. SFUN [IMS87a, IMS87c, IMS87j, IMS87h, IMS87i, IMS89d, IMS89e, Lib89b, IMS89l, IMS89a]. SFUN/ [Lib89b]. SFUN/LIBRARY [IMS87a, IMS87c, IMS89d, IMS89e, IMS89l, IMS87j, IMS87h, IMS87i, IMS89a]. SHADOW [She89b]. shapes [Iwa84]. SHARE [Noh84]. shared [AP87, Jor86]. sharing [Fon85]. shell [Jam86a, Jam86b]. Shelley [Edm86]. Shih [RkC84, Cha83, mCaLjH84, CwL83, hHtM81, sKcH81, mM84, fTBcL7 , mT82b, fY84]. ship [NM85]. shock [Tho81, Tho82a]. Short [Web88, Jan84]. shortest [Uni88]. show [Hig86]. SICEDR [Don82a, Don82b]. side [CK88]. side-effect [CK88]. sided [Dun88b]. Siemens [Ano82a, Ehi82]. Signal [LCMM88, RMFG85]. significance [Hym82]. Significant [Ber85b, JC82]. SIGPLAN [ACM82, ACM84, ACM87, Van84a, Wex81, Wex87, For82b]. SIGSAM [ACM89a, Gon89]. silica [KD84a, KD84b]. silicates [Bod87]. Silicon [MF84]. SIMD [Sol89, ZBG88]. similarity [BS84]. Simple [Ess88, Lan88, And84a, CV88a, Nag81a, Sch89a, vM84f, EA87]. Simplex [Hel85b, GG88, Hel85a]. Simplified [CSD83, Hua82, Min88]. Simplifying [Dix85]. SIMULA [Bro81b]. simulate [PS82, PS83, Wyl86, vM84e]. Simulating [sT85, ZGS89]. Simulation [BCM87, Dah81, Fis82, Gab89, Hel85b, KBRM^{+}86, LW88b, Mar84b, OM82, Rey80, Sch82d, AHU81, Ant81, BK84, Ber84a, CT88, Dav86, Gri82, Hur82, JK82, Lee84a, LOU86, Sch88b, Uni84a, Uni84b, VMS81, Hof84, Sch82c]. simulations [GDK89, NE81].

Simulationslaufes [Lep86]. Simulationsprogramm [Hei83]. Simulationsrechner [Hel85a, Hel85b]. Simulationstechnik [Gol82a]. Simulator [Ber84b, Dun86, HS81, MS84b, Sch84b, Roc86, Sch87b]. Simulators [HS81, Lag85, Mul85, PF85]. Simultaneous [HL86a, LK88, HL86b, Hof87, HP89, IZP81, dR87]. SIN [Nor83]. single [DGNP88a, DGNP88b, SDH84]. single-program-multiple-data [DGNP88a, DGNP88b]. Sintez [MK86]. SIR/DBMS [SIR82a, SIR82b]. SIR/HOST [SIR82a, SIR82b]. Sistem [MK86]. sitisiti [SS84]. situ [SPS84, SS87a, Sav87]. situations [Tho84a]. six [Hon81b, Lam89, You82]. six-component [Lam89]. SLAC [MH82]. SLACINPT [MH82]. SLAM [Lee84a]. SLATEC [JK83]. SLECT [Kle89b]. sliding [Sav87]. slip [HS86, Sav87]. slopes [Hua82]. slow [KS81b, KS81a]. slowly [Bar89]. small [Art81, Col82]. SMOOFF [Gru88]. Smoothing [Dur80, Ano82b, Ano85c, Gru88]. Social [Leh86, Cor81, Cor82c]. sock [DS82]. sock-free [DS82]. softback [RB82]. Software [Air77, Ano85a, Ano88c, BS83, Sof87, Cow84, Don84a, Don84d, Don85a, Don87b, Don88b, Don89, Fen87a, GGLM88, GL90, Ger83, Gla83, Gro82, HG82a, Hey85, JRS88a, JRS88b, JW86, Jur86, KMN89, KODG^{+}87, Law83, Mar84a, MGH81b, MGH81a, Mos88, Ott81, PP82a, Rey80, Sof83b, Sof84, SAN^{+}81, Smi85a, Smi88c, Usr82, Adv86, Ada89, ACG^{+}88, Boo81, CKT85, CM81c, Don83b, Don84b, Don85b, Don87a, Don88c, Fen87b, Gre88, Gul86, IMS82, Lib84a, Lib84b, Lib87, Jus88, KRW88, MMS88, Not89, Obl85, Par84, Pet88, PP82b, PRL^{+}85, PP85, RS89, Rob83, Sel83, Smi83c, Tho86a, Voe89, Zho89, ZGK88, CSD82, Ger83]. Soil [SG88]. SOILMOP [RMS82]. SOILWAT [BMS84]. Solid [Ano88b, Vor89, KD84a, KD84b]. Solid-State [Ano88b]. solubilities [KD84a, KD84b]. Solution [Abd80, Bur81a, Cal86, Gaf84, Hop81, MC80, Ste79, SS79, Ada89, BD80, GL81, Mel88, Pet89, RAKK88, SPS84]. Solutions [JSW85b, Mil87b, NL85d, Rei84b, Tel82, Adm85, RO85, RO86]. Solve [Sch88a, HS86, HK83]. Solver [LK88, Sha87, Sha89]. Solvers [ET86, RRS88]. Solving [BS81a, BS81b, BS81c, BL83, BS81e, Boi84, Boi87, BA85b, BA85c, BA85a, Cas89a, Cas89b, Col87a, CC82, CMM^{+}88, DFK83a, DFK83b, DFK88, Dil85, DM84, DS84, DR82, DLS84a, Edg89b, Ett84a, Ett84b, Gre86, GKRY82, Hah87, Hig91, HB83, Hon81a, JW86, KW87b, KW87a, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Kre86a, MHK86, Nan81b, NL83, Rei84b, BZ85, BS81f, BW84, BW87b, DFK81, DLS84b, Edg89a, Ett85, FKSS81, FK81, FK82, HB84, HB81, IMS82, Lib84a, Lib84b, Lib87, IZP81, KF87, KF88, Lew81a, Num84d, Nan81a, Rei84a, Sch86, Smi85d, SFKS81, Tod85, VC89, Zho89, Lew81b, Smi85e]. Some [Gaf84, GQ88, JW86, LR89, MR86, Mul83, Pal86, Bar89, Car88b, Chi85b, Rei89a, Rei89b, RS87, TMS88a, TMS88b, VC89]. sort [ZDS81a]. sorting [Car88b, Hig86, Hou83, RS87]. sound [Jam86a, Jam86b]. Source [Bod87, KK89a, FE82, Kir89, Mye83a, Mye83b, Sis85, Wal81, ZGS89]. Sources [Cow84, Dal88b]. sous [LB89, TR84]. sous-ensemble [LB89]. sous-programmes [TR84]. SP [DW83b, Int88b]. Space [Pal86, BHK^{+}85, Zak84]. Space-filling [Pal86]. Spaces [SSS84c]. spall [RS84, RS81]. spallation [RS81, RS84]. Sparse [CGM84b, CGM85a, DR82, Gre86, GKRY82, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, She78, BHK^{+}85, CGM84a, CGM85b, GL81, Mel88, Num88a, Par84]. Spath [Eve85]. Spearman [Kem85]. special [IMS87a, Lib87, IMS87c, IMS87j, IMS87h, IMS87i, IMS89d, IMS89e, Lib89b, IMS89l, Kem87]. species [Mal85]. Specifically [BB82, DB82a]. Specification [BBG^{+}82, BBB^{+}83, BBG^{+}84, Gab89, SAS86]. Specifications [Dix85, RW89]. Specified [PCK84, Van82, BT83, Tho84a]. spectra [CSC^{+}86, Kie86]. spectral [Her88, MD88]. Spectre [Oli81]. spectroscopic [Wyl86]. spectroscopy [Abe89]. Spectrum [PCK84, Van82, vM84a]. Speed [Hus84]. Sperry [Uni84d, Uni84e]. Spezialprozessors [Wis81]. Spezifischen [RS82]. Spezifischer [San82]. spherical [Jam86a, Jam86b, MW84b, Rap82a, dB82b, dB84]. spin [SDH84, vMF81, vMF84, vM84d]. spin-echo [vM84d]. SPINC2 [Uni88]. SPINC2/SPINS4 [Uni88]. SPINS4 [Uni88]. spiral [KS82b]. Spline [Ano82b, Ano85c, Dur80]. splines [Ano87a, PS84]. SPOC [Lag85, Mul85, PF85]. spring [Art81, Iwa84]. spring-mass [Iwa84]. SQL [Rel88, Rat89, Rel89a, Rel89b, Rel89d]. SQSIMUL [BSdlT87]. Square [BBF^{+}82, VVV89b, VVV89a]. Square-Well [BBF^{+}82]. squared [Wat82a, Wat82b]. squares [GHM^{+}86, TU81]. Squeezing [DE84, BSdlT87]. sravnenie [Osi82a, Osi82b]. St. [Wex87]. Stability [PFF83, EH81, Hua82, Red86, Thu86]. Standard [Bee85b, Com89, DH82, DH83, Don84a, Don84d, Don85a, Don87b, Don88b, Don89, Fat82a, Fat82b, For82a, FR82, Sne88, Ass82, Ass83a, AR87, BP81b, CR84, Don83b, Don84b, Don85b, Don87a, Don88c, Kne81, Lah87a, Met87a, RB82, Sal84, Wie86a, Wie86b, Ame85c, Ame85a, Ame87a, AC87, Ame87c, Ame87d, BS81g, BP81a, DH84a, EmR84, Ame85b, Ame89b, Ame87e]. Standard-Fortran-programmen [EmR84]. Standardization [Gre84, Int87a, Met89c]. standardized [ISJ85]. standards [Uni83b]. Stanford [Wri89]. stark [Nai84]. start [Rin83, ASM89]. START/PAT [ASM89]. Started [Dav81b]. Stat [IMS87m, Lib89c, IMS87a, IMS87d, IMS87l, IMS87k, IMS89f, IMS89g, IMS89m, IMS89n, IMS89a]. STAT/ [IMS87l, IMS89m]. Stat/library [IMS87m, Lib89c, IMS87a, IMS87d, IMS87k, IMS89f, IMS89g, IMS89n, IMS89a]. State [Ano88b, Vor89, Wyl86]. statement [Bur86a]. statements [Dun85b, Sal84, Wil87a]. states [Sav87]. static [McD89, VH87]. Statischer [Jac82]. Statistical [OM82, SB83, Stu81a, IMS82, IMS87a, IMS87d, IMS87l, IMS87m, IMS87k, IMS89a, IMS89f, IMS89g, IMS89m, IMS89n, IA89, Ras84]. statistics [Ano84, BSdlT87, Chi85c, IMS84, IMS87a, IMS89a, LOU86, Mar81]. Status [Smi83b, Wag84, BS86b, BT83]. Staveren [Pem83]. steady [Dul81, Kee88b, Wyl86]. steam [KD84a, KD84b]. Stellar [Moo85a]. step [MS83]. Stevenson [Pem83]. Stiff [Gaf84]. stirred [Gla88]. stochastic [CGQS89, Fra84a]. Stock [Sim86, Sim85, Sim88b]. Stokes [Gro87]. Storage [Jai84, Sch88a, Con81b, Con83a, Con85a, GS81, BJ81a, BJ84a]. store [Jac85b]. Stored [AW82]. Stored-Program [AW82]. strain [Lam89]. strange [Jan84]. Strategies [CT88]. stream [ZSD82a, ZSD82b]. stream-sediment [ZSD82b]. Stress [Fen87a, Fen87b, LN89, Sav87]. stresses [SPS84]. strike [HS86]. string [GS81]. string-matching [GS81]. striped [Mel88]. Strömgren [Moo85a]. structural [Bod87, Per81, Sas83a]. Structure [GMPW79, HS81, Jai84, Sch89a, SP82, And84a, FS86b, Joh81, Nai84, Nai86, Tel82].

Structured [AM81, BS81e, Ber88a, Boi84, Boi87, Con82d, Con82e, Col83, Col87a, CM83a, CM87, DH88b, DLS84a, Ell83, Ett83a, Ett83b, Ett84a, Ett84b, Ett87, Fis83, Gol82b, Gri85, Hil81, HB83, HB81, Hon81a, Law83, LHP87, LH87, Mas83, Mas87, MSM84, McK85b, McK85a, Mei84, MM81, Moo85b, Nic85b, Pad85, Pol81, Pol82, Pol83, Sas81, TWI88, AM84, BS81f, BW84, BW87b, CwL83, Cou85a, CM83b, DH82, DH83, DKG89a, DKG89b, DLS84b, Ear85, Ell81a, Ell82e, Ett85, FR82, FKSS81, FK81, FK82, Gol81, Gui81, HB84, Kha81, KF87, KF88, LH81, MO82, MO84, Mer81, Nic85a, Nic85c, Sas83b, SFKS81, Wie83, Zwa81, Cou85b]. structurée [Ain89]. Structures [DR86, Pou87, HS83, Mat83a, Mat83b, Wat86]. Structuring [Jor86, See81a, See81b]. Strukturanalyse [Mey84]. Strukturierte [Els82, Weh85]. strumming [Iwa84]. Student [CM84b, RWA84, Rod84, WAD81, Con82e, Cal85, Lee85]. Students [Mil82b, Baj81, MP81, Bis81]. Studies [DM87c, Rod87, PP85]. Study [BRK^{+}87a, BRK^{+}87b, BRK^{+}88, NE89, RV8 , ZM86, CDHP86, Der82, GGJ^{+}89, LCH^{+}88, Red86, Spe83, Wie82, Wyl86]. Style [BGG85a, BGG86, DH88b, FGGF86, Fuo86b, Fuo86c, BGG85b, DH82, DH83, FR82, Mer81, Mer85, Ros87]. suan [mCaLjH84, hH82, mM84]. subcritical [kK89c]. subject [Der82, SD89]. submitted [Ame87d]. subprogram [Bru86]. Subprograms [MAT89a, MAT89b, Ste79, SS79, Boo82, DCHH85, LN87, DG82, DDHH84, DCHH87, DCHH88a, DCHH88b, DCHH88c, HK87a, HK87b, LHKK79a, LHKK79b, LN88b, LN88a]. Subroutine [Abd80, Amo83a, Amo83b, BA85b, BA85c, BA85a, Cas89a, Cas89b, CL83, Chu88, Cod88, Don82a, Don82b, Gaf83b, Gaf83c, Gaf83a, Hig91, MS88a, MP86a, MP86c, MP86b, MC80, Pee84a, Ser85, Ser89, She78, SSS84a, SSS84b, SSS84c, BS86b, Bow82, CF85, CHPS85, Col84b, FE82, Int82f, IZP81, IA89, KP86, Pee84b, Pee84c, Pee84d, Pee85d, Pee85e, Pee86, Pee89, Sch86, Sun88a, TU81, Vu 89]. Subroutines [AC86, CGM84b, CGM84a, CGM85a, DM87b, DM87a, DR82, EK87b, EK87a, FW82, Gaf84, GN89, Hop81, IEC81, MGH81b, MGH81a, Mor81b, Mor90, Mor93, PCK84, SR86, Ste76, Van82, VVV89b, VVV89a, Alb86, Ano84, CGM85b, Dun85a, IMS84, IMS87a, Lib87, IMS87c, IMS87b, IMS87d, IMS87g, IMS87e, IMS87f, IMS87j, IMS87h, IMS87i, IMS87l, IMS87m, IMS87k, IMS87n, IMS89a, IMS89b, IMS89c, Lib89a, IMS89d, IMS89e, IMS89f, IMS89g, Lib89c, Lib89b, IMS89h, IMS89m, IMS89i, IMS89j, IMS89k, IMS89l, IMS89n, Joh87b, Joh87c, Num88d, Pee85a, Pee85b, Pee85c, RV89, SC83, TSU88, Uni82b, Uni85a, Uni84d, Uni84e, Uni88, Dod83]. Subset [Dur83b, Par86a, Par86b, Rom81]. subsonic [Wal85]. Subspaces [PCK84, Van82]. successive [PJ84]. Successors [Ros84]. sugar [PM87]. suggestions [Dun85b]. suite [CDL88, Kip82, Nag85]. sum [Bar89]. Sumador [RL81]. SUMLIST [Kle89b, Kle89a]. summarizing [BK89]. Summary [AW89, Mei89c, Dig82d, Dig84g, Dun85b, Fed81, Gri82, Int82d, IBM85, Int88p, Kle89b, Kle89a, Sof83a]. summation [HM81, HM84]. Summer [Sof83b, Sof84]. sums [HM81, HM84]. Sun [Sun88b, GMW86]. SunINGRES [Sun87]. super [Hey85]. super-mini [Hey85]. SUPERB [ZBG88]. Supercomputer [Kow85, PZA86, Wat87]. Supercomputers [Hwa84, LW88a, LW89, PW86, ZM86, RS89, hTDT88]. Supercomputing [ACM89b, IEE88b, ML88, Car89a]. Supermap [Gue86]. Supervector [MMM85]. supplement [HF81, Int83g, Num81, Num85a, Nag81a, Nag85, Tan85b, Tri89]. Support [MBP^{+}85b, Lee85, MBP^{+}85a]. supporting [Ber88b]. suppression [Col82]. Suprenum [Sol89]. Suprenum-Fortran [Sol89]. surface [Ano87a, BDJ^{+}89, Gra84b, Gre88, NM85]. surface-level [Gra84b]. surfaces [FS86a, Wal85]. survey [EIT85]. Surveys [MF84]. survival [CCHT89, Mul88]. Survive [Ros84]. susceptibility [vM84b]. SVDTJP [dR87]. SX [LtW88, Wat87]. SX-2 [LtW88]. Sylvester [KW87b, KW87a]. symbol [SL82, ZSD82b]. Symbolic [ACM89a, AB89, CV88a, DR86, Gia89, Gon89, Her88, Red86, RT85, Vor89, Wan85, AM89b, DFD81, DFD84, Fra84a, Hil82a, Hil82b, Pet88, VC89, Wan86]. Symbolics [Sym88]. symbols [BT83]. Symmetric [CL83, Cra86a, DR82, HPB82, CHPS85, MT84a, MW84a, SPS84]. Symposium [ACM89a, Ano88b, Gia89, Gon89, IEE81, KM83, RW86, POP82, ACM82, ACM84, ACM87, Wex87, Van84a]. Synchronization [FJS85]. Synchrotron [GKKY89]. Syntax [BS86a, Bro82a, Can81, Ell81b, Tay84]. synthesis [Gin82, She89b, Van84c, Van85]. Synthetic [Tha89a, Tha89b]. System [Ame85d, AKLS88, Bur84c, DNV81, DM85, Fon85, GBJ81, HS81, Int83b, Int86b, IBM89b, IEC88, ISO88, KBRM^{+}86, MBP^{+}85b, Miz83, Rud83, Ame85a, Ack84, ABC^{+}88, Ano82c, Ant81, BG82, BGCS82, BR89a, BJ81a, BJ81b, BJ84a, BJ84b, Bur85b, Con83a, Con85a, CGQS89, CDW82, CDW84, Cre89, Dob85, DJM87, Dun85b, Fed82a, Gic88, Gui81, Gul86, Hew85, Hig86, Hue83, Int83g, Int86c, Int88c, IBM88, Int88d, IMS82, Lib84a, Lib84b, Int88q, Ame85b, Jac85b, Jac85a, JC82, Joh85b, KS81b, KS81a, KRW88, Lah88b, Lah88c, Le83, Lee85, MBP^{+}85a, Mic84c, Mic85c, Mic85b, Mic87c, MSG86, Nag81a, Obl85, OO86, Ols83, Owe87, Pay84a, Pay84b, Pet89, Rom81, Sch82b, Tan81a, Tan82, Int88r, VLV^{+}86, Wan86, Wat87, YHKM89, ZSD82a, ZSD82b, Ame85c, ACG^{+}86, HWS^{+}88, Int83c, SAS86, WD81a, WD81b]. system-Harray [YHKM89]. System/2 [IBM89b]. System/34 [Int83b, Int86b]. System/36 [Int86b]. System/360 [Int83c]. System/370 [ACG^{+}86, Int83c]. Systemanalyse [Sch84c, Sch82e, Sch85a]. Systematic [JRS88a, JRS88b, LG86]. systématique [Str82, Str85]. Systeme [MF84, Jac82]. Systemen [Hel85b]. Systems [Ame85d, Abd80, Com89, Cas89a, Cas89b, Dav84b, DFK83a, DFK83b, DFK88, DK84, DS84, Gre86, GKRY82, Hig91, Ame89b, KGRY81, KRYG82b, KRYG82c, KRYG82a, KRYG82d, Mar84b, MSA86, Sch82d, Sch83b, Sch83a, Sch88a, Ame85c, Ame85a, AC87, AM89b, Bur81b, Bur84b, Bur84a, Col82, CMM^{+}88, Cor81, Cor82c, CBS81, CB82, Cra83, Cre89, Dig86b, Dig86c, Dig86a, DFK81, Fra84a, GGJ^{+}89, Gal89, GL81, Int81d, Int83c, IEC88, Lib87, ISO88, IZP81, Ame85b, Int85a, Mel88, RRS88, Tod85, VP84]. sytle [Dun85b].

T [Cou85a, Gro82]. T. [BD80]. Table [Car87, Car88a, Car89b, Kah80]. Tables [DDH84, MP86a, MP86c, MP86b, Dal88b, WF85]. TADS [Bur85b]. Talmi [Zoh84]. Talwani [For85]. Talwani-Ewing [For85]. Tandem [Sha87, Sha89]. Tanenbaum [Pem83, Tan83b]. tao [LcY83]. tape [BJ81a, BJ84a]. tapes [Gri85, HM82]. task [BS84]. taut [Iwa84]. taxes [Tew81, TMjC81]. TAYLOR [Gra84b, CC82]. Teach [Rad81, Rad83b, Le83, YS84d, Mic84f]. Teaching [Rag86, Bel84, Ric82a, Wie83]. Technical [Aha85d, For82b, Pem83, Sam81, Tan83b, Adv86, Int85i]. Technik [McC85a, McC85b]. technique [BK89, MRS84, PM87]. Techniques [AB83, ASU86, Con82a, Rey80, ACM87, SP84a, Wan85, Wex87, SP85a, SP85b]. Technological [All84, All82]. Technologists [Val85]. Technology [Voe89, Baj81]. tectonic [SPS84]. Tekmar [Adv86]. telescope [La 87]. temperature [Cza83, Rin83]. Temperaturen [RS82]. tension [Ano87a]. TEP [Joh81]. TEP-II [Joh81]. Teresa [Noh84]. terrain [Gra86b, SP84b, SP84c]. terrain-correction [Gra86b]. Tessellations [Bow81]. Test [Cod86b, Cod86a, Cod89, DCHH87, DCHH88a, MP86a, MP86c, MP86b, Sch81a, Sch84f, Wis81, Bur86a, Bur85b, CDL88, Dav82, JC82, Mat83a, Mat83b, McM86, Tho84b, Tho84c, WP84]. Testing [Sof87, vC87, DAG^{+}88, MGH81b, MGH81a, JK82, Tan81a, Tan82, TBM85]. Tests [JW86, vdV85b, vdV85a, CM81a, Dal88a]. Testschnittstelle [Kna84]. Texas [Mor82]. Text [MM81, Moo85b, OO86]. textual [BS81g]. their [BH89, DM87b, DM87a, GQ88, RW85]. them [Ber85b]. theorem [Ril83]. theoretic [GQ88]. Theorie [VPH82, LPJ83, LP87]. Theories [LCMM88]. Theory [Ano88b, AB89, Ban88, EW87, Eve85, IEE88a, Spa85b, Spa85a, Kip82, SH88]. there [AS89a, AS89b]. thermal [Joh81]. thermal-ellipsoid [Joh81]. thermodynamic [Kie86, McC81]. thermophysical [McC86, You82]. thesis [Chi85a]. thin [SDC82]. Third [Car88a]. Thomas [KM83, Rit89]. those [Bro81b]. Threads [Doe88]. Three [Arn82, Bur87, PP82a, PP82b, PRL^{+}85, CM86, DM89b, EH81, LW88b, Tho81, Tho82a, Tho84a, Tho84c, Tho84e, ZGS89, dB82b, dB84]. three-beam [Tho84a, Tho84c, Tho84e]. Three-Dimensional [PP82a, PP82b, PRL^{+}85, CM86, DM89b, EH81, Tho81, Tho82a, ZGS89]. ti [cT81]. TIDY [Ell81b, Wal81]. Tiefen [RS82]. Tien [mCaLjH84, iL82, fS82, Cha83, hK85]. Time [Cro85b, Fon85, Gla83, HL82c, Hol87a, Hol87b, HG83, MBP^{+}85b, CK88, Cro85a, Dig85e, Dun85a, GS81, IEC85, Joh85b, Kne81, Kri83, Mai87, MBP^{+}85a, Min88, MSG86, Ols83, PMBK82a, PMBK82b]. time-dependent [Joh85b, PMBK82a, PMBK82b]. time-invariant [MSG86]. Time-sharing [Fon85]. Timing [Car88b, RS87]. Titan [Cod89]. Toeplitz [Hop81]. toil [MMM85]. too [Wag85]. Tool [BA86, DS87b, DBFK89, JRS88a, JRS88b, Sym85, Sym86, Sym88, BKK^{+}89, BDS89, BCF^{+}88, HM82, KK89b, Sch89f, Smi88d, ZBG88]. Toolbox [Cha86b]. Toolkit [ASM89, Cul88, Lah88d]. TOOLPACK [Kim86]. Tools [ASU86, Ano85a, BB83, BL83, BDS88b, BDS88a, BDS88c, DS86b, Gro82, Gus84, JC89, PBB^{+}88, Usr82, BK88, DS86a, JC88]. top [Leh83b]. top-down [Leh83b]. topics [Con85d, Mic89b]. TOPPER [PP85]. tops [MW84a, MW84b, Dig83]. TOPS-10 [Dig83]. TOPS-10/ [Dig83]. Toronto [Sof83b]. total [SMD84, SR84a]. Tour [Hor83b]. TQ [hH82]. TQ-16 [hH82]. TR [Ano82a, Ehi82]. traces [CDW82, CDW84]. TRACK [Dal89]. tracking [Dal89]. traditional [MB81]. traffic [NJLB81, NJLB81]. Traitement [Ass83b]. Trajectories [GKKY89]. transcendental [JBJ84]. transcribing [Rub83]. transfer [PB84]. transferability [Can81]. transferring [Hue83]. Transform [AC86, GGLM88, GL90, Bai87, Joh86]. Transformation [Boy84b, NEM84, Gro89, RW89]. Transformations [Bos88, BK89]. Transforming [CT86a, CT86b]. transforms [Bun86, GQ88, Pra89]. transition [Bro89a, Bro89b, Bro89c, Bro89d, Nai84, Nai86]. Translation [AK87, Ano85a, Fut78, HK87a, HK87b, LN87, LN88b, LN88a, MT82a, Mon89a, PBB^{+}88, Pra89, Sch89e, Sim76, BNZ87, Sul88, Wil87a]. translations [SMD84]. Translator [CCN^{+}79, DO86, Fre81, Goo89, Bai86, Bro82a, Chi85a]. transmission [Wyl86]. transonic [Dul81]. transparency [Nic85c].

Transport

[And84b, Kee88a, McC81, Tat87]. transportable [KJM89]. Transputer [GH87]. Tree [SW83, Uni88, Win85]. trees [Zak84]. TREESOLVE [Rei84a, Rei84b]. trend [TBM85]. trials [WM85a]. Triangular [Gre86]. Triangulation [CY89, Ren84]. triaxial [Sav87]. Triebwerken [RAKK88]. Triodic [Gol82b]. TRS [Bel84, McN83]. TRS-80 [Bel84, McN83]. Truncated [DM89a]. trust [MS83]. ts’ao [hHtM81]. tso [hHtM81, Int82b, Int82c, Int82d]. tsung [hC83]. Tulsa [BBB^{+}83]. t’ung [hHtM81]. Tuning [HMB^{+}88, CF85]. turbomachinery [HL82a]. Turbulent [MHK86, HL82a]. Tutor [Pet89]. Tutorial [BDR87, Dir81, Hwa84, LO85a, LO85b, Pri89, Wor84b, DW83a, DW84, Doe88, Kir85]. twelve [Gri85, McC86]. twin [SDH84]. TWO [VMS81, Cla89, Cra86a, EIT85, EP87a, EBS88, KWWK86, Wal3 , Dal88b, GRB88, HS86, kK89c, LCH^{+}88, EP89]. Two-dimensional [KWWK86, GRB88, HS86, kK89c]. Two-electron [EBS88]. two-way [Dal88b]. type [SM87, SMD84, SR84a, Whe84a, Whe84b]. tzu [mCaLjH84].

U [Gru88]. U.W. [Dun85b]. Uberblicke [MF84]. UCSD [Ano82c, BG82, BGCS82]. Ueberarbeitung [Wie85]. Ueberblicke [Mau84]. Ueberpruefung [Wid88]. Ueberwachung [Lep86]. Ultracomputer [Gre86]. ULTRIX [Dig86b, Dig86c, Dig86a]. unabhaengige [Sto84a, Sto84b]. UNAERO [Dun85a]. Unconstrained [MGH81b, MGH81a]. Understanding [Boi81, BW84, Boi84, Boi85, BW87b, Boi87, DBFK89, Lud81, WB85]. unfolding [YHKM89]. ungaged [Uni81a]. UNICOS [Cra89b]. Uniform [Gui89, MR83, DZ88, Gui88]. Uniformly [PLR85]. Unimate [Fra84b]. unimodal [MT84a]. Uniprocessor [Cal86]. Unit [CHH81, CHH83, Sch81a, Lag85, Min88, Mul85, PF85]. UNIVAC [SDC82]. univariate [TSU88, TFH86, Wal3 ]. University [IEE81, Pie85, Lee85]. UNIX [Ber85a, Cod89, Pay84a, Pay84b, Uni82b]. Unordered [MP86a, MP86c, MP86b]. unsteady [Dun85a, Wal85]. Untersuchung [RAKK88, San82]. Unusual [DR86]. Update [FCG83, KKK89, Hof87, Jac85b, Kir89]. Updating [Buc81a, Buc81b, Buc81c, Buc82]. Upper [FW82, Ste76]. Uranium [LZ82]. uravnenii [Sko88]. uravneniia [Sko88]. Urbana [Rei89c]. Ure [PFF83]. USA [Wri89, Sof84]. Usage [DG82, Dod83, HK87a, HK87b, LHKK79a, LHKK79b, LN88b, LN88a, Con83c, Con84, Con85b, Con85d, Con86, Con87c, Con87a, Con87b, Con88a, LN87, Ols83, Wie82]. usager [TR84]. Use [GSZ88, Gol82b, LO85a, LO85b, BK89, Con83a, Con85a, Col82, Hos88, Int86c, IS84a, IS84b, Nag81a, Nag85, RW85, SL82, Sul88, VC89]. used [Per81]. useful [RH84b, TMS88a, TMS88b]. USENIX [Sof83b, Sof84, Usr82]. User [AHU81, Ano84, ADP88, BMS84, Bre81, Bri84, Sof87, Dig85b, Den82, DPA87, GRB88, Gic88, GHM^{+}86, Gil86, Hua82, IMS82, Lib84a, IMS84, IMS87n, IMS89h, IMS89m, IMS89i, IMS89j, IMS89k, IMS89l, IMS89n, Kle89b, Kle89a, MSG86, Obl85, RMS82, Sym85, Sym86, Sym88, TMjC81, Tew81, Uni84a, Uni84b, Uni81a, VMS81, Zho89, AD84, Ano82b, Ano85c, Ano87c, Apo83, Bur85b, Con81a, Con82a, Con82c, CB86, CBS81, CB82, CDW83a, CDW83b, DW85, Dig82e, Dig84h, Dig84i, Dig86c, Dig88c, DW83b, Dir84, Fed82b, Fed82a, GC84, Gue86, Int88c, Lib84b, Lib87, IMS87c, IMS87b, IMS87d, IMS87g, IMS87j, IMS87l, IMS87m, IMS89c, IMS89e, IMS89g, IS84a, IS84b, ISJ85, Int84a, Int85a, Mar81, Mic84c, Mic85c, Mic87f, Nor83, Rel83, Rel86, Rel88, Rel89b, Tho81, Tho82a, Uni83c, Uni84c, Uni86b, Uni84d]. user-interface [Nor83]. Users [Usr82, Ber85a, Int84b, Iwa84]. Usersin [Nor83]. Using [AG87c, AEV89, Bar89, Cod89, CC82, DAG^{+}88, Don84a, Don84d, Don85a, Don87b, Don88b, Don89, GY82, Gre86, Gro83, HG83, McK85b, Mos88, Rei87a, Ros87, Sch82d, Sch88a, SP84a, SW83, SR84b, SR86, TvSS82, dEV89, Ack84, AM89a, AG87a, AG87b, CK86a, Don83b, Don84b, Don85b, Don87a, Don88c, DS82, GS81, Gra84b, Gro89, Hig86, HP88, HPR81, Kem87, LW88b, Rod87, Num84d, Red86, SH88, Sch88b, SP85a, SP85b, Sul2 , Tan81a, Tan82, Tat87, Tur86, VSH83, Wat86, Wie86a, Wie86b, Pem83]. USSAERO [Wie86a, Wie86b]. USU [Gic88]. Utah [Sof84, Ket85b, Ket84, Ket85a]. utilisation [Ano85b]. utilities [Int85b, Mic87e, Pee85a, Pee85b, Pee85c]. utility [BS86c, CR84, GC84, Mai87]. utilization [CT88].

V [Fra84a, Lav83, Lig82a, Lig82b, Lig84, Lig85a, MSG86, SPS84, Tho86a]. V1.1 [Pre87a]. V2.5.3 [Cod86b]. V3.0 [Dig82f]. V4.1 [Dig84j]. Validation [BSP83, How82, Sof83a, Tho86a, Fed81, Fed82a]. Valley [MF84]. valleys [SPS84]. Value [Cas89a, Cas89b, EP87a, EP89, Hig91, CR84]. Variable [SP82, Dav86, Gra86b, PM87, RMS82]. variable-magnetization [Gra86b]. Variables [Maa89]. Varian [Moo81, Moo83]. Variates [HPB82]. VARIATM [LN89]. varied [BMS84]. variety [Alc82, BT83]. Various [Don84a, Don84d, Don85a, Don87b, Don88b, Don89, Don83b, Don84b, Don85b, Don87a, Don88c, ZGK88]. VARMAG [Gra86b]. VAST [Bro82a, LCH^{+}88]. VAST-2 [LCH^{+}88]. VAX [Fed81, Ano87c, Ber85a, Cal85, Chi85a, Dig82a, Dig82b, Dig82d, Dig82c, Dig82e, Dig82f, Dig84a, Dig84d, Dig84f, Dig84e, Dig84g, Dig84h, Dig84i, Dig84j, Dig85a, Dig85c, Dig85d, Dig86b, Dig86c, Dig86a, Dig86d, Dig88b, Dig88a, Dig88c, Dun85b, Gre88, hHtM81, Joh87b, Joh87c, JC88, JC89, Kle89b, Kle89a, Mid84, Moo81, Moo83, Sul88, sT85, Uni88, Wat86, Wei86a, Wei86b, Wei89b, Wie82]. VAX-11 [Dig82a, Dig82b, Dig82d, Dig82c, Dig82e, Dig82f, Dig84e, Dig88a, Sul88, sT85, Wie82]. VAX-11/ [sT85]. VAX-II [Fed81]. VAX-VMS [hHtM81]. VAX. [Ano85b]. VAX/VMS [Dig84a, Dig86d, Gre88, Joh87b, Joh87c, JC88, JC89, Moo81, Moo83]. VAXELN [Dig85e]. VAXIMA [SR84b]. VE [Con83c, Con84, Con85b, Con85d, Con85c, Con86, Con87c, Con87a, Con87b, Con88a]. Vector [ACG^{+}86, AK87, Ano88a, Blu78, CT86b, DD86, GPHL90, LS85, LS88, MS88a, Pet83, Ric84, Riz85, vdV86, ZM86, BK88, Bro82a, CT88, CMM^{+}88, CT86a, GF89, GSZ88, Guz88, GPHL88, Int86a, Lec89, LS87, LCH^{+}88, MMM85, RRS88, RS87, Sch89a, Sor84, Swa84, hTDT88, VSH83, VH87, van86, AC86, RV89]. vectorizable [hTD88]. Vectorization [Bos88, VSH83, AJ88]. Vectorized [GDK89, Col89a, Col89b, VH87]. Vectorizer [Arn82]. Vectorizing [ACK86a, CDL88, EBS88, NE89, ACK86b, SK86]. vegetation [Mal85]. Vektorisieren [MW83]. vented [WP84]. Verbindungslisten [Wie85]. Vergleich [Hof84, Hah81]. Vergleichende [San82]. verification [KS88, Tho86a, Var85]. versatile [Nag85]. Version [Ano85b, Ber84b, Con81b, Con88a, Cod86a, HV83, Int87f, Int88i, Int88j, Int88k, Int88l, Int88m, Int88n, Int88o, Int88p, IBM89a, Int89c, MS84b, PFTV86, Sch84c, Sch84d, Sch84e, Sch84b, Sch87a, (??87, AI88, And89, BGCS82, Bru86, Con81a, Con82b, Con82c, Con83a, Con83b, Con83d, Con85a, CBS81, CB82, Dig85b, Doe88, Fed81, GRB88, Gic88, GHM^{+}86, Gil86, Gre88, HL82b, Hua82, Hud88a, Hud88b, Int86e, Int86f, Int86g, Int87c, Int87d, Int87e, Int87g, Int87h, Int87i, Int88e, Int88f, Int88h, Int89a, Int89b, Int89d, Int89e, IBM89b, KM89, LB89, LOU86, Met89a, Mai87, Mon82a, Mon82b, MT84b, NM85, NJLB81, Sch87b, SDC82, (??84, Tri84, Tri89, TR84, Uni83b, Van85, Wie86a, Wie86b, Zho89, Obl85]. versus [Joh84]. vertical [HS86, The88]. very [Gui88, Jan84]. Vespan [Mal85]. VF [ALPC88]. VI.1 [Num85c]. viable [LD87]. vibration [Jam86a, Jam86b]. vibrational [Kie86]. vicinity [Tho84a]. video [Gri85]. VII [Sav87]. Vindicated [KTW84]. Virtual [IBM88, Int88d, Con83a, Con85a]. viscosities [KWM88]. Visionaries [Tay86]. VM [Ber89, DW83b, Hew85, Int82e, Int85b, Int88b, Sto85a, Sto85b, Uni83c, Uni84c, Uni86b]. VM/ [Int82e]. VM/370 [Int85b]. VM/370-III [Int85b]. VM/CMS [Ber89, Uni83c, Uni84c, Uni86b]. VM/EPEX [Sto85a, Sto85b]. VM/SP [DW83b]. VM/XA [Int88b]. VM/XA-SP [Int88b]. VME [Uni87]. VMS [Ano87c, Dig84a, Dig86d, Gre88, hHtM81, Joh87b, Joh87c, JC88, JC89, Moo81, Moo83, TR84, Uni88, Wat86]. volnovykh [BZ85]. Volterra [Bow82, BA85b, BA85c, BA85a]. Volume [Coc83, Owe87, Riz85]. Vortex [HL82b]. Voruebersetzer [Kal85a]. vous [Ain89]. VP [MMM85]. VP-100 [MMM85]. VP-100/200 [MMM85]. VS [Int88a, Int88c, Pag87, hTDT88, Ano82d, Dat84, Int81b, Int81c, Int81d, Int82f, Int82g, Int82h, Int82i, Int82j, Int83d, Int83e, Int83f, Int83g, Int83h, Int83i, Int84c, Int84d, IBM85, Int85c, Int85d, Int85e, Int85f, Int85g, Int85h, Int85i, Int86d, Int86e, Int86f, Int86g, Int87c, Int87d, Int87e, Int87f, Int87g, Int87h, Int87i, Int88e, Int88f, Int88h, Int88g, Int88i, Int88j, Int88k, Int88l, Int88m, Int88n, Int88o, Int88p, Int89a, Int89b, IBM89a, Int89c, Int89d, Int89e, IBM89b, Kir89, LS87, LS88, Sof83a, Wan84]. VSPC [Har86b]. VX [Ess88].

W [Mil82b, Rzy84]. WADA1 [And89]. Waerme [RS82]. wakes [HL82a]. Walsh [NEM84]. Walsh-Transformation [NEM84]. WATCOM [CDW83a, CDW83b, DW85, DW83a, DW83b, DW84, Dir84]. WATEQ2 [BNZ87]. WATEQ4F [BNZ87]. water [BMS84, KD84a, KD84b, KWWK86, RMS82]. Waterloo [CDW83b, Dir81]. watersheds [Uni81a]. Watfiv [Ett84b, McK85a, AM84, BS81d, DH82, DH83, DH84a, Ett85, FK82, KS82b, Mer85, Moo85b, Sas83a, Sas83b, Sin81, Tel82, BGG86, MM81]. WATFIV-S [DH82, DH83, DH84a, BGG86]. WATFOR [BS81d, DH82, DH83, DH84a, KS82b, Ler83]. WATFOR/WATFIV [KS82b]. Watson [KM83]. Wave [EW87, And89, BZ85, LW88b, ZGS89]. Wave-Current [EW87]. wave-propagation [LW88b]. wavefront [Slo88]. waveguide [Bai89]. waveguide-fed [Bai89]. waves [Tho81, Tho82a]. Way [LP85b, Dal88b, LP85a]. weakness [Sav87]. Weeks [GGLM88, GL90]. Weibull [Szy87]. Weights [EK87b, EK87a]. Weirel [Joh85b]. Weiterentwicklung [Ull84]. Well [BBF^{+}82, BT83, Gla88, Hig86, WD81a, WD81b]. well-stirred [Gla88]. WELLMAP [BT83]. Wesley [Cou85a, Rid82a]. WG [Wri89]. WG5 [Rei89c]. WHCS [WD81a, WD81b]. Where [WM85b]. whether [Bur86a]. white [Fra84a]. who [Bro81b]. Whole [AL81b, AL81a]. Whole-Rock [AL81b, AL81a]. width [LKM88]. WIENER [Chu88]. Wigner [AD84]. WILDMAP [Boy84a]. Wiley [Edm86]. window [Owe87]. Wise [Per83a, Sho85]. without [MMM85]. WITS [Ler83]. wolli [mK84]. work [Wag84]. Workbook [CM84b, SFKS81, Wor84b, RWA84, Rod84, WAD81]. Working [Int86c, Wri89]. Workstation [Whi89, RMFG85, Sun84, Sun85]. Workstations [Wan85]. World [ML87, VV86]. Write [SR84b, SR86, AS89a, AS89b]. writer [Bar84]. writeups [FE82]. Writing [Ano88c, Oni85, Smi85c, Smi88c, Int86a]. Written [BP81a, BP81b, LB86]. Wsing [McK85a]. WY [Rei89c].

X [Abe89, AGS88, Hoc85, JL81a, JL81b, MT84b, RS85, hTD88, TPR85, Tem89a, Tem89b, Tho84e, ZM86, vM84f]. X-MP [AGS88, Hoc85, RS85, hTD88, Tem89a, Tem89b, ZM86]. X-MP/24 [AGS88]. X-ray [Abe89, MT84b, Tho84e, JL81b, JL81a, TPR85, vM84f]. X.25 [And84b]. X3.124 [Ame85d]. X3.124-1985 [Ame85d]. X3.9 [AC87, Ame87b, Ano82a, Com89, AC87, Com89, Ehi82]. X3.9- [Ehi82]. X3.9-1978 [Ano82a, AC87, Com89]. X3.9-198X [AC87, Com89]. X3J3 [BBB^{+}83, Rei89c, Ame87d, Mei89d, Rei87b, Rei87d, Rei87c, Rei87e, Rei89c]. X3J3-Meetings [BBB^{+}83]. XA-SP [Int88b]. XERROR [JK83]. xii [Nad86]. XRDPLT [JL81a, JL81b]. XREF [TR84]. XSRAIN [Uni81a]. XT [Fuo86a, Joh86].

yen [Cha83, mCaLjH84, hH82, sKcH81, iL82, fS82, fTBcL7 , cT81, cTcT84, mT82b, yW85]. Yes [McG84]. yin [LcY83]. yonsup [hYsA82]. Yorktown [DO86]. Yourself [Rad81, Rad83b]. yu [Cha83, mCaLjH84, CwL83, hH82, sKcH81, iL82, fS82, fTBcL7 , cT81, cTcT84, mT82b, yW85]. yung [Cha83, hH82, RkC84].

Z [Ass83b, Kah80]. Z80 [Hei83]. Z80cpu [Hei83]. zadach [BZ85, Sko88]. zeitdiskreten [Hel85b]. zero [Alb86]. zerophase [Alb86]. zitureisyu [SS84]. zone [SP84b, SP84c]. zur [CLW81, Dah81, EmR84, Fis82, Hel85b, Hof84, Kna84, Mey84, RAKK88, RS82, San82, Sch84f, Sch87c, Ste87, Wie85]. zusammenhängende [KS81c]. Zweier [Hof84].

References

Tricot:1984:MBFa

[(??84] Laurence Tricot (??). Modulad: Bibliothèque FORTRAN pour l’analyse des données: brochure de documentation: version 1.0. INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, 1984. ISBN 2-7261-0374-X. 268 pp.

Tricot:1987:MBF

[(??87] Laurence Tricot (??). MOD-ULAD: Bibliothèque Fortran 77 pour l’analyse des données. Version 2.1. INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, 1987. ISBN 2-7261-0469-X. 547 pp.

xxx:1988:PC

[??88]  ??, editor.  Parallel Comput-

ing 9. North-Holland Publishing Co., Amsterdam, The Netherlands, 1988. ISBN ???? LCCN ????

Ashcroft:1981:PF

[A^{+}81]  J. Ashcroft et al.  Program-

ming With Fortran 77. Sheridan House: Distributed exclusively in North America by Renouf/USA, London, UK; New York, NY, USA: Granada; Brookfield, VT, June 1981. ISBN 0-246-11573-4 (paperback). 294 pp. LCCN QA76.73.F25 P76 1981. US$29.50; UK£6.95 (US$19.95

U.S.). URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0246115734. Dis-

tributed exclusively in North America by Renouf/USA.

Adey:1983:BCT

[AB83] R. A. Adey and C. A. Brebbia. Basic Computational Techniques for Engineers. John Wiley and Sons, New York, London, Sydney, June 1983. ISBN 0-471-88970-9. 208 pp. LCCN TA345 .A33 1983. US$34.95.

Autin:1989:SEI

[AB89] B. Autin and J. Bengtsson. Symbolic evaluation of integrals occurring in accelerator orbit theory. Journal of Symbolic Computation, 7(2):183–188 (or 183–187??), February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Allen:1988:OPA

[ABC^{+}88] F. Allen, M. Burke, P. Charles, R. Cytron, and J. Ferrante. An overview of the PTRAN analysis system for multiprocessing. Journal of Parallel and Distributed Computing, 5(5):617–640, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abdelmalek:1980:AFS

[Abd80] Nabih N. Abdelmalek. Algorithm 551: A FORTRAN subroutine for the L_{1} solution of overdetermined systems of linear equations [F4]. ACM Transactions on Mathematical Software, 6(2):228–230, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Abel:1989:FPX

[Abe89] Phillip Benjamin Abel. Fortran program for X-ray photoelectron spectroscopy data reformatting. NASA technical paper 2957, National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Division, Washington, DC, USA, 1989. 7 pp. For sale by the National Technical Information Service.

Abstreiter:1988:PPF

[Abs88] F. Abstreiter. PARFOR — Paralleles FORTRAN: Benutzungsanleitung unter BS2000. Technical report, Institut für Informatik, TU München, München, Germany, December 1988.

Agarwal:1986:FTC

[AC86] Ramesh C. Agarwal and James W.

Cooley. Fourier transform and convolution subroutines for the IBM 3090 Vector Facility. IBM Journal of Research and Development, 30 (2):145–162, March 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

ANSI:1987:ANSb

[AC87] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language Fortran: S8 (X3.9-198X): Revision of ANSI X3.9-1978. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1987. various pp.

Agarwal:1986:NSV

[ACG^{+}86] Ramesh C. Agarwal, James W. Cooley, Fred G. Gustavson, James B. Shearer, Gordon Slishman, and Bryant Tuckerman. New scalar and vector elementary functions for the IBM System/370. IBM Journal of Research and Development, 30(2):126–144, March 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Argon:1988:MSP

[ACG^{+}88] Elvira Argon, I.-Lok Chang, Gamini Gunaratna, David K. Kahaner, and Martin A. Reed. Mathematical software: Plod. IEEE Micro, 8(4):56–61, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ackeret:1984:IFP

[Ack84] James R. Ackeret. An interactive FORTRAN program for determining reliability of pseudorange geodetic point positioning using the global positioning system. Thesis (M.S.), Ohio State University, Columbus, OH, USA, 1984. x + 125. pp.

Allen:1986:IIA

[ACK86a] Randy Allen, David Callahan, and Ken Kennedy. An implementation of interprocedural analysis in a vectorizing Fortran compiler. Technical Report TR86-38, Department of Computer Science, Rice University (??), Houston, TX, USA, May 1986.

Allen:1986:IID

[ACK86b] Randy Allen, David Callahan, and Ken Kennedy. An implementation of interprocedural data flow analysis in a vectorizing Fortran compiler. Computer science technical report COMP TR86-38, Rice University, Dept. of Computer Science, Houston, TX, USA, 1986. 13 pp.

SIGPLAN:1982:PSS

[ACM82] ACM Press. Proceedings of the SIGPLAN ’82 Symposium on Compiler Construction. ACM Press, New York, NY 10036, USA, 1982. ISBN ???? LCCN ???? Available as SIGPLAN Notices 17(6) June 1982.

SIGPLAN:1984:PSS

[ACM84] ACM Press. Proceedings of the SIGPLAN ’84 Symposium on Compiler Construction. ACM Press, New York, NY 10036, USA, 1984. ISBN ???? LCCN ???? Available as SIGPLAN Notices 19(6) June 1984.

SIGPLAN:1987:PSS

[ACM87] ACM Press. Proceedings SIG-PLAN ’87 Symposium on Interpreters and Interpretive Techniques. ACM Press, New York, NY 10036, USA, June 1987. ISBN ????

LCCN ???? Also available as SIG-PLAN Notices 22(7) July 1987.

ACM:1989:PAI

[ACM89a] ACM, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation, ISSAC ’89. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN ????

ACM:1989:PSN

[ACM89b] ACM, editor. Proceedings, Supercomputing ’89: November 13–17, 1989, Reno, Nevada. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Akiyama:1984:UGF

[AD84] Yoshimi Akiyama and J. P. Draayer. A user’s guide to Fortran programs for Wigner and Racah coefficients of SU(3). Computer Physics Communications, 35 (1–3):C–194, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/S001046558482439X.

Adams:1984:IF

[Ada84] Jeanne Adams. Institutionalization of FORTRAN. Annals of the History of Computing, 6(1): 28, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1028.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1028abs.htm.

Adams:1989:MMP

[Ada89] John C. Adams. mudpack:

Multigrid portable Fortran software for the efficient solution of linear elliptic partial differential equations. Applied Mathematics and Computation, 34 (2 (part 2)):113–146, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0096300389900106.

Agrawal:1989:DMO

[ADH^{+}89] H. Agrawal, R. Demillo, R. Hathaway, Wm. Hsu, Wynne Hsu, E. Krauser, R. J. Martin, A. Mathur, and E. Spafford. Design of mutant operators for the C programming language. Technical Report SERC-TR-41-P, Software Engineering Research Centre, Utrecht, The Netherlands (??), March 20, 1989.

Adman:1984:FN

[Adm84] Peter Adman. Fortran 77 for nonscientists. A Chartwell-Bratt student text. Chartwell-Bratt, Bromley, Kent, UK, 1984. ISBN 0-86238-074-X (paperback), 91-44-23061-3 (Sweden). 101 pp.

Adman:1985:FSN

[Adm85] Peter Adman. Fortran 77: solutions to non-scientific problems. Chartwell-Bratt student text. Chartwell-Bratt, Bromley, Kent, UK, 1985. ISBN 0-86238-087-1 (paperback). 150 pp.

Antoniewicz:1988:UMI

[ADP88] Robert F. Antoniewicz, Eugene L. Duke, and Brian P. Patterson. User’s manual for interactive LIN-EAR, a FORTRAN program to derive linear aircraft models. NASA technical paper 2835, National Aeronautics and Space Administration, Scientific and Technical Information Division, Washington, DC, USA, 1988. ???? pp. For sale by the National Technical Information Service.

ASC:1986:TGL

[Adv86] Advanced Systems Consultants. Tekmar graphics library and Tekmar technical graphics software and high-level Fortran and graphics library, 1986.

Arvind:1987:FSPb

[AE87a] Arvind and K. Ekanadham. Future scientific programming on parallel machines. Computation Structures Group Memo 272, Laboratory for Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1987.

Arvind:1987:FSPa

[AE87b] Arvind and Kattamuri Ekanadham. Future Scientific Programming on Parallel Machines. Computation Structures Group Memo 272, Laboratory for Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, March 1987.

Arvind:1988:FSP

[AE88] Arvind and K. Ekanadham. Future scientific programming on parallel machines. Journal of Parallel and Distributed Computing, 5(5): 460–493, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Arnold:1986:CPD

[AEL^{+}86] J. Arnold, V. K. Eckenrode, K. Lemke, G. J. Phillips, and S. W. Schaeffer. A comprehensive package for DNA sequence analysis in FORTRAN IV for the PDP-11. Nucleic Acids Research, 14(1):239– 254, January 10, 1986. CODEN NARHAD. ISSN 0305-1048.

Antonio:1989:UPL

[AEV89] R. Di Antonio, J. Eilert, and M. Vitaletti. Using PAGE-AHEAD for large FORTRAN programs. In Proceedings of the Supercomputing ’89 (Reno, NV, USA), pages 511–520. ACM Press, New York, NY 10036, USA, 1989.

AFNOR:1983:LPF

[AFN83] AFNOR. Le langage de programmation FORTRAN. Norme ISO 1539 (norme NF Z 65-110), Association Française de Normalisation, Tour Europe, Cedex 7, F-92080 Paris La Défense Cedex, Paris, France, 1983.

Averbukh:1994:RA

[AFS94] Victoria Z. Averbukh, Samuel Figueroa, and Tamar Schlick. Remark on algorithm 566. ACM Transactions on Mathematical Software, 20(3):282–285, September 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [MGH81b].

Audin:1985:FGP

[AFV85] Eugene Audin and Françoise Ficheux-Vapne. FORTRAN 77: guide pour l’écriture de programmes portables. Collection de la Direction des études et recherches d’Electricité de France, 60 0399-4198. Eyrolles, Paris, France, 1985. ISBN ???? 145 pp. LCCN ????

Angell:1987:HCGa

[AG87a] Ian O. Angell and Gareth Griffith. High-resolution computer graphics using Fortran 77. Macmillan computer science series. MacMillan Publishing Company, New York, NY, USA, 1987. ISBN 0-333-40398-3 (cased), 0-333-40399-1 (paperback). ix + 355 pp.

Angell:1987:HCGb

[AG87b] Ian O. Angell and Gareth Griffith. High-resolution computer graphics using FORTRAN 77. John Wiley and Sons, New York, London, Sydney, 1987. ISBN 0-333-40398-3 (cased), 0-333-40399-1 (paperback), 0-470-20773-6 (paperback). ix + 355 pp. LCCN QA76.73.F25 A54 1987.

Angell:1987:HRC

[AG87c] Ian O. Angell and Gareth Griffith. High-Resolution Computer Graphics Using Fortran 77. Halsted Press, New York, USA, October 1987. ISBN 0-470-20773-6. 355 pp. LCCN ???? US$55.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0470207736.

Anderson:1988:PCC

[AGS88] Richard E. Anderson, Roger G.

Grimes, and Horst D. Simon. Performance comparison of the CRAY X-MP/24 with SDD and the CRAY-2. The Journal of Supercomputing, 1(4):409–419, August 1988. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=1&issue=4&
spage=409.

Aharonian:1985:AFC

[Aha85a] G. Aharonian. Automated Fortran conversion. ACM SIGPLAN FOR-TRAN Forum, 4(4):14–19, December 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Aharonian:1985:LKF

[Aha85b] G. Aharonian. Let’s kill future Fortrans. ACM SIGPLAN FOR-TRAN Forum, 4(3):21–23, October 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Aharonian:1985:MLB

[Aha85c] G. Aharonian. Multi-language BLAS: a proposal. ACM SIG-PLAN Notices, 20(11):11–13, November 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aharonian:1985:TCA

[Aha85d] Gregory Aharonian. Technical contributions: Automated Fortran conversion. ACM SIGPLAN FOR-TRAN Forum, 4(4):13–19 (or 14– 19??), December 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Husayni:1983:AFA

[aHH83] Abd al Hasan Husayni. al-Musaid fi al-barmajah: FORTRAN IV. Dar al-Qalam, Bayrut, Lebanon, al-Tabah 1 edition, 1983. ISBN ???? 167 pp. LCCN ????

Atkinson:1989:NMF

[AHH89] Laurence Atkinson, P. J. Harley, and J. D. Hudson. Numerical Methods With Fortran 77: a Practical Introduction. International Computer Science Series. Addison-Wesley, Reading, MA, USA, March 1, 1989. ISBN 0-201-17430-8. x + 395 pp. LCCN QA76.73.F25 A84 1989. US$39.76; US$35.50; UK£18.00

(est.). URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0201174308.

ANW:1981:HSP

[AHU81] Anderson-Nichols/West, Hydrocomp Incorporated, and United States. Environmental Protection Agency. Office of Research and Development. Hydrological simulation program — FORTRAN (HSPF): User’s manual — Release 7.0. U.S. Environmental Protection Agency, Athens, GA, USA, 1981. 677 pp.

Aldea:1988:FAE

[AI88] N. Aldea and E. Indrea. Fourier analysis of EXAFS data — a selfcontained FORTRAN programpackage — a second version. Computer Physics Communications, 51(3):451–462, November 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465588901580.

Ain:1989:SPF

[Ain89] Maryse Ain. Savez-vous parler Fortran?: cours progressif de programmation structurée en Fortran 7. Publications universitaires scientifiques. Orsay Plus, Orsay, France, 1989. ISBN 2-87800-000-

5. 482 pp.

Aird:1977:PMS

[Air77] Thomas J. Aird. Portability of mathematical software coded in Fortran. ACM Transactions on Mathematical Software, 3(2):113– 127, June 1977. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Allen:1988:CCV

[AJ88] R. Allen and S. Johnson. Compiling C for vectorization, parallelization, and inline expansion. ACM SIGPLAN Notices, 23(7):241–249, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:80/
pubs/citations/proceedings/
pldi/53990/p241-allen/.

Allen:1981:PPC

[AK81] John R. Allen and Ken Kennedy.

PFC: a program to convert Fortran to parallel form. Technical report, Rice University, Department of Mathematical Sciences, Houston, TX, USA, 1981. 63 pp.

Allen:1982:PPC

[AK82] J. R. Allen and K. Kennedy. PFC: a program to convert Fortran to parallel form. Technical Report MASC-TR82-6, Rice University, Houston, TX, USA, 1982.

Allen:1984:ALI

[AK84] John R. Allen and Ken Kennedy. Automatic loop interchange. In Proceedings of the SIGPLAN ’84 Symposium on Compiler Construction [ACM84], pages 233–246. ISBN ???? LCCN ???? Available as SIGPLAN Notices 19(6) June 1984.

Allen:1985:PPE

[AK85] John R. Allen and Ken Kennedy. A parallel programming environment. IEEE Software, 2(4):21– 29, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Allen:1987:ATF

[AK87] Randy Allen and Ken Kennedy. Automatic translation of Fortran programs to vector form. ACM Transactions on Programming Languages and Systems, 9 (4):491–542, October 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

ANSSSR:1988:MPM

[Aka88] Materialy po matematicheskomu obespecheniiu EVM, page various, 1988. Nauch. tsentr biologicheskikh issledovanii AN SSSR v Pushchine, Pushchino, USSR.

Albert:1988:CFA

[AKLS88] Eugene Albert, Kathleen Knobe, Joan D. Lukas, and Guy L. Steele, Jr. Compiling Fortran 8x array features for the Connection Machine computer system. ACM SIGPLAN Notices, 23(9):42–56, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:80/
pubs/citations/proceedings/
plan/62115/p42-albert/.

Andrew:1981:PAF

[AL81a] A. S. Andrew and J. Linde. PRP:

A Fortran IV interactive plotting program (for whole-rock analysis). Computers and Geosciences, 7:3– 20, 1981. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Andrew:1981:PFI

[AL81b] A. S. Andrew and J. Linde. PRP:

a Fortran IV interactive plotting program (for whole-rock analysis). Computers and Geosciences, 7:3– 20, 1981. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Albert:1986:FSZ

[Alb86] Donald G. Albert. FORTRAN subroutines for zero-phase digital frequency filters. Special report 86-4, US Army Corps of Engineers, Cold Regions Research and Engineering; Available from National Technical Information Service, Hanover, NH, USA, March 1986. iii + 26 pp.

Alcock:1982:IFP

[Alc82] Donald Alcock. Illustrating Fortran (the portable variety). Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24598-2, 0-521-28810-X (paperback). ix + 134 pp. LCCN QA76.73.F25 A4 1982.

Alcock:1983:IFa

[Alc83a] Donald. Alcock. Illustrating Fortran. Cambridge University Press, Cambridge, UK, July 1983. ISBN 0-521-24598-2. ???? pp. LCCN ???? US$49.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0521245982.

Alcock:1983:IFb

[Alc83b] Donald Alcock. Illustrating Fortran. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-28810-X. ???? pp. LCCN ???? US$24.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
052128810X.

Allen:1982:TRF

[All82] F. E. Allen. A technological review of the FORTRAN I compiler. In Morgan [Mor82], pages 805– 809. ISBN 0-88283-035-X. LCCN TK7885.A1 J6 1982. URL http:

//community.computerhistory.
org/scc/projects/FORTRAN/paper/
Allen-technological_review_
FORTRAN-1982.pdf.

Allen:1984:TRE

[All84] Frances E. Allen. A technological review of the early FORTRAN compilers. Annals of the History of Computing, 6(1):22–25, January/March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1015.pdf;
http://www.computer.org/annals/
an1984/a1015abs.htm.

Alliant:1987:FFL

[All87] Alliant Computer Systems Corporation. FX/FORTRAN language manual. Alliant Computer Systems Corp., Littleton, MA, USA, 1987. various pp.

Allison:1990:IMC

[All90] Bob Allison. Interfacing Microsoft C and FORTRAN. The Journal of C Language Translation, 1(4):300– 304, March 1990. ISSN 1042-5721.

ACS:1983:F

[Alp83] Alpha Computer Service. Forlib-Plus, 1983. 1 computer floppy disk.

Angeleri:1988:PCI

[ALPC88] P. Angeleri, D. F. Lozupone, F. Piccolo, and J. Clinckemaillie. PAM-CRASH on the IBM 3090/ VF: an integrated environment of crash analysis. IBM Systems Journal, 27(4):541–560, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Ageloff:1981:AFF

[AM81] Roy Ageloff and Richard Mojena. Applied Fortran 77: Featuring Structured Programming.

Wadsworth, Pacific Grove, CA, USA, May 1981. ISBN 0-534-00961-1. xvi + 604 pp. LCCN QA 76.73 F25 A33 1981. US$33.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534009611.

Ageloff:1984:ASW

[AM84] Roy Ageloff and Richard Mojena. Applied structured WAT-FIV. Wadsworth, Pacific Grove, CA, USA, March 1984. ISBN 0-534-03079-3 (paperback). xvii + 619 pp. LCCN QA76.73.F25A336 1984. US$31.75.

Ahn:1989:PFP

[AM89a] C. W. Ahn and J. E. Mezzich. PROPOV-K: a FORTRAN program for computing a kappa coefficient using a proportional overlap procedure. Computers in Biomedical Research, 22(5):415–423, October 1989. CODEN CPBRAF. ISSN 0010-4809 (print), 1090-2368 (electronic).

Ashrafiuon:1989:ASC

[AM89b] H. Ashrafiuon and N. K. Mani. Applications of symbolic computing for numerical analysis of mechanical systems. American Society of Mechanical Engineers, Design Engineering Division (Publication) DE, 19-3(pt 3):141–149, 1989. CO-DEN AMEDEH.

ANSI:1985:ANSb

[Ame85a] American National Standards Institute. American National Standard for information systems: computer graphics — graphical kernal system (GKS) FORTRAN binding. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? 106 pp. LCCN ????

Institute:1985:ANS

[Ame85b] American National Standards Institute. American National Standard for information systems: computer graphics — graphical kernal system (GKS) FORTRAN binding. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? 106 pp. LCCN ????

ANSI:1985:ANSa

[Ame85c] American National Standards Institute. American National Standard for information systems: computer graphics — Graphical Kernel System (GKS) functional description. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? 268 + 106 pp. LCCN ????

ANSI:1985:ISC

[Ame85d] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems—Computer Graphics— Graphical Kernel System (GKS). ANSI X3.124-1985, 1985. Includes Fortran bindings to GKS.

ANSI:1987:ANSa

[Ame87a] American National Standards Institute. American National Standard programming language, FOR-TRAN: draft proposed. American National Standard; ANSI X3.9-1978 American National Standards Institute. American National Standard; ANSI X3.9-1987. Global Engineering Documents, Washington, DC, USA, revised edition, 1987. ISBN ???? various pp. LCCN ????

ANSI:1987:DPA

[Ame87b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed ANSI Fortran X3.9–198x, September 18, 1987. See also [MR87].

ANSI:1987:DPR

[Ame87c] American National Standards Institute. Technical Committee X3J3 — Fortran. Draft proposed revised American National Standard programming language Fortran. X3 Secretariat/CBEMA, Washington, DC, USA, version 104 edition, 1987. ISBN ???? various pp. LCCN ????

ANSI:1987:XDF

[Ame87d] American National Standards Institute. Technical Committee X3J3 — Fortran. X3J3 draft Fortran Standard submitted document. X3 Secretariat/Computer and Business Equipment Manufacturers Association, Washington, DC, USA, 1987. ISBN ???? 166 columns. pp. LCCN ????

X3J3:1987:DPR

[Ame87e] American National Standards Institute. Technical Committee X3J3. Draft proposed revised American National Standard programming language Fortran. X3 Secretariat/CBEMA, Washington, DC, USA, version 104 edition, 1987. various pp.

ANSI:1989:F

[Ame89a] American National Standard for Information Systems programming language: Fortran. Fortran. Fortran Forum; 8, no. 4. ACM Press, New York, NY 10036, USA, revision of X3.9-1978 edition, 1989. ISBN ???? various pp. LCCN ????

Institute:1989:ANS

[Ame89b] American National Standards Institute. American National Standard for Information Systems programming and language: Fortran 8X draft, volume 8(4) of Fortran Forum. ACM Press, New York, NY 10036, USA, May 1989. various pp. Revision of X3.9-1978.

Amos:1983:APFa

[Amo83a] D. E. Amos. Algorithm 609:

a portable FORTRAN subroutine for the Bickley functions Ki_{n}(x). ACM Transactions on Mathematical Software, 9(4):480–493, December 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Amos:1983:APFb

[Amo83b] D. E. Amos. Algorithm 610:

a portable FORTRAN subroutine for derivatives of the Psi function. ACM Transactions on Mathematical Software, 9(4):494–502, December 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Anantharaman:1987:APE

[Ana87] Th. S. Anantharaman. Algorithms for parallel execution of programs with memory aliasing. Report CMU-CS-87-109, Carnegie-Mellon University, Department of Computer Science, Pittsburgh, PA, USA, March 1987.

Anderson:1984:SFR

[And84a] B. Edward Anderson. A simple FORTRAN ring data structure for representing building envelope geometries. Thesis (M.A.), University of California at Los Angeles, Architecture and urban planning, Los Angeles, CA, USA, 1984. vii + 205 pp.

Andreoni:1984:FIX

[And84b] Gaetano Andreoni. Fortran interface to X.25 and transport service. Computer Networks: The International Journal of Distributed Informatique, 8(1):17–22, February 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Ando:1989:WFP

[And89] Samon Ando. WADA1: FOR-TRAN program for wave diffraction analysis: version 1. Technical memorandum 89/204, Defence Research Establishment Atlantic, Dartmouth, NS, Canada, 1989. v + 20 pp.

Anonymous:1981:MFD

[Ano81a] Anonymous. Miscellaneous: Fortran data base facility (Journal of development 1980). Software —Practice and Experience, 11(1): 100, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anonymous:1981:PF

[Ano81b] Anonymous. Programming in FORTRAN. Heath Co., Benton Harbor, MI, USA, 1981. 6 sound cassettes kit + program coding forms. pp.

Anonymous:1982:FAB

[Ano82a] Anonymous. Fortran-Lexikon: Anweisungen und Begriffe: Fortran IV — DIN 60027, TR 440, IBM 360, 370, CDC FTN 4, DEC IV-PLUS: Fortran 77–ANSI X3.9-1978, Siemens FORI, CDC FTN5. Walter de Gruyter, New York, NY, USA, 1982. ISBN 3-11-008359-0. xix + 492 pp. LCCN QA76.73.F25F677 1982.

Anonymous:1982:SSR

[Ano82b] Anonymous. Spline smoothing routines for ASCII Fortran: user manual for the 1100. Mathematical routines series 1420, University of Wisconsin — Madison, Academic Computing Center, Madison, WI, USA, 1982. 136 pp.

Anonymous:1982:UPF

[Ano82c] Anonymous. UCSD p-system FORTRAN. IBM Corporation, Boca Raton, FL, USA, version IV.0 edition, 1982. 1 program file on 6 computer disks pp.

Anonymous:1982:VFC

[Ano82d] Anonymous. VS FORTRAN compiler and library: general information: program product. IBM Corporation, San Jose, CA, USA, fourth edition, 1982. ix + 45 pp.

Anonymous:1983:NFP

[Ano83] Anonymous. Nevada Fortran; programmer’s reference manual. Ellis Computing, San Francisco, CA, USA, 1983. various pp.

Anonymous:1984:ILF

[Ano84] Anonymous. IMSL library: FOR-TRAN subroutines for mathematics and statistics: User’s manual. IMSL Problem-Solving Software Systems, Houston, TX, USA, ed. 9.2 edition, 1984. ???? pp.

Anonymous:1985:ATA

[Ano85a] Anonymous. AI tools automate software translation. Electronics, 58(38):59–61, September 23, 1985. ISSN 0883-4989.

Anonymous:1985:QIP

[Ano85b] Anonymous. Quelques instructions pour l’utilisation du FORTRAN 77 sur le VAX. Version 3.5. ????, ????, 1985. 26 pp.

Anonymous:1985:SSR

[Ano85c] Anonymous. Spline smoothing routines for ASCII Fortran: user manual for the 1100. Mathematical routines series ????, University of Wisconsin — Madison, Academic Computing Center, Madison, WI, USA, 1985. various pp.

Anonymous:1987:FRM

[Ano87a] Anonymous. FITLIB reference manual: a PC Scientific IBM PC and PS/2 Fortran Library product for curve and surface fitting employing splines under tension. McGraw-Hill Book Company, New York, NY, USA, 1987. ISBN 0-07-852424-5. xiv + 427 pp.

Anonymous:1987:IF

[Ano87b] Anonymous. IBM Fortran/2. IBM Corporation, Boca Raton, FL, USA, 1987. various pp.

Anonymous:1987:IFU

[Ano87c] Anonymous. IPMIXD for Fortran 77: user manual for the VMS VAX cluster. Technical report, University of Wisconsin — Madison, MACC Academic Computing Center, Madison, WI, USA, 1987. various pp.

Anonymous:1987:MPL

[Ano87d] Anonymous. Manuel pratique de l’informatique. Ed. Sciences et techniques, ????, 1987. ISBN 2-86840-014-0. 14 + [3] (env. 918 p.) pp.

Anonymous:1987:MCG

[Ano87e] Anonymous. Model Construction With GPSS-Fortran. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., April 1987. ISBN 3-540-96503-3. ???? pp. LCCN ???? US$39.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
3540965033.

Anonymous:1988:ICV

[Ano88a] Anonymous. International conference on vector and parallel processors in computational science III.

Parallel Computing, 8(1–3):??, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Anonymous:1988:ISQ

[Ano88b] Anonymous, editor. International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods. John Wiley and Sons, New York, London, Sydney, 1988. CODEN IJQSDI. ISSN 0161-3642.

Anonymous:1988:WGS

[Ano88c] Anonymous. Writing Good Software in Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1988. ISBN 0-7248-1281-4. ???? pp. LCCN ????

Anonymous:1989:MHP

[Ano89] Anonymous. MacFortran high performance ANSI FORTRAN 77 compiler and debugger, 1989.

X3J3:1989:FD

[ANS89] ANSI X3J3. Fortran 8X draft. ACM SIGPLAN Notices, 8(4), December 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Also published by Global Engineering.

Antonelli:1981:OSS

[Ant81] Albert Eugene Antonelli. An operating system simulation program in FORTRAN. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1981. iv + 84 pp.

Allen:1987:DPF

[AP87] Todd R. Allen and David A.

Padua. Debugging parallel Fortran on a shared memory machine. Technical Report CSRD 624, University of Illinois, Center for Supercomputing Research and Development, Urbana, IL, USA, 1987. 17 pp.

Alpert:1986:IMA

[APD86] Elizabeth Alpert, Rex L. Page, and Richard L. Didday. Instructor’s manual to accompany Fortran 77 for humans, 3rd edition. West Publishing Company, St. Paul, MN, USA, 1986. ISBN 0-314-97145-9. iii + 116 pp.

Apollo:1983:DFU

[Apo83] Apollo Computer. DOMAIN FORTRAN user’s guide. Apollo Computer Inc., Chelmsford, MA, USA, revised 04 edition, 1983. various pp.

Apollo:1986:DFL

[Apo86] Apollo Computer. DOMAIN FORTRAN language reference. Apollo Computer Inc., Chelmsford, MA, USA, revised 05 edition, 1986. various pp.

Adams:1987:FDF

[AR87] J. Adams and J. Reid. Fortran 8X — the draft Fortran standard revision. ACM SIGPLAN FORTRAN Forum, 6(2):27–38, October 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Arnold:1982:PET

[Arn82] C. N. Arnold. Performance evaluation of three automatic vectorizer packages. ICPP82, pages 235–242, 1982.

Artley:1981:DFI

[Art81] J. A. Artley. Description of the FORTRAN implementation of the spring small grains planting date distribution model. Report 17414 JSC, Lyndon B. Johnson Space Center, NASA, Houston, TX, USA, 1981. 55 pp. For sale by National Technical Information Service.

Alam:1988:CCFa

[AS88a] S. S. Alam and S. K. Sen. Computer and computing with Fortran 77. IIT, Kharagpur-Oxford and IBH series in engineering and technology. Oxford and IBH Publishing, New Delhi, India, 1988. ISBN ???? 594 pp. LCCN ????

Alam:1988:CCFb

[AS88b] S. S. Alam and S. K. Sen. Computer and Computing With Fortran 77. IIT, Kharagpur-Oxford & IBH Series in Engineering & Technology. South Asia Books, ????, December 1988. ISBN 81-204-0301-0. ???? pp. LCCN ???? US$14.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
8120403010.

Asbury:1989:FOIa

[AS89a] Raymond K. Asbury and David S. Scott. FORTRAN I/O on the iPSC/2: Is there read after write?

Technical report, Intel Corporation, Santa Clara, CA, USA, 1989. conference proceedings unknown.

Asbury:1989:FOIb

[AS89b] Raymond K. Asbury and David S.

Scott. FORTRAN I/O on the iPSC/2: Is there read after write? In Fourth Conference on Hypercube Concurrent Computers and Applications, pages 129–132. ????, ????, 1989.

Ashworth:1981:PPF

[Ash81a] R. Ashworth. Prettyprinting for power (FORTRAN). ACM SIGPLAN Notices, 16(2):16–17, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashworth:1981:PP

[Ash81b] Robert Ashworth. Prettyprinting for power (FORTRAN). ACM SIGPLAN Notices, 16(2):16–17, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashby:1985:CAF

[Ash85a] Steven F. Ashby. CHEBY-CODE: A FORTRAN implementation of Manteuffel’s adaptive Chebyshev algorithm. Technical Report UIUCDCS-R-85-1203, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, May 1985. 105 pp.

Ashby:1985:CFIa

[Ash85b] Steven F. Ashby. ChebyCode, a FORTRAN implementation of Manteuffel’s adaptive Chebyshev algorithm. Report UIUCDCS-R-85-1203, Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1985. iv + 100 pp.

Ashby:1985:CFIb

[Ash85c] Steven F. Ashby. ChebyCode: a FORTRAN, implementation of Manteuffel’s adaptive Chebyshev algorithm. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1985. iv + 100 pp.

Appelbe:1989:SPT

[ASM89] Bill Appelbe, Kevin Smith, and Charlie McDowell. START/PAT: a toolkit for developing parallel programs. IEEE Software, 6 (4):29–38, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

ATC:1982:FCS

[Ass82] Associated Technology Company. A FORTRAN coding standard. Associated Technology Co., Estill Springs, TN, USA, 1982. vi + 40 pp.

ATC:1983:FCS

[Ass83a] Associated Technology Company. A FORTRAN coding standard. Associated Technology Co., Estill Springs, TN, USA, 1983. 47 pp.

AFN:1983:NAN

[Ass83b] Association française de normalisation. Norme AFNOR NF Z 65-110: Traitement de l’information, langage de programmation Fortran. AFNOR, Paris, France, 1983. ISBN ???? 444 pp. LCCN ????

ACM:1984:FF

[Ass84] Fortran forum, 1984. ACM Press, New York, NY 10036, USA.

ICPR:1986:EIC

[Ass86] Association française de cybernetique économique et technique, editor. Eighth International Conference on Pattern Recognition, Paris, France, October 27–31, 1986: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0742-4 (paperback), 0-8186-4742-6 (microfiche), 0-8186-8742-8 (hardcover). LCCN Q 327 I615 1986. IEEE Computer Society order number 742. IEEE catalog number 86CH2342-4. On spine: 1986 International Conference on Pattern Recognition. Includes bibliographies and index.

Aho:1986:CPC

[ASU86] Alfred V. Aho, Ravi Sethi, and Jeffrey D. Ullman. Compilers: Principles, Techniques, and Tools. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10088-6 (hardcover), 0-201-10194-7 (paperback). x + 796 pp. LCCN QA76.76.C65 A371 1986. See [AU77].

Aho:1977:PCD

[AU77] Alfred V. Aho and Jeffrey D. Ullman. Principles of Compiler Design. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-00022-9 (hardcover), 0-201-10073-8. x + 604 pp. LCCN QA76.6 .A285 1977. See also the much expanded subsequent book [ASU86].

Aspray:1982:MRP

[AW82] William F. Aspray and Virginia C. Walker. Meetings in retrospect: Pioneer Day, NCC ’82: History of the Stored-Program Concept; The 25th Anniversary of FOR-TRAN. Annals of the History of Computing, 4(4):358–367, October/December 1982. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1982/pdf/a4358.pdf;
 http://www.computer.org/annals/
an1982/a4358abs.htm.

Adams:1989:SCF

[AW89] Jeanne C. Adams and Jerrold L. Wagener. Summary of changes to Fortran draft. ACM SIGPLAN FORTRAN Forum, 8(3):13–19, August 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Ayatey:1984:EFI

[Aya84] Siegfried B. Y. Ayatey. Elementary Fortran IV Microeconomics Programs. University Press of America, Lanham, MD, USA, July 1984. ISBN 0-8191-3950-5 (paperback). 232 pp. LCCN HB172 .A875 1984. US$22.00; US$11.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0819139505.

Bownds:1985:FSS

[BA85a] J. M. Bownds and L. Applebaum. A Fortran subroutine for solving Volterra integral equations. ACM Transactions on Mathematical Software, 11(1):58–65, March 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bownds:1985:AAF

[BA85b] John M. Bownds and Lee Applebaum. Algorithm 627: A FORTRAN subroutine for solving Volterra integral equations. ACM Transactions on Mathematical Software, 11(1):58–65, March 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bownds:1985:AFS

[BA85c] John M. Bownds and Lee Applebaum. Algorithm 627: A FORTRAN subroutine for solving Volterra integral equations. ACM Transactions on Mathematical Software, 11(1):58–65, March 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/214314.html.

Ben-Ari:1986:FTD

[BA86] Mordechai Ben-Ari. FOREET: a tool for design and documentation of Fortran programs. Software— Practice and Experience, 16(10): 915–924, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Backus:1981:HFI

[Bac81] J. Backus. The history of FOR-TRAN I, II, and III. In Wexelblat [Wex81], pages 25–74. ISBN 0-12-745040-8. LCCN QA76.7 .H56 1978. US$45.

Backus:1984:A

[Bac84a] John Backus. Afterword. Annals of the History of Computing, 6(1):26–27, January/ March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1015.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1015abs.htm.

Backus:1984:EDF

[Bac84b] John Backus. Early days of FOR-TRAN. Annals of the History of Computing, 6(1):15, January/ March 1984. CODEN AHCOE5. ISSN 0164-1239.

Baillie:1986:GFC

[Bai86] C. F. Baillie. A general FOR-TRAN to C translator. Computer Physics Communications, 41 (2–3):409–414, August 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465586900810.

Bailey:1987:HPF

[Bai87] David H. Bailey. A highperformance fast Fourier transform algorithm for the Cray-2. The Journal of Supercomputing, 1(1): 43–60, March 1987. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=

0920-8542&volume=1&issue=1&
spage=43.

Bailey:1989:CFP

[Bai89] M. C. Bailey. CWG, a FORTRAN program for mutual coupling in a planar array of circular waveguidefed apertures. NASA technical memorandum 101614, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, June 1989. 11 + 13 pp.

Bajpai:1981:FAP

[Baj81] Avinash Chandra Bajpai. Fortran and Algol: a programmed course for students of science and technology. A series of programmes on mathematics for scientists and technologists. John Wiley and Sons, New York, London, Sydney, 1981. ISBN ???? 273 pp. LCCN ????

Baldwin:1984:BDG

[Bal84] R. R. Baldwin. The backwarddirected GO TO in FORTRAN. ACM SIGPLAN Notices, 19(8): 62–64, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balagurusamy:1985:FBI

[Bal85] E. Balagurusamy. Fortran for beginners: including Fortran 77. Tata McGraw-Hill, New Delhi, India, 3eme reimpr. (1989) edition, 1985. ISBN 0-07-451527-6. x + 196 pp.

Bank:1978:AFI

[Ban78] Randolph E. Bank. Algorithm 527: A Fortran implementation of the generalized marching algorithm [D3]. ACM Transactions on Mathematical Software, 4(2):165– 176, June 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Banerjee:1988:IFT

[Ban88] Utpal Banerjee. An introduction to a formal theory of dependence analysis. The Journal of Supercomputing, 2(2):133– 149, October 1988. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=2&issue=2&
spage=133.

Barnard:1984:FFF

[Bar84] A. J. Barnard. FURI — a Fortran function writer. Computer Physics Communications, 35(1–3):C–429, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584826387.

Barnett:1989:UPF

[Bar89] M. P. Barnett. Using partial fraction formulas to sum some slowly convergent series analytically for molecular integral calculations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (3):13–18, July 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bates:1985:GAC

[Bat85] B. L. Bates. GENTRAN: an automatic code generation facility for REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(3):24–42, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bauer:1982:F

[Bau82] Charles R. Bauer. Fortran 77. Four Pi, Inc., Norridge, IL, USA, 1982. ISBN ???? 304 pp. LCCN ????

Baughman:1988:FFB

[Bau88] A. L. Baughman. A FORTRAN function for the bivariate normal integral. Computer Methods and Programs in Biomedicine, 27 (2):169–174, September–October 1988. CODEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Blackwood:1982:AFD

[BB82] Brian D. Blackwood and George H.

Blackwood. Apple Fortran: Detailed Language Instructions Specifically for the Apple Computer. Howard W. Sams, Indianapolis, IN 46268, USA, November 1982. ISBN 0-672-21911-5 (paperback). 236 pp. LCCN QA76.8.A66 B58 1982. US$14.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0672219115.

Bader:1983:NRT

[BB83] Morris Bader and William Allan Bader. New Ratfor tools for numeric and graphics processing. In Software Tools Users Group [Sof83b], pages 411–417. ISBN ???? LCCN QA76.8.U65 U74 1983. “Sponsored by USENIX Association in cooperation with Software Tools Users Group.”–p. iii.

Bechlars:1986:GP

[BB86] J. Bechlars and R. Buhtz. GKS in der Praxis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-16139-2.

Bohlender:1983:AMS

[BBB^{+}83] G. Bohlender, H. Boehm, K. Braune, K. Gruener, E. Kaucher, R. Klatte, W. Kraemer, U. Kulisch, W. L. Miranker, Ch. Ullrich, and J. Wolff Von Gudenberg. Application Module: Scientific-Computation for Fortran 8x. Modified Proposal for Arithmetic Specification According to Guidelines of the X3J3-Meetings in Tulsa and Chapel Hill. Interner Bericht, Instituts F. Angew. Math. Universität Karlsruhe, Karlsruhe, Germany, 1983.

Baker:1982:HSF

[BBF^{+}82] G. A. Baker, L. P. Benofy, M. Fortes, M. de Llano, S. M. Peltier, and A. Plastino. Hardcore square-well Fermion. Phys. Rev. A, 26:3575–3588, 1982. CO-DEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Bohlender:1982:PAS

[BBG^{+}82] G. Bohlender, H. Boehm, K. Gruener, E. Kaucher, R. Klatte, W. Kraemer, U. Kulisch, S. M. Rump, Ch. Ullrich, and J. Wolff Von Gudenberg. Proposal for arithmetic specification in Fortran 8x. Interner Bericht Des Instituts F. Angew. Math., Universität Karlsruhe, Karlsruhe, Germany, 1982.

Bohlender:1984:ASF

[BBG^{+}84] G. Bohlender, H. Boehm, K. Gruener, E. Kaucher, R. Klatte, W. Kraemer, U. Kulisch, W. L. Miranker, S. M. Rump, Ch. Ullrich, and J. Wolff Von Gudenberg. Arithmetic specification in Fortran 8x. In B. Ford, J. C. Rault, and F. Thomasset, editors, Tools, Methods and Languages For Scientific and Engineering Computation, pages 213–243. Elsevier Sci. Publ., North Holland, Amsterdam, The Netherlands, 1984.

Blackwood:1984:AFK

[BBuC84] Brian D. Blackwood, George H.

Blackwood, and So uk Cha. Apple FORTRAN kipon maenyuol. Kanamsa, Seoul, Korea, 1984. ISBN ???? 248 pp. LCCN ????

Burke:1988:ADP

[BCF^{+}88] Michael Burke, Ron Cytron, Jeanne Ferrante, Wilson Hsieh, Vivek Sarkar, and David Shields. Automatic discovery of parallelism: a tool and an experiment (extended abstract). ACM SIGPLAN Notices, 23(9):77–84, September 1988. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1989:CHR

[BCKT89] P. Briggs, K. D. Cooper, K. Kennedy, and L. Torczon. Coloring heuristics for register allocation. ACM SIGPLAN Notices, 24(7):275–284, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bagrodia:1987:MBA

[BCM87] R. L. Bagrodia, K. M. Chandy, and J. Misra. A message-based approach to discrete-event simulation. IEEE Transactions on Software Engineering, SE-13(6): 654–665, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702271.

Burkard:1980:AMP

[BD80] Rainer E. Burkard and Ulrich Derigs. Assignment and matching problems: solution methods with FORTRAN programs. With the assistance of T. Bonniger and G. Katzakidis, volume 184 of Lecture Notes in Economics and Mathematical Systems. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-10267-1. iii + 148 pp.

Bischof:1989:LAL

[BD89] C. H. Bischof and J. J. Dongarra. A linear algebra library for highperformance computers. In Carey [Car89a], pages 45–56. ISBN 0-471-92436-9. LCCN M89.E02452; QA76.6.

Booth:1989:EMC

[BDJ^{+}89] S. P. Booth, R. W. Dobinson, D. R. N. Jeffery, W. Lu, K. M. Storr, and A. Thornton. An evaluation of the Meiko computing surface for HEP Fortran farming. Computer Physics Communications, 57(1–3):486–491, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465589902695.

Bolmarich:1987:TEF

[BDR87] T. Bolmarich and Frederica Darema-Rogers. Tutorial for EPEX/FORTRAN program parallelization and execution. Research Report RC 12515, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1987. 8 pp.

Byrne:1984:PFR

[BDS84] G. D. Byrne, A. J. DeGregoria, and D. E. Salane. A program for fitting rate constants in gas phase chemical kinetics models. SIAM Journal on Scientific and Statistical Computing, 5(3): 642–657, September 1984. CO-DEN SIJCD4. ISSN 0196-5204.

Brewer:1988:TAAb

[BDS88a] O. Brewer, J. Dongarra, and D. Sorensen. Tools to aid in the analysis of memory access patterns for FORTRAN programs. LA-PACK Working Note 06, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June

1988. URL http://www.netlib.
org/lapack/lawns/lawn06.ps;
 http://www.netlib.org/lapack/
lawnspdf/lawn06.pdf.  ANL,

MCS-TM-120, June 1988.

Brewer:1988:TAAa

[BDS88b] Orlie Brewer, Jack Dongarra, and Danny Sorenson. Tools to aid in the analysis of memory access patterns for Fortran programs. Mathematics and Computer Science Division Report ANL-MCS-TM-119, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June 1988. LAPACK Working Note #6.

Brewer:1988:TAAc

[BDS88c] Orlie Brewer, Jack Dongarra, and Danny Sorenson. Tools to aid in the analysis of memory access patterns for FORTRAN programs. Parallel Computing, 9(1): 25–35, December 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brewer:1989:GTA

[BDS89] O. Brewer, J. Dongarra, and D. Sorensen. A graphics tool to aid in the generation of parallel FOR-TRAN programs. In Stanley Y. W. Su and George J. Knafl, editors, Proceedings: the thirteenth annual International Computer Software & Applications Conference, pages 89–93. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1964-3. LCCN QA76.6.C6295 1989. Computer Society order number 1964. IEEE catalog number 89CH2743-3.

Beebe:1981:ICF

[Bee81] Nelson H. F. Beebe. Ideas for a combined FORTRAN 77/ SFTRAN3 prettyprinter. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, March 30, 1981. 2 pp.

Beebe:1982:PPF

[Bee82] Nelson H. F. Beebe. PRETTY— A Portable FORTRAN Program Prettyprinter, second edition. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, June 10 1982. 46 pp.

Bathie:1985:FER

[BEE^{+}85a] William W. Bathie, Hotten A.

Elleby, J. C. Even, Richard E. Horton, James William Nilsson, Allen H. Pulsifer, William F. Riley, and Donald F. Young. Fundamentals of engineering review, 1985.

Beebe:1985:PSS

[Bee85b] Nelson H. F. Beebe. Proposal for a Standard Set of Primitives for Machine-Independent Character Manipulation in FORTRAN 77. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, May 23 1985. 31 pp.

Beebe:1988:FSP

[Bee88] Nelson H. F. Beebe. A Fortran 77 and SFTRAN3 Prettyprinter. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 8 1988. 18 pp.

Bell:1984:TPL

[Bel84] Frederick H. Bell. TRS-80 programming for learning and teaching: 40+ applications programs. Reston Publishing Co., Inc., Reston, VA, USA, 1984. ISBN 0-8359-7863-X (hard) : US$19.95, 0-8359-2096-8 (paperback) : US$16.95, 0-8359-7862-1 (paperback : cover) : US$16.95. xviii + 301 pp. LCCN QA76.8.T18 B44 1984.

Bellamy:1989:CPF

[Bel89] Clifford John Bellamy. Computer programming in FORTRAN 77. Longman Cheshire, Melbourne, Australia, 1989. ISBN 0-582-71185-1 (paperback). xii + 316 pp.

Bemer:1984:CPF

[Bem84] Robert W. Bemer. Computing prior to FORTRAN. Annals of the History of Computing, 6(1):16– 18, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1015.pdf;

http://www.computer.org/annals/
an1984/a1015abs.htm.

Berche:1982:LPF

[Ber82a] Stephane Berche. Langages de programmation: Fortran, LSE, Basic, Pascal, Cobol, PL/1, APL, Assembleur. Ed. du P.S.I., Paris, France, 1982. ISBN ???? 134 pp. LCCN ????

BerzosaValencia:1982:PF

[Ber82b] Jose Ramon Berzosa Valencia. Programacion FORTRAN. Technical report, Departamento de Publicaciones, E.T.S.I. de Telecommunicacion, Madrid, Spain, 1982. ISBN 84-7402-081-6. 330 pp.

Bermudez:1984:EFS

[Ber84a] Victor M. Bermudez. Ellips — a Fortran simulation of a polarization-modulation ellipsometer. Computer Physics Communications, 35(1–3):C–426, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584826351.

Bernd:1984:SGF

[Ber84b] Schmidt Bernd. The Simulator Gpss-Fortran Version 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., June 1984. ISBN 0-387-96504-

1. ???? pp. LCCN ????

Berkman:1985:GVU

[Ber85a] Jerry Berkman. A guide for VAX UNIX Fortran users. Unx; 2.3.3 unx; 2.3.3., University of California, Berkeley, Academic Computing Services, Berkeley, CA, USA, September 19, 1985. iii + 62 pp.

Berns:1985:SFR

[Ber85b] G. M. Berns. Significant Fortran reliability problems and what Fortran 8X can do about them. ACM SIGPLAN FORTRAN Forum, 4 (1):25–32, March 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Berz:1987:DAF

[Ber87] Martin Berz. The differential algebra FORTRAN precompiler DAFOR. Technical Report AT– 3 TN–87–32, Los Alamos National Laboratory, Los Alamos, NM, USA, 1987.

Berke:1988:PAS

[Ber88a] Wayne Berke. ParFOR — A structured environment for parallel FORTRAN. Ultracomputer Note 137, New York University, New York, NY, USA, April 1988.

Bernstein:1988:PSN

[Ber88b] David Bernstein. PREFACE-2: supporting nested parallelism in Fortran. Research report RC 14160 (#63397), IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1988. 21 pp.

Berkman:1989:FUV

[Ber89] Jerry Berkman. Fortran under VM/CMS. IBM 2.3.1, Computing Services, University of California, Berkeley, CA, USA, 1989. 39 pp.

Bezner:1988:F

[Bez88] Hart C. Bezner. Fortran 77. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, October 1988. ISBN 0-13-329509-5. ???? pp. LCCN ???? US$41.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0133295095.

Bezner:1989:F

[Bez89] Hart C. Bezner. Fortran 77.

Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1989. ISBN 0-13-329087-5 (paperback), 0-13-329509-5. xvi + 395 pp. LCCN QA 76.73 F25 B493 1989.

Barth:1982:FRU

[BG82] Jeffrey Barth and R. Steven Glanville. Fortran 77 reference for the UCSD p-system, 1982.

Bryant:1984:ILD

[BG84] Barrett R. Bryant and A. A. Grau.

An intermediate language to define dynamic semantics. Computer Languages, 9(3-4):149–159, ???? 1984. CODEN COLADA. ISSN 0096-0551.

Barth:1982:UPF

[BGCS82] Jeffrey Barth, R. Steven Glanville, Randy Clark, and Stan Stringfellow. UCSD p-system FORTRAN version IV.0, 1982.

Brainerd:1985:FFS

[BGG85a] Walter S. Brainerd, Charles H.

Goldberg, and Jonathan L. Gross. Fortran 77: Fundamentals and Style. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, July 1985. ISBN 0-87835-143-4 (paperback). xvi + 423 + 33 + 3 pp. LCCN QA76.73.F25 B728 1985. US$38.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0878351434.

Brainerd:1985:IMA

[BGG85b] Walter S. Brainerd, Charles H. Goldberg, and Jonathan L. Gross. Instructor’s manual to accompany Fortran 77 fundamentals and style. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, 1985. ISBN 0-87835-146-9 (paperback). 238 pp.

Brainerd:1986:WFS

[BGG86] Walter S. Brainerd, Charles H. Goldberg, and Jonathan L. Gross. Watfiv-S Fundamentals and Style. The Boyd and Fraser programming language series. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, April 1986. ISBN 0-87835-174-4 (paperback). xiii + 370 + 33 + 3 pp. LCCN QA76.73.F25B74 1986. US$35.95.

Barth:1983:FLR

[BGM83] Jeffrey Barth, R. Steven Glanville, and Henry McGilton. FORTRAN language reference manual. Silicon Valley Software, Inc., Cupertino, CA, USA, 1983. vi + 180 pp.

Brooks:1982:OCL

[BGS82] Rodney A. Brooks, Richard P. Gabriel, and Guy L. Steele, Jr.

An optimizing compiler for lexically scoped Lisp. In Proceedings of the SIGPLAN ’82 Symposium on Compiler Construction [ACM82], pages 261–275. ISBN ???? LCCN ???? Available as SIGPLAN Notices 17(6) June 1982.

Brown:1985:FFP

[BH85] James W. Brown and Stanford B. Hooker. FIXSRC: a Fortran preprocessor. Computer Physics Communications, 38(3): 435–440, December 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465585901110.

Bomans:1989:AGM

[BH89] Luc Bomans and Rolf Hempel.

The Argonne/GMD macros in FORTRAN for portable parallel programming and their implementation on the Intel iPSC/

2. Arbeitspapiere der GMD 406 0723-0508, Gesellschaft fur Mathematik und Datenverarbeitung mbH, Sankt Augustin, Germany, 1989. 18 pp.

Berry:1985:ACS

[BHK^{+}85] M. W. Berry, M. T. Heath, I. Kaneko, M. Lawo, R. J. Plemmons, and R. C. Ward. An algorithm to compute a sparse basis of the null space. Numerische Mathematik, 47(4):483–504, 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brent:1980:AIB

[BHY80] Richard P. Brent, Judith A. Hooper, and J. Michael Yohe. An AUGMENT interface for Brent’s multiple precision arithmetic package. ACM Transactions on Mathematical Software, 6(2):146–149, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Bre78b, Bre79].

Binder:1985:ADM

[Bin85] Robert Binder. Application debugging: an MVS abend handbook for COBOL, Assembly, PL/I, and FORTRAN programmers. Prentice-Hall Software Series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, February 1985. ISBN 0-13-039348-7. xi + 366 pp. LCCN QA76.6.B56 1985. US$65.00; US$30.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0130393487.

Bishop:1981:BRBa

[Bis81] J. M. Bishop. Book review: Fortran for students, Roger Hutty, Macmillan, London, 1980. Software—Practice and Experience, 11 (1):101, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burke:1981:FSMa

[BJ81a] V. M. Burke and C. Jackson. A Fortran system to maintain a program library: 1. Storage of the program decks in magnetic tape files. Computer Physics Communications, 22(1):59–75, February/ March 1981. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465581900795.

Burke:1981:FSMb

[BJ81b] V. M. Burke and C. Jackson.

A Fortran system to maintain a program library: 2. Retrieval of program decks from the library files. Computer Physics Communications, 22(1):77–84, February/ March 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465581900801.

Burke:1984:FSMa

[BJ84a] V. M. Burke and C. Jackson.

A Fortran system to maintain a program library: 1. Storage of the program decks in magnetic tape files. Computer Physics Communications, 35(1–3):C–685, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584828362.

Burke:1984:FSMb

[BJ84b] V. M. Burke and C. Jackson.

A Fortran system to maintain a program library: 2. Retrieval of program decks from the library files. Computer Physics Communications, 35(1–3):C–686, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584828374.

Barnwell:1984:HSP

[BK84] Thomas O. Barnwell and John L. Kittle. Hydrologic simulation program — Fortran development, maintenance and applications. Technical report, US Environmental Protection Agency, Athens, GA, USA, 1984. 14 pp.

Braswell:1988:EVF

[BK88] R. N. Braswell and M. S. Keech. An evaluation of vector Fortran 200 generated by Cyber 205 and ETA-10 pre-compilation tools. In IEEE [IEE88b], pages 106–113. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Balasundaram:1989:TSD

[BK89] Vasanth Balasundaram and Ken Kennedy. A technique for summarizing data access and its use in parallelism enhancing transformations. ACM SIGPLAN Notices, 24(7):41–53, July 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org:80/pubs/
                                        citations/proceedings/pldi/
                                        73141/p41-balasundaram/.

Bohlender:1981:FCN

[BKK^{+}81] G. Bohlender, E. Kaucher, R. Klatte, U. Kulisch, W. L. Miranker, Ch.

Ullrich, and J. Wolff v. Gundenberg. Fortran for contemporary numerical computation. Computing, 26(4):277–314, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Balasundaram:1989:PEI

[BKK^{+}89] V. Balasundaram, K. Kennedy, U. Kremer, K. McKinley, and J. Subhlok. The ParaScope editor: an interactive parallel programming tool. In ACM [ACM89b], pages 540–550. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Borisiuk:1989:PDB

[BKL89] R. M. Borisiuk, G. P. Kreitser, and V. V. Levitin. Programmirovanie dlia biologov: osnovy iazyka Fortran. Nauch. tsentr biologicheskikh issledovanii AN SSSR, Pushchino, USSR, 1989. ISBN ???? 92 pp. LCCN ????

Beakley:1983:CCC

[BL83] George C. Beakley and Robert E.

Lovell. Computation, Calculators, and Computers: Tools for Engineering Problem Solving, Including Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, January 1983. ISBN 0-02-307150-8. xii + 334 pp. LCCN TA345.B35 1983. US$30.07. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0023071508.

Blatt:1987:CNH

[Bla87] David W. E. Blatt. Comparison of network and hierarchical data management in large programming projects. The Journal of Systems and Software, 7(4):297–309, December 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bliss:1989:IFP

[Bli89] Brian Eugene Bliss. Instrumentation of Fortran programs for automatic roundoff error analysis and performance evaluation. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. xiii + 132 pp.

Blue:1978:PFP

[Blu78] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. ACM Transactions on Mathematical Software, 4(1):15–23, March 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Boyer:1981:CPC

[BM81] R. S. Boyer and J. S. Moore. The Correctness Problem in Computer Science. Academic Press, New York, NY, USA, 1981. ISBN 0-12-122920-3. xiii + 279 pp. LCCN QA76.6 .C68 1981.

Billica:1984:UMS

[BMS84] Judith A. Billica and Hubert J. Morel-Seytoux. User’s manual for SOILWAT: a finite difference Fortran 77 program for prediction of infiltration rates and water content profiles under varied initial and boundary conditions. Technical report, Colorado State University, Colorado Water Resources Research Institute, Hydrowar Program, Fort Collins, CO, USA, 1984. iv + 178 pp.

Ball:1987:WPC

[BNZ87] J. W. Ball, Darrell Kirk Nordstrom, and Dieter W. Zachmann. WATEQ4F: a personal computer FORTRAN translation of the geochemical model WATEQ2 with revised data base. Open-file report 87-50, U.S. Geological Survey, Menlo Park, CA, USA, 1987. iii + 108 pp.

Bodine:1987:SCC

[Bod87] Marc W. Bodine. Source codes of CLAYFORM and associated file of data constants, a FORTRAN 77 computer program for calculating structural formulae of clay minerals and other silicates. Openfile report/ Department of the Interior, U.S. Geological Survey 87-0004-A-B, U.S. Geological Survey, Denver, CO, USA, 1987. ii + 53 pp.

Boehm:1987:CRI

[Boe87] Hans-Juergen Boehm. Constructive real interpretation of numerical programs. In Proceedings SIGPLAN ’87 Symposium on Interpreters and Interpretive Techniques [ACM87], pages 214–221. ISBN ???? LCCN ???? Also available as SIGPLAN Notices 22(7) July 1987.

Boillot:1981:UF

[Boi81] Michel H. Boillot. Understanding FORTRAN. West Publishing Company, St. Paul, MN, USA, second edition, 1981. ISBN 0-8299-0355-0. xii + 505 pp.

Boillot:1984:UFS

[Boi84] Michel H. Boillot. Understanding Fortran 77 With Structured Problem Solving. West Publishing Company, St. Paul, MN, USA, 1984. ISBN 0-314-77845-4 (paperback). xvi + 592 pp. LCCN QA 76.73 F25 B643 1984. US$17.95.

Boillot:1985:UF

[Boi85] Michel H. Boillot. Understanding Fortran. West Information Publishing Group, ????, third edition, January 1985. ISBN 0-314-85219-0 (paperback). xvii + 592 pp. LCCN QA76.73.F25B64 1985. US$50.75.

Boillot:1987:UFS

[Boi87] Michel H. Boillot. Understanding Fortran 77 With Structured Problem Solving. West Publishing Company, St. Paul, MN, USA, second edition, February 1987. ISBN 0-314-27031-0 (paperback). xx + 525 + 63 pp. LCCN QA76.73.F25 B643 1987. US$43.75.

Bolmarcich:1989:IME

[Bol89] A. S. Bolmarcich. An introduction to Mach/EPEX. Research report RC 14369 (#64355), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, 1989. 13 pp.

Booch:1981:DSD

[Boo81] Grady Booch. Describing software design in Ada. ACM SIGPLAN Notices, 16(9):42–47, September 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Booch:1982:NSC

[Boo82] Grady Booch. Naming subprograms with clarity. ACM SIG-PLAN Notices, 17(1):18–22, January 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borse:1985:FNM

[Bor85a] Garold J. (Garold J.) Borse. Fortran 77 and Numerical Methods for Engineers. PWS Engineering, Boston, MA, USA, second edition, 1985. ISBN 0-534-04638-X. xiv + 656 pp. LCCN TA345 .B67 1985. US$17.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0-534-04638-X.

Borse:1985:FE

[Bor85b] Garold J. (Garold J.) Borse. Fortran 77 for Engineers. PWS Engineering, Boston, MA, USA, March 1985. ISBN 0-534-04650-9 (paperback). xvii + 420 pp. LCCN TA345 .B68 1985. US$43.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534046509.

Borse:1989:PFC

[Bor89] Garold J. Borse. Programacion en FORTRAN 77: con aplicaciones de calculo numerico en ciencias e ingenieria. Informatica profesional y universitaria. Anaya Multimedia, Madrid, Spain, 1989. ISBN 84-7614-185-8. xviii + 668 pp.

Bose:1988:HRP

[Bos88] Pradip Bose. Heuristic, rule-based program transformations for enhanced vectorization. PROC of the 1988 ICPP, II, Software:63–66, August 1988. IBM TJW.

Bowyer:1981:CDT

[Bow81] A. Bowyer. Computing Dirichlet tessellations. The Computer Journal, 24(2):162–166, May 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Bownds:1982:TIE

[Bow82] John M. Bownds. Comments on the performance of a Fortran subroutine for certain Volterra equations. In Treatment of integral equations by numerical methods (Durham, 1982), pages 163–167. Academic Press, New York, NY, USA, 1982.

Boyce:1984:WFC

[Boy84a] E. G. Boyce. WILDMAP, a FOR-TRAN computer program for the plotting of digital boundaries and mineral locations for the assessment of a given area’s resources. Open-file report 84-651, U.S. Geological Survey, Reston, VA, USA, 1984. ii + 25 pp.

Boyle:1984:LFP

[Boy84b] J. M. Boyle. LISP to FOR-TRAN — program transformation applied. In P. Pepper, editor, Program Transformation and Programming Environments. Proceedings of the NATO Advanced Research Workshop, volume 8 of NATO Advanced Studies Institutes Series F: Computer and Systems Sciences, pages 291–298. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-12932-4.

Boyle:1985:FP

[Boy85a] Thomas A. Boyle. FORTRAN 77 P.D.Q. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1985. ISBN 0-534-04938-9 (paperback). xii + 164 pp. LCCN QA76.73.F25B69 1985.

Boyle:1985:FPD

[Boy85b] Thomas A. Boyle. Fortran 77 P.D.Q. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1985. ISBN 0-534-04938-9. xii + 164 pp. LCCN QA76.73.F25B69 1985. US$32.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534049389.

Boyle:1989:FP

[Boy89a] Thomas A. Boyle. FORTRAN 77 PDQ. Brooks/Cole brief programming guides. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, second edition, 1989. ISBN 0-534-09936-X. ix + 140 pp. LCCN QA76.73.F25 B69 1989.

Boyle:1989:FPD

[Boy89b] Thomas A. Boyle. Fortran 77 P.D.Q. Brooks/Cole Brief Programming Guides. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, second edition, January 1989. ISBN 0-534-09936-X. ix + 140 pp. LCCN QA76.73.F25B69 1989. US$15.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
053409936X.

Butler:1981:FDI

[BP81a] G. F. Butler and J. Pike. Freeformat data input scheme written in Standard FORTRAN (ANSI 66). The Computer Journal, 24(3): 256–257, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Butler:1981:FFD

[BP81b] G. F. Butler and J. Pike. A freeformat data input scheme written in standard Fortran (ANSI 66). The Computer Journal, 24 (3):256–257, August 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
24/3/256.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/
256.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_03/tiff/257.
tif.

Beebe:1989:PCP

[BR89a] Nelson H. F. Beebe and R. P. C. Rodgers. (PLOT79): a comprehensive portable Fortran scientific line graphics system, as applied to biomedical research. Computers in Biology and Medicine, 19(6):385– 402, 1989. CODEN CBMDAW. ISSN 0010-4825.

Brainerd:1989:FJ

[BR89b] Fortran Journal, 1989. ISSN 1060-

0221. URL http://www.fortran.
com/fortran/fug_fj.html. ????,

????

Brent:1978:FMP

[Bre78a] Richard P. Brent. A Fortran multiple-precision arithmetic package. ACM Transactions on Mathematical Software, 4(1):57– 70, March 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brent:1978:AMF

[Bre78b] Richard P. Brent. Algorithm 524: MP, A Fortran multipleprecision arithmetic package [A1]. ACM Transactions on Mathematical Software, 4(1):71–81, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Bre79, BHY80].

Brent:1979:RMF

[Bre79] R. P. Brent. Remark on “Algorithm 524: MP, A Fortran multiple-precision arithmetic package [A1]”. ACM Transactions on Mathematical Software, 5(4): 518–519, December 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Bre78b, BHY80].

Brent:1981:MUG

[Bre81] Richard P. Brent. MP user’s guide. Technical Report TR-CS-81-08, Department of Computer Science, Australian National University, Canberra, ACT, Australia, June 1981. 73 pp.

Bright:1984:EFU

[Bri84] Herbert S. Bright. Early FOR-TRAN user experience. Annals of the History of Computing, 6(1):28– 30, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1028.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1028abs.htm.

Brich:1985:FE

[Bri85] Zinaida Sergeevna Brich. Fortran ES EVM. Finansy i statistika, Moskva, USSR, 2-e izd., perer. i dop. edition, 1985. ISBN ???? 286 pp. LCCN ????

Bleher:1987:FSS

[BRK^{+}87a] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. FORTRAN-SC. a study of a Fortran extension for engineering scientific computation with access to ACRITH. Computing, 39(2):93–110, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bleher:1987:SFE

[BRK^{+}87b] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. A study of a FORTRAN extension for engineering/scientific computation with access to ACRITH. Computing, 39: 93–110, 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bleher:1988:FSS

[BRK^{+}88] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. FORTRAN-SC: A study of a FORTRAN extension for engineering/scientific computation with access to ACRITH. In Kulisch and Stetter [KS88], pages 227–244. CO-DEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Brode:1981:PFP

[Bro81a] B. Brode. Precompilation of Fortran programs to facilitate array processing. Computer, 14(9):46– 51, September 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1981:SWK

[Bro81b] Willard A. Brown. SIMULA for those who know FORTRAN, PL/I or BASIC. Norwegian computing center/simula information publication, Norwegian Computing Center, Oslo, Norway, 1981. 47 pp.

Brode:1982:RFP

[Bro82a] Brian Quincy Brode. Restructuring Fortran programs for VAST [the vector and array syntax translator]. Technical report, Pacific-Sierra Research Corp., Santa Monica, CA, USA, 1982. various pp.

Brown:1982:CIF

[Bro82b] Channing Brown. Code improvement in a Fortran 77 compiler. Master of science, plan ii, Dept. of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1982. various pp.

Browne:1982:FP

[Bro82c] Lewis Wallace Bertram Browne. A FORTRAN primer. Prentice-Hall of Australia, Sydney, NSW, Australia, 1982. ISBN 0-7248-0454-4. xi + 92 pp.

Brown:1983:FPD

[Bro83a] Gary DeWard Brown. FORTRAN to PL/I dictionary, PL/I to FOR-TRAN dictionary. Krieger Pub. Co., Malabar, FL, USA, 1983. ISBN 0-89874-587-X. xi + 204 pp. LCCN QA76.73.F25 B76 1983.

Browne:1983:FHB

[Bro83b] L. W. B. Browne. FORTRAN for human beings. Interface, London, UK, 1983. ISBN 0-907563-52-X (paperback). xi + 92 pp. LCCN QA76.73.C15B774 1990.

Brooks:1984:MKC

[Bro84a] E. D. Brooks, III. A multitasking kernel for the C and FORTRAN programming languages. Technical Report UCID-20167, Lawrence Livermore National Laboratory, Livermore, CA, USA, September 1984.

Brown:1984:FPD

[Bro84b] Gary D. Brown. Fortran to Pl-I Dictionary: Pl-I to Fortran Dictionary. Robert E. Krieger Publishing Company, Huntington, NY, USA, June 1984. ISBN 0-89874-587-X. ???? pp. LCCN ???? US$19.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
089874587X.

Brown:1985:FPC

[Bro85] M. W. Brown. A Fortran program to calculate several diversity indices. ARS / United States Department of Agriculture, Agricultural Research Service 27, U.S. Dept. of Agriculture, Agricultural Research Service; [Available from] National Technical Information Service, Washington, DC, USA, 1985. 21 pp.

Brophy:1989:LLTa

[Bro89a] Carolyn Elizabeth Brophy. Lessons learned in the transition from Ada to FORTRAN at NASA/ goddard. Software Engineering Laboratory series; SEL-89-005 NASA-TM 103311, National Aeronautics and Space Administration, Goddard Space Flight Center, Greenbelt, MD, USA, 1989. ???? pp. For sale by the National Technical Information Service.

Brophy:1989:LLTb

[Bro89b] Carolyn Elizabeth Brophy. Lessons learned in the transition to Ada from FORTRAN at NASA/ goddard. Thesis (M.S.), University of Maryland, College Park, MD, USA, 1989. xiii + 131 pp.

Brophy:1989:LLTc

[Bro89c] Carolyn Elizabeth Brophy. Lessons learned in the transition to Ada from FORTRAN at NASA/ goddard. Computer science technical report series CS-TR-2305, University of Maryland, College Park, MD, USA, August 1989. vii + 90 pp. Supported in part by NASA. UMIACS-TR-89-84.

Brophy:1989:LLTd

[Bro89d] Carolyn Elizabeth Brophy. Lessons learned in the transition to ADA from FORTRAN at NASA/ goddard. NASA CR-186458, Dept. of Computer Science, University of Maryland, College Park, MD, USA, 1989. ???? pp.

Brusencov:1982:BF

[Bru82] Nikolaj Petrovic Brusencov. Bazisnyj Fortran. Izd. Moskovskogo universiteta, Moskva, USSR, 1982. ISBN ???? 199 pp. LCCN ????

Bruey:1984:BF

[Bru84] Alfred J. Bruey. From Basic to Fortran. Tab Books, Blue Ridge Summit, PA, USA, October 1984. ISBN 0-8306-0753-6 (hardcover), 0-8306-1753-1 (paperback). vi + 138 pp. LCCN QA76.73.F25B78 1984. US$17.95; US$13.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0830607536;   http://www.cbooks.

com/sqlnut/SP/search/gtsumt?
source=&isbn=0830617531.

Brut:1986:FVF

[Bru86] F. Brut. A FORTRAN 77 version of “A function subprogram in order to calculate the matrix elements of rotation operators”. Computer Physics Communications, 39(2):297–299, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465586901396.

Barrodale:1981:AAF

[BS81a] I. Barrodale and G. F. Stuart. Algorithm 576: A FOR-TRAN program for solving Ax = b. ACM Transactions on Mathematical Software, 7(3):391–397, September 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barrodale:1981:AFP

[BS81b] I. Barrodale and G. F. Stuart.

Algorithm 576: A FORTRAN program for solving Ax = b [F4]. ACM Transactions on Mathematical Software, 7(3):391–397, September 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barrodale:1981:FPS

[BS81c] I. Barrodale and G. F. Stuart. A Fortran program for solving Ax = b. ACM Transactions on Mathematical Software, 7(3):391–397, September 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Basso:1981:PFW

[BS81d] David T. Basso and Ronald D. Schwartz. Programming with FORTRAN/WATFOR/WATFIV. Winthrop computer systems series. Winthrop Publishers, Cambridge, MA, USA, 1981. ISBN 0-316-08315-1, 0-87626-638-3. vii + 407 pp. LCCN QA76.73.F25 B373. US$14.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0316083151.

Bent:1981:FPS

[BS81e] Robert J. Bent and George C. Sethares. Fortran With Problem Solving: a Structured Approach. Brooks/Cole series in computer science. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, August 1981. ISBN 0-8185-0436-6. x + 374 pp. LCCN QA76.73.F25B46. US$29.70. URL http://www.

                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=0818504366.

Bent:1981:IMF

[BS81f] Robert J. Bent and George C. Sethares. Instructor’s manual for FORTRAN with problem solving: a structured approach. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, 1981. ISBN ???? 108 pp. LCCN ????

Bielecki:1981:TCP

[BS81g] Jan Bielecki and Marek A. Suchenek. A textual criticism of Polish Standard of programming language FORTRAN. Research Reports 22/81, Institute of Computer Science, Warsaw Technical University, Nowowiejska 15/19, 00-665 Warszawa, Poland, 1981. In Polish.

Brooks:1983:SMF

[BS83] Barbara Brooks and George Sutherland. Software maintenance on Fortran libraries at a large computer installation. Technical Report UCRL 89913. CONF-831247– 1, National Technical Information Service, Springfield, VA, USA, 1983. 6 pp.

Berghel:1984:MPS

[BS84] H. L. Berghel and D. L. Sallach.

Measurements of program similarity in identical task environments. ACM SIGPLAN Notices, 19(8): 65–76, August 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Behforooz:1986:FS

[BS86a] Ali Behforooz and Onkar P.

Sharma. Fortran 77 Syntax. Reston Publishing Co., Inc., Reston, VA, USA, January 1986. ISBN 0-8359-3273-7 (paperback). vii + 136 pp. LCCN QA 76.73 F25 B44 1986. US$14.00; US$8.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835932737.

Bohning:1986:FSC

[BS86b] D. Böhning and F. P. Schelp.

A Fortran subroutine for computing indicators of the nutritional status of children and adolescents. Statistical Papers, 27(1):

141–150, December 1986. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http:

//link.springer.com/article/
10.1007/BF02932563.

Boyd:1986:FUP

[BS86c] K. M. H. Boyd and N. S. Scott. FTIDY — a utility program for FORTRAN 77 programs. Computer Physics Communications, 39 (3):421–430, April 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465586900986.

Barnard:1988:EFEb

[BS88a] David T. Barnard and David B. Skillcorn. Effective Fortran 77 for Engineers and Scientists. Wm. C. Brown Publishers, Dubuque, IA, USA, January 1988. ISBN 0-697-06754-8. ???? pp. LCCN ???? US$39.15. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0697067548.

Barnard:1988:EFEa

[BS88b] David T. Barnard and David B. Skillicorn. Effective Fortran 77 for engineers and scientists. Wm. C. Brown Publishers, Dubuque, IA, USA, 1988. ISBN 0-205-11098-3. xxi + 501 pp. LCCN QA76.73.F25 B366 1988. US$29.95.

Bielecki:1988:FAP

[BS88c] Jan Bielecki and Marek A. Suchenek. FORTRAN for advanced programmers. Polish Scientific Publishers, Warsaw, Poland, third edition, 1988. ISBN ???? LCCN ???? Textbook in Polish.

Bermejo:1987:SFC

[BSdlT87] F. J. Bermejo, J. Santoro, and L. Sainz de los Terreros. SQSIMUL: a FORTRAN code for the computation of squeezing properties and photon statistics in multiphoton processes. Computer Physics Communications, 43 (2):245–256, January 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465587902098.

Basili:1983:MAD

[BSP83] V. R. Basili, R. W. Selby, and T. Phillips. Metric analysis and data validation across Fortran projects. IEEE Transactions on Software Engineering, SE-9(6):652–663, November/ December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703112.

Boyce:1983:WFC

[BT83] E. G. Boyce and Robert W.

Turner. WELLMAP, a FOR-TRAN computer program for the generation of maps of specified geographic areas with a variety of well symbols available to depict the current status of given oil and gas drilling. Open-file report 83-155, U.S. Geological Survey, Reston, VA, USA, 1983. ii + 24 pp.

Bertsekas:1988:RCL

[BT88] Dimitri P. Bertsekas and P. Tseng. The relax codes for linear minimum cost network flow problems. In B. Simeone et al., editors, FOR-TRAN Codes for Network Optimization, Annals of Operations Research, vol. 13, pages 125–190. ????, ????, 1988.

Buckley:1981:AQA

[Buc81a] A. Buckley. Algorithm 580: QRUP: a set of FORTRAN routines for updating QR factorizations. ACM Transactions on Mathematical Software, 7(4):548–549, December 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Buckley:1981:AQS

[Buc81b] A. Buckley. Algorithm 580: QRUP: a set of FORTRAN routines for updating QR factorizations [F5]. ACM Transactions on Mathematical Software, 7(4): 548–549, December 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Buc82].

Buckley:1981:QAS

[Buc81c] A. Buckley. QRUP: a set of Fortran routines for updating QR factorizations. ACM Transactions on Mathematical Software, 7(4): 548–549, December 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Buckley:1982:RQS

[Buc82] A. Buckley. Remark on “Algorithm 580: QRUP: a set of FOR-TRAN routines for updating QR factorizations [F5]”. ACM Transactions on Mathematical Software, 8(4):405, December 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Buc81b].

BCS:1984:PGF

[Buc84] Bucknell Computer Services. Programmer’s guide for FORTRAN. Bucknell Computer Services, Lewisburg, PA, USA, 1984. various pp.

Buneman:1986:CFF

[Bun86] Oscar Buneman. Conversion of FFTs to fast Hartley transforms. SIAM Journal on Scientific and Statistical Computing, 7(2):624– 638, April 1986. CODEN SIJCD4. ISSN 0196-5204.

Burkard:1981:AMP

[Bur81a] Rainer E. Burkard. Assignment and Matching Problems: Solution Methods With Fortran-Programs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., April 1981. ISBN 0-387-10267-1. ???? pp. LCCN ???? US$18.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0387102671.

Burroughs:1981:BBS

[Bur81b] Burroughs Corporation. B 6000/ B 7000 systems: FORTRAN 77 reference manual. Burroughs Corporation, Detroit, MI, USA, 1981. various pp.

Burroughs:1984:BSF

[Bur84a] Burroughs Corporation. B 1000 systems, FORTRAN 77: language manual. Burroughs Corp., Detroit, MI, USA, 1984. xviii + [265] pp.

Burroughs:1984:BBB

[Bur84b] Burroughs Corporation. B5000/ B6000/B7000 systems FORTRAN 77 reference manual. Burroughs Corporation, Detroit, MI, USA, 1984. various pp.

Burton:1984:RPKa

[Bur84c] C. E. Burton. RSA: a public key cryptography system, part I. Dr. Dobb’s Journal of Software Tools, 9(3):16–??, March 1984. CODEN DDJSDM. ISSN 1044-789X.

Burns:1985:FRG

[Bur85a] Evelyn Burns. FORTRAN 77 reference guide. Prime Computer, Natick, MA, USA, 4th, rev. 19.4 edition, 1985. ???? pp.

Burroughs:1985:AFT

[Bur85b] Burroughs Corporation. A-series FORTRAN 77 test and debug system (TADS): (relative to release level 3.6): user’s guide. Burroughs, Detroit, MI, USA, 1985. various pp.

Burroughs:1985:SFRb

[Bur85c] Burroughs Corporation. A series: FORTRAN 77 (relative to release level 3.5): reference manual. Burroughs, Detroit, MI, USA, 1985. various pp.

Burroughs:1985:SFRa

[Bur85d] Burroughs Corporation. A series FORTRAN (relative to release level 3.5): reference manual. Burroughs, Detroit, MI, USA, 1985. various pp.

Burke:1986:TDW

[Bur86a] Michael Burke. A test for determining whether a Fortran 8x IDENTIFY statement is many-to-one. Research report RC 12150, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1986. 2 pp.

Burky:1986:DIP

[Bur86b] John J. Burky. The design and implementation of a portable Fortran debugger. Kent State University honors papers: Department of Mathematical Sciences, Kent State University, Kent, OH, USA, 1986. iv + 132 pp.

Burroughs:1986:SFR

[Bur86c] Burroughs Corporation. A series:

FORTRAN 77 (relative to release level 3.6): reference manual. Burroughs, Detroit, MI, USA, 1986. various pp.

Burleigh:1987:TFC

[Bur87] D. W. Burleigh. Three Fortran 77 compilers. BYTE Magazine, 12: 187–194, 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Boillot:1984:IGA

[BW84] Michel H. Boillot and David C.

Whitney. Instructor’s guide to accompany Understanding FOR-TRAN 77 with structured problem solving. West Publishing Company, St. Paul, MN, USA, 1984. ISBN 0-314-77846-2. vii + [various] pp.

Ballmann:1987:FCP

[BW87a] J. S. Ballmann and J. F. Wilkerson. FORTRAN compiler performance benchmarks for the IBM Personal Computer. Technical report, Los Alamos National Laboratory; available from National Technical Information Service, Los Alamos, NM, USA, 1987. 31 pp.

Boillot:1987:IGA

[BW87b] Michel H. Boillot and David C. Whitney. Instructor’s guide to accompany Understanding FOR-TRAN 77 with structured problem solving. West Publishing Company, St. Paul, MN, USA, second edition, 1987. ISBN 0-314-34720-8. vii + 394 pp.

Buell:1989:MIA

[BW89] Duncan A. Buell and Robert L. Ward. A multiprecise integer arithmetic package. The Journal of Supercomputing, 3(2):89– 107, July 1989. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

http://www.springerlink.com/
openurl.asp?genre=article&issn=
0920-8542&volume=3&issue=2&
spage=89.

Bakhvalov:1985:CAF

[BZ85] N. S. Bakhvalov and Ya. M. Zhileĭkin. Chislennyĭanaliz na FORTRANe. (Russian) [Numerical analysis in FORTRAN] Metody i programmy resheniya zadach fiziki volnovykh protsessov. [Methods and programs for solving problems in the physics of wave processes]. Technical Report 76-08, Moskov. Gos. Univ., Moscow, USSR, 1985. 96 pp.

Cooke:1986:IFD

[CA86] D. J. Cooke and F. Abdollahzadeh.

Insecurities in FORTRAN DOloops. Software—Practice and Experience, 16(3):201–216 (or 201– 215??), March 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cadete:1984:LF

[Cad84] Maria Odete Rodrigues Cadete. A linguagem FORTRAN 77. Number 8 in Estudos matematica e informatica. Fundacão Calouste Gulbenkian, Instituto Gulbenkian de Ciencia Centro de Calculo Cientifico, Oeiras, Portugal, 1984. ISBN ???? 197 + [1] pp. LCCN ????

Calderbank:1983:CPF

[Cal83] Valerie Joyce Calderbank. A Course on Programming in Fortran, volume 65 of Science paperbacks; 65. Chapman and Hall, Ltd., London, UK, second (revised to incorporate FORTRAN 77) edition, July 1983. ISBN 0-412-24270-2 (hardcover), 0-412-23790-3 (paperback). xi + 183 pp. LCCN QA76.73.F25 C35 1983. US$17.95 (hardcover), US$8.95 (paperback); US$18.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=

0412237903. Rev. ed. of: A course on programming in FORTRAN IV. 1969. Includes index.

Calderbank:1985:VFS

[Cal85] Valerie Joyce Calderbank. VAX FORTRAN: student guide. VAX/VMS education series. Digital Equipment Corporation, Educational Services, Bedford, MA, USA, 1985. ISBN ???? various pp. LCCN ????

Calahan:1986:BOL

[Cal86] D. A. Calahan. Block-oriented, local-memory-based linear equation solution on the Cray-2: Uniprocessor algorithms. Proceedings of the International Conference on Parallel Processing, pages 375–378, 1986. CODEN PCPADL. ISBN 0-8186-0724-6. ISSN 0190-3918. IEEE Service Cent. Piscataway, NJ, USA.

Calderbank:1989:PFCb

[Cal89a] V. J. Calderbank. Programming in Fortran. Chapman and Hall Computing Series. Chapman and Hall, Ltd., London, UK, third edition, August 1, 1989. ISBN 0-412-30510-0. 235 pp. LCCN ???? US$28.50. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0412305100.

Calderbank:1989:PF

[Cal89b] Valerie Joyce Calderbank. Programming in Fortran. Chapman and Hall computing series. Chapman and Hall, Ltd., London, UK, third edition, July 1989. ISBN 0-412-30510-0 (paperback), 0-412-30500-3. x + 235 pp. LCCN QA76.73.F25 C35 1989. US$55.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0412305003. Rev. ed. of: A course on programming in FORTRAN. 2nd ed., rev. to incorporate FOR-TRAN 77. c1983.

Calderbank:1989:PFCa

[Cal89c] Valerie Joyce Calderbank. Programming in Fortran. Chapman and Hall Computing Series. Chapman and Hall, Ltd., London, UK, July 1989. ISBN 0-412-30500-3. ???? pp. LCCN ???? US$55.00.

Cannon:1981:TLS

[Can81] Francis Robert Cannon. The transferability of learning of syntax between COBOL and FORTRAN. Thesis (Ed.D.), Temple University, Philadelphia, PA, USA, 1981. vii + 134 pp.

Carlson:1987:TEI

[Car87] B. C. Carlson. A table of elliptic integrals of the second kind. Mathematics of Computation, 49 (180):595–606, S13–S17, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1988:TEI

[Car88a] B. C. Carlson. A table of elliptic integrals of the third kind. Mathematics of Computation, 51(183): 267–280, S1–S5, July 1988. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carnevali:1988:TRS

[Car88b] P. Carnevali. Timing results of some internal sorting algorithms on the IBM 3090. Parallel Computing, 6(1):115–117, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Carey:1989:PSM

[Car89a] Graham F. Carey, editor. Parallel supercomputing: methods, algorithms and applications. Wiley series in parallel computing. John Wiley and Sons, New York, London, Sydney, 1989. ISBN 0-471-92436-9. x + 287 pp. LCCN M89.E02452; QA76.6.

Carlson:1989:TEI

[Car89b] B. C. Carlson. A table of elliptic integrals: Cubic cases. Mathematics of Computation, 53(187):327– 333, July 1989. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cassel:1981:ICI

[Cas81] Don Cassel. Introduction to Computers and Information Processing: With Basic, Cobol, Fortran, Pascal. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1981. ISBN 0-8359-3150-1. ???? pp. LCCN ???? US$28.67. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0835931501.

Cash:1989:ABF

[Cas89a] J. R. Cash. Algorithm 669: BRKF45: A FORTRAN subroutine for solving first-order systems of nonstiff initial value problems for ordinary differential equations. ACM Transactions on Mathematical Software, 15(1):29–30, March 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/214388.html.  See also

[Hig91].

Cash:1989:BAF

[Cas89b] J. R. Cash. BRK45: A Fortran subroutine for solving first-order systems of nonstiff initial value problems for ordinary differential equations. ACM Transactions on Mathematical Software, 15(1):29– 30, March 1989. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Casimir:1989:FGP

[Cas89c] Rommert J. Casimir. Fourth generation problems. ACM SIGPLAN Notices, 24(5):83–86, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cosgrove:1982:LUM

[CB82] J. H. Cosgrove and E. T. Bayliss.

LFP user’s manual (Lincoln FORTRAN preprocessor): version 02.01 for IBM/Amdahl systems. Technical report 623, Massachusetts Institute of Technology, Lincoln Laboratory, Lexington, MA, USA, 1982. ix + 71 pp.

Cheng:1986:CMF

[CB86] Yi-Wen Cheng and Jerry J. Broz.

Cycle-counting methods for fatigue analysis with random load histories: a FORTRAN user’s guide. NBSIR 86-3055, U.S. Dept. of Commerce, National Bureau of Standards; Order from National Technical Information Service, Boulder, CO, USA, 1986. v + 45 pp.

Cosgrove:1981:LUM

[CBS81] J. H. Cosgrove, E. T. Bayliss, and J. M. Sivak. LFP user’s manual (Lincoln FORTRAN preprocessor): version 02.01 for MOD-COMP systems. Technical report 570, Massachusetts Institute of Technology, Lincoln Laboratory, Lexington, MA, USA, 1981. viii + 61 pp. Also Defense Advanced Research Projects Agency, U.S.D.O.A. ESD/TR-81/150.

Corliss:1982:SOD

[CC82] George F. Corliss and Y. F. Chang. Solving ordinary differential equations using Taylor series. ACM Transactions on Mathematical Software, 8(2):114–144, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chivers:1984:IFH

[CC84a] Ian David Chivers and Malcolm W. Clark. Interactive Fortran 77: a hands-on approach. Ellis Horwood series in computers and their applications; 28 Computers and their applications; 28. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, 1984. ISBN 0-85312-775-1, 0-470-20101-0 (U.S.). 231 pp.

Chen:1984:AFC

[cC84b] Sung chu Chen. APPLE FOR-TRAN chieh hsi. Ti 3 po tien noa tsung shu; 119. Ti san po wen hua shih yeh ku fen yu hsien kung ssu, Tai-pei shih, chu pan edition, 1984. ISBN ???? 259 pp. LCCN ????

Chen:1987:ILP

[CC87] Zen Chen and Chih-Chi Chang.

Iteration-level parallel execution of DO loops with a reduced set of dependence relations. Journal of Parallel and Distributed Computing, 4(5):488–504, October 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ciampi:1989:GFP

[CCHT89] A. Ciampi, J. A. Chapman, S. A.

Hogg, and J. Thiffault. GENCOV: a Fortran program that generates randomly censored survival data with covariates. Computer Methods and Programs in Biomedicine, 29(1):51–57, May 1989. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Callahan:1986:ICP

[CCKT86] David Callahan, Keith D. Cooper, Ken Kennedy, and Linda Torczon. Interprocedural constant propagation. In SIGPLAN ’86 Symposium on Compiler Construction, pages 152–161. ACM Press, New York, NY 10036, USA, June 1986.

URL http://www.acm.org:80/
pubs/citations/proceedings/
plan/12276/p152-callahan/.

Castaneda:1979:PTP

[CCN^{+}79] Fernando Castaneda, Frederick Chow, Peter Nye, Dan Sleator, and Gio Wiederhold. PCFORT — a Fortran-to-Pcode translator. Technical Report STAN-CS-79-714, Computer System Laboratories, Stanford University, Stanford, CA, USA, January 1979.

Castle:1984:EF

[CD84] P. T. (Peter T.) Castle and G. Doherty. Elements of Fortran. Dept. of Mathematics, Wollongong, NSW, Australia, 1984. ISBN 0-86418-005-5. iv + 124 pp.

Coleman:1986:CSA

[CDHP86] M. Coleman, A. Douglas, C. Hermon, and J. Peto. Cohort study analysis with a FORTRAN computer program. International Journal of Epidemiology, 15(1): 134–137, March 1986. CODEN IJEPBF. ISSN 0300-5771.

Callahan:1988:VCT

[CDL88] David Callahan, Jack Dongarra, and David Levine. Vectorizing compilers: A test suite and results. In IEEE [IEE88b], pages 98–105. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Chang:1982:SGF

[CDW82] B. D. Chang, J. P. Draayer, and S. S. M. Wong. A system to generate Fortran programs for calculating configuration traces of angular momentum coupled product operators. Computer Physics Communications, 28(1):41–60, November 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465582900613.

Cowan:1983:WFUa

[CDW83a] Donald D. Cowan, Paul Dirksen, and J. W. (James William) Welch. WATCOM FORTRAN user’s guide for the IBM Personal Computer with IBM DOS. WAT-COM / WATFAC Series in computer science and computer applications. WATCOM Publications, Waterloo, Ontario, Canada, 1983. ISBN 0-919884-34-2. vi + 52 pp.

Cowan:1983:WFUb

[CDW83b] Donald D. Cowan, Paul Dirksen, and James William Welch. WAT-COM FORTRAN users’ guide for the IBM Personal Computer with Waterloo microNet. WATCOM Publications, Waterloo, Ontario, Canada, 1983. ISBN 0-919884-46-

6. vi + 42 pp.

Chang:1984:SGF

[CDW84] B. D. Chang, J. P. Draayer, and S. S. M. Wong. A system to generate Fortran programs for calculating configuration traces of angular momentum coupled product operators. Computer Physics Communications, 35(1–3):C–833–C–834, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0010465584829458.

Casagrande:1985:FFS

[CF85] J. T. Casagrande and N. R. Froelich. Fine-tuning a Fortran subroutine [letter]. M.D. computing: computers in medical practice, 2(1):8–??, January–February 1985. CODEN MDCOE7. ISSN 0724-6811.

Coleman:1984:FSE

[CGM84a] T. F. Coleman, B. S. Garbow, and J. J. Moré. Fortran subroutines for estimating sparse Jacobian matrices. ACM Transactions on Mathematical Software, 10(3): 346–347, September 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1984:AFS

[CGM84b] Thomas F. Coleman, Burton S. Garbow, and Jorge J. Moré. Algorithm 618: Fortran subroutines for estimating sparse Jacobian matrices. ACM Transactions on Mathematical Software, 10(3):346–347, September 1984. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1985:FSE

[CGM85a] T. F. Coleman, B. S. Garbow, and J. J. Moré. Fortran subroutines for estimating sparse Hessian matrices. ACM Transactions on Mathematical Software, 11(4):378, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1985:AFS

[CGM85b] Thomas F. Coleman, Burton S.

Garbow, and Jorge J. Moré. Algorithm 636: FORTRAN subroutines for estimating sparse Hessian matrices. ACM Transactions on Mathematical Software, 11(4):378, December 1985. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/6193.html.

Chastain:1988:CCA

[CGMW88] M. Chastain, G. Gostin, J. Mankovich, and S. Wallach. The Convex C240 architecture. In IEEE [IEE88b], pages 321–329. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Chancelier:1989:ESS

[CGQS89] J.-P. Chancelier, C. Gomez, J.-P.

Quadrat, and A. Sulem. An expert system for stochastic control problems: automatic report generation. Computer Science in Economics and Management, 2(1):65– 82, ???? 1989. CODEN CSMNEF. ISSN 0921-2736.

Chang:1983:KCT

[Cha83] Shi Kuo Chang. Kung cheng tien nao: ko hsueh kung cheng yung cheng shih yu yen FORTRAN. Mei-kuo chih shih hsi tung hsueh yuan juan ti kung cheng hsueh chiao tsai: ti 3 tan yuan; CS2.2 Chang, S. K. (Shi Kuo), 1944- Meikou chih shih hsi tung hsueh yuan juan ti kung cheng hsueh chiao tsai; CS2.2. Chih shih hsi tung chu pan yu hsien kung ssu, Tai-pei shih, Taiwan, 1983. ISBN ???? 156 pp. LCCN ????

Chang:1986:ILG

[Cha86a] Wei-Te Chang. An implementation of level 1a GKS in FORTRAN 77. Thesis (M.S.), University of Texas at El Paso, El Paso, TX, USA, 1986. ix + 190 pp.

Chang:1986:AT

[Cha86b] Y. F. Chang. The ATOMCC toolbox. BYTE Magazine, 11(4):215– 224, April 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chang:1987:OF

[Cha87a] Wen-Zen Chang. An overview of Fortran 8X. Thesis (M.S.), University of Houston — University Park, Dept. of Computer Science, Houston, TX, USA, 1987. viii + 152 pp.

Charley:1987:HEC

[Cha87b] William J. Charley. The hydrologic engineering center FORTRAN compatible library: HECLIB programmers manual.

Technical report, The Hydrologic Engineering Center, Davis, CA, USA, 1987. various pp.

Cohen:1981:CAU

[CHH81] M. Cohen, V. C. Hamacher, and T. E. Hull. CADAC: An arithmetic unit for clean decimal arithmetic and controlled precision. In IEEE CA5 ’81 [IEE81], pages 106– 112. LCCN QA 76.6 S985t 1981.

URL http://www.acsel-lab.
com/arithmetic/arith5/papers/
ARITH5_Cohen_Hamacher_Bull.

pdf. IEEE catalog number 81CH1630-C.

Cohen:1983:CCP

[CHH83] Marty S. Cohen, T. E. Hull, and V. Carl Hamacher. CADAC: a controlled-precision decimal arithmetic unit. IEEE Transactions on Computers, C-32(4):370–377, April 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chirlian:1981:MF

[Chi81] Paul M. Chirlian. Microsoft Fortran. Dilithium Press, Forest Grove, OR, May 1981. ISBN 0-918398-46-0 (paperback). ix + 333 pp. LCCN QA76.73.F25 C483. US$15.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0918398460.

Chico:1985:VFF

[Chi85a] Stephen B. Chico. A VAX FOR-TRAN to FORTRAN 77 translator.(part I): a thesis. Thesis (M.S.), Clarkson University, Potsdam, NY, USA, 1985. 3 v. in 1 pp. Pt.1, p.17, Pt.2, p.61, Pt.3, p.70. Pt.2. Solving the convection-diffusion equation using Galerkin’s method and Hermitian finite elements–Pt.3. Multitasking algorithms in finite element analysis.

Ching:1985:ACS

[Chi85b] Wai-Mee Ching. An APL/370 compiler and some performance comparisons with APL interpreter and Fortran. Research report RC 11676, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1985. 7 pp.

Chino:1985:FPE

[Chi85c] Sadako Chino. FORTRAN programmes in the education of statistics. Number 21 in Computer science monographs(Tokei Suri Kenkyujo (Tokyo Japan)). Institute of Statistical Mathematics, Tokyo, Japan, February 1985. ISBN ???? 85 (or 86??) pp. LCCN QA276.4 .C551 1985.

Ching:1986:PAC

[Chi86a] Wai-Mee Ching. Program analysis and code generation in an APL/ 370 compiler. IBM Journal of Research and Development, 30(6): 594–602, November 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Chirlian:1986:MF

[Chi86b] Paul M. Chirlian. Microsoft Fortran. Crown Publishing, ????, June 1986. ISBN 0-517-56382-7. ???? pp. LCCN ???? US$15.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0517563827.

Chiu:1988:FPP

[Chi88] Min-Bing Chiu. Fortran programs for pressure flow in polymer processing. Thesis (M.S.), University of Louisville. Department of Chemical Engineering, Louisville, KY, USA, 1988. xiii + 132 pp.

Clint:1985:ADF

[CHPS85] M. Clint, C. Holt, R. H. Perrott, and A. Stewart. Algorithm 120: A DAP FORTRAN subroutine for the eigensolution of real symmetric matrices. The Computer Journal, 28(3):340–342, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/
340.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_28/Issue_03/tiff/341.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
28/Issue_03/tiff/342.tif.

Christopher:1984:RCG

[Chr84] Thomas W. Christopher. Reference count garbage collection. Software—Practice and Experience, 14(6):503–507, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chung:1988:WPF

[Chu88] Chang-Jo F. Chung. WIENER PACK: a FORTRAN 77 subroutine package for computing probabilities associated with Wiener (Brownian motion) and Brownian bridge processes. The American Statistician, 42(3):226–227, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        jstor.org/stable/2685015.

Chen:1986:ALE

[CK86a] T. Y. Chen and S. C. Kwan. An analysis of length equation using a dynamic approach. ACM SIGPLAN Notices, 21(4):42–47, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choora:1986:FIP

[CK86b] M. G. Choora and Ram Kumar. Fortran IV Programming. Advent Books, ????, June 1986. ISBN 0-7069-3008-8. ???? pp. LCCN ???? US$27.50. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0706930088.

Cooper:1988:ISA

[CK88] K. D. Cooper and K. Kennedy. Interprocedural side-effect analysis in linear time. ACM SIGPLAN Notices, 23(7):57–66, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org:80/pubs/
                                        citations/proceedings/pldi/
                                        53990/p57-cooper/.

Cooper:1985:IIA

[CKT85] Keith D. Cooper, Ken Kennedy, and Linda Torczon. The impact of interprocedural analysis and optimization on the design of a software development environment. ACM SIGPLAN Notices, 20 (7):107–116, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chan:1983:AFS

[CL83] N. N. Chan and Kim-Hung Li. Algorithm 115: A FORTRAN subroutine for finding a real symmetric matrix with prescribed diagonal elements and eigenvalues. The Computer Journal, 26(2):184–186, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/
184.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_26/Issue_02/tiff/185.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
26/Issue_02/tiff/186.tif.

Clark:1986:PF

[Cla86a] M. W. (Malcolm W.) Clark. Pcportable Fortran. Ellis Horwood series in computers and their applications. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, 1986. ISBN 0-7458-0005-X, 0-470-20756-6 (Halsted Press). 228 pp. LCCN QA76.73.F25 C55 1986. UK£18.50. Laser-printed using TEX.

Clark:1986:PPFa

[Cla86b] M. W. (Malcolm W.) Clark. PC-Portable Fortran. Ellis Horwood series in computers and their applications. Ellis Horwood, New York, NY, USA, November 1986. ISBN 0-470-20756-6. 228 pp. LCCN QA76.73.F25C55 1986. US$21.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0470207566.

Clark:1986:PPFb

[Cla86c] Malcolm W. Clark. PC-Portable Fortran. Ellis Horwood Series in Computers and Their Applications. Ellis Horwood, New York, NY, USA, November 1986. ISBN 0-13-654385-5. ???? pp. LCCN ???? US$52.50. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0136543855.

Clayton:1989:TFP

[Cla89] P. M. Clayton. Two Fortran programs for the production of isarithmic maps on a PDP-11 computer. Divisional report / Division of Soils, no. 98 0725-8526; Divisional report (Commonwealth Scientific and Industrial Research Organization (Australia). Division of Soils); no. 98. CSIRO, Canberra, Australia, 1989. ISBN 0-643-04786-7. 25 pp.

Clinch:1984:PRR

[Cli84] Simon Clinch. Programming With RT-11. The RT-11 Technical User’s Series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1984. ISBN 0-932376-32-0. ???? pp. LCCN ???? US$29.00.

Claus:1981:EBA

[CLW81] V. Claus, U. Lichtblau, and Frank Wankmueller. Erster Bericht zur Aufwaertsuebersetzung. Technical report, Universität Dortmund (??), Dortmund, Germany, 1981.

Couger:1981:ATA

[CM81a] J. Daniel Couger and Fred R. Mc-Fadden. Alternate tests to accompany first course in data processing with BASIC, COBOL, FORTRAN and RPG/2E: and first course in data processing with BASIC. John Wiley and Sons, New York, London, Sydney, 1981. ISBN ???? 57 pp. LCCN ????

Couger:1981:FCD

[CM81b] J. Daniel Couger and Fred R.

McFadden. First course in data processing with BASIC, COBOL, FORTRAN, RPG II. John Wiley and Sons, New York, London, Sydney, second edition, 1981. ISBN 0-471-05581-6 (paperback). xix + 532 pp. LCCN HF5548.2 .C68 1981b. Published in 1977 under title: A first course in data processing. Includes bibliographical references and index.

Cowley:1981:FCS

[CM81c] Ann D. Cowley and Greg McArthur.

FORTRAN compilers and software libraries at NCAR. NCAR technical note ????, Scientific Computing Division, NCAR, Boulder, CO, USA, October 1981. 10 pp.

Crawley:1983:SAFa

[CM83a] J. Winston Crawley and Charles E. Miller. A Structured Approach to Fortran. Reston Publishing Co., Inc., Reston, VA, USA, 1983. ISBN 0-8359-7092-2, 0-8359-7091-4 (paperback). xvi + 750 pp. LCCN QA76.73.F25C725 1983. US$33.92. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835970922.

Crawley:1983:SAFb

[CM83b] J. Winston Crawley and Charles E. Miller. A structured approach to FORTRAN: instructor’s manual. Reston Publishing Co., Inc., Reston, VA, USA, 1983. ISBN 0-8359-7093-0 (guide). 74 pp.

Couger:1984:FCD

[CM84a] J. Daniel Couger and Fred R. McFadden. First course in data processing with BASIC, COBOL, FORTRAN, and RPG. Wiley series in computers and information processing systems in business. John Wiley and Sons, New York, London, Sydney, third edition, January 1984. ISBN 0-471-86946-5. xxi + 640 + [30] pp. LCCN ???? US$51.47. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471869465.

Couger:1984:SWA

[CM84b] J. Daniel Couger and Fred R. McFadden. Student Workbook to Accompany First Course in Data Processing With Basic, Cobol, Fortran, and RPG. John Wiley and Sons, New York, London, Sydney, January 1984. ISBN 0-471-86951-1. ???? pp. LCCN ???? US$24.43. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471869511.

Creutz:1986:FCT

[CM86] Michael Creutz and K. J. M.

Moriarty. FORTRAN code for the three-dimensional Ising model. Computer Physics Communications, 39(2):173–180, February/ March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465586901281.

Crawley:1987:SAF

[CM87] J. Winston Crawley and Charles E.

Miller. A Structured Approach to Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1987. ISBN 0-13-854183-3 (paperback). xxii + 632 pp. LCCN QA76.73.F25 C725 1987. US$55.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0138541833.

Couger:1989:IAP

[CM89] J. Daniel Couger and Fred R. Mc-Fadden. Introduccion al procesamiento de datos con BASIC, COBOL, FORTRAN y RPG. Trillas, Mexico, DF, Mexico, 1a edition, 1989. ISBN 968-24-2557-3. 760 pp.

Corona:1988:SLE

[CMM^{+}88] A. Corona, C. Martini, M. Morando, S. Ridella, and C. Rolando. Solving linear equation systems on vector computers with maximum efficiency. Parallel Computing, 8(1– 3):133–139, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cockrell:1983:NGP

[Coc83] P. R. Cockrell. A new general purpose method for large volume production of contour charts. Computer Graphics Forum, 2(1):35–47, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cody:1986:ETRc

[Cod86a] W. J. Cody. ELEFUNT test results under FX/Fortran version 1.0 on the Alliant FX/8. Technical Report MCS-TM-78, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1986. ?? pp.

Cody:1986:ETRb

[Cod86b] W. J. Cody. ELEFUNT test results under NS32000 Fortran V2.5.3 on the Sequent Balance. Technical Report MCS-TM-80, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1986. 12 pp.

Cody:1988:AMS

[Cod88] W. J. Cody. Algorithm 665. MACHAR: A subroutine to dynamically determine machine parameters. ACM Transactions on Mathematical Software, 14(4): 303–311, December 1988. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/51907.html.

Cody:1989:ETR

[Cod89] W. J. Cody. ELEFUNT test results using Titan Fortran under Ardent UNIX 2.0 on the Titan. Technical Report MCS-TM-129, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, March 1989. iii + 14 pp.

Coleman:1982:FIP

[Col82] H. P. Coleman. A Fortran IV plot routine with hidden line suppression for use with small computer systems. NRL memorandum report 4776, Naval Research Laboratory; National Technical Information Service, Washington, DC, USA, 1982. iii + 12 pp.

Cole:1983:AFIb

[Col83] J. W. Perry Cole. ANSI Fortran IV With Fortran 77 Extensions: a Structured Programming Approach. Wm. C. Brown Publishers, Dubuque, IA, USA, second edition, March 1983. ISBN 0-697-08172-9 (paperback). xv + 534 + 117 + [41] pp. LCCN QA76.73.F25C64 1983. US$53.65. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0697081729.

Cole:1984:RIK

[Col84a] Allen Cole. A Ratfor implementation of KERMIT. In Software Tools Users Group [Sof84], pages 355–367. ISBN ???? LCCN QA76.8.U65 U83 1984.

Coleman:1984:FSB

[Col84b] J. P. Coleman. A Fortran subroutine for the Bessel function J_{n}(x) of order 0 to 10. Computer Physics Communications, 35 (??):C–654, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Cole:1987:AFS

[Col87a] J. W. Perry Cole. ANSI Fortran 77: a Structured Problem Solving Approach. Wm. C. Brown Publishers, Dubuque, IA, USA, February 1987. ISBN 0-697-00442-2 (paperback). xxviii + 638 + [169] pp. LCCN ???? US$53.65. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0697004422.

Coll:1987:HSR

[Col87b] Richard Coll. How to set up and run your micro: everything you need to know. Compsit Corp., Plainfield, NJ, USA, fourth edition, 1987. ISBN ???? 63 pp. LCCN ????

Colquitt:1989:CFVa

[Col89a] Walter Colquitt. Creation of fully vectorized FORTRAN code for integrating the movement of dust grains in interplanetary environments. Technical report, Research Institute for Computing and Information Systems University of Houston–Clear Lake, Houston, TX, USA, 1989. 10 pp.

Colquitt:1989:CFVb

[Col89b] Walter Colquitt. Creation of fully vectorized FORTRAN code for integrating the movement of dust grains in interplanetary environments. NASA contractor report CR-187263, Research Institute for Computing and Information Systems University of Houston-Clear Lake; National Aeronautics and Space Administration, Houston, TX, USA, 1989. ???? pp. National Technical Information Service, distributor.

CBEMA:1989:FD

[Com89] Computer and Business Equipment Manufacturers Association (CBEMA). Fortran 8X draft, May 1989: American National Standard for Information Systems Programming Language Fortran, S8 (X3.9-198x), Revision of X3.9-1978. ACM SIGPLAN FOR-TRAN Forum, 8(4):various, December 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

CDC:1981:FVU

[Con81a] Control Data Corporation. FOR-TRAN version 5 user’s guide. The Corporation, Sunnyvale, CA, USA, revision A edition, 1981. 142 pp.

CDC:1981:MSF

[Con81b] Control Data Corporation. Mass storage Fortran Version 3A / B reference manual. Control Data Corporation, LaJolla, CA, USA, 1981. various pp.

CDC:1982:CSE

[Con82a] Control Data Corporation. Cyber 205 service efficient Fortran Techniques: user guide. CDC, Minneapolis, MN, USA, 1982. ???? pp.

CDC:1982:FVR

[Con82b] Control Data Corporation. FOR-TRAN version 5 reference manual. CDC, Sunnyvale, CA, USA, revision F edition, 1982. 345 pp.

CDC:1982:FVU

[Con82c] Control Data Corporation. FOR-TRAN version 5 user’s guide. The Corporation, Sunnyvale, CA, USA, revised edition, 1982. 138 pp.

CDC:1982:SPFa

[Con82d] Control Data Corporation. Structured programming with FOR-TRAN 77. Control Data Corporation, Minneapolis, MN, USA, 1982. program files (Control Data 110) on 9 computer disks student guide + administration guide. pp.

CDC:1982:SPFb

[Con82e] Control Data Corporation. Structured programming with FOR-TRAN 77: student guide. Control Data Corporation, Minneapolis, MN, USA, 1982. various pp.

CDC:1983:CCF

[Con83a] Control Data Corporation. CDC Cyber 200 Fortran version 2, for use with CDC Cyber 200 virtual storage operating system version 2: reference manual. CDC, Sunnyvale, CA, USA, 1983. ???? pp.

CDC:1983:FEV

[Con83b] Control Data Corporation. FOR-TRAN extended version 4 reference manual. CDC, Sunnyvale, CA, USA, 1983. ???? pp.

CDC:1983:FNV

[Con83c] Control Data Corporation. FOR-TRAN for NOS/VE, language definition: usage. Control Data Corporation, Sunnyvale, CA, USA, 1983. ???? pp.

CDC:1983:FVR

[Con83d] Control Data Corporation. FOR-TRAN version 5 reference manual. Control Data Corporation, Sunnyvale, CA, USA, revised edition, 1983. 342 pp.

ConvergentTechnologies:1983:MFM

[Con83e] Convergent Technologies, Inc.

MegaFrame FORTRAN manual. Convergent Technologies, Inc., Santa Clara, CA, USA, 1983. various pp. B-09-00316-01.

CDC:1984:FNV

[Con84] Control Data Corporation. FOR-TRAN for NOS/VE language definition: usage. CDC, St. Paul, MN, USA, 1984. ???? pp.

CDC:1985:FVU

[Con85a] Control Data Corporation. Fortran 200 version 1, for use with CDC Cyber 200 virtual storage operating system version 2: reference manual. CDC, Sunnyvale, CA, USA, 1985. ???? pp.

CDC:1985:FNVa

[Con85b] Control Data Corporation. FOR-TRAN for NOS/VE language definition: usage. CDC, Sunnyvale, CA, USA, 1985. ???? pp.

CDC:1985:FNVc

[Con85c] Control Data Corporation. FOR-TRAN for NOS/VE, quick reference. Control Data Corporation, Sunnyvale, CA, USA, 1985. ???? pp.

CDC:1985:FNVb

[Con85d] Control Data Corporation. FOR-TRAN for NOS/VE: topics for FORTRAN programmers: usage. Control Data Corporation, St. Paul, MN, USA, revised B edition, 1985. various pp.

CDC:1986:FNV

[Con86] Control Data Corporation. FOR-TRAN for NOS/VE LIB99: usage. CDC, Sunnyvale, CA, USA, 1986. ???? pp.

CDC:1987:FNVb

[Con87a] Control Data Corporation. FOR-TRAN for NOS/VE language definition: usage. CDC, Sunnyvale, CA, USA, revised G edition, 1987. ???? pp.

CDC:1987:FNVc

[Con87b] Control Data Corporation. FOR-TRAN for NOS/VE language definition: usage. CDC, Sunnyvale, CA, USA, revised H edition, 1987. ???? pp.

CDC:1987:FNVa

[Con87c] Control Data Corporation. FOR-TRAN for NOS/VE LIB99: usage. CDC, Sunnyvale, CA, USA, revised B edition, 1987. ???? pp.

CDC:1988:FVN

[Con88a] Control Data Corporation. FOR-TRAN Version 1 for NOS/VE language definition: usage. Control Data Corporation, St. Paul, MN, USA, revised J edition, 1988. various pp.

Convex:1988:CF

[Con88b] Convex Computer Corporation.

CONVEX FORTRAN. CONVEX Computer Corp., Richardson, TX, USA, 1988. ???? pp. 720-000033-200.

Correa:1981:ESM

[Cor81] Hector Correa. Elements of systems mathematics and FORTRAN programming for social scientists. ????, Pittsburgh, PA, USA, 1981. ISBN ???? v + 496 pp. LCCN ????

Corbett:1982:EAF

[Cor82a] R. P. Corbett. Enhanced arithmetic for Fortran. ACM SIG-PLAN Notices, 17(12):41–48, December 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cornyn:1982:GCF

[Cor82b] John Cornyn. Guidelines for coding Fortran programs. NORDA report 41, Naval Ocean Research and Development Activity; National Technical Information Service, NSTL Station, MS, USA, 1982. vi + 54 pp.

Correa:1982:ESS

[Cor82c] Hector Correa. Elements of social systems mathematics and FOR-TRAN programming. ????, Pittsburgh, PA, USA, 1982. ISBN ???? 560 pp. LCCN ????

Corbett:1983:EAF

[Cor83] Robert Paul Corbett. Enhanced arithmetic for Fortran. ACM SIGNUM Newsletter, 18(1):24–28, January 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Corliss:1988:ADA

[Cor88] George F. Corliss. Applications of differentiation arithmetic. In Moore [Moo88b], pages 127– 148. ISBN 0-12-505630-3. LCCN QA76.9.E94 R45 1988.

Counihan:1985:BRBa

[Cou85a] Martin Counihan. Book review: A structured approach to Fortran 77 programming, T. M. R. Ellis, Addison-Wesley International Computer Science Series, 1982. no. of pages: 350. price: £8.95. Software—Practice and Experience, 15 (3):319, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Counihan:1985:BRBb

[Cou85b] Martin Counihan. Book review: Structured Fortran 77 for engineers and scientists, D. M.

Etter, Benjamin/Cummings Publishing Company, 1983. No. of pages: 357. Price: £9.95. Software—Practice and Experience, 15 (3):319, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cowell:1984:SDM

[Cow84] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984.

Clinch:1984:PR

[CP84] Simon Clinch and Stephen Peters.

Programming with RT-11. The RT-11 technical user’s series. Digital DECbooks, Burlington, Mass., 1984. ISBN 0-932376-32-0 (paperback: v. 1). v. ¡1–2 ¿ pp. LCCN QA76.6.C55 1984. US$24.00 (v. 1), US$100.00 (set). Vol. ¡2 ¿ by Stephen Peters and others. Contents: v. 1. Program development facilities – v. 2. Callable system facilities.

Christiansen:1984:OSC

[CR84] J. P. Christiansen and K. V.

Roberts. Olympus: a standard control and utility package for initial-value Fortran programs. Computer Physics Communications, 35(1–3):C–243–C–244, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0010465584824819.

Cray:1983:CCS

[Cra83] Cray Research, Inc. Cray-1 computer systems: Fortran (CFT) internal reference manual. Number SM-0017 in Publication. Cray Research, Inc., Minneapolis, MN, revised B edition, 1983. ISBN ???? various pp. LCCN ????

Cray:1984:FCR

[Cra84] Cray Research, Inc. Fortran (CFT) reference manual. Number SR-0009 in Publication. Cray Research, Inc., Minneapolis, MN, revision K edition, 1984. ISBN ???? ???? pp. LCCN ????

Crawford:1986:APR

[Cra86a] Charles R. Crawford. Algorithm 646: PDFIND: a routine to find a positive definite linear combination of two real symmetric matrices. ACM Transactions on Mathematical Software, 12(3):278–282, September 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/citations/journals/
                                        toms/1986-12-3/p278-crawford/

.

Cray:1986:FCR

[Cra86b] Cray Research, Inc., Minneapolis, MN. Fortran (CFT) reference manual, 1986. ???? pp.

Cray:1989:FCR

[Cra89a] Cray Research, Inc. Fortran (CFT) reference manual. Number SR-0009 in Publication. Cray Research, Inc., Minneapolis, MN, revised M edition, 1989. ISBN ???? various pp. LCCN ????

Cray:1989:UFL

[Cra89b] Cray Research, Inc. UNICOS Fortran library reference manual. Number SR-2079 5.0 in Publication. Cray Research, Inc., Minneapolis, MN, 1989. ISBN ???? various pp. LCCN ????

Cress:1989:DES

[Cre89] Jill Janene Cress. Development of an expert system to convert knowledge-based geological engineering systems into Fortran. Computer science thesis (M.S.), University of Missouri– Rolla, Rolla, MO, USA, 1989. viii + 80 pp.

Crowl:1985:RFE

[Cro85a] Daniel A. Crowl. A real-time FOR-TRAN executive. IEEE Micro, 5 (4):48–66, July/August 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Crowl:1985:RTF

[Cro85b] Daniel A. Crowl. A real-time Fortran executive. IEEE Micro, 5(4): 48–66, July/August 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Crockett:1987:PFFa

[Cro87a] T. Crockett. Performance of Fortran floating-point operations on the Flex/32 multicomputer. ICASE Interim Report 4, ICASE, NASA Langley Research Center, Hampton, VA, USA, 1987.

Crockett:1987:PFFb

[Cro87b] Thomas W. Crockett. Performance of FORTRAN floatingpoint operations on the Flex/32 multicomputer. NASA contractor report NASA CR-178364, ICASE interim report 178364, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1987. ???? pp.

Chandra:1989:AAF

[CRV^{+}89] U. Chandra, G. Riccardi, J. Vagi, J.-L. Dekeyser, and F. Hannedouche. Aftran: Array Fortran programming language. Computer Physics Communications, 57(1–3):263– 267, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465589902257.

Cassel:1983:FME

[CS83] Don Cassel and Richard Swanson. Fortran Made Easy. Reston Publishing Co., Inc., Reston, VA, USA, May 1983. ISBN 0-8359-2090-9 (hardcover), 0-8359-2089-5 (paperback). xii + 258 pp. LCCN QA76.73.F25C38 1983. US$19.95 (hardcover), US$14.95; US$29.93. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835920895;   http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0835920909.

Couger:1984:IAL

[CS84] James Daniel Couger and Loren E.

Shannon. Introduccion al lenguaje FORTRAN. Serie Computacion Serie de Ensenanza Personal. El Ateneo, Buenos Aires, Argentina, 2., réimpresion edition, 1984. ISBN 950-02-5012-8. xiii + 161 pp.

Ciarcia:1986:DFP

[CSC^{+}86] C. A. Ciarcia, W. A. Schier, G. P.

Couchell, D. J. Pullen, R. S. Tanczyn, M. H. Haghighi, and Q. Sharfuddin. DENTS: a FOR-TRAN program for analysing continuous neutron spectra. Computer Physics Communications, 39(2):233–243, February/March 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465586901347.

Conte:1982:EDC

[CSD82] S. D. Conte, V. Y. Shen, and K. Dickey. On the effect of different counting rules for control flow operators on Software Science metrics in Fortran. ACM SIG-METRICS Performance Evaluation Review, 11(2):118–126, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Couger:1983:PLA

[CSD83] J. Daniel Couger, Loren E. Shannon, and Charles B. Drebes. Personal Learning Aid for Fortran: a Simplified Approach. Dow Jones-Irwin personal learning aid series. R. D. Irwin, Homewood, IL, USA, third edition, July 1983. ISBN 0-87094-327-8 (paperback). xi + 201 pp. LCCN ???? US$11.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0870943278.

Carnevali:1986:MIM

[CSZ86] Paolo Carnevali, Piero Sguazzero, and Vittorio Zecca. Microtasking on IBM multiprocessors. IBM Journal of Research and Development, 30(6):574–582, November 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Tan:1981:FYY

[cT81] Hao chiang Tan. FORTRAN yu yen hsi ti yu chieh ta. Ching-hua ta hsueh chu pan she: Hsin hua shu tien Pei-ching fa hsing so fa hsing, Pei-ching, ti 1 pan edition, 1981. ISBN ???? 56 pp. LCCN ????

Cowell:1986:TFDa

[CT86a] Wayne R. Cowell and Christopher P. Thompson. Transforming Fortran DO loops to improve performance on vector architectures. Numerical Algorithms Group technical report NP1168, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1986. iv + 28 pp.

Cowell:1986:TFDb

[CT86b] Wayne R. Cowell and Christopher P. Thompson. Transforming Fortran DO loops to improve performance on vector architectures. ACM Transactions on Mathematical Software, 12(4):324–353, December 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/
0098-3500/24035.html.

Cochrane:1988:SPN

[CT88] David L. Cochrane and Donald G.

Truhlar. Strategies and performance norms for efficient utilization of vector pipeline computers as illustrated by the classical mechanical simulation of rotationally inelastic collisions. Parallel Computing, 6(1):63–85, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tan:1984:FYY

[cTcT84] Hao chiang Tan and Shu ching Tien. FORTRAN yu yen. Ching hua ta hsueh chu pan she: Hsin hua shu tien Pei-ching fa hsing so fa hsing, Pei-ching, tseng ting pen Ti 3 pan edition, 1984. ISBN ????

[2] + 3 + 248 pp. LCCN ????

Culloch:1988:PPT

[Cul88] A. D. Culloch. Parallel programming toolkit for 3L — C FOR-TRAN and Pascal. In Jon Kerridge, editor, Developments Using Occam, pages 23–30. IOS Press, Amsterdam, The Netherlands, 1988.

Cizek:1988:SCQ

[CV88a] J. Cizek and F. Vinette. Symbolic computation in quantum mechanics. several simple examples.

In Anonymous [Ano88b]. CODEN IJQSDI. ISSN 0161-3642.

Coleman:1988:HMC

[CV88b] Thomas F. Coleman and Charles F. Van Loan. Handbook for Matrix Computations, volume 4 of Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1988. ISBN 0-89871-227-0 (paperback), 1-61197-104-7 (ebook). vii + 264 pp. LCCN QA188 .C651 1988. US$34.00.

Carnahan:1985:FMI

[CW85] Brice Carnahan and James O. Wilkes. FORTRAN 77 with MTS and the IBM PC. Technical report, College of Engineering, University of Michigan, Ann Arbor, MI, USA, 1985. 306 pp.

Carnahan:1988:FMI

[CW88] Brice Carnahan and James O. Wilkes. FORTRAN 77 with MTS and the IBM PC. Technical report, Collegeof Engineering, University of Michigan, Ann Arbor, MI, USA, 1988. various pp.

Carnahan:1989:FMI

[CW89] Brice Carnahan and James O. Wilkes. FORTRAN 77 with MTS and the IBM PS/ 2. Technical report, Collegeof Engineering, University of Michigan, Ann Arbor, MI, USA, 1989. various pp.

Cole:1983:AFIa

[CwL83] J. W. Perry Cole and Kuang wu Lai. ANSI FORTRAN IV yu FORTRAN 77 cheng shih she chi: a structured programming approach. Ju lin tu shu kung ssu, Tai-pei shih, chu pan edition, 1983. ISBN ???? ???? pp. LCCN ????

Chen:1989:PTA

[CY89] Xiangping Chen and Daoning Ying. Polygon triangulation algorithm as a powerful core processor of PLAN-I. Computer Graphics Forum, 8(3):193–198, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Czarnecki:1983:FCP

[Cza83] John B. Czarnecki. Fortran computer programs to plot and process aquifer pressure and temperature data. U.S. Geological Survey water-resources investigations report 83-4051, U.S. Dept. of the Interior, Geological Survey [Open-File Service Section, Western Distribution Branch], 1, St. Paul, MN, USA, 1983. iii + 50 pp.

Demillo:1988:UMA

[DAG^{+}88] R. A. Demillo, W. F. Appelbe, D. S. Guindi, K. N. King, and W. M. Mccracken. Using mutation analysis for testing Ada programs. Technical Report SERC-TR-9-P, Software Engineering Research Centre, Utrecht, The Netherlands (??), February 3, 1988.

Dahmen:1981:FFP

[Dah81] N. Dahmen. FORCASD — ein FORTRAN-orientiertes programmsystem zur modellbil dung und simulation. Informatik Fachberichte, 41:133–147, 1981. ISSN 0343-3005.

Dallal:1988:PFP

[Dal88a] G. E. Dallal. PITMAN: a FOR-TRAN program for exact randomization tests. Computers in Biomedical Research, 21(1):9–15, February 1988. CODEN CP-BRAF. ISSN 0010-4809 (print), 1090-2368 (electronic).

Dallal:1988:RFP

[Dal88b] G. E. Dallal. RCJOIN: a FORTRAN program for isolating sources of interaction in two-way tables of counts. Computers in Biomedical Research, 21(2):129– 136, April 1988. CODEN CP-BRAF. ISSN 0010-4809 (print), 1090-2368 (electronic).

Dallal:1989:TFP

[Dal89] G. E. Dallal. TRACK: a FOR-TRAN program for calculating the Foulkes-Davis tracking index. Computers in Biology and Medicine, 19(5):367–371, ???? 1989. CODEN CBMDAW. ISSN 0010-4825.

DGC:1981:FRM

[Dat81] Data General Corporation. FOR-TRAN 77: reference manual. Data General Corporation, Southboro, MA, USA, 1981. 274 pp.

DGC:1984:FEM

[Dat84] Data General Corporation. FOR-TRAN 77 environment manual (AOS/VS). Data General Corporation, Southboro, MA, USA, revised10 edition, 1984. various pp.

DGC:1985:FEM

[Dat85a] Data General Corporation. FOR-TRAN 77 environment manual (AOS). Data General Corporation, Southboro, MA, USA, revision 01, Jan. 1985 edition, 1985. various pp.

DGC:1985:FR

[Dat85b] Data General Corporation. Fortran 77 reference. Data General Corporation, Southboro, MA, USA, 1985. various pp.

David:1981:PF

[Dav81a] Daniel Jean David. Programmer en Fortran. Editests, Paris, France, 1981. ISBN ???? 127 pp. LCCN ????

Davis:1981:FGS

[Dav81b] William S. Davis. Fortran, Getting Started. Addison-Wesley, Reading, MA, USA, February 1981. ISBN 0-201-03104-3. vii + 168 pp. LCCN ???? US$8.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201031043.

Davidson:1982:FIF

[Dav82] S. A. Davidson. FORTRAN implementation of friedman’s test for several related samples. JSC 8, Lyndon B. Johnson Space Center, Houston, TX, USA, 1982. various pp.

David:1984:LF

[Dav84a] Daniel Jean David. Le Langage Fortran. Editests, Paris, France, 1984. ISBN 2-86699-020-X. 141 pp.

Davidson:1984:ELG

[Dav84b] Charles Davidson. The emergence of load-and-go systems for FOR-TRAN. Annals of the History of Computing, 6(1):35–37, January/ March 1984. CODEN AHCOE5. ISSN 0164-1239.

Davis:1985:IF

[Dav85] William S. Davis. Iniciacion en FORTRAN. Fondo Educativo Interamericano, Mexico, DF, Mexico, 1985. ISBN 968-858-011-2. xii + 207 pp.

Davis:1986:FSM

[Dav86] Christopher Paul Davis. A Fortran simulation of multi-effect evaporation with provision for boilingpoint rise and variable feed configuration. Thesis (M.S.), University of Arkansas, Fayetteville, Fayetteville, AR, USA, 1986. vi + 84 pp.

Davenport:1989:EEC

[Dav89] James Harold Davenport, editor. EUROCAL ’87: European Conference on Computer Algebra, Leipzig, GDR, June 2–5, 1987: proceedings, volume 378 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. CO-DEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

D:1982:AFD

[DB82a] Brian D. and George H. Blackwood. Apple Fortran: Detailed Language Instructions Specifically for the Apple Computer. Howard W. Sams, Indianapolis, IN 46268, USA, November 1982. ISBN 0-672-21911-5. ???? pp. LCCN ????

deBrito:1982:FPI

[dB82b] Adelsindo Liberato de Brito. FOR-TRAN program for the integral of three spherical harmonics. Computer Physics Communications, 25(1):81–85, January 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465582900467.

deBrito:1984:FPI

[dB84] Adelsindo Liberato de Brito. Fortran program for the integral of three spherical harmonics. Computer Physics Communications, 35 (1–3):C–770, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584828982.

Fineberg:1989:TAD

[DBFK89] J. Dongarra, O. Brewer, S. Fineberg, and J. A. Kohl. A tool to aid in the design, implementation, and understanding of matrix algorithms for parallel processors. Technical Report MCS-P115-1189, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1989.

Devon:1989:FAK

[DcF89] Richard F. Devon and Wu chun Feng. FORTRAN at the keyboard: an introduction to FORTRAN for engineers. Kendall/Hunt Pub., Dubuque, IA, USA, 1989. ISBN 0-8403-5461-4 (paperback). 78 + [26] pp.

Dongarra:1985:PES

[DCHH85] J. J. Dongarra, J. Du Croz, Sven Hammarling, and R. J. Hanson. A proposal for an extended set of Fortran basic linear algebra subprograms. ACM SIGNUM Newsletter, 20(1):2–18, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1987:ESF

[DCHH87] J. J. Dongarra, J. Du Croz, S. Hammarling, and R. J. Hanson. An extended set of Fortran Basic Linear Algebra Subprograms: Model implementation and test programs. Technical Report MCS-TM-81, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, January 1987.

Dongarra:1988:AES

[DCHH88a] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. Algorithm 656: An extended set of Basic Linear Algebra Subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 14(1):18–32, March 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1988:CES

[DCHH88b] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. Corrigenda: “An extended set of FOR-TRAN Basic Linear Algebra Subprograms”. ACM Transactions on Mathematical Software, 14 (4):399, December 1988. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [DCHH88c].

Dongarra:1988:ESF

[DCHH88c] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. An extended set of FORTRAN Basic Linear Algebra Subprograms. ACM Transactions on Mathematical Software, 14(1):1–17, March 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/42291.html.  See also

[DCHH88b].

Dongarra:1986:PVC

[DD86] J. Dongarra and I. Duff. Performance of vector computers for direct and indirect addressing in Fortran. Harwell report, AERE Harwell Laboratory, Chilton, Oxon, England, ?? 1986.

Demmel:1989:PDL

[DDDG89] J. Demmel, J. J. Dongarra, J. DuCroz, and A. Greenbaum. A project for developing a linear algebra library for high-performance computers. In Wright [Wri89], pages 87–92. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Dencker:1984:OPT

[DDH84] Peter Dencker, Karl Dürre, and Johannes Heuft. Optimization of parser tables for portable compilers. ACM Transactions on Programming Languages and Systems, 6(4):546–572, October 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dongarra:1984:PES

[DDHH84] J. Dongarra, J. DuCroz, S. Hammarling, and R. Hanson. A proposal for an extended set of Fortran Basic Linear Algebra Subprograms. Technical Memo 41, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, December 1984.

Dongarra:1984:SMA

[DE84] Jack J. Dongarra and Stanley C. Eisenstat. Squeezing the most out of an algorithm in CRAY FORTRAN. ACM Transactions on Mathematical Software, 10(3): 219–230, September 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Delannoy:1982:APF

[Del82] Claude Delannoy. Apprendre à programmer en Fortran. Eyrolles, Paris, France, 1982. ISBN ???? xii + 187 pp. LCCN ????

Delannoy:1983:APF

[Del83] Claude Delannoy. Apprendre à programmer en FORTRAN. Eyrolles, Paris, France, 2e edition, 1983. ISBN ???? xii + 187 pp. LCCN ????

Delannoy:1985:APF

[Del85] Claude Delannoy. Apprendre à programmer en Fortran. Eyrolles, Paris, France, 3e edition, 1985. ISBN ???? xii + 188 pp. LCCN ????

Delannoy:1988:PF

[Del88] Claude Delannoy. Programmer en Fortran. Eyrolles, Paris, France, 4e revue et corrigée edition, 1988. ISBN ???? xii + 187 pp. LCCN ????

Denelcor:1982:HFU

[Den82] Inc. Denelcor. HEP FORTRAN user’s guide. Publication Number 9000006, February 1982.

Dence:1984:FCa

[Den84a] Thomas P. Dence. The FORTRAN cookbook. Tab Books, Blue Ridge Summit, PA, USA, second edition, October 1984. ISBN 0-8306-0737-4, 0-8306-1737-X (paperback). viii + 232 pp. LCCN ????

Dence:1984:FCb

[Den84b] Thomas P. Dence. The Fortran Cookbook. Tab Books, Blue Ridge Summit, PA, USA, October 1984. ISBN 0-8306-1737-X. ???? pp. LCCN ????

Derbyshire:1982:IFS

[Der82] Elisabeth N. Derbyshire. Introductory FORTRAN [a self study subject], 1982.

Desai:1989:FPN

[Des89] R. C. Desai. FORTRAN programming and numerical methods. Tata McGraw-Hill, New Delhi, India, 1989. ISBN 0-07-451579-9. xi + 309 pp.

diAntonio:1989:UPA

[dEV89] R. di Antonio, J. Eilert, and M. Vitaletti. Using PAGE-AHEAD for large FORTRAN programs. In ACM [ACM89b], pages 511–520. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Devon:1989:FKI

[DF89] Richard Devon and Wu Chun Feng. Fortran at the Keyboard: An Introduction to Fortran for Engineers. Kendall/Hunt Pub., Dubuque, IA, USA, July 1989. ISBN 0-8403-5461-4. ???? pp. LCCN ???? US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0840354614.

DeRaedt:1981:MCF

[DFD81] H. De Raedt, J. Fivez, and B. De Raedt. Manual for commute, a FORTRAN program for symbolic evaluation of commutators and correlation functions. Computer Physics Communications, 23(2):209–220, July 1, 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465581900345.

DeRaedt:1984:MCF

[DFD84] H. De Raedt, J. Fivez, and B. De Raedt. Manual for commute, a FORTRAN program for symbolic evaluation of commutators and correlation functions. Computer Physics Communications, 35 (1–3):C–721, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584828611.

Diaz:1981:FPS

[DFK81] J. C. Diaz, Graeme Fairweather, and Patrick Keast. FOR-TRAN packages for solving almost block diagonal linear systems by modified alternate row and column elimination. Technical report 148/81, University of Toronto, Dept. of Computer Science, Toronto, Ontario, Canada, 1981. 34 + 16 + 15 pp.

Diaz:1983:ACA

[DFK83a] J. C. Díaz, G. Fairweather, and P. Keast. Algorithm 603: COL-ROW and ARCECO: FORTRAN packages for solving certain almost block diagonal linear systems by modified alternate row and column elimination. ACM Transactions on Mathematical Software, 9 (3):376–380, September 1983. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [DFK88].

Diaz:1983:FPS

[DFK83b] J. C. Díaz, G. Fairweather, and P. Keast. FORTRAN packages for solving certain almost block diagonal linear systems by modified alternate row and column elimination. ACM Transactions on Mathematical Software, 9(3):358–375, September 1983. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Diaz:1988:RCA

[DFK88] J. C. Diaz, G. Fairweather, and P. Keast. Remark on “Algorithm 603: COLROW and ARCECO: FORTRAN packages for solving certain almost block diagonal linear systems by modified alternate row and column elimination”. ACM Transactions on Mathematical Software, 14(2):196, June 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [DFK83a].

Dodson:1982:RBL

[DG82] David S. Dodson and Roger G. Grimes. Remark on “Algorithm 539: Basic Linear Algebra Subprograms for Fortran usage [F1]”. ACM Transactions on Mathematical Software, 8(4):403–404, December 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [LHKK79a, Dod83, HK87a, LN88b].

Darema:1988:SCM

[DGNP88a] F. Darema, D. A. George, V. A. Norton, and G. F. Pfister. A single-program-multiple-data computational model for EPEX/ FORTRAN. Parallel Computing, 7(1):11–24, April 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Darema:1988:SPM

[DGNP88b] F. Darema, D. A. George, V. A.

Norton, and G. F. Pfister. A single-program-multiple-data computational model for EPEX/ FORTRAN. Parallel Computing, 7(1):11–24, April 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Davis:1982:IMA

[DH82] Gordon Bitter Davis and Thomas Russell Hoffmann. Instructor’s manual to accompany FORTRAN 77: a structured, disciplined style; based on 1977 American National Standard FORTRAN and compatible with WATFOR, WATFIV, WATFIV-S, and M77 Fortran compilers. McGraw-Hill, New York, NY, USA, second edition, 1982. ISBN 0-07-015904-1. iv + 383 pp.

Davis:1983:FSD

[DH83] Gordon Bitter Davis and Thomas Russell Hoffmann. FORTRAN 77: a structured, disciplined style: based on 1977 American National Standard FORTRAN and compatible with WATFOR, WAT-FIV, WATFIV-S, and M77 FOR-TRAN compilers. McGraw-Hill, New York, NY, USA, second edition, June 1983. ISBN 0-07-015903-3 (paperback), 0-07-015904-1 (hard). x + 400 pp. LCCN QA76.73.F25 D385 1983. US$43.95; US$16.95

                                        (est.). URL http://www.cbooks.
                                        com/sqlnut/SP/search/gtsumt?
                                        source=&isbn=0070159033.

Davis:1984:FEE

[DH84a] Gordon Bitter Davis and Thomas Russell Hoffmann. Fortran 77: un estilo estructurado y disciplinado basado en Fortran 1977 American National Standard y compatible con los compiladores Fortran, WATFOR, WATFIV, WATFIV-S y M 77. Libros McGraw-Hill de Mexico, Mexico, DF, Mexico, 1a edition, 1984. ISBN 968-451-562-6. x + 416 pp.

Dongarra:1984:CPL

[DH84b] J. J. Dongarra and R. E. Hiromoto. A collection of parallel linear equations routines for the Denelcor HEP. Parallel Computing, 1 (2):133–142, December 1984. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Davis:1988:F

[DH88a] Gordon B. Davis and Thomas R. Hoffmann. Fortran 77. Mc-Graw-Hill, New York, NY, USA, third edition, February 1988. ISBN 0-07-015904-1. x + 400 pp. LCCN QA76.73.F25D385 1983. US$43.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0070159041.

Davis:1988:FSD

[DH88b] Gordon Bitter Davis and Thomas Russell Hoffmann. Fortran 77: a Structured, Disciplined Style. Mc-Graw-Hill, New York, NY, USA, third edition, February 1988. ISBN 0-07-100194-8 (paperback), 0-07-015905-X (paperback). ix + 434 pp. LCCN QA76.73.F25 D36 1988. US$46.65. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
007015905X.

Agrawal:1988:REF

[Dha88] Dharma P. Agrawal. Retargeting of existing FORTRAN program and development of parallel compilers final report. NASA contractor report NASA CR-182806, Dept. of Electrical and Computer Engineering, North Carolina State University, Raleigh, NC, USA, 1988. ???? pp. National Aeronautics and Space Administration; National Technical Information Service, distributor.

Didriksen:1986:DIP

[Did86] T. Didriksen. A database interface for a prototype programming environment. EPOS 3, Division of Computer Science, University of Trondheim, Norwegian Institute of Technology, Trondheim, Norway, November 1986. 1–25 pp.

DEC:1982:VF

[Dig82a] Digital Equipment Corporation.

VAX-11 Fortran. Digital Equipment Corporation, Maynard, MA, USA, 1982. various pp.

DEC:1982:VFI

[Dig82b] Digital Equipment Corporation.

VAX-11 FORTRAN installation guide/release notes. Digital Equipment Corporation, Maynard, MA, USA, 1982. vi + 17 pp.

DEC:1982:VFLb

[Dig82c] Digital Equipment Corporation. VAX-11 FORTRAN language reference manual. Digital Equipment Corp., Maynard, MA, USA, microfiche edition, 1982. various pp.

DEC:1982:VFLa

[Dig82d] Digital Equipment Corporation. VAX-11 Fortran: language summary. Digital Equipment Corporation, Maynard, MA, USA, 1982. 40 pp.

DEC:1982:VFU

[Dig82e] Digital Equipment Corporation. VAX-11 FORTRAN user’s guide. Digital Equipment Corporation, Maynard, MA, USA, 1982. xv + 238 pp.

DEC:1982:VFV

[Dig82f] Digital Equipment Corporation. VAX-11 FORTRAN V3.0. Digital Equipment Corp., Maynard, MA, USA, 1982. ???? pp.

DEC:1983:TFL

[Dig83] Digital Equipment Corporation. TOPS-10/ 20 FORTRAN language manual: preliminary draft. Digital Equipment Corp., Maynard, MA, USA, 1983. ???? pp.

DEC:1984:GPV

[Dig84a] Digital Equipment Corporation. Guide to programming on VAX/ VMS (FORTRAN edition). Digital Equipment Corp., Maynard, MA, USA, revised edition, 1984. various pp.

DEC:1984:MFPa

[Dig84b] Digital Equipment Corporation.

MicroVMS FORTRAN programmer’s manual. Digital Equipment Corp., Maynard, MA, USA, 1984. various pp.

DEC:1984:MFPb

[Dig84c] Digital Equipment Corporation.

MicroVMS FORTRAN programmer’s primer. Digital Equipment Corp., Maynard, MA, USA, revised edition, 1984. various pp.

DEC:1984:PVF

[Dig84d] Digital Equipment Corporation.

Programming in VAX FORTRAN. Digital Equipment Corporation, Maynard, MA, USA, 1984. various pp.

DEC:1984:VFI

[Dig84e] Digital Equipment Corporation.

VAX-11 FORTRAN installation guide/release notes. Digital Equipment Corporation, Maynard, MA, USA, 1984. vi + 20 pp.

DEC:1984:VF

[Dig84f] Digital Equipment Corporation.

VAX FORTRAN. Digital Equipment Corp., Maynard, MA, USA, 1984. ???? pp.

DEC:1984:VFL

[Dig84g] Digital Equipment Corporation.

VAX FORTRAN language summary. Digital Equipment Corporation, Maynard, MA, USA, revised edition, 1984. 50 pp.

DEC:1984:VFUa

[Dig84h] Digital Equipment Corporation. VAX FORTRAN user’s guide. Digital Equipment Corporation, Maynard, MA, USA, microfiche edition, 1984. various pp.

DEC:1984:VFUb

[Dig84i] Digital Equipment Corporation. VAX FORTRAN user’s guide. Digital Equipment Corporation, Maynard, MA, USA, revised edition, 1984. various pp.

DEC:1984:VFV

[Dig84j] Digital Equipment Corporation. [VAX FORTRAN V4.1 documentation kit.]. Digital Equipment Corp., Maynard, MA, USA, revised edition, 1984. 10 pp.

DigitalResearch:1984:F

[Dig84k] Digital Research. FORTRAN 77, 1984.

DEC:1985:FVF

[Dig85a] Digital Equipment Corporation. FORTRAN-10/ 20 and VAX FORTRAN compatibility manual. Digital Equipment Corp., Marlboro, MA, USA, revised edition, 1985. various pp.

DEC:1985:MPR

[Dig85b] Digital Equipment Corporation. MicroVMS primers and release notes: includes User’s primer, Release notes, version 4.0 … 4.1 … 4.1M, FORTRAN programmer’s primer. Digital Equipment Corp., Merrimack, NH, USA, 1985. 5 v. in 1 pp.

DEC:1985:VFI

[Dig85c] Digital Equipment Corporation.

VAX FORTRAN installation guide/release notes. Digital Equipment Corporation, Maynard, MA, USA, 1985. various pp.

DEC:1985:VLE

[Dig85d] Digital Equipment Corporation.

VAX language-sensitive editor; VAX FORTRAN pocket guide. Digital Equipment Corporation, Maynard, MA, USA, 1985. various pp.

DEC:1985:VCR

[Dig85e] Digital Equipment Corporation.

VAXELN: C run-time library reference manual, FORTRAN programmer’s guide. Digital Equipment Corporation, Maynard, MA, USA, 1985. various pp.

DEC:1986:VFUc

[Dig86a] Digital Equipment Corporation.

VAX FORTRAN for ULTRIX systems. Digital Equipment Corp., Maynard, MA, USA, 1986. 2 v. pp.

DEC:1986:VFUa

[Dig86b] Digital Equipment Corporation.

VAX FORTRAN for ULTRIX systems language reference manual. Digital Equipment Corporation, Maynard, MA, USA, 1986. various pp.

DEC:1986:VFUb

[Dig86c] Digital Equipment Corporation.

VAX FORTRAN for ULTRIX systems user manual. Digital Equipment Corporation, Maynard, MA, USA, 1986. various pp.

DEC:1986:VVF

[Dig86d] Digital Equipment Corporation. VAX/VMS (FORTRAN edition) guide to programming. Digital Equipment Corporation, Maynard, MA, USA, microfiche edition, 1986. 8 microfiche. pp.

DEC:1988:VFL

[Dig88a] Digital Equipment Corporation. VAX-11 FORTRAN language reference manual. Digital Equipment Corporation, Maynard, MA, USA, revised edition, 1988. various pp.

DEC:1988:VF

[Dig88b] Digital Equipment Corporation. VAX FORTRAN. Digital Equipment Corp., Maynard, MA, USA, 1988. 2 v. pp.

DEC:1988:VFU

[Dig88c] Digital Equipment Corporation. VAX FORTRAN user manual. Digital Equipment Corporation, Maynard, MA, USA, revised edition, 1988. various pp.

Dillman:1985:PSF

[Dil85] Richard W. Dillman. Problem Solving With Fortran. Holt, Reinhart, and Winston, New York, NY, USA, January 1985. ISBN 0-03-063734-1 (paperback). xiii + 354 pp. LCCN QA76.73.F25 D53 1985. US$27.93. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0030637341.

Dirksen:1981:WMT

[Dir81] P. H. Dirksen. Waterloo Microfortran: Tutorial and Reference Manual. Howard W. Sams, Indianapolis, IN 46268, USA, December 1981. ISBN 0-672-21904-2. ???? pp. LCCN ????

Dirksen:1984:WFU

[Dir84] Paul Dirksen. WATCOM FOR-TRAN user’s guide for the CEM-CORP ICON with QNX. WAT-COM Publications, Waterloo, Ontario, Canada, 1984. ISBN 0-919884-66-0. vi + 32 pp.

VPISUDEF:1985:NII

[Div85] Division of Engineering Fundamentals. Notes on the IBM-PC and IBM DOS: with EDLIN, DVED, and MS FORTRAN 77. Technical report, Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1985. 65 pp.

Dixon:1985:SSS

[Dix85] F. J. Dixon. Simplifying screen specifications — the ‘Full Screen Manager’ interface and ‘Screen Form’ generating routines. The Computer Journal, 28(2):117–127, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Domich:1987:IRS

[DJM87] Paul D. Domich, Richard Henry Frymuth Jackson, and Marjorie A. McClain. The Internal Revenue Service post-of-duty location modeling system: programmer’s manual for Fortran driver. NBSIR 86-3473, U.S. Dept. of Commerce, National Bureau of Standards,; National Technical Information Service, Gaithersburg, MD, USA, 1987. 50 pp.

Dolk:1984:KRM

[DK84] D. R. Dolk and B. R. Konsynski. Knowledge representation for model management systems. IEEE Transactions on Software Engineering, SE-10 (6):619–628, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010291.

Dhaliwal:1989:PFSa

[DKG89a] R. S. (Ranjit S.) Dhaliwal, S. (Sudhir) Kumar, and S. K. (Subodh K.) Gupta. Programming With Fortran 77: a structured approach. Wiley Eastern Limited, ????, 1989. ISBN 81-224-0094-9. xii + 408 + 5 pp. LCCN ???? US$45.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
8122400949.

Dhaliwal:1989:PFSb

[DKG89b] Ranjit S. Dhaliwal, Sudhir Kumar, and Subodh K. Gupta. Programming with FORTRAN 77: a structured approach. Halsted Press, New York, USA, December 1989. ISBN 0-470-21356-6, 81-224-0094-9 (India). xii + 408 + 5 pp. LCCN ???? US$47.95. URL http:

//www.cbooks.com/sqlnut/SP/

search/gtsumt?source=&isbn=
0470213566.

Dyck:1984:FISa

[DLS84a] V. A. Dyck, John Douglas Lawson, and J. A. Smith. Fortran/77: An Introduction to Structured Problem Solving. Reston Publishing Co., Inc., Reston, VA, USA, February 1984. ISBN 0-8359-3163-3. xxii + 696 pp. LCCN QA76.73.F25 D91 1984. US$35.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835931633.

Dyck:1984:FISb

[DLS84b] V. A. Dyck, John Douglas Lawson, and J. A. Smith. FORTRAN/ 77 an introduction to structured problem solving, 1984.

Dongarra:1984:EPS

[DM84] J. J. Dongarra and C. B. Moler.

EISPACK — A package for solving matrix eigenvalue problems. In Cowell [Cow84], pages 68– 87. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

Delves:1985:DDS

[DM85] L. M. Delves and S. C. Mawdsley.

DAP-Algol: a development system for parallel algorithms. The Computer Journal, 28(2):148–153, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

DiDonato:1987:FSC

[DM87a] A. R. DiDonato and A. H. Morris Jr. Fortran subroutines for computing the incomplete gamma function ratios and their inverse. ACM Transactions on Mathematical Software, 13(3):318–319, September 1987. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

DiDonato:1987:AFS

[DM87b] Armido R. DiDonato and Alfred H. Morris Jr. Algorithm 654: FORTRAN subroutines for computing the incomplete gamma function ratios and their inverse. ACM Transactions on Mathematical Software, 13(3):318–319, September 1987. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/pdf/
10.1145/29380.214348.

Dorn:1987:NMF

[DM87c] William S. Dorn and Daniel D. McCracken. Numerical Methods With Fortran IV: Case Studies. Robert E. Krieger Publishing Company, Huntington, NY, USA, January 1987. ISBN 0-89874-982-4. xii + 447 pp. LCCN ???? US$56.65. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0898749824.

Dixon:1989:PTN

[DM89a] Lawrence C. W. Dixon and Z. Maany. The performance of the truncated Newton, conjugate gradient algorithm in Fortran and Ada. Technical Report NOC TR210, The Numerical Optimisation Center, Hatfield Polytechnic, Hatfield, UK, June 1989. ?? pp.

Drouffe:1989:FCT

[DM89b] J.-M. Drouffe and K. J. M. Moriarty. FORTRAN code for the three-dimensional Ising model on bcc and fcc lattices. Computer Physics Communications, 52(2): 249–259, January/February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465589900088.

Deen:1981:DCD

[DNV81] S. M. Deen, D. Nikodem, and A. Vashishta. Design of a canonical database system (PRECI). The Computer Journal, 24(3):200–209, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Driscoll:1986:CAY

[DO86] Graham C. Driscoll, Jr. and Donald L. Orth. Compiling APL: the Yorktown APL translator. IBM Journal of Research and Development, 30(6):583–593, November 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Dobes:1985:FFM

[Dob85] K. Dobes. Formal — a Fortran memory allocation system. Computer Physics Communications, 36(2):147–159, April 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465585901201.

Dodson:1983:CRB

[Dod83] David S. Dodson. Corrigendum: Remark on “Algorithm 539: Basic Linear Algebra Subroutines for FORTRAN usage”. ACM Transactions on Mathematical Software, 9(1):140, March 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [LHKK79a, DG82, HK87a, LN88b].

Doeppner:1988:TTF

[Doe88] Thomas W. Doeppner, Jr. A threads tutorial: Fortran version. Technical report CS-88-19, Brown University, Dept. of Computer Science, Providence, RI, USA, 1988. 29 pp.

Doherty:1982:EF

[Doh82] G. Doherty. Elements of FOR-TRAN. Technical report, Dept. of Mathematics, University of Wollongong, Wollongong, NSW, Australia, 1982. ISBN 0-909716-79-X. 78 + A34 pp.

Dolan:1988:FD

[Dol88] Kent P. Dolan. Fortran 8X discussion. ACM SIGPLAN FOR-TRAN Forum, 7(1):17–27, April 1988. ISSN 1061-7264 (print), 1931-1311 (electronic).

Don:1981:ICI

[Don81] Cassel Don. Introduction to Computers and Information Processing: With Basic, Cobol, Fortran, Pascal. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1981. ISBN 0-8359-3150-1. ???? pp. LCCN ???? US$28.67.

Dongara:1982:ASF

[Don82a] J. J. Dongara. Algorithm 589 SICEDR: A FORTRAN subroutine for improving the accuracy of computed matrix eigenvalues. ACM Transactions on Mathematical Software, 8:371–375, 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1982:ASF

[Don82b] Jack J. Dongarra. Algorithm 589:

SICEDR: A FORTRAN subroutine for improving the accuracy of computed matrix eigenvalues. ACM Transactions on Mathematical Software, 8(4):371–375, December 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1983:RLA

[Don83a] J. J. Dongarra. Redesigning linear algebra algorithms. Bull. Dir. Etud. Rech. C (France), Bulletin de la Direction des Etudes et Recherches, Serie C, C(1):51–60 (or 51–59??), ???? 1983. CODEN EDBCAA. ISSN 0013-4511.

Dongarra:1983:PVC

[Don83b] Jack J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ACM SIGARCH Computer Architecture News, 11(5):22–27, December 1983. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Donigian:1983:HPA

[Don83c] Anthony S. Donigian. HSPF parameter adjustments to evaluate the effects of agricultural best management practices. Technical report, U.S. Environmental Protection Agency, Office of Research and Development, Environmental Research Laboratory, Athens, GA, USA, 1983. 2 microfiches pp.

Dongarra:1984:PVCa

[Don84a] J. J. Dongarra. Performance of various computers using standard linear equations software in a FORTRAN environment. Technical Memorandum 23, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1984. 8 pp.

Dongarra:1984:PVCc

[Don84b] J. J. Dongarra. Performance of various computers using standard linear equations software in a FORTRAN environment. Technique et Science Informatiques, 3(5):317–320, 1984. CODEN TTSIDJ, TTSIEK. ISSN 0752-4072, 0264-7419. URL http:

//www.netlib.org/utk/people/
JackDongarra/PAPERS/Performance-
of-Various-Computers-Using-
Linear-Equations-Software.pdf.

Dongarra:1984:PCO

[Don84c] J. J. Dongarra. Performances comparèes de 80 ordinateurs sur des programmes Fortran. Technique et Science Informatiques, 3(5):355–360, 1984. CODEN TTSIDJ. ISSN 0752-4072, 0264-

7419. URL http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/Performances-Comparees-
de-80-Ordinateurs-sur-des-Programmes-
Fortran.pdf.

Dongarra:1984:PVCb

[Don84d] Jack J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ACM SIGNUM Newsletter, 19(1):23–26, January 1984. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1985:PVCa

[Don85a] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ANL/MCS-TM–23. Argonne National Laboratory. Mathematics and Computer Science Division. Technical memorandum; no. 23. MCA-TM-23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1985. 14 pp. For sale by the National Technical Information Service.

Dongarra:1985:PVCc

[Don85b] J. J. Dongarra. Performance of various computers using standard linear equations software in a FORTRAN environment. ACM SIGARCH Computer Architecture News, 13(1):3–11, March 1985. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Dongarra:1987:PVCa

[Don87a] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. In Walter J. Karplus, editor, Multiprocessors and array processors: proceedings of the Third Conference on Multiprocessors and Array Processors: 14–16 January 1987, San Diego, California, pages 15–32. Society for Computer Simulation, San Diego, CA, USA, 1987. ISSN 0735-9276. LCCN QA76.5 .C61923.

Dongarra:1987:PVCb

[Don87b] Jack J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Simulation, 49 (2):51–62, August 1987. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Dongarra:1988:LBE

[Don88a] J. J. Dongarra. The LIN-PACK benchmark: an explanation. In Evaluating Supercomputers: Strategies for Exploiting, Evaluating and Benchmarking Computers with Advanced Architecture, pages 150–167. Unicom Seminars, Uxbridge, UK, ???? 1988. ISBN ???? LCCN ????

Dongarra:1988:PVCa

[Don88b] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Technical Report MCS-TM-23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, January 1988. ???? pp.

Dongarra:1988:PVCb

[Don88c] Jack J. Dongarra. Performance of various computers using standard linear equations software in a FORTRAN environment. ACM SIGARCH Computer Architecture News, 16(1):47–69, March 1988. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Dongarra:1989:PVC

[Don89] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Technical Memorandum 23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June 4, 1989.

Didday:1981:FH

[DP81] Richard L. Didday and Rex L.

Page. Fortran for Humans. West Publishing Company, St. Paul, MN, USA, third edition, March 1981. ISBN 0-8299-0356-9. xiv + 480 pp. LCCN QA76.73.F25D52 1981. US$32.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0829903569.

Didday:1984:FHb

[DP84a] Rich Didday and Rex Page. Fortran for Humans. West Information Publishing Group, ????, fourth edition, March 1984. ISBN 0-8299-0356-9. xiv + 480 pp.

LCCN ???? US$32.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0829903569.

Didday:1984:FHa

[DP84b] Richard L. Didday and Rex L. Page. FORTRAN for humans. West Publishing Company, St. Paul, MN, USA, fourth edition, March 1984. ISBN 0-314-77887-X (paperback), 0-314-77891-8 (International edition: paperback). xvi + 488 pp. LCCN QA76.73.F25 D52 1984. US$16.95; US$38.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
031477887X.

Didday:1984:IMA

[DP84c] Richard L. Didday and Rex L. Page. Instructor’s manual to accompany FORTRAN for humans. West Publishing Company, St. Paul, MN, USA, fourth edition, 1984. ISBN 0-314-77889-6 (paperback). iii + 65 pp.

Duke:1987:UML

[DPA87] Eugene L. Duke, Brian P. Patterson, and Robert F. Antoniewicz. User’s manual for LINEAR, a FORTRAN program to derive linear aircraft models. NASA technical paper 2768, National Aeronautics and Space Administration, Scientific and Technical Information Division, Washington, DC, USA, 1987. ???? pp.

Duff:1982:MSF

[DR82] I. S. Duff and J. K. Reid. MA27— A set of Fortran subroutines for solving sparse symmetric sets of linear equations. Technical Report R.10533, AERE, Harwell, England, Oxon, UK, 1982. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from math.liu.se

in pub/references.

Duncan:1986:RUM

[DR86] Anthony Duncan and Ralph Roskies. Representations of unusual mathematical structures in scientific applications of symbolic computation. Journal of Symbolic Computation, 2(2):201–206 (or 201–207??), June 1986. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Rijk:1987:NFL

[dR87] P. P. M. de Rijk. NUMVEC FORTRAN library manual: chapter, simultaneous linear equations: routine, SVDTJP and LSQUMNS. Centrum voor Wiskunde en Informatica Report NM-R8719, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, 1987. 4 + 5 pp.

Dreyfus:1981:FI

[Dre81] Michel Dreyfus. Fortran IV.

Dunod informatique: phase formation. Dunod, Paris, France, 5e edition, 1981. ISBN 2-04-007726-X. xii + 228 pp.

Dulikravich:1982:CFP

[DS82] Djordje Stevo Dulikravich and Helmut Sobieczy. CAS22 — FOR-TRAN program for fast design and analysis of sock-free airfoil cascades using fictitious-gas concept. NASA contractor report 3507, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1982. 56 pp. For sale by the National Technical Information Service.

Dongarra:1984:LPS

[DS84] J. J. Dongarra and G. W. Stewart. LINPACK — A package for solving linear systems. In Cowell [Cow84], pages 20–48. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

Dongara:1986:STD

[DS86a] Jack J. Dongara and Danny C. Sorensen. SCHEDULE; tools for developing and analyzing parallel Fortran programs. Argonne National Laboratory. Mathematics and Computer Science Division. Technical Memorandum 86, Argonne National Laboratory, Mathematics and Computer Science Division, Argonne, IL, USA, 1986. 32 pp.

Dongarra:1986:STD

[DS86b] J. Dongarra and D. Sorensen. SCHEDULE: Tools for developing and analyzing parallel Fortran programs. Technical Memorandum ANL-MCS-TM-86, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1986.

Dongarra:1987:PED

[DS87a] J. J. Dongarra and D. C. Sorensen. A portable environment for developing parallel FORTRAN programs. Parallel Computing, 5(1– 2):175–186, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dongarra:1987:STD

[DS87b] J. J. Dongarra and D. C. Sorensen.

Schedule: A tool for developing and analyzing parallel Fortran programs. In Jamieson et al. [JGD87], pages 363–394 (??). ISBN 0-262-10036-3. LCCN QA76.6 .C42981 1987.

Dongarra:1988:PMP

[DSCP88] J. J. Dongarra, D. C. Sorensen, K. Connolly, and J. Patterson. Programming methodology and performance issues for advanced computer architectures. Parallel Computing, 8(1–3):41–58, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dubois:1984:IAF

[Dub84] Philippe Dubois. Initiation au Fortran par l’exemple. Publications de l’Institut français du petrole. Ed. Technip, Paris, France, 2e entièrement revue et corrigée edition, 1984. ISBN ???? 136 pp. LCCN ????

Dulikravich:1981:CFP

[Dul81] Djordje S. Dulikravich. CAS2D — FORTRAN program for nonrotating blade-to-blade, steady potential transonic cascade flows. NASA technical paper 1705, National Aeronautics and Space Administration, Scientific and Technical Information Office, Washington, DC, USA, 1981. 33 pp. For sale by the National Technical Information Service.

Dunham:1980:FPD

[Dun80] Charles B. Dunham. A Fortran program for discrete nonlinear Chebyshev approximation. Journal of Computational and Applied Mathematics, 6(3):241–245, 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Dunn:1985:UPF

[Dun85a] H. J. Dunn. UNAERO, a package of FORTRAN subroutines for approximating unsteady aerodynamics in the time domain. NASA technical memorandum 86392, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1985. ???? pp. For sale by the National Technical Information Service.

Dunn:1985:EFM

[Dun85b] Walter L. Dunn. Engr. 141 FOR-TRAN manual: example computer problems processed on the U.W. VAX system, suggestions for good FORTRAN sytle and readability, and summary of FORTRAN statements. ASUW Publishing, Seattle, WA, USA, 1985. ISBN ???? 180 pp. LCCN ????

Dunigan:1986:DHM

[Dun86] T. W. Dunigan. Denelcor HEP multiprocessor simulator. Report ORNL/TM-9971, Engineering Physics and Mathematics Division, Mathematical Sciences Section, Oak Ridge National Laboratory, Oak Ridge, TN 37831, USA, June 1986. iv + 22 pp. URL http:

//web.ornl.gov/info/reports/
1986/3445600639931.pdf.

Dunham:1987:FPN

[Dun87a] Charles Burton Dunham. A FORTRAN program for nonlinear Chebyshev approximation on an interval. Technical report 179, Dept. of Computer Science, University of Western Ontario, London, UK, 1987. ISBN 0-7714-0918-4. 6 + [6] pp.

Dunham:1987:PDF

[Dun87b] Charles Burton Dunham. A public domain Fortran program for the Fraser–Hart rational Remez algorithm. Technical report 187, Dept. of Computer Science, University of Western Ontario, London, UK, 1987. ISBN 0-7714-0942-7. 12 pp.

Dunham:1988:FPBb

[Dun88a] Charles Burton Dunham. Fortran program for biased rational approximation. Technical report; #202. Dept. of Computer Science, University of Western Ontario, London, UK, 1988. ISBN 0-7714-0978-8. 13 pp.

Dunham:1988:FPO

[Dun88b] Charles Burton Dunham. FOR-TRAN program for one-sided (from above) rational approximation. Technical report 201, Dept. of Computer Science, University of Western Ontario, London, UK, 1988. ISBN 0-7714-0976-1. 15 pp.

Duris:1980:AFR

[Dur80] Charles S. Duris. Algorithm 547: FORTRAN routines for discrete cubic spline interpolation and smoothing [E1], [E3]. ACM Transactions on Mathematical Software, 6(1):92–103, March 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Durgin:1983:F

[Dur83a] Max W. Durgin. Fortran 77. National Publishers of the Black Hills, Rapid City, SD, USA, 1983. ISBN 0-935920-04-8. 316 pp.

Durgin:1983:SF

[Dur83b] Max W. Durgin. Subset Fortran 77. National Publishers of the Black Hills, Rapid City, SD, USA, 1983. ISBN 0-935920-11-0. 316 pp.

Dodrill:1981:FMC

[DV81] William H. Dodrill and Gretchen L. Van Meer. FORTRAN multiple choice and programming exercises. Computing Technology Associates, Morgantown, WV, USA, 1981. ISBN ???? v + 224 pp. LCCN ????

Dirksen:1983:WFT

[DW83a] Paul Dirksen and James William Welch. WATCOM FORTRAN: tutorial and reference manual. WAT-COM / WATFAC Series in computer science and computer applications. WATCOM Publications, Waterloo, Ontario, Canada, revised Waterloo microFORTRAN edition, 1983. ISBN 0-919884-24-5. ix + 196 pp.

Dirksen:1983:WFU

[DW83b] Paul Dirksen and James William Welch. WATCOM FORTRAN users’ guide for the IBM 370 with VM/SP CMS. WATCOM Publications, Waterloo, Ontario, Canada, 1983. ISBN 0-919884-07-5. v + 25 pp.

Dirksen:1984:WFT

[DW84] Paul Dirksen and James William Welch. WATCOM FORTRAN: tutorial and reference manual. WAT-COM series in computer science and computer applications. WAT-COM Publications, Waterloo, Ontario, Canada, 1984. ISBN ???? ix + 196 pp. LCCN ????

Cowan:1985:WFU

[DW85] Donald D. Cowanand Paul Dirksen and James William Welch. WAT-COM FORTRAN user’s guide: IBM PC with DOS. WATCOM series in computer science and computer applications. WATCOM Publications, Waterloo, Ontario, Canada, 1985. ISBN 0-919884-14-8. vi + 52 pp.

Dyck:1981:IC

[Dyc81] V. A. Dyck. Introduction to Computing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, October 1981. ISBN 0-8359-3158-7. ???? pp. LCCN ???? US$27.67.

Zeeuw:1986:NFL

[dZ86] P. M. de Zeeuw. NUMVEC FOR-TRAN library manual: chapter, elliptic PDEs; routine, MGD1V and MGD5V. Report. Centrum voor Wiskunde en Informatica; NM-R8624 Report. Centre for Mathematics and Computer Science. Dept. of Numerical Mathematics; NM-R8624 Centrum voor Wiskunde en Informatica (Amsterdam, Netherlands). Report; NM-R8624. Stichning Mathematisch Centrum, Amsterdam, The Netherlands, 1986. ISBN ???? 14 pp. LCCN ????

Dunham:1987:FPD

[DZ87] Charles Burton Dunham and C. Z. Zhu. A FORTRAN program for discrete nonlinear biased approximation. Technical report 181, Dept. of Computer Science, University of Western Ontario, London, UK, 1987. ISBN 0-7714-0919-2. 11 pp.

Dunham:1988:FPBa

[DZ88] Charles Burton Dunham and C. Z. Zhu. A FORTRAN program for biased uniform approximation on an interval. Technical report, #197 0707-235X. Dept. of Computer Science, University of Western Ontario, London, UK, 1988. ISBN 0-7714-0971-0. 21 pp.

Ekanadham:1987:SPE

[EA87] Kattamuri Ekanadham and Arvind. SIMPLE: Part 1 — An Exercise in Future Scientific Programming. Computation Structures Group Memo 273, Laboratory for Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, July 1987.

Earnest:1985:GDS

[Ear85] B. M. Earnest. Guidelines for developing structured FORTRAN programs. NASA technical memorandum 86407, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1985. ???? pp.

Ernenwein:1988:VSC

[EBS88] René Ernenwein, Marc Benard, and Isaiah Shavitt. Vectorizing a sequence of conditional branches: The calculation of the class index of two-electron repulsion integrals on Cray computers. Computer Physics Communications, 48(2):175–180, February 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465588900380.

Edgar:1989:APSa

[Edg89a] Stacey L. Edgar. Advanced problem solving with FORTRAN 77, including a preview of FORTRAN 8X. Science Research Associates, Chicago, IL, USA, 1989. ISBN 0-574-18735-9. xv + 445 pp. LCCN QA 76.73 F25 E34 1989. US$28.00.

Edgar:1989:APSb

[Edg89b] Stacey L. Edgar. Advanced Problem Solving With Fortran 77: Including a Preview of Fortran 8X. MacMillan Publishing Company, New York, NY, USA, January 1989. ISBN 0-02-331484-2. ???? pp. LCCN ???? US$39.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0023314842.

Edmunds:1986:BRB

[Edm86] G. Edmunds. Book review: Essentials of Fortran 77, John Shelley, Wiley, Chichester. 1984. no of pages: 218. price: £8.25. Software —Practice and Experience, 16(1): 99–100, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ekman:1984:PF

[EE84] Torgil Ekman and Göran Eriksson. Programmering i Fortran 77. Studentlitteratur, Lund, Sweden, 1984. ISBN 91-44-16663-X. ???? pp. LCCN ????

Egger:1983:ALG

[Egg83] Thomas J. Egger. Assembly language graphics library (EGR1.LB) with FORTRAN interfacing. NOAA, National Weather Service, Central Region computer programs and problems NWS CRCP 9, U.S. Dept. of Commerce, National Oceanic and Atmospheric Administration, National Weather Service, Washington, DC, USA, 1983. 91 pp.

Etlinger:1981:FSP

[EGP81] Henry A. Etlinger, Gordon I. Goodman, and Charles Plummer. FORTRAN: a self-paced, masterybased course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 13 (1):62–73, February 1981. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th SIGCSE Symposium on Computer Science Education.

El-Hady:1981:HFP

[EH81] Nabil M. El-Hady. HADY-1, a FORTRAN program for the compressible stability analysis of three-dimensional boundary layers. NASA contractor report 3467, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1981. 114 pp. For sale by the National Technical Information Service.

Ehinger:1982:FAB

[Ehi82] Gerd Ehinger. FORTRAN-Lexikon: Anweisungen und Begriffe: FORTRAN IV, DIN 60,027, TR 440, IBM/ 360,/370, CDC FTN 4, DEC IV-PLUS: FORTRAN 77, ANSI X3.9- 1978, SIEMENS FOR 1, CDC FTN 5. Walter de Gruyter, New York, NY, USA, 1982. ISBN 3-11-008359-0. xix + 492 pp.

Endo:1985:TFP

[EIT85] Elliot T. Endo, Gene Y. Iwatsubo, and Lyn J. Topinka. Two FOR-TRAN programs for the reduction of atmospheric data collected by aircraft along EDM survey lines. Open-file report 85-279, U.S. Dept. of the Interior, Geological Survey, Reston, VA, USA, 1985. 39 pp.

Elhay:1987:IFS

[EK87a] S. Elhay and J. Kautsky. IQPACK:

Fortran subroutines for the weights of interpolatory quadratures. ACM Transactions on Mathematical Software, 13(4):399–415, December 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Elhay:1987:AIF

[EK87b] Sylvan Elhay and Jaroslav Kautsky. Algorithm 655: IQPACK: FORTRAN subroutines for the weights of interpolatory quadratures. ACM Transactions on Mathematical Software, 13(4): 399–415, December 1987. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/toc/Abstracts/0098-
                                        3500/214351.html.

Eisenstat:1981:BAF

[EL81] S. C. Eisenstat and J. W. Lewis. Band arrays in Fortran: an elegant representation. ACM SIGNUM Newsletter, 16(3):9–11, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Elliott:1981:FSD

[Ell81a] D. G. Elliott. FORTRAN 77 and structured design. ACM SIG-PLAN Notices, 16(12):7–9, December 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ellis:1981:SAT

[Ell81b] Keith Charles Ellis. A syntax analyzer for TIDY 81 — a FORTRAN 77 preprocessor. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1981. vi + 234 pp.

Elliott:1982:HDP

[Ell82a] B. Elliott. A high-level debugger for PL/I, FORTRAN and BA-SIC. Software—Practice and Experience, 12(4):331–340, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Elliott:1982:HLD

[Ell82b] Brig Elliott. A high-level debugger for PL/I, FORTRAN and BA-SIC. Software—Practice and Experience, 12(4):331–340, April 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ellis:1982:GDFa

[Ell82c] Margaret A. Ellis. Global data flow analysis in a Fortran 77 compiler. Electrical engineering and computer sciences master of science report., University of California, Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1982. 18 pp.

Ellis:1982:GDFb

[Ell82d] Margaret A. Ellis. Global data flow analysis in a Fortran 77 compiler. Master of science, plan ii., University of California, Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1982. ???? pp.

Ellis:1982:SAF

[Ell82e] T. M. R. Ellis. A structured approach to FORTRAN 77 programming. International computer science series. Addison-Wesley, Reading, MA, USA, 1982. ISBN 0-201-13790-9 (paperback). x + 350 pp. LCCN QA76.73 F25 E44 1982 Sci-Eng. US$7.95.

Ellis:1983:SAF

[Ell83]  T. Ellis.  A Structured Ap-

proach to Fortran 77 Programming. Addison-Wesley, Reading, MA, USA, January 1983. ISBN 0-201-13790-9. ???? pp. LCCN ???? US$23.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201137909.

Elsner:1982:FSP

[Els82] Manfred Elsner. Fortran-preprozessor für Strukturierte Programme. Master’s thesis, Technische Universität Braunschweig (??), Braunschweig, Germany, 1982.

Eoyang:1988:BSG

[EML88] C. Eoyang, R. H. Mendez, and O. M. Lubeck. The birth of the second generation: the Hitachi S-820/ 80. In IEEE [IEE88b], pages 296– 303. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Engeln-muellges:1984:FNM

[EmR84] G. Engeln-muellges and F. Reutter. Formelsammlung zur Numerischen Mathematik mit Standard-Fortran-programmen. Bibliographisches Institut, Mannheim, Germany, 4 edition, 1984. ISBN 3-411-01677-9.

Encore:1987:FM

[Enc87] Encore Computer Corporation.

Fortran-77 manual. Encore Computer Corp., Marlboro, MA, USA, 1987. 192 pp.

Evett:1981:FP

[EP81a] Jack B. Evett and Richard P.

Pinckney. Fortran Programming. Engineering Press, San Jose, CA, USA, second edition, 1981. ISBN 0-910554-32-3. xi + 209 pp. LCCN QA76.73.F25E93.

US$9.95.  URL http://www.
cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0910554323.

Evett:1981:FPA

[EP81b] Jack B. Evett and Richard P.

Pinckney. Fortran programming with applications to engineering. Engineering Press, San Jose, CA, USA, 1981. ISBN 0-910554-32-3. xi + 209 pp. LCCN QA 76.73 F25 E93 1981. US$4.95.

Enright:1987:TFP

[EP87a] Wayne H. Enright and John D.

Pryce. Two FORTRAN packages for assessing initial value methods. ACM Transactions on Mathematical Software, 13(1):1–27, March 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/27645.html.  See also

[EP89].

Evett:1987:FP

[EP87b] Jack B. Evett and Richard P. Pinckney. Fortran Programming. Engineering Press, San Jose, CA, USA, second edition, January 1987. ISBN 0-910554-65-X (paperback). xi + 209 pp. LCCN QA76.73.F25E93 1987. US$13.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        091055465X.

Enright:1989:CFP

[EP89] W. H. Enright and J. D. Pryce. Corrigenda: “Two FORTRAN packages for assessing initial value methods”. ACM Transactions on Mathematical Software, 15 (3):287, September 1989. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [EP87a].

Ess:1988:SCI

[Ess88] M. Ess. Simple convolution on the Intel iPSC/1 and iPSC/1-VX. Supercomputer, 5(4):22–30, July 1988. CODEN SPCOEL. ISSN 0168-7875.

Estes:1982:DPO

[Est82] George E. Estes. Distinguishing the potential operands in FOR-TRAN programs. ACM SIG-METRICS Performance Evaluation Review, 11(2):115–117, Summer 1982. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Efrat:1986:PIL

[ET86] Ilan Efrat and Miron Tismenetsky.

Parallel iterative linear solvers for oil reservoir models. IBM Journal of Research and Development, 30 (2):184–192, March 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

EtaSystems:1988:FRM

[ETA88] ETA Systems, Inc. FORTRAN 77 reference manual. Publication PUB-1255, ETA Systems, Inc., St. Paul, MN, USA, 1988. ???? pp.

Etter:1983:IGA

[Ett83a] D. M. Etter. Instructor’s guide to accompany Structured Fortran 77 for engineers and scientists. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1983. ISBN 0-8053-2521-2. 365 pp.

Etter:1983:SFE

[Ett83b] D. M. Etter. Structured FOR-TRAN 77 for engineers and scientists. The Benjamin/Cummings series in computing and information sciences. Benjamin/ Cummings Pub. Co., Menlo Park, CA, USA, 1983. ISBN 0-8053-2520-4. xv + 357 pp.

Etter:1984:PSS

[Ett84a] D. M. Etter. Problem Solving With Structured Fortran 77. The Benjamin/Cummings series in structured programming. Benjamin/ Cummings Pub. Co., Menlo Park, CA, USA, March 1984. ISBN 0-8053-2522-0 (paperback). xxii + 416 pp. LCCN QA 76.73 F25 E848 1984. US$33.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0805325220.

Etter:1984:WSP

[Ett84b] D. M. Etter. Watfiv: Structured Programming and Problem Solving. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, August 1984. ISBN 0-8053-2502-6. ???? pp. LCCN ???? US$31.25.

Etter:1985:WSP

[Ett85] D. M. Etter. WATFIV: structured programming and problem solving. The Benjamin/Cummings series in structured programming. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1985. ISBN 0-8053-2502-6 (paperback). xv + 367 pp. LCCN QA 76.73 F25 E87 1985.

Etter:1987:SFE

[Ett87] D. M. Etter. Structured FOR-TRAN 77 for engineers and scientists. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, second edition, 1987. ISBN 0-8053-2495-X (paperback). xxi + 519 pp.

Evans:1981:FCP

[Eva81] John Richard Evans. FORTRAN computer programs for running median filters and a general despiker. Open-file report 81-1091, U.S. Dept. of the Interior, Geological Survey, Reston, VA, USA, 1981. 14 pp.

Evenden:1984:R

[Eve84] Gerald I. Evenden. RATFOR [letter]. Dr. Dobb’s Journal of Software Tools, 9(6):10–??, June 1984.

CODEN DDJOEB. ISSN 1044-789X.

Everitt:1985:BRB

[Eve85] B. Everitt. Book review: Cluster Dissection and Analysis: Theory, Fortran Programs and Examples, by H. Spath. Journal of the Royal Statistical Society. Series A (General), 148(3):285–286, ???? 1985. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.
org/stable/2981981.

Eastwood:1987:ATW

[EW87] James W. Eastwood and Christopher J. H. Watson. An analytic theory of wave-current interactions. Plasma Physics Note 87/7, Culham Laboratory, Theory and Optics Division, Abingdon, Berkshire, UK, February 1987.

Fateman:1982:HLI

[Fat82a] Richard J. Fateman. High-level language implications of the proposed IEEE floating point standard. ACM Transactions on Programming Languages and Systems, 4(2):239–257, April 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fateman:1982:HLL

[Fat82b] Richard J. Fateman. High-level language implications of the proposed IEEE floating-point standard. ACM Transactions on Programming Languages and Systems, 4(2):239–257, April 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fellows:1983:UFR

[FCG83] David M. Fellows, Brian P. Cassidy, and Uday G. Gujar. Update to ‘FORTRAN routines with optional arguments’. Software— Practice and Experience, 13(11): 1093–1094, November 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fondrat:1986:PCQ

[FDL86] C. Fondrat, P. Dessen, and P. Le Beux. Principle of codification for quick comparisons with the entire biomolecule databanks and associated programs in FORTRAN 77. Nucleic Acids Research, 14(1):197– 204, January 10, 1986. CODEN NARHAD. ISSN 0305-1048.

Frane:1982:BPG

[FE82] James W. Frane and Laszlo Engelman. BMDP programmer’s guide and subroutine writeups: documentation of the FORTRAN source. Bmdp technical report; no. 55, BMDP Statistical Software, Dept. of Biomathematics, University of California, Los Angeles, CA, USA, 1982. 449 pp.

FCTC:1981:FCV

[Fed81] Federal Compiler Testing Center (U.S.). FORTRAN compiler validation summary report: Digital Equipment Corporation VAX-II FORTRAN version 2. Validation number FCVS78-VSR806, General Services Administration, Washington, DC, USA, 1981. v + 5 pp.

FSTC:1982:FCV

[Fed82a] Federal Software Testing Center. 1978 Fortran compiler validation system user’s guide. Report OSD/FSTC-82/021 (FCVS 78), Federal Software Testing Center; National Technical Information Service, Falls Church, VA, USA, 1982. various pp.

FSTC:1982:FAU

[Fed82b] Federal Software Testing Center (U.S.). FORTRAN 77 analyzer: user’s manual. Report OSD/FCTC-82/017, Federal Software Testing Center, U.S. General Services Administration, Falls Church, VA, USA, July 1982. 118 pp.

Fenner:1987:ESAb

[Fen87a] D. N. Fenner. Engineering Stress Analysis: a Finite Element Approach With Fortran 77 Software. Ellis Horwood Series in Mechanical Engineering. Ellis Horwood, New York, NY, USA, December 1987. ISBN 0-13-275835-0. ???? pp. LCCN ???? US$64.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0132758350.

Fenner:1987:ESAa

[Fen87b] D. N. (David N.) Fenner. Engineering stress analysis: a finite element approach with FOR-TRAN 77 software. Ellis Horwood series in mechanical engineering. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, 1987. ISBN 0-7458-0246-X, 0-7458-0302-4 (paperback), 0-470-20895-3 (US). 250 pp. LCCN TA347.F5 F458 1987. UK£30.00 (hardcover), UK£14.50 (paperback).

Fishwick:1984:PPG

[FF84] Paul A. Fishwick and Stefan Feyock. PROFGEN: a procedure for generating machine independent high-level language profilers. ACM SIGMETRICS Performance Evaluation Review, 12(2):27–31, Spring-Summer 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fuori:1986:FEPa

[FGGF86] William M. Fuori, Louis Gioia, Stephen Gaughran, and Michael Fuori. Fortran 77: Elements of Programming Style. The Art of programming IBM personal computers. Hayden Book Co., Rochelle Park, NJ, USA, March 1986. ISBN 0-8104-6397-0 (paperback). 358 pp. LCCN QA76.73.F25 F67155 1986. US$22.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0810463970.

Friedmann:1981:FI

[FGH81] Jehosua Friedmann, Philip Greenberg, and Alan Hoffberg. Fortran IV. A Self-teaching guide. John Wiley and Sons, New York, London, Sydney, second edition, 1981. ISBN 0-471-07771-2 (paperback). xii + 499 pp.

Fischer:1982:EPS

[Fis82]  P. Fischer.  Entwicklung des

Programmpakets zur Simulation eines Rdbm-prozessors (datenkonverter). Master’s thesis, Technische Universität Braunschweig (??), Braunschweig, Germany, 1982.

Fisher:1983:SAF

[Fis83] D. L. Fisher. A structured approach to Fortran 77 programming. The Computer Journal, 26(4):381a–381, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
26/4/381-a.full.pdf+html.

Frederickson:1985:SCP

[FJS85] Paul O. Frederickson, Rondall E.

Jones, and Brian T. Smith. Synchronization and control of parallel algorithms. Parallel Computing, 2 (3):255–264, November 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Friedman:1981:PSSb

[FK81] Frank L. Friedman and Elliot B.

Koffman. Problem solving and structured programming in FOR-TRAN. Addison-Wesley series in computer science and information processing. Addison-Wesley, Reading, MA, USA, second edition, 1981. ISBN 0-201-02461-6. 530 pp. LCCN QA76.73.F25F74 1981.

Friedman:1982:PSS

[FK82] Frank L. Friedman and Elliot B.

Koffman. Problem solving and structured programming in WAT-FIV. Addison-Wesley series in computer science and information processing. Addison-Wesley, Reading, MA, USA, February 1982. ISBN 0-201-10482-2 (paperback). xvi + 483 + [56] pp. LCCN QA 76.73 F25 F743 1982. US$37.75; US$15.95.

Friedman:1984:FIA

[FK84] Frank L. Friedman and Elliott B. Koffman. FORTRAN: introduccion al lenguaje y resolucion de problemas con programacion estructurada. Fondo Educativo Interamericano, Mexico, DF, Mexico, 1984. ISBN 968-858-006-6. xvi + 537 pp.

Friedman:1981:PSSa

[FKSS81] Frank L. Friedman, Eliot B. Koffman, Robert Soloman, and Judith O’shea Stebulis. Problem solving and structured programming in Fortran. Addison-Wesley, Reading, MA, USA, second edition, 1981. ISBN ???? 212 pp. LCCN ????

Floyd:1989:MCF

[Flo89] Michael A. Floyd. Making the Cto-Fortran connection. Dr. Dobb’s Journal of Software Tools, 14(8): 22–23, 25–27, 102–104, August 1989. CODEN DDJOEB. ISSN 1044-789X.

Freeborn:1985:MFP

[FMH85] W. Phelps Freeborn, E. S. McGee, and J. S. Huebner. MINCLC: a FORTRAN program for recalculating mineral analyses. Open-file report 85-257, Dept. of the Interior, U.S. Geological Survey, Reston, VA, USA, 1985. i + 45 pp.

Foulk:1985:APN

[FN85] Patrick W. Foulk and Salwa M.

Nassar. Analysis of parallelism in nested do loops. The Journal of Systems and Software, 5(1):73–80, February 1985. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fogiel:1985:HGC

[Fog85] Max Fogiel. Handbook and guide for comparing and selecting computer languages BASIC, FOR-TRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, 1985. ISBN 0-87891-561-3 (paperback). v + 122 pp.

Fogiel:1987:HGC

[Fog87] Max Fogiel. Handbook and guide for comparing and selecting computer languages BASIC, FOR-TRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, revised print. edition, 1987. ISBN ???? v + 122 pp. LCCN ????

Fogiel:1988:HGC

[Fog88] Max Fogiel. Handbook and guide for comparing and selecting computer languages BASIC, FOR-TRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, 1988. ISBN ???? v + 122 pp. LCCN ????

Fong:1985:NCT

[Fon85] Kirby W. Fong. NMFECC Cray time-sharing system. Software— Practice and Experience, 15(1): 87–103, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ford:1982:SFPa

[For82a] Donald H. Ford. Standard Fortran Programming. R. D. Irwin, Homewood, IL, USA, fourth edition, July 1982. ISBN 0-256-01998-3. xv + 330 pp. LCCN QA76.73.F25F664 1978. US$11.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0256019983.

ForTec:1982:FFQ

[For82b] ForTec forum: a quarterly publication of the SigPlan Technical Committee on FORTRAN, page ????, 1982. ISSN 0735-3731. ACM Press, New York, NY 10036, USA.

Forrest:1985:DTF

[For85] Michael Robin Forrest. Development of a Talwani-Ewing Fortran program for gravity interpretation. Thesis (M.S.), California State University, Long Beach, Long Beach, CA, USA, 1985. v + 64 pp.

FUG:1989:FJ

[For89] Fortran journal, page various, 1989. ISSN 1060-0221. Fortran Users Group, Fullerton, CA, USA.

Ford:1982:SFPb

[FR82] Donald H. Ford and Joseph Rue, Ph.D. Standard FORTRAN programming: a structured style. The Irwin series in information and decision sciences. R. D. Irwin, Homewood, IL, USA, fourth edition, July 1982. ISBN 0-256-02608-4 (paperback). xvii + 319 pp. LCCN QA76.73.F25F664 1982. US$20.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0256026084.

Franke:1984:FVS

[Fra84a] Michael Martin Franke. A FOR-TRAN V symbolic derivation of first and second stochastic moment coefficient matrices for linear systems with parametric white noise excitations. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1984. vi + 76 pp.

Franzmeier:1984:FLC

[Fra84b] Nathan Victor Franzmeier. A FORTRAN library for control of the unimate PUMA 600. Computer science thesis (M.S.), Texas A and M University, College Station, TX, USA, 1984. vii + 164 pp.

Freak:1981:FPT

[Fre81] R. A. Freak. A Fortran to Pascal translator. Software—Practice and Experience, 11(7):717– 732, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

FRCC:1983:PGF

[Fre83] Freas-Rooke Computer Center. Programmer’s guide to FOR-TRAN. Frcc user’s manual; section c, Freas-Rooke Computer Center, Lewisburg, PA, USA, 1983. 13 + A1–11 + B1–6 + C1–10 + D1–8 + E1–5 + F1–22 pp.

Fritz:1984:ALC

[Fri84] Jane Manning Fritz. An analysis of the 198O CODASYL report on a FORTRAN data base facility. Thesis (M.Sc.C.S.), University of New Brunswick, Ottawa, Ontario, Canada, 1984. 2 microfiches (125 fr.) pp.

Shih:1982:TNY

[fS82] Po feng Shih. Tien nao yu yen. Hsing yeh tu shu ku fen yu hsien kung ssu, Tai-nan shih, 3 pan edition, 1982. ISBN ???? viii + 12 + 794 + [173] pp. LCCN ????

Fabrikant:1986:AGM

[FS86a] V. I. Fabrikant and T. S. Sankar. An algorithm for geometrical modelling of surfaces of revolution. Computers and Graphics, 10(3):245–255, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Faroult:1986:FSM

[FS86b] Stephane Faroult and Didier Simon. Fortran structure et méthodes numériques. Dunod informatique. Bo-Pre, Saint-Jean-sur-Richelieu, Québec, 1986. ISBN 2-89315-009-8, 2-04-016483-9 (Dunod). xv + 326 pp.

Fosdick:1989:BFA

[FSO89] Lloyd Dudley Fosdick, C. J. C.

Schauble, and Kurt M. Olender. Benchmarking Fortran and Ada programs on parallel machines. Technical report CU-CS-420-89, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, January 1989. 47 pp.

Tai:1987:FCC

[fTBcL7 ] Cheng fang Tai, Michel H. Boillot, and Chi chang Liu. Fu chuan cheng shih yu yen. Hsiang-kang ko chi chu pan she, Chiu-lung, 1987 (??). ISBN ???? 483 pp. LCCN ????

Fullerton:1987:ADF

[Ful87] J. Fullerton. An alternate design for Fortran 8X. ACM SIG-PLAN FORTRAN Forum, 6(3): 31–40, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Fung:1986:DIB

[Fun86] Kin M. Fung. A data interface between Prolog and Fortran. Thesis (M.S.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1986. viii + 93 pp.

Fuori:1986:FPI

[Fuo86a] Bill Fuori. Fortran 77 Programming for the IBM PC and XT. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, April 1986. ISBN 0-8359-2096-8. ???? pp. LCCN ???? US$19.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835920968.

Fuori:1986:FEPb

[Fuo86b] William Fuori. Fortran 77: Elements of Programming Style. Howard W. Sams, Indianapolis, IN 46268, USA, March 1986. ISBN 0-8104-6597-3. ???? pp. LCCN ???? US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0810465973.

Fuori:1986:FEPc

[Fuo86c] William Fuori. Fortran 77: Elements of Programming Style. Howard W. Sams, Indianapolis, IN 46268, USA, teacher’s edition, March 1986. ISBN 0-8104-6697-X. ???? pp. LCCN ???? US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
081046697X.

Futrell:1978:RTA

[Fut78] R. Futrell. Remark on “Fortran translation of Algorithm 409: Discrete Chebychev curve fit [E2]”. ACM Transactions on Mathematical Software, 4(1):95, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Sim76].

Flamm:1982:RHE

[FW82] David S. Flamm and Robert A. Walker. Remark on “Algorithm 506: HQR3 and EXCHNG: Fortran subroutines for calculating and ordering the eigenvalues of a real upper Hessenberg matrix [F2]”. ACM Transactions on Mathematical Software, 8(2):219– 220, June 1982. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Ste76].

Feldman:1983:PFCb

[FW83a] S. I. Feldman and P. J. Weinberger. A portable Fortran 77 compiler. In Unix Programmer’s Manual, Volume II. Holt, Reinhart, and Winston, New York, NY, USA, 1983. See [FW83b].

Feldman:1983:PFCa

[FW83b] Stuart I. Feldman and P. J. Weinberger. A portable Fortran 77 compiler. Technical report, University of California, Berkeley, Berkeley, CA, USA, 1983. 21 pp. Also available in [FW83a].

Yu:1984:CKH

[fY84] Shu feng Yu. Chieh kou hua FOR-TRAN cheng shih she chi. “Tzu hsun chiao yu” hsi lieh; 5. Tzu hsun chiao yu tui kuang chung hsin, Tai-pei shih, ti 1 pan edition, 1984. ISBN ???? 278 pp. LCCN ????

Gabriel:1989:AGD

[Gab89] J. R. Gabriel. Automated generation of databases from a specification and simulation input data. In in Yu(ed) 1989 Workshop on heterogeneous databases (??). ????, ????, December 1989.

Gaffney:1983:FSC

[Gaf83a] P. W. Gaffney. A Fortran subroutine for computing the optimal estimate of f(x). ACM Transactions on Mathematical Software, 9(1):98–116, March 1983. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gaffney:1983:AAF

[Gaf83b] Patrick W. Gaffney. Algorithm 592: A FORTRAN subroutine for computing the optimal estimate of f(x). ACM Transactions on Mathematical Software, 9(1):98– 116, March 1983. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gaffney:1983:AFS

[Gaf83c] Patrick W. Gaffney. Algorithm 592: A FORTRAN subroutine for computing the optimal estimate of f(x). ACM Transactions on Mathematical Software, 9(1):98– 116, March 1983. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gaffney:1984:PES

[Gaf84] Patrick W. Gaffney. A performance evaluation of some FOR-TRAN subroutines for the solution of stiff oscillatory ordinary differential equations. ACM Transactions on Mathematical Software, 10(1):58–72, March 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gallivan:1989:PPL

[Gal89] Kyle A. Gallivan. Performance prediction of loop constructs on multiprocessor hierarchicalmemory systems. Technical Report CSRD 853, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1989. 19 pp.

Gatewood:1982:EFI

[Gat82] Walter Patrick Gatewood. An experimental FORTRAN implementation of HCPRVR. Thesis (M.A.), University of Texas at Austin, Austin, TX, USA, 1982. 115 pp.

Gates:1985:GAC

[Gat85] Barbara L. Gates. GENTRAN: an automatic code generation facility for REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(3):24–42, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gourdin:1983:MNA

[GB83] A. Gourdin and M. Boumahrat.

Méthodes numériques appliquées: (avec nombreux problémes résolus en Fortran IV). Lavoisier/ Technique et Documentation, Paris, France, 1983. ISBN 2-85206-198-8. 422 pp.

Gourdin:1989:MNA

[GB89] A. Gourdin and M. Boumahrat.

Méthodes numériques appliquées: (avec nombreux problèmes résolus en Fortran 77). Technique et Documentation-Lavoisier, Paris, France, 2e edition, 1989. ISBN 2-85206-554-1. xiii + 440 pp.

Gregory:1981:ACR

[GBJ81] W. Gregory, R. Bell, and R. R.

Jackson. An automated clinical research and data analysis system.

Computers and Biomedical Research, 16(5), October 1981. CO-DEN CBMRB7. ISSN 0010-4809 (print), 1090-2368 (electronic).

Gergely:1984:UGD

[GC84] Peter J. Gergely and D. J. Chiasson. A user’s guide to DE-CLIB 2.0: a library of FORTRAN callable utility routines. Technical memorandum 84/R, Defence Research Establishment Atlantic, Dartmouth, NS, Canada, 1984. vi + 26 pp.

Grest:1989:VLC

[GDK89] Gary S. Grest, Burkhard Dünweg, and Kurt Kremer. Vectorized link cell Fortran code for molecular dynamics simulations for a large number of particles. Computer Physics Communications, 55(3):269–285, October 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465589901252.

Gray:1985:PRI

[GE85] P. M. D. Gray and P. Esslemont. Performance of a relational interface to a Codasyl database. The Computer Journal, 28(5):501–507, November 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gear:1986:CAAb

[Gea86a] C. William Gear. Computer Applications and Algorithms. Science Research Associates, Chicago, IL, USA, February 1986. ISBN 0-02-341206-2. ???? pp. LCCN ????

Gear:1986:CAAa

[Gea86b] C. William (Charles William)

Gear. Computer Applications and Algorithms. Science Research Associates, Chicago, IL, USA, February 1986. ISBN 0-574-21970-6 (paperback). x + 244 pp. LCCN QA76.6 .G3751 1986. US$11.95. Revision of the applications and algorithms modules from the Introduction to computers, structured programming, and applications series.

Geenen:1986:LAF

[Gee86] Donald J. Geenen. Learning Apple FORTRAN. Computers and math series; [9]. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, 1986. ISBN 0-88175-024-7 (paperback). xiii + 264 pp. LCCN QA76.8.A6623 G44 1986 Sci-Eng.

Genz:1982:LEA

[Gen82] Alan C. Genz. A Lagrange extrapolation algorithm for sequences of approximations to multiple integrals. SIAM Journal on Scientific and Statistical Computing, 3 (2):160–172, June 1982. CODEN SIJCD4. ISSN 0196-5204.

Gerlach:1983:GSI

[Ger83] H. Gerlach. Graphische Software im CMS GR-Software. Technical report, Interner Bericht KfA Jülich ZENTRAL Zentralinstitut für angewandte Mathematik, Jülich, Germany, 1983. 1–37 pp.

Gujar:1981:FRO

[GF81] Uday G. Gujar and David M. Fellows. FORTRAN routines with optional arguments. Software— Practice and Experience, 11(2): 187–193, February 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gary:1989:OPF

[GF89] J. Gary and L. Fosdick. An optimizing precompiler for finitedifference computations on a vector computer. Parallel Computing, 10(1):51–64, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Goldfarb:1988:CCD

[GG88] D. Goldfarb and M. D. Grigoriadis. A computational comparison of the Dinic and network simplex methods for maximum flow. In B. Simeone et al., editors, FOR-TRAN Codes for Network Optimization, Annals of Operations Research, vol. 13, pages 83–124. ????, ????, 1988.

Gallivan:1989:BCM

[GGJ^{+}89] K. Gallivan, D. Gannon, W. Jalby, A. Malony, and H. Wijshoff. Behavioral characterization of multiprocessor memory systems: a case study. ACM SIGMETRICS Performance Evaluation Review, 17 (1):79–88, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garbow:1988:AFS

[GGLM88] B. S. Garbow, G. Giunta, J. N. Lyness, and A. Murli. Algorithm 662: A FORTRAN software package for the numerical inversion of the Laplace transform based on Weeks’ method. ACM Transactions on Mathematical Software, 14(2):171–176, June 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/214375.html.  See also

[GL90].

Glendinning:1987:TAF

[GH87] Ian Glendinning and Anthony Hey. Transputer arrays as FOR-TRAN farms for particle physcs. Computer Physics Communications, 45(1–3):367–371, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Gill:1986:UGL

[GHM^{+}86] P. E. Gill, S. J. Hammarling, W. Murray, M. A. Saunders, and M. H. Wright. User’s guide for LSSOL (version 1.0): a Fortran package for constrained linear least-squares and convex quadratic programming. Report SOL ??, Dept. of Operations Research, Stanford University, Stanford, CA, USA, 1986.

Gianni:1989:SAC

[Gia89] P. Gianni, editor. Symbolic and Algebraic Computation. International Symposium ISSAC ’88. Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN ????

Gichuki:1988:UMF

[Gic88] Francis N. Gichuki. User’s manual for the Fortran version of USU main system hydraulic model. Wms report; 73, Utah State University, Agricultural and Irrigation Engineering Dept., Logan, UT, USA, 1988. viii + 71 pp.

Gill:1986:UGN

[Gil86] Philip E. Gill. User’s guide for NPSOL (version 4.0): a Fortran package for nonlinear programming. Technical report SOL 86-2, Stanford University, Dept. of Operations Research, Systems Optimization Laboratory, Stanford, CA, USA, 1986. 54 pp.

Gini:1982:ASI

[Gin82] Guiseppina Gini. The automatic synthesis of iterative programs. Information Processing Letters, 14 (2):67–73, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghezzy:1982:PLC

[GJ82] C. Ghezzy and M. Jazayeri. Programming Language Concepts. John Wiley and Sons, New York, London, Sydney, 1982. ISBN 0-471-86482-X, 0-471-08755-6. xvi + 327 pp. LCCN QA76.7 .G48 1982.

Gerdt:1989:ACP

[GKKY89] V. P. Gerdt, Z. T. Kostova, N. A. Kostov, and I. P. Yudin. Algebraic-numeric calculations of proton trajectories in bending magnets of synchrotron accelerator. Preprint E11-89-755, J.I.N.R., Dubna, USSR, 1989.

Grimes:1982:AIF

[GKRY82] Roger G. Grimes, David R. Kincaid, John Respess, and David M. Young. Algorithm 586, ITPACK 2C: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8(3):302– 322, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

George:1981:CSL

[GL81] Alan George and Joseph W. H.

Liu. Computer solution of large sparse positive definite systems. Prentice-Hall series in computational mathematics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1981. ISBN 0-13-165274-5. xii + 324 pp. LCCN QA188 .G46 1980.

Grob:1986:APP

[GL86] L. S. Grob and J. Lipkis. Approaches to parallel programming on multiprocessors. EUUG Autumn ’86, pages 283–294, September 1986.

Garbow:1990:RFS

[GL90] B. S. Garbow and J. N. Lyness. Remark on “Algorithm 662: A FORTRAN software package for the numerical inversion of the Laplace transform based on Weeks’ method”. ACM Transactions on Mathematical Software, 16(4):405–406, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.

                                        org/pubs/toc/Abstracts/0098-
                                        3500/98302.html. See [GGLM88].

Glass:1983:RS

[Gla83] Robert L. Glass. Real-Time Software. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1983. ISBN 0-13-767103-2. vii + 456 pp. LCCN QA76.54 .R425 1983. US$24.95.

Glarborg:1988:PFP

[Gla88] Peter Glarborg. PSR: a FOR-TRAN program for modeling wellstirred reactors. Sandia report; sand86-8209, Sandia National Laboratories, Albuquerque, NM, USA, 1988. 54 pp.

Goldberg:1983:INP

[GM83] Jeffrey L. Goldberg and Thomas C. Miller. Interactive name placement for provisional maps. Technical Papers of the American Congress of Surveying and Mapping, pages 314–321, 1983. CO-DEN TPAMDF.

Gill:1979:DSF

[GMPW79] Philip E. Gill, Walter Murray, Susan M. Picken, and Margaret H. Wright. The design and structure of a Fortran program library for optimization. ACM Transactions on Mathematical Software, 5 (3):259–283, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ghodssi:1986:GOS

[GMW86] Vida Ghodssi, Steven S. Muchnick, and Alex Wu. Global optimizer for Sun FORTRAN, C & Pascal. In USENIX Conference Proceedings, pages 318–334. USENIX Association, Berkeley, CA, USA, Summer 1986.

Gragg:1989:FSE

[GN89] W. Gragg and B. Neta. Fortran subroutines for the evaluation of the confluent hypergeometric functions. Technical Report NPS-MA-89-014, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943,

USA, 1989. URL http://math.

nps.navy.mil. Software available URL http://math.nps.navy.mil.

Goldstein:1981:INS

[Gol81] Ira Goldstein. Integrating a network-structured database into an object-oriented programming language. ACM SIGPLAN Notices, 16(1):124–125, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goller:1982:S

[Gol82a] M. Goller, editor. Simulationstechnik. ????, Berlin, Germany, 1982. ISBN 3-540-11605-2.

Goller:1982:TLU

[Gol82b] N. E. Goller. Triodic logic and its use in structured program design. The Computer Journal, 25(2):218–226, May 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
25/2/218.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/

hdb/Volume_25/Issue_02/tiff/
218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_02/tiff/219.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
25/Issue_02/tiff/220.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/221.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/
222.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_02/tiff/223.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
25/Issue_02/tiff/224.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
02/tiff/225.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/
226.tif.

Goldberg:1984:RAF

[Gol84] Richard Goldberg. Register allocation in FORTRAN I. Annals of the History of Computing, 6(1):19– 20, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1015.pdf;
 http://www.computer.org/annals/
an1984/a1015abs.htm.

Gonnet:1989:PAI

[Gon89] Gaston H. Gonnet, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89 / July 17– 19, 1989, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Goodman:1989:DFC

[Goo89] Goodman. Design of a FORTRAN to C translator, Part I. The Journal of C Language Translation, 1 (3):166–179, December 1989. ISSN 1042-5721.

Gottfried:1984:PFI

[Got84] Byron S. Gottfried. Programming With Fortran IV. QPI series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, January 1984. ISBN 0-13-729699-1 (paperback). x + 373 pp. LCCN QA76.73.F25 G67 1984. US$32.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0137296991.

Guzzi:1988:CFOb

[GPHL88] M. D. Guzzi, D. A. Padua, J. P.

Hoeflinger, and D. H. Lawrie. Cedar Fortran and other vector and parallel Fortran dialects. In IEEE [IEE88b], pages 114–121. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Guzzi:1990:CFO

[GPHL90] Mark D. Guzzi, David A. Padua, Jay Hoeflinger, and Duncan H. Lawrie. Cedar Fortran and other vector and parallel Fortran dialects. The Journal of Supercomputing, 4(1):37–62, March 1990. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        0920-8542&volume=4&issue=1&
                                        spage=37.

Gajski:1982:SOD

[GPKK82] D. D. Gajski, D. A. Padua, D. J. Kuck, and R. H. Kuhn. A second opinion on data flow machines and languages. Computer, 15(2):15–25, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1984:SOD

[GPKK84] D. D. Gajski, D. A. Padua, D. J. Kuck, and R. H. Kuhn. A second opinion on data flow machines and languages. In Hwang [Hwa84], pages 489–500. ISBN 0-8186-0581-2. LCCN TK 7888.3 H82 1984. Also published in/as: Computer Magazine of the Computer Group News of the IEEE Computer Group Society, Vol. 15 No. 2, Feb. 1982, pp. 15–25.

Gu:1988:SIA

[GQ88] Jia Chang Gu and Lin Jun Qian. Some iterative algorithms for the construction of number-theoretic transforms and their FORTRAN programs. (Chinese). J. Nanjing Inst. Tech., 18(5):71–78, 1988.

Gray:1981:FIC

[Gra81a] Robert F. Gray. A Fortran IV computer program for designing and implementing recursive digital filters. Technical report HDL-TR-1942, U.S. Army Material Development and Readiness Command; Available from the National Technical Information Service, Alexandria, VA, USA, 1981. 63 pp.

Grayson:1981:RKF

[Gra81b] Martin Grayson. RNFREE — keyword free-format input. Software —Practice and Experience, 11(9): 907–911, September 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Graham:1984:IFS

[Gra84a] N. Graham. Introduction to Fortran Seventy-Seven. Holt, Reinhart, and Winston, New York, NY, USA, May 1984. ISBN 0-03-059559-2. ???? pp. LCCN ???? US$18.55. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0030595592.

Grauch:1984:TFP

[Gra84b] V. J. S. Grauch. TAYLOR a FORTRAN program using Taylor series expansion for level-surface or surface-level continuation of potential-field data. Open-file report 84-501, U.S. Dept. of the Interior, Geological Survey, Denver, CO, USA, 1984. i + 31 pp.

GSS:1985:GFB

[Gra85] Graphic Software Systems, Inc. GSS-DRIVERS Fortran binding, 1985.

Gras:1986:DND

[Gra86a] Georges Gras. Détermination numérique du groupe d’Artin des extensions cycliques de Q à ramification donnée (programme FOR-TRAN IV-GALCYCL). (French) [Numerical determination of the Artin group of cyclic extensions of Q with given ramification (FOR-TRAN IV-GALCYCL program)]. Technical Report Exp. No. 7, Publ. Math. Fac. Sci. Besançon, Univ. Franche-Comté, Besançon, France, 1986. 34 pp. Théorie des nombres, Années 1984/85–1985/86, Fasc. 2.

Grauch:1986:VFP

[Gra86b] V. J. S. Grauch. VARMAG: a FORTRAN program to implement the variable-magnetization terrain-correction method for aeromagnetic data. Open-file report 86-268, U.S. Dept. of the Interior, Geological Survey, Denver, CO, USA, 1986. ii + 52 pp.

Gran:1988:HAF

[Gra88] Pamela K. Gran. A hybrid archetype for FORTRAN to ADA amelioration. Thesis (M.S.), National University, Vista, CA, USA, 1988. vii + 76 pp.

Garbinski:1988:UME

[GRB88] Charles Garbinski, Paul C. Redin, and Gerald D. Budd. User’s manual for EZPLOT version 5.5 a FORTRAN program for twodimensional graphic display of data. NASA technical memorandum 88293, National Aeronautics and Space Administration, Ames Research Center, Dryden Flight Research Facility, Edwards, CA, USA, 1988. iii + 41 pp. For sale by the National Technical Information Service.

Greenwood:1981:FDB

[Gre81] Stephen Greenwood. FORTRAN data base facility. Software—Practice and Experience, 11(1):100, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Letter to the editor.

Greenfield:1984:IFS

[Gre84] Martin N. Greenfield. The impact of FORTRAN standardization. Annals of the History of Computing, 6(1):33, January/March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1028.pdf;
 http://www.computer.org/annals/
an1984/a1028abs.htm.

Greenen:1985:LAF

[Gre85] Donald J. Greenen. Learning Apple Fortran. Computers and Math Series. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, October 1985. ISBN 0-88175-024-7. ???? pp. LCCN ???? US$32.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0881750247.

Greenbaum:1986:SST

[Gre86] Anne Greenbaum. Solving sparse triangular linear systems using Fortran with parallel extensions on the NYU Ultracomputer prototype. Ultracomputer Note 99, New York University, New York, NY, USA, April 1986.

Green:1988:DFP

[Gre88] Gregory N. Green. DLG2ISM, a Fortran program to read DLG-3 optional format digital data files into the VAX/VMS version of the interactive surface modeling software package. Open-file report 88-258-A, U.S. Dept. of the Interior, Geological Survey: [Books and Open-File Reports Section, distributor], Denver, CO, USA, 1988. 18 pp.

Grimsrud:1982:ESH

[Gri82] G. P. Grimsrud. Executive summary for the hydrological simulation program FORTRAN (HSPF). Technical report, U.S. Environmental Protection Agency, Research and Development Environmental Research Laboratory; Center for Environmental Research Information [distributor], Athens, GA, USA, 1982. 2 pp.

Griffiths:1985:SFV

[Gri85] Lloyd J. Griffiths. Structured Fortran 77: video course manual to accompany twelve color video tapes. Technical report, Interactive Instructional Television Program, School of Engineering, University of Southern California, Los Angeles, CA, USA, 1985. 253 pp.

Groundwater:1982:NSD

[Gro82] Neil P. Groundwater. Navy software development with ratfor-T and software tools. In Usr Group [Usr82], pages 342–?? ISBN ???? LCCN ???? Abstract only.

Grout:1983:FCP

[Gro83] Jarrell C. Grout. Fundamental Computer Programming: Using Fortran 77. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, February 1983. ISBN 0-13-335141-6 (paperback). xiv + 384 pp. LCCN QA76.73.F25 G76 1983 Sci-Eng. US$27.67. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0133351416.

Grosch:1987:ANS

[Gro87] C. E. Grosch. Adapting a Navier– Stokes code to the ICL-DAP. SIAM Journal on Scientific and Statistical Computing, 8(1):S96– 117, January 1987. CODEN SI-JCD4. ISSN 0196-5204.

Grotendorst:1989:MPC

[Gro89] J. Grotendorst. MAPLE programs for converting series expansions to rational functions using the Levin transformation. automatic generation of FORTRAN functions for numerical applications. Computer Physics Communications, 55(3): 325–335, October 1, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Grundy:1988:SFS

[Gru88] I. H. Grundy. SMOOFF — a Fortran smoothing program (U).

Aircraft structures technical memorandum; 490, Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratory, Melbourne, Victoria, Australia, 1988. ISBN 0-7316-5838-8 (not printed in book). 17 pp.

Galil:1981:LTS

[GS81] Z. Galil and J. Seireras. Lineartime string-matching using only a fixed number of local storage locations. Theoretical Computer Science, 13(3):331–336, March 1981. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gentzsch:1988:UPF

[GSZ88] W. Gentzsch, F. Szelenyi, and V. Zecca. Use of parallel FOR-TRAN for engineering problems on the IBM 3090 vector multiprocessor. Parallel Computing, 9(1):107– 115, December 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeGoede:1989:NFL

[GtTB89] E. D. De Goede and J. H. M. ten Thije Boonkkamp. NUMVEC FORTRAN library manual: chapter, partial differential equations, routine, OEHPC. Report. Centrum voor Wiskunde en Informatica NM-R8915, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, 1989. 10 pp.

Guest:1986:SCF

[Gue86] Clayton J. Guest. Supermap for Cray Fortran a user’s guide.

NASA contractor report NASA CR-177410, National Aeronautics and Space Administration, Ames Research Center, Moffett Field, CA, USA, 1986. ???? pp. For sale by the National Technical Information Service.

Guida:1981:EPS

[Gui81] G. Guida. An effective preprocessor for structured FORTRAN: the HENTRAN system. International Journal of Computer and Information Sciences, 10(4):283–297, August 1981. CODEN IJCIAH. ISSN 0091-7036.

Guinier:1987:FPP

[Gui87] D. Guinier. Fortran programs on a personal computer. ACM SIGPLAN FORTRAN Forum, 6 (2):39–47, October 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Guinier:1988:FPU

[Gui88] D. Guinier. A fast and portable uniform quasi-random generator of very large period based on a generalized multi-moduli congruential method. ACM SIMULETTER (SIGSIM) jointly with IEEE TC-SIM, 19, 3:27–33, 1988.

Guinier:1989:FUA

[Gui89] Daniel Guinier. A fast uniform ‘astronomical’ random number generator. SIGSAC Review, 7(1):1–13, Spring 1989. CODEN SSARE7. ISSN 0277-920X (print), 1558-0261 (electronic).

Gulve:1986:FFE

[Gul86] Samir S. Gulve. Fortran front end for a software evaluation system. Project (M.S.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1986. viii + 89 pp.

Gust:1984:LPB

[Gus84] H. Gust. Logic programming for beginners: An introduction to the tools of artificial intelligence. Elektronik, 33(24):64–9, November 30, 1984. CODEN EKRKAR. ISSN 0013-5658.

Guzzi:1987:CFP

[Guz87] Mark David Guzzi. Cedar Fortran programmer’s handbook. Technical Report CSRD 601, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. iv + 40 pp.

Guzzi:1988:CFOa

[Guz88] Mark David Guzzi. Cedar Fortran and other vector and parallel Fortran dialects. Technical Report CSRD 731, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 8 pp.

Griffin:1988:DPP

[GWM88] J. H. Griffin, H. J. Wasserman, and L. P. McGavran. A debugger for parallel processes. Software —Practice and Experience, 18(12): 1179–1190, December 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gibson:1982:IPU

[GY82] Glenn A. Gibson and James R.

Young. Introduction to Programming Using Fortran 77. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1982. ISBN 0-13-493551-9. xv + 461 pp. LCCN QA76.73.F25 G48 1982. US$42.20. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0134935519.

Hanna:1983:FSS

[HA83] Christine B. Hanna and Milton H.

Aronson. Fortran seventy seven: a professional course. Measurements and Data Corp., Pittsburgh, PA, USA, 1983. ISBN ???? 70 pp. LCCN ????

An:1984:FCK

[hA84] Po hui An. Flowchart chaksong kipop: FORTRAN, COBOL chungsim. Taeun Chulpansa, Soul Tukpyolsi, Korea, chopan edition, 1984. ISBN ???? 259 pp. LCCN ????

Haas:1987:MPR

[Haa87] Alexander Haas. The multiple prime random number generator. ACM Transactions on Mathematical Software, 13(4): 368–381, December 1987. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://doi.acm.org/10.
1145/35078.214349;   http://
www.acm.org/pubs/citations/
journals/toms/1987-13-4/p368-
haas/.

Hahn:1981:HPI

[Hah81] R. Hahn. Höhere Programmiersprachen im Vergleich. Akademische Verlagsgesellschaft, Wiesbaden, Germany, 1981. ISBN 3-400-00433-2.

Hahn:1987:PSF

[Hah87] Brian D. Hahn. Problem Solving With Fortran 77. Edward Arnold, London, UK, April 1987. ISBN 0-7131-3592-1 (paperback). viii + 247 pp. LCCN ???? US$22.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0713135921.

Hamilton:1985:RRK

[Ham85] Dennis E. Hamilton. Remark on “Algorithm 620: References and keywords for Collected Algorithms of the ACM”. ACM Transactions on Mathematical Software, 11(3): 305–307, September 1985. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RH84a, HM90].

Harris:1981:HFR

[Har81] Harris Corporation. Harris FOR-TRAN reference manual. Harris Corporation, Fort Lauderdale, FL, USA, 1981. 270 pp.

Harris:1985:HFR

[Har85] Harris Corporation. Computer Systems Division. Harris FOR-TRAN reference manual. Harris Corp., Fort Lauderdale, FL, USA, 1985. various pp.

Harada:1986:FP

[Har86a] Ken’ichi Harada. Fortran 77 puroguramingu. Information and computing; 9. Saiensusha, Tokyo, Japan, shohan edition, 1986. ISBN 4-7819-0461-0. xiv + 305 pp.

Harrigan:1986:IGP

[Har86b] Kevin A. (Kevin Ambrose) Harrigan. An interactive graphics package in VSPC FORTRAN. Thesis (M.Sc.C.S.), University of New Brunswick, Ottawa, Ontario, Canada, 1986. 2 microfiches (134 fr.) pp.

Harrison:1989:IAA

[Har89] W. Ludwell Harrison III. The Interprocedural Analysis and Automatic Parallelization of Scheme Programs. Technical Report CSRD No. 860, UILU-ENG-89-8003, Center of Supercomputing Research and Development, University of Illinois, Urbana, IL, USA, February 1989.

Hayes:1986:ABF

[Hay86] Manny Hayes. Accessing bit fields in FORTRAN-77. Comm. ACM, 29(10):991–995, October 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Honess:1981:IMA

[HB81] C. Brian Honess and C. Brian.

Instructor’s manual to accompany Structured business problem solving with FORTRAN. Allyn and Bacon, Needham Heights, MA, USA, 1981. ISBN 0-205-07333-6. 85 pp.

Holoien:1983:PSS

[HB83] Martin O. Holøien and Ali Behforooz. Problem Solving and Structured Programming With Fortran 77. Brooks/Cole series in computer science. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, March 1983. ISBN 0-534-01275-2 (paperback). xii + 522 (or xii + 514??) pp. LCCN QA76.73.F25 H63 1983. US$33.95; US$18.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0534012752.

Holoien:1984:PSS

[HB84] Martin O. Holøien and Ali Behforooz. Problem solving and structured programming with For-Tran 77, 1984.

Chou:1983:FCT

[hC83] Wei hsin Chou. Fu chuan tsung heng = FORTRAN IV, 77, F-80. Han Tung-mei: Tsung ching hsiao Sung kang tien nao tu shu tzu liao yu hsien kung ssu, Tai-pei shih, 1983. ISBN ???? 466 + [14] pp. LCCN ????

Heath:1981:FP

[Hea81] Heath Company (Benton Harbor, Mich.). FORTRAN programming. Heath Company, Benton Harbor, MI, USA, 1981. various pp.

Heinsohn:1983:SZM

[Hei83] J. Heinsohn. Ein Simulationsprogramm für die Z80-cpu mit Monitor, Lader und Binder in Fortran 4. Studienarbeit, 1983.

Heising:1984:EFI

[Hei84] William P. Heising. The emergence of FORTRAN IV from FOR-TRAN II. Annals of the History of Computing, 6(1):31–32, January/March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1028.pdf;
 http://www.computer.org/annals/
an1984/a1028abs.htm.

Heller:1983:RFI

[Hel83] M. Heller. RASC FORTRAN IV computer program for ranking and scaling of biostratigraphic events. Open file 922, Geological Survey of Canada, Ottawa, Ontario, Canada, 1983. 54 pp.

Hellmold:1985:SS

[Hel85a] K.-U. Hellmold. Der Simulationsrechner SIMPLEX. Dissertation, Universität Erlangen-Nürnberg, Arbeitsberichte des IMMD, Erlangen, Germany, 1985. 316 pp.

Hellmold:1985:SSM

[Hel85b] K.-U. Hellmold. Der Simulationsrechner Simplex, ein Multiprozessorsystem zur Simulation von zeitdiskreten Systemen auf GPSS-Fortran-Basis. PhD thesis, Universität Erlangen (??), Erlangen, Germany, 1985.

Hemmes:1986:FR

[Hem86] David Hemmes. FORTRANSIT recollections. Annals of the History of Computing, 8(1):70–73, January/March 1986. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1986/pdf/a1070.pdf;
                                         http://www.computer.org/annals/
                                        an1986/a1070abs.htm.

Herbold:1981:PPD

[Her81] Robert John Herbold. Prime program decomposition of Fortran programs. Thesis (M.S.), University of Maryland, College Park, MD, USA, 1981. 47 pp.

Herbert:1988:SCS

[Her88] T. Herbert. Symbolic computations with spectral methods. American Society of Mechanical Engineers, Heat Transfer Division, (Publication) HTD, 105:25– 31, 1988. CODEN ASMHD8. ISSN 0272-5673.

HP:1985:FRM

[Hew85] Hewlett–Packard Company. FOR-TRAN 77 reference manual: RTE-6/VM and RTE-A, HP 1000 computer system. Hewlett–Packard, Cupertino, CA, USA, 1985. various pp.

HP:1986:FR

[Hew86] Hewlett–Packard Company. FOR-TRAN/ 9000 reference. Hewlett– Packard, Fort Collins, CO, USA, 1986. various pp.

Heyman:1985:SCR

[Hey85] J. Heyman. Software conversion: retargeting from a super-mini to a micro. ACM SIGPLAN Notices, 20(2):54–57, February 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoffberg:1981:ACS

[HF81] Alan Hoffberg and Jehosua Friedmann. Apple computer supplement to FORTRAN IV: 2nd edition. John Wiley and Sons, New York, London, Sydney, 1981. ISBN 0-471-09813-2. vi + 30 pp.

Hamlin:1982:EDG

[HG82a] G. Hamlin and J. E. George. Experiences with distributing graphic software between processors (Fortran). In Proceedings of the 3RD International Conference on Distributed Computing Systems, pages 486–492. ????, ????, 1982.

Hennessy:1982:CGR

[HG82b] J. L. Hennessy and T. R. Gross.

Code generation and reorganization in the presence of pipeline constraints. In POPL ’82 [POP82], pages 120–127. ISBN ???? LCCN ????

Horning:1983:PTU

[HG83] R. R. Horning and W. E.

Goode. Processing time using EASYTRIEVE-11, clunks, and FORTRAN. In Proc. DECUS, Las Vegas, NV, USA. ????, ????, October 1983.

Hsu:1982:TCF

[hH82] Chih hung Hsu. TQ-16 chi FOR-TRAN yu yen chang yung suan fa cheng hsu chi. Hua hsueh kung yeh chu pan she: Hsin hua shu tien Pei-ching fa hsing so fa hsing, Peiching, ti 1 pan edition, 1982. ISBN ???? 452 pp. LCCN ????

Hume:1985:FSE

[HH85] J. N. P. Hume and R. C. (Richard C.) Holt. Fortran 77 for Scientists and Engineers. Reston Publishing Co., Inc., Reston, VA, USA, second edition, February 1985. ISBN 0-8359-2065-8 (paperback). xvi + 364 pp. LCCN QA76.73.F25 H85 1985. US$33.92. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835920658.

Hsieh:1981:RVC

[hHtM81] Ying hsiung Hsieh and Pao t’ien Mei. RT-11 VAX-VMS ch’eng shih chi ch’iao chi hsi t’ung ts’ao tso. Sung kang tien nao t’u shu tzu liao yu hsien kung ssu, T’ai-pei shih, ch’u pan edition, 1981. ISBN ???? 2 + 475 pp. LCCN ????

Higley:1986:FSP

[Hig86] Debra K. Higley. Fortran sorting program to code hydrocarbon production and show data using well data from petroleum information’s well history control system. Openfile report 86-437, U.S. Dept. of the Interior, Geological Survey, Denver, CO, USA, 1986. i + 28 pp.

Higham:1988:FCE

[Hig88a] N. J. Higham. Fortran codes for estimating the one-norm of a real or complex matrix, with applications to condition estimation. ACM Transactions on Mathematical Software, 14(4):381–396, December 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from math.liu.se

in pub/references.

Higham:1988:AFC

[Hig88b] Nicholas J. Higham. Algorithm 674: FORTRAN codes for estimating the one-norm of a real or complex matrix, with applications to condition estimation. ACM Transactions on Mathematical Software, 14(4):381–396, December 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/214386.html.  See also

[Hig89].

Higham:1989:CFC

[Hig89] Nicholas J. Higham. Corrigendum:

“Algorithm 674: FORTRAN codes for estimating the one-norm of a real or complex matrix, with applications to condition estimation”. ACM Transactions on Mathematical Software, 15(2):168, June 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/214391.html. See [Hig88b].

Higham:1991:RBF

[Hig91] Desmond J. Higham. Remark on “Algorithm 669: BRKF45: A FORTRAN subroutine for solving first-order systems of nonstiff initial value problems for ordinary differential equations”. ACM Transactions on Mathematical Software, 17(3):424–426, September 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/toc/
                                        Abstracts/0098-3500/116814.

html. See [Cas89a].

Hill:1981:SPF

[Hil81] Louis A. Hill. Structured Programming in Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1981. ISBN 0-13-854612-6 (paperback). xvi + 526 pp. LCCN ???? US$30.80. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0138546126.

Hillstrom:1982:JAP

[Hil82a] Kenneth E. Hillstrom. JAKEF — A portable symbolic differentiator of functions given by algorithms. Technical Report ANL– 82–48, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, 1982.

Hillstrom:1982:JPS

[Hil82b] Kenneth E. Hillstrom. JAKEF — A portable symbolic differentiator of functions given by algorithms. Technical Report ANL– 82–48, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Ave., Argonne, IL 60439– 4801, 1982.

Houtman:1983:FPR

[HK83] Hubert Houtman and C. J. (Cornelis Jan) Kost. A Fortran program (RELAX3D) to solve the 3 dimensional Poisson (Laplace) equation. Technical Report TRI-PP-83-95, Triumf, Vancouver, BC, Canada, 1983. 6 pp.

Hood:1984:PEF

[HK84] R. T. Hood and K. Kennedy. A programming environment for Fortran. Technical Report TR84-1, Rice University, Houston, TX, USA, June 1984.

Kuo:1985:KJT

[hK85] Ting ho Kuo. Ko jen tien nao (Apple): FORTRAN. Fu wen shu chu, Tai-nan shih, chu pan edition, 1985. ISBN ???? 212 pp. LCCN ????

Hanson:1987:ATA

[HK87a] R. J. Hanson and F. T. Krogh.

Algorithm 653: Translation of Algorithm 539: PC-BLAS Basic Linear Algebra Subprograms for FORTRAN usage with the INTEL 8087, 80287 numeric data processor. ACM Transactions on Mathematical Software, 13(3):311–317, September 1987. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/toc/
Abstracts/0098-3500/214346.

html. See [LHKK79a, DG82, Dod83, LN88b].

Hanson:1987:TAP

[HK87b] R. J. Hanson and F. T. Krogh.

Translation of Algorithm 539: PC-BLAS Basic Linear Algebra Subprograms for Fortran usage with the INTEL8087 80287 numeric data processor. ACM Transactions on Mathematical Software, 13(3): 311–317, September 1987. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Husmann:1988:ACF

[HKP88] Harlan E. Husmann, David J. Kuck, and David A. Padua. Automatic compound function definition for multiprocessors. In Proceedings of the 1988 International Conference on Parallel Processing, volume II, Software, pages 33–41. Penn State, University Park, PA, USA, August 1988. CSRD, U. Ill.

Hah:1982:NAF

[HL82a] C. Hah and B. Lakshminarayana. Numerical analysis and FOR-TRAN program for the computation of the turbulent wakes of turbomachinery rotor blades, isolated airfoils and cascade of airfoils. NASA contractor report 3509, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1982. [v] + 164 pp. For sale by the National Technical Information Service.

Herbert:1982:PVE

[HL82b] Henry E. Herbert and John E. Lamar. Production version of the extended NASA-Langley vortex lattice FORTRAN computer program. NASA technical memorandum 83304, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1982. various pp.

Heuft:1982:ITP

[HL82c] R. W. Heuft and W. D. Little. Improved time and parallel processor bounds for Fortranlike loops. IEEE Transactions on Computers, C-31(1):78–81, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675890.

Hoffman:1986:NFL

[HL86a] W. Hoffman and W. M. Lioen.

NUMVEC Fortran library manual, chapter: Simultaneous equations. Technical report, Report NM-R8614, Centre for Mathematics and Computer Science, Amsterdam, ????, 1986. ???? pp.

Hoffmann:1986:NFL

[HL86b] W. Hoffmann and W. M. Lioen. NUMVEC FORTRAN library manual: Chapter: simultaneous linear equations. Report NM 8614., CWI, Amsterdam, The Netherlands, 1986. various pp.

Hiromoto:1984:EDH

[HLM84] R. E. Hiromoto, O. M. Lubeck, and J. Moore. Experiences with the Denelcor HEP. Parallel Computing, 1(3–4):197–206, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hernando:1981:PFP

[HM81] J. A. Hernando and V. Massidda. Plattsum: a Fortran program that evaluates electrostatic lattice sums by the planewise summation method. Computer Physics Communications, 22(1): 13–31, February/March 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/001046558190076X.

Harris:1982:FPC

[HM82] A. G. Harris and J. R. Montgomery. A FORTRAN program to check numerical control machine tapes and graphically display cutting tool paths. Report 1982/IE/1, School of Mechanical and Industrial Engineering, University of New South Wales, Kensington, NSW, Australia, 1982. ISBN 0-909287-19-8. 11 + [6] pp.

Hernando:1984:PFP

[HM84] J. A. Hernando and V. Massidda. Plattsum: a Fortran program that evaluates electrostatic lattice sums by the planewise summation method. Computer Physics Communications, 35(1–3): C–680–C–681, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584828337.

Hopkins:1990:RRK

[HM90] Tim Hopkins and David Morse. Remark on “Algorithm 620: References and keywords for Collected Algorithms of the ACM”. ACM Transactions on Mathematical Software, 16(4):401–403, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RH84a, Ham85].

Hessel:1988:TAK

[HMB^{+}88] R. E. Hessel, M. Myszewski, G. Brussino, J. A. Swanson, and L. Wagner. Tuning the ANSYS kernel LSOLVE for a parallel computer. In Martin and Lundstrom [ML88], pages 191–201. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Houstis:1985:AGC

[HMR85a] E. N. Houstis, W. F. Mitchell, and J. R. Rice. Algorithm 637: GENCOL: Collocation of general domains with bicubic Hermite polynomials. ACM Transactions on Mathematical Software, 11(4): 413–415, December 1985. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Houstis:1985:AIH

[HMR85b] E. N. Houstis, W. F. Mitchell, and J. R. Rice. Algorithm 638: INTCOL and HERMCOL: Collocation on rectangular domains with bicubic Hermite polynomials. ACM Transactions on Mathematical Software, 11(4):416–418, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Haase:1988:FRG

[HO88] Camilla B. Haase and Jerry W. Ornstein. FORTRAN 77 reference guide. Prime Computer, Natick, MA, USA, fifth edition, 1988. various pp.

Hocking:1989:BIS

[HO89] Stephan A. Hocking and Gerard J. O’Neill. A butterfly implementation of SCHEDULE — an environment for developing and analysing parallel Fortran programs. In Jesshope and Reinartz [JR89], pages 108–115. ISBN 0-521-37177-5. LCCN QA76.5 .C6194 1988.

Hockney:1985:MCC

[Hoc85] Roger W. Hockney. (r_{∞}, n_{1/2}, s_{1/2}) measurements on the 2-CPU CRAY X-MP. Parallel Computing, 2(1):1–14, March 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hofmann:1984:SRV

[Hof84] M. Hofmann. Simulation eines Rechnernetzes — Vergleich Zweier Formeln zur Paketzerlegung in Netzen mit Paketvermittlung. Studienarbeit, Technische Universität Braunschweig (??), Braunschweig, Germany, 1984.

Hoffmann:1987:NFL

[Hof87] W. Hoffmann. NUMVEC FOR-TRAN library manual: chapter, simultaneous linear equations update,# 1. Report NM. Centrum voor Wiskunde en Informatica; 8712 Report NM. Centre for Mathematics and Computer Science. Department of Numerical Mathematics; 8712 Centrum voor Wiskunde en Informatica (Amsterdam, Netherlands). Report NM; 8712. NM-R8712. CWI, Amsterdam, The Netherlands, 1987. ISBN ???? 8 pp. LCCN ????

Holt:1987:DDA

[Hol87a] R. C. Holt. Data descriptors:

a compile-time model of data and addressing. ACM Transactions on Programming Languages and Systems, 9(3):367–389, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Holt:1987:DDC

[Hol87b] Richard C. Holt. Data descriptors: a compile-time model of data and addressing. ACM Transactions on Programming Languages and Systems, 9(3):367–389, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Honess:1981:SBP

[Hon81a] C. Brian Honess. Structured Business Problem Solving With Fortran. Allyn and Bacon, Needham Heights, MA, USA, March 1981. ISBN 0-205-07332-8. viii + 244 pp. LCCN HF5548.5.F2 H66. US$24.24; US$13.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0205073328.

HoneywellIS:1981:CPC

[Hon81b] Honeywell Information Systems, Inc. Control program-six (CP-6) FORTRAN programmer guide: level 66. Honeywell Information Systems, Inc., Waltham, MA, USA, 1981. various pp.

Honarvar:1982:IFC

[Hon82] Hassanali Honarvar. Improving FORTRAN character manipulation. Thesis (M.S.), Texas Woman’s University, Denton, TX, USA, 1982. v + 93 pp.

HoneywellIS:1985:DFP

[Hon85] Honeywell Information Systems, Inc. DM-IV (FORTRAN 77) programmer reference manual: large systems. Honeywell Information Systems, Inc., Waltham, MA, USA, 1985. various pp.

Hopkins:1981:CSS

[Hop81] T. R. Hopkins. Corrections to “Fortran subroutines for the solution of Toeplitz sets of linear equations”. IEEE Trans. Acoustics, Speech, and Signal Processing, 29 (6):1212, 1981. CODEN IETABA. ISSN 0096-3518.

Horowitz:1983:FPL

[Hor83a] E. Horowitz. Fundamentals of Programming Languages. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-914894-37-4. xiv + 450 pp. LCCN QA76.7 .H67 1983.

Horowitz:1983:PLG

[Hor83b] E. Horowitz, editor. Programming Languages, a Grand Tour: a collection of papers. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-914894-67-6, 3-540-11931-0. viii + 664 pp. LCCN QA76.7 .P78 1983.

Hosking:1988:FRU

[Hos88] J. R. M. Hosking. Fortran routines for use with the method of L-moments. Research report RC 13844, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1988. 80 pp.

Houlsby:1983:ESR

[Hou83] G. T. Houlsby. Efficient sorting routines in FORTRAN 77. Ouel report; 1529/84, Oxford University, Dept. of Engineering Science, Oxford, UK, 1983. [ii] + 20 pp. Available from National Technical Information Service.

Howden:1982:VSP

[How82] William E. Howden. Validation of scientific programs. ACM Computing Surveys, 14(2):193–227, June 1982. CODEN CMSVAN. ISSN 0010-4892.

Hopkins:1988:NMP

[HP88] Tim Hopkins and Chris Phillips.

Numerical methods in practice: using the NAG library. International Computer Science Series. Addison-Wesley, Reading, MA, USA, December 1988. ISBN 0-201-19248-9. xi + 308 pp. LCCN QA297.H589 1988. UK£14.95; US$26.95.

Hoffmann:1989:NFL

[HP89] W. Hoffmann and K. Potma.

NUMVEC FORTRAN library manual: chapter, simultaneous linear equations. Report. Centrum voor Wiskunde en Informatica NM-R8903, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, January 1989. various pp.

Huba:1982:OFP

[HPB82] G. J. Huba, A. L. Palisoc, and P. M. Bentler. ORSIM2: a FORTRAN program for symmetric and asymmetric orthogonal rotation of canonical variates and interbattery factors. The American Statistician, 36(1):62, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.
org/stable/2684100.

Henstridge:1984:ABO

[HPB84] J. D. Henstridge, R. W. Payne, and R. J. Baker. Algorithm 117: Buffered output in FORTRAN. The Computer Journal, 27(2):179– 184, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/
179.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_02/tiff/180.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
27/Issue_02/tiff/181.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/182.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/
183.tif;   http://www3.oup.

co.uk/computer_journal/hdb/
Volume_27/Issue_02/tiff/184.

tif. See note [WH87].

Hughes:1981:MPK

[HPR81] Charles Edward Hughes, Charles P.

Pfleeger, and Lawrence L. Rose. Metody programmirovnija kurs na osnoveFortrana (A second course in programming using Fortran). Mir, Moscow, Russia, 1981. ISBN ???? 336 pp. LCCN ????

Hammond:1983:IFI

[HR83a] Robert H. Hammond and William B.

Rogers. Introduction to Fortran IV. McGraw-Hill, New York, NY, USA, third edition, 1983. ISBN 0-07-025908-9 (paperback). vi + 330 pp. LCCN QA76.73.F25 H36 1983. US$10.50 (est.); US$27.45. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070259089.

Hughes:1983:OFC

[HR83b] M. H. Hughes and K. V. Roberts.

The OLYMPUS Fortran compositor. Computer Physics Communications, 29(1):45–57, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465583900267.

Hughes:1983:OFG

[HR83c] M. H. Hughes and K. V. Roberts.

The OLYMPUS Fortran generator. Computer Physics Communications, 29(1):59–71, March 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465583900279.

Hughes:1984:OFC

[HR84a] M. H. Hughes and K. V. Roberts. The Olympus Fortran compositor. Computer Physics Communications, 35(1–3):C–862, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584829690.

Hughes:1984:OFG

[HR84b] M. H. Hughes and K. V. Roberts. The Olympus Fortran generator. Computer Physics Communications, 35(1–3):C–863, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584829707.

Hammond:1987:IFP

[HRC87] Robert H. Hammond, William B. Rogers, and John B. Crittenden. Introduction to Fortran 77 and the Personal Computer. McGraw-Hill, New York, NY, USA, February 1987. ISBN 0-07-025912-7 (paperback). xvi + 413 pp. LCCN QA76.73.F25 H365 1987. US$35.95; US$26.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070259127.

Hammond:1989:IAF

[HRC89] Robert H. Hammond, William B. Rogers, and John B. Crittenden.

Introduccion al FORTRAN 77 y la PC. McGraw-Hill, New York, NY, USA, 1a en espanol edition, 1989. ISBN 968-422-505-9. xvi + 415 pp.

Hammond:1981:IAF

[HRH81] Robert H. Hammond, William B.

Rogers, and Byard Houck. Introduccion al FORTRAN IV. Mc-Graw-Hill, New York, NY, USA, second edition, 1981. ISBN ???? 203 pp. LCCN ????

Hellmold:1981:PSG

[HS81] K. U. Hellmold and B. Schmidt.

Parallelstruktur des Simulators GPSS-FORTRAN — Entwurf für ein angepaßtes Multiprozessorsystem. (German) [Parallel structure of the simulator GPSS-FORTRAN — A design for a matching multiprocessor system]. In Knödel and Schneider [KS81c], pages 25–39. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Haring:1983:REC

[HS83] G. Haring and O. Schechtner. On the realization of extended control structures in FORTRAN. Software—Practice and Experience, 13 (5):431–445, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harris:1986:IFP

[HS86] R. A. Harris and P. Segall. IN-VERSE, a FORTRAN program to solve for two-dimensional slip on a vertical strike slip fault by inversion of line-length data. Open-file report 86-451, U.S. Geological Survey, Menlo Park, CA, USA, 1986. 58 pp.

Hsiao:1983:FCP

[Hsi83] James K. Hsiao. A FORTRAN computer program to compute the radiation pattern of an array-fed paraboloid reflector. Nrl report; 8740, Naval Research Laboratory, Washington, DC, USA, 1983. iii + 13 pp.

Hinxman:1982:PEC

[HT82] A. I. Hinxman and Austin Tate. Parsing an extension to CODA-SYL FORTRAN DML. Software —Practice and Experience, 12(3): 205–209, March 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tang:1988:ECC

[hTD88] Ju ho Tang and Edward S. Davidson. An evaluation of Cray-1 and Cray X-MP performance on vectorizable Livermore Fortran loops. Technical Report CSRD 785, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1988. 9 pp.

Tang:1988:PVS

[hTDT88] Ju ho Tang, E. S. Davidson, and J. Tong. Polycyclic vector scheduling vs. chaining on 1-port vector supercomputers. In IEEE [IEE88b], pages 122–129. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Huang:1982:UMR

[Hua82] Yang H. (Yang Hsien) Huang.

User’s manual [for] REAMES, a simplified version of REAME in both BASIC and FORTRAN for the stability analysis of slopes. Technical report IMMR82/067, Institute for Mining and Minerals Research, University of Kentucky, Lexington, KY, USA, 1982. i + 135 pp.

Huddleston:1988:ICF

[Hud88a] John V. Huddleston. Introduction to computers — FORTRAN version. Exchange Pub. Division, Buffalo, NY, USA, 1988. ISBN 0-945261-01-2. vi + 170 pp. LCCN QA76.6.H817 1988.

Huddleston:1988:ICV

[Hud88b] John V. Huddleston. Introduction to computers — FORTRAN version. Exchange Pub. Division, Buffalo, NY, USA, 1988. ISBN 0-945261-01-2. vi + 170 pp.

Huebner:1983:RFI

[Hue83] J. S. Huebner. RDARL4, a FORTRAN interface for transferring chemical analytical data from an Applied Research Laboratories electron microprobe to a PDP-11 computing system. Open-file report 83-713, U.S. Geological Survey, Reston, VA, USA, 1983. 38 pp.

Hughes:1984:EFL

[Hug84] Robert A. Hughes. Early FOR-TRAN at Livermore. Annals of the History of Computing, 6(1): 30, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1028.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1028abs.htm.

Hurst:1982:GIC

[Hur82] Nicholas Richard Hurst. GASP IV, a combined continuous/discrete Fortran based simulation language. Thesis (Ph.D.), Purdue University, Lafayette, IN, USA, 1982. xii + 247. pp.

Husmann:1984:HFC

[Hus84] Harlan Edward Husmann. Highspeed format conversion and parallel I/O in numerical programs. Master’s thesis, Department of Computer Science, Univ. of Illinois at Urbana-Champaign, Urbana, IL, USA, January 1984. Available as TR number UIUCDCS-R-84-1152.

Hildebrandt:1983:EGV

[HV83] J. Hildebrandt and M. Vering. Einführung in GPSS-Fortran, Version 2. Technical report, Technische Universität Braunschweig (??), Braunschweig, Germany, 1983.

Hochman:1986:FCP

[HW86] Murray Hochman and Oliver Watson. FORTRAN 77: the complete primer. Hochman Watson Pub. Co., ??, ??, USA, 1986. ISBN ???? iii + 257 pp. LCCN QA76.73.F25 H62 1986.

Hwang:1984:TSD

[Hwa84] Kai Hwang, editor. Tutorial—Supercomputers: Design and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0581-2. viii + 640 pp. LCCN TK 7888.3 H82 1984.

Hatton:1988:SKS

[HWS^{+}88] Les Hatton, Andy Wright, Stuart Smith, Gregg Parkes, Paddy Bennett, and Robert Laws. The Seismic Kernel System—A large-scale exercise in Fortran 77 portability. Software—Practice and Experience, 18(4):301–329, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Describes portability issues in a 500,000+ line system.

Hybl:1987:C

[Hyb87] Albert Hybl. Correspondence.

ACM SIGPLAN FORTRAN Forum, 6(3):10–11, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Hyman:1982:FEF

[Hym82] James M. Hyman. FORSIG: an extension of FORTRAN with significance arithmetic. Technical report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1982. iv + 101 pp.

Hromadka:1987:BAM

[HYP87] Theodore V. Hromadka, Ch.-Ch. Yen, and George Francis Pinder. The Best Approximation Method: An Introduction, volume 27 of Lecture Notes in Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., March 1987. ISBN 0-387-17572-5. xii + 168 pp. LCCN TA335 .H761 1987. US$27.00.

Yu:1982:FIY

[hYsA82] Ki han Yu and Chung su An. Fortran IV yonsup. Tonga Hagyonsa, Soul, Korea, 1982. ISBN ???? 317 pp. LCCN ????

Indrea:1984:FAE

[IA84] E. Indrea and N. Aldea. Fourier analysis of exafs data, a selfcontained Fortran program-package. Computer Physics Communications, 35(1–3):C–652, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S001046558482812X.

Ishiguro:1989:DDA

[IA89] M. (Makio) Ishiguro and Hirotsugu Akaike. DALL: Davidons algorithm for log likelihood maximization: a FORTRAN subroutine for statistical model builders. Number 25 in Computer science monographs (Tokei Suri Kenkyujo (Tokyo, Japan)). Institute of Statistical Mathematics, Tokyo, Japan, 1989. ISBN ???? 29 pp. LCCN ????

IBM:1985:VFCa

[IBM85] IBM Corporation. VS FOR-TRAN compiler and library reference summary. IBM Corporation, New York, NY, USA, sixth edition, 1985. 23 pp.

IBM:1986:DGR

[IBM86] IBM Corporation, San Jose, CA, USA. Development Guide, Relational Applications, 1986.

IBMUK:1987:IF

[IBM87] IBM United Kingdom. IBM Fortran/2. IBM United Kingdom, PO Box 41, Portsmouth, UK, 1987. 3 books + 4 diskettes + 1 booklet. pp.

IBM:1988:VMSa

[IBM88] IBM Corporation. Virtual machine / extended architecture system product: application development guide for FORTRAN and COBOL. IBM Corporation, New York, NY, USA, 1988. iii–x + 263 pp.

IBM:1989:VFVa

[IBM89a] IBM Corporation. VS FORTRAN Version 2: installation and customization for MVS, Release 4. IBM Corporation, New York, NY, USA, third edition, 1989. xiv + 94 pp.

IBMUK:1989:IAI

[IBM89b] IBM United Kingdom. IBM advanced interactive executive for the personal System/2 (AIX PS/ 2): AIX PS/ 2 VS Fortran: version 1.1. IBM United Kingdom, PO Box 41, Portsmouth, UK, 1989. various pp.

IEC:1981:ISC

[IEC81] IEC. IEC 60713 (1981-01): Subroutines for CAMAC. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://www.

iec.ch, 1981. 57 pp. US$98.00.

URL http://www.iec.ch/cgi-
bin/procgi.pl/www/iecwww.p?
wwwlang=E&wwwprog=cat-det.p&
wartnum=017001.

IEC:1985:III

[IEC85] IEC. ISO/IEC 7846 (1985-12): Industrial real-time FORTRAN — Application for the control of industrial processes. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://www.

iec.ch, 1985. 32 pp. US$63.00.

URL http://www.iec.ch/cgi-
bin/procgi.pl/www/iecwww.p?
wwwlang=E&wwwprog=cat-det.p&
wartnum=018044.

IEC:1988:III

[IEC88] IEC. ISO/IEC 8651-1 (1988-12): Information processing systems — Computer Graphics — Graphical Kernel System (GKS) language bindings — Part 1: FORTRAN. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://www.

iec.ch, 1988. 116 pp. US$156.00.

URL http://www.iec.ch/cgi-
bin/procgi.pl/www/iecwww.p?
wwwlang=E&wwwprog=cat-det.p&
wartnum=018128.

IEEE:1981:PSC

[IEE81] Proceedings: 5th Symposium on Computer Arithmetic: May 18– 19, 1981, University of Michigan, Ann Arbor, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN QA 76.6 S985t 1981. IEEE catalog number 81CH1630-C.

IEEE:1988:PIC

[IEE88a] IEEE, editor. Proceedings. International Computer Science Conference ’88. Artificial Intelligence: Theory and Applications. IEEE Comput. Soc. Hong Kong Chapter, Hong Kong, 1988.

IEEE:1988:PSN

[IEE88b] IEEE, editor. Proceedings, Supercomputing ’88: November 14– 18, 1988, Orlando, Florida, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes.

Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Li:1982:TNY

[iL82] Chung i Li. Tien nao yu yen FOR-TRAN. Fu wen shu chu, Tai-nan, 23rd edition, 1982. ISBN ???? 532 pp. LCCN ????

IMSL:1982:UMI

[IMS82] IMSL, Inc. User’s manual, IMSL library: problem-solving software system for mathematical and statistical FORTRAN programming. IMSL, Houston, TX, USA, nineth edition, 1982. 4 v. pp.

IMSL:1984:UMI

[IMS84] IMSL, Inc. User’s manual, IMSL library: FORTRAN subroutines for mathematics and statistics. IMSL, Houston, TX, USA, ed. 9.2 edition, 1984. 4 v. pp.

IMSL:1987:ILF

[IMS87a] IMSL, Inc. The IMSL libraries: FORTRAN subroutines for mathematics and statistics: MATH/ LIBRARY — mathematical applications; STAT/LIBRARY — statistical analysis; SFUN/LIBRARY — special functions. IMSL, Houston, TX, USA, 1987. 32 pp.

IMSL:1987:IUMb

[IMS87b] IMSL, Inc. IMSL user’s manual: MATH/LIBRARY: FOR-TRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, 1987. ???? pp.

IMSL:1987:IUMa

[IMS87c] IMSL, Inc. IMSL user’s manual: SFUN/LIBRARY: FOR-TRAN subroutines for evaluating special functions. IMSL, Houston, TX, USA, version 2.0 edition, 1987. v + 60 + [19] pp.

IMSL:1987:IUMc

[IMS87d] IMSL, Inc. IMSL user’s manual: STAT/LIBRARY: FOR-TRAN subroutines for statistical analysis. IMSL, Houston, TX, USA, 1987. ???? pp.

IMSL:1987:MLFb

[IMS87e] IMSL, Inc. MATH/library: FOR-TRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, version 1.0, April 1987 edition, 1987. v + 1152 + [80] (doubled–up) pp.

IMSL:1987:MLFc

[IMS87f] IMSL, Inc. MATH/LIBRARY:

FORTRAN subroutines for mathematical applications: quick reference. IMSL, Houston, TX, USA, 1987. 91 pp.

IMSL:1987:MLFa

[IMS87g] IMSL, Inc. MATH/ library: FOR-TRAN subroutines for mathematical applications: user’s manual. IMSL, Houston, TX, USA, 1987. 3 v. pp.

IMSL:1987:SLFb

[IMS87h] IMSL, Inc. SFUN/library: FOR-TRAN subroutines for evaluating special functions. IMSL, Houston, TX, USA, version 2.0, April 1987 edition, 1987. v + 60 + [23] pp.

IMSL:1987:SLFc

[IMS87i] IMSL, Inc. SFUN/library: FOR-TRAN subroutines for evaluating special functions: quick reference. IMSL, Houston, TX, USA, 1987. 29 pp.

IMSL:1987:SLFa

[IMS87j] IMSL, Inc. SFUN/library: FOR-TRAN subroutines for evaluating special functions: user’s manual. IMSL, Houston, TX, USA, 1987. various pp.

IMSL:1987:SLFf

[IMS87k] IMSL, Inc. STAT/LIBRARY: FORTRAN subroutines for statistical analysis: quick reference. IMSL, Houston, TX, USA, 1987. 76 pp.

IMSL:1987:SLFd

[IMS87l] IMSL, Inc. STAT/ library: FOR-TRAN subroutines for statistical analysis: user’s manual. IMSL, Houston, TX, USA, 1987. 3 v. pp.

IMSL:1987:SLFe

[IMS87m] IMSL, Inc. Stat/library: FOR-TRAN subroutines for statistical analysis: user’s manual. IMSL, Houston, TX, USA, version 1.0, April 1987 edition, 1987. vi + 1232 + [75] (doubled–up) pp.

IMSL:1987:UMM

[IMS87n] IMSL, Inc. User’s manual: MATH/LIBRARY: FORTRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, 1987. various pp.

IMSL:1989:ILF

[IMS89a] IMSL, Inc. The IMSL libraries:

FORTRAN subroutines for mathematics and statistics. Math/ Library for mathematical applications, Stat/Library for statistical analysis, Sfun/Library. IMSL, Houston, TX, USA, 1989. 60 pp.

IMSL:1989:MLFa

[IMS89b] IMSL, Inc. MATH/LIBRARY:

FORTRAN subroutines for mathematical applications: quick reference. IMSL, Houston, TX, USA, 1989. 97 pp.

IMSL:1989:MLFb

[IMS89c] IMSL, Inc. MATH/LIBRARY:

FORTRAN subroutines for mathematical applications: user’s manual. IMSL, Houston, TX, USA, software (v.1.1, Jan. 1989) edition, 1989. ???? pp.

IMSL:1989:SLFa

[IMS89d] IMSL, Inc. SFUN/LIBRARY:

FORTRAN subroutines for evaluating special functions: quick reference. IMSL, Houston, TX, USA, 1989. 32 pp.

IMSL:1989:SLFb

[IMS89e] IMSL, Inc. SFUN/LIBRARY:

FORTRAN subroutines for evaluating special functions: user’s manual. IMSL, Houston, TX, USA, softcover (v.2.1, Jan. 1989) edition, 1989. v + 60 + [23] pp.

IMSL:1989:SLFc

[IMS89f] IMSL, Inc. STAT/LIBRARY:

FORTRAN subroutines for statistical analysis: quick reference.

IMSL, Houston, TX, USA, 1989. 82 pp.

IMSL:1989:SLFd

[IMS89g] IMSL, Inc. STAT/LIBRARY: FORTRAN subroutines for statistical analysis: user’s manual. IMSL, Houston, TX, USA, softcover (v.1.1, Jan. 1989) edition, 1989. ???? pp.

IMSL:1989:UMIa

[IMS89h] IMSL, Inc. User’s manual, IMSL MATH/LIBRARY: FORTRAN subroutines for mathematical applications. IMSL, Inc., Houston, TX, USA, version 1.1; softcover 1.1 edition, 1989. vi + 1152 columns pp.

IMSL:1989:UMIc

[IMS89i] IMSL, Inc. User’s manual: IMSL MATH/LIBRARY: FORTRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, software 1.1 edition, 1989. 1152 columns pp.

IMSL:1989:UMId

[IMS89j] IMSL, Inc. User’s manual: IMSL MATH/LIBRARY: FORTRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, version 1.1 edition, 1989. 3 v. pp.

IMSL:1989:UMIe

[IMS89k] IMSL, Inc. User’s manual: IMSL MATH/LIBRARY: FORTRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, 1989. ???? pp.

IMSL:1989:UMIf

[IMS89l] IMSL, Inc. User’s manual:

IMSL SFUN/LIBRARY: FOR-TRAN subroutines for evaluating special functions. IMSL, Houston, TX, USA, version 2.1 edition, 1989. various pp.

IMSL:1989:UMIb

[IMS89m] IMSL, Inc. User’s manual, IMSL STAT/ Library: FORTRAN subroutines for statistical analysis. IMSL, Inc., Houston, TX, USA, version 1.1; softcover 1.1 edition, 1989. vii + 1232 columns pp.

IMSL:1989:UMIg

[IMS89n] IMSL, Inc. User’s manual:

IMSL STAT/LIBRARY: FOR-TRAN subroutines for statistical analysis. IMSL, Houston, TX, USA, version 1.1 edition, 1989. 3 v. pp.

IBM:1981:UFA

[Int81a] International Business Machines Corporation. US Fortran application programming: guide: program product. IBM Corporation, White Plains, NY, USA, 1981. xiv + 253 pp.

IBM:1981:VFAa

[Int81b] International Business Machines Corporation. VS Fortran application programming: guide: program product. IBM Corporation, White Plains, NY, USA, 1981. xiv + 253 pp.

IBM:1981:VFAb

[Int81c] International Business Machines Corporation. VS Fortran application programming: library reference: program product. IBM Corporation, White Plains, NY, USA, 1981. 139 pp.

IBM:1981:VFAc

[Int81d] International Business Machines Corporation. VS Fortran application programming: library reference: systems. IBM Corporation, White Plains, NY, USA, 1981. 251 pp.

ICL:1981:DFL

[Int81e] International Computers, Ltd. DAP: FORTRAN language. International Computers, Ltd., London, UK, third edition, 1981. 141 pp.

IBM:1982:FC

[Int82a] International Business Machines Corporation. Fortran compiler, 1982.

IBM:1982:FIDa

[Int82b] International Business Machines Corporation. FORTRAN interactive debug for CMS and TSO. IBM Corporation, New York, NY, USA, 1982. 2 pp.

IBM:1982:FIDb

[Int82c] International Business Machines Corporation. FORTRAN interactive debug for CMS and TSO: guide and reference. IBM Corporation, New York, NY, USA, fourth edition, 1982. 151 pp.

IBM:1982:FIDc

[Int82d] International Business Machines Corporation. FORTRAN interactive debug for CMS and TSO: reference summary. IBM Corporation, New York, NY, USA, second edition, 1982. 1 pp.

IBM:1982:IFP

[Int82e] International Business Machines Corporation. IBM FORTRAN program products for OS and the CMS component of VM/ 370: general information. IBM Corporation, New York, NY, USA, third edition, 1982. 38 pp.

IBM:1982:OVG

[Int82f] International Business Machines Corporation. OS/VS graphic subroutine package for FORTRAN IV, COBOL, and PL/I. IBM Corporation, New York, NY, USA, second edition, 1982. viii + 179 pp.

IBM:1982:VFAa

[Int82g] International Business Machines Corporation. VS FORTRAN application programming: guide: program product. IBM Corporation, San Jose, CA, USA, third edition, 1982. xix + 271 pp.

IBM:1982:VFAb

[Int82h] International Business Machines Corporation. VS FORTRAN application programming: language reference. IBM Corporation, New York, NY, USA, third edition, 1982. ???? pp.

IBM:1982:VFC

[Int82i] International Business Machines Corporation. VS FORTRAN compiler and library: diagnosis. IBM Corporation, New York, NY, USA, third edition, 1982. ???? pp.

IBM:1982:VFM

[Int82j] International Business Machines Corporation. VS FORTRAN master index. IBM Corporation, New York, NY, USA, 1982. ???? pp.

IBM:1983:F

[Int83a] International Business Machines Corporation. Fortran, 1983.

IBM:1983:ISF

[Int83b] International Business Machines Corporation. IBM System/34 FORTRAN IV reference manual: program no. 5726-FO1. IBM Corporation, New York, NY, USA, third edition, 1983. various pp.

IBM:1983:ISS

[Int83c] International Business Machines Corporation. IBM System/360 and System/370 FORTRAN IV language: systems. IBM Corporation, Programming Publishing Division, San Jose, CA, USA, 12th, September 1983 edition, 1983. ???? pp.

IBM:1983:VFAa

[Int83d] International Business Machines Corporation. VS FORTRAN application programming: guide. IBM Corporation, New York, NY, USA, fourth edition, 1983. xxi + 284 pp.

IBM:1983:VFAb

[Int83e] International Business Machines Corporation. VS FORTRAN application programming: language reference. IBM Corporation, New York, NY, USA, fourth edition, 1983. xvii + 291 pp.

IBM:1983:VFAc

[Int83f] International Business Machines Corporation. VS FORTRAN application programming: library reference. IBM Corporation, New York, NY, USA, third edition, 1983. ???? pp.

IBM:1983:VFAd

[Int83g] International Business Machines Corporation. VS FORTRAN application programming: system services reference supplement. IBM Corporation, New York, NY, USA, third edition, 1983. ???? pp.

IBM:1983:VFC

[Int83h] International Business Machines Corporation. VS FORTRAN compiler and library: general information. IBM Corporation, New York, NY, USA, fifth edition, 1983. ???? pp.

IBM:1983:VFI

[Int83i] International Business Machines Corporation. VS FORTRAN installation and customization. IBM Corporation, New York, NY, USA, third edition, 1983. ???? pp.

Intel:1984:FUG

[Int84a] Intel Corporation. Fortran-86 user’s guide. Intel Corp., Santa Clara, CA, USA, revised-003 edition, 1984. ISBN ???? various pp. LCCN ????

IBM:1984:EER

[Int84b] International Business Machines Corporation. Extended exponent range for FORTRAN users: program description/operations manual. IBM Corporation, New York, NY, USA, 1984. viii + 54 pp.

IBM:1984:VFCa

[Int84c] International Business Machines Corporation. VS FORTRAN compiler and library: diagnosis. IBM Corporation, New York, NY, USA, fifth edition, 1984. ???? pp.

IBM:1984:VFCb

[Int84d] International Business Machines Corporation. VS FORTRAN compiler, library, and interactive debug general information. IBM Corporation, San Jose, CA, USA, second edition, 1984. x + 72 pp.

Intel:1985:FUG

[Int85a] Intel Corporation. FORTRAN-86 user’s guide for DOS systems. Intel Corporation, Santa Clara, CA, USA, 1985. ISBN 0-917017-70-6. xix + 274 pp. LCCN QA76.73.F25F67 1985. US$24.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0917017706. Cover title: Software development tools. Includes index.

IBM:1985:FUV

[Int85b] International Business Machines Corporation. FORTRAN utilities for VM/370-III: program description/operations manual. IBM Corporation, San Jose, CA, USA, fourth edition, 1985. vi + 94 pp.

IBM:1985:VFCb

[Int85c] International Business Machines Corporation. VS FORTRAN compiler and library. IBM Corporation, New York, NY, USA, 1985. 6 pp.

IBM:1985:VFCc

[Int85d] International Business Machines Corporation. VS FORTRAN compiler and library: installation and customization. IBM Corporation, New York, NY, USA, sixth edition, 1985. xiii + 108 pp.

IBM:1985:VFCd

[Int85e] International Business Machines Corporation. VS FORTRAN compiler, library, and interactive debug: general information. IBM Corporation, New York, NY, USA, fourth edition, 1985. xii + 80 pp.

IBM:1985:VFI

[Int85f] International Business Machines Corporation. VS FORTRAN installation and customization. IBM Corporation, New York, NY, USA, fifth edition, 1985. ???? pp.

IBM:1985:VFL

[Int85g] International Business Machines Corporation. VS FORTRAN language and library reference. IBM Corporation, New York, NY, USA, second edition, 1985. xxi + 561 pp.

IBM:1985:VFP

[Int85h] International Business Machines Corporation. VS FORTRAN programming guide. IBM Corporation, New York, NY, USA, second edition, 1985. xxviii + 455 pp.

IBM:1985:VFT

[Int85i] International Business Machines Corporation. VS FORTRAN technical bulletin. IBM Corporation, New York, NY, USA, second edition, 1985. 14 pp.

IBM:1986:DWF

[Int86a] International Business Machines Corporation. Designing and writing FORTRAN programs for vector and parallel processing. IBM Corporation, Kingston, NY, USA, 1986. x + 116 pp.

IBM:1986:ISP

[Int86b] International Business Machines Corporation. IBM System/36: programming with FORTRAN IV (System/34 compatible). IBM Corporation, ??, Ontario, Canada, fourth edition, 1986. various pp.

IBM:1986:LUY

[Int86c] International Business Machines Corporation. Learning to use your system: Working with FORTRAN and advanced graphics. IBM Corporation, Armonk, NY, USA, 1986. various pp.

IBM:1986:VFC

[Int86d] International Business Machines Corporation. VS FORTRAN compiler and library. IBM Corporation, New York, NY, USA, 1986. 5 pp.

IBM:1986:VFI

[Int86e] International Business Machines Corporation. VS FORTRAN interactive debug guide and reference version 2. IBM Corporation, New York, NY, USA, 1986. viii + 254 pp.

IBM:1986:VFL

[Int86f] International Business Machines Corporation. VS FORTRAN language and library reference version 2. IBM Corporation, New York, NY, USA, 1986. xiv + 604 pp.

IBM:1986:VFP

[Int86g] International Business Machines Corporation. VS FORTRAN programming guide version 2. IBM Corporation, New York, NY, USA, 1986. xix + 557 pp.

IBM:1987:FS

[Int87a] International Business Machines Corporation. Fortran standardization. IBM Corporation, San Jose, CA, USA, 1987. 43 pp.

IBM:1987:IFL

[Int87b] International Business Machines Corporation. IBM Fortran/2: language reference. IBM Corporation, Boca Raton, FL, USA, 1987. various pp.

IBM:1987:VFI

[Int87c] International Business Machines Corporation. VS FORTRAN interactive debug guide and reference version 2. IBM Corporation, New York, NY, USA, second edition, 1987. ix + 286 pp.

IBM:1987:VFL

[Int87d] International Business Machines Corporation. VS FORTRAN language and library reference version 2. IBM Corporation, New York, NY, USA, third edition, 1987. xvi + 595 pp.

IBM:1987:VFP

[Int87e] International Business Machines Corporation. VS FORTRAN programming guide version 2. IBM Corporation, New York, NY, USA, third edition, 1987. xix + 606 pp.

IBM:1987:VFVa

[Int87f] International Business Machines Corporation. VS FORTRAN Version 2 diagnosis guide. IBM Corporation, New York, NY, USA, second edition, 1987. xi + 39 pp.

IBM:1987:VFVb

[Int87g] International Business Machines Corporation. VS FORTRAN version 2 interactive debug guide and reference, release 2. IBM Corporation, San Jose, CA, USA, second edition, 1987. ix + 286 pp.

IBM:1987:VFVc

[Int87h] International Business Machines Corporation. VS FORTRAN version 2 language and library reference, release 4. IBM Corporation, San Jose, CA, USA, third edition, 1987. xvi + 595 pp.

IBM:1987:VFVd

[Int87i] International Business Machines Corporation. VS FORTRAN version 2 programming guide, release 2. IBM Corporation, San Jose, CA, USA, third edition, 1987. xix + 606 pp.

IBM:1988:AVF

[Int88a] International Business Machines Corporation. AIX VS Fortran reference manual. IBM Corporation, Danbury, CT, USA, 1988. various pp.

IBM:1988:ADG

[Int88b] International Business Machines Corporation. Application development guide for FORTRAN and COBOL: VM/XA-SP release 1 and release 2. IBM Corporation, Kingston, NY, USA, 1988. x + 263 pp.

IBM:1988:IAV

[Int88c] International Business Machines Corporation. IBM AIX VS Fortran user’s guide: AIX operating system: RT, PS/2. IBM Corporation, Danbury, CT, USA, 1988. various pp.

IBM:1988:VMSb

[Int88d] International Business Machines Corporation. Virtual machine / system product: application development guide for FORTRAN and COBOL. IBM Corporation, New York, NY, USA, fourth edition, 1988. x + 280 pp.

IBM:1988:VFI

[Int88e] International Business Machines Corporation. VS FORTRAN interactive debug guide and reference version 2. IBM Corporation, New York, NY, USA, third edition, 1988. xiii + 276 pp.

IBM:1988:VFL

[Int88f] International Business Machines Corporation. VS FORTRAN language and library reference version 2. IBM Corporation, New York, NY, USA, fourth edition, 1988. xv + 466 pp.

IBM:1988:VFPb

[Int88g] International Business Machines Corporation. VS FORTRAN programming guide. IBM Corporation, San Jose, CA, USA, third edition, 1988. xxvii + 477 pp.

IBM:1988:VFPa

[Int88h] International Business Machines Corporation. VS FORTRAN programming guide version 2. IBM Corporation, New York, NY, USA, fourth edition, 1988. xi + 470 pp.

IBM:1988:VFVa

[Int88i] International Business Machines Corporation. VS FORTRAN Version 2 directory of programming interfaces for customers, Release 3. IBM Corporation, New York, NY, USA, 1988. v + 7 pp.

IBM:1988:VFVb

[Int88j] International Business Machines Corporation. VS FORTRAN Version 2 directory of programming interfaces for customers, Release 3. IBM Corporation, New York, NY, USA, 1988. v + 7 pp.

IBM:1988:VFVc

[Int88k] International Business Machines Corporation. VS FORTRAN Version 2 installation and customization for MVS, Release 3. IBM Corporation, New York, NY, USA, second edition, 1988. xiv + 96 pp.

IBM:1988:VFVd

[Int88l] International Business Machines Corporation. VS FORTRAN Version 2 installation and customization for MVS, Release 3. IBM Corporation, New York, NY, USA, second edition, 1988. xiv + 96 pp.

IBM:1988:VFVe

[Int88m] International Business Machines Corporation. VS FORTRAN Version 2 interactive debug guide and reference, Release 3. IBM Corporation, New York, NY, USA, third edition, 1988. xiii + 276 pp.

IBM:1988:VFVf

[Int88n] International Business Machines Corporation. VS FORTRAN Version 2 language and library reference, Release 3. IBM Corporation, New York, NY, USA, fourth edition, 1988. xv + 466 pp.

IBM:1988:VFVg

[Int88o] International Business Machines Corporation. VS FORTRAN Version 2 programming guide, Release 3. IBM Corporation, New York, NY, USA, fourth edition, 1988. xxi + 470 pp.

IBM:1988:VFVh

[Int88p] International Business Machines Corporation. VS FORTRAN Version 2 reference summary, Release 3. IBM Corporation, New York, NY, USA, fourth edition, 1988. iii + 45 pp.

ISO:1988:IPS

[Int88q] International Organization for Standardization. Technical and Information Processing Systems Committee ISO/TC 97. Information processing systems: computer graphics, graphical kernel system (GKS) language bindings. International standard; ISO 8651-1. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 116 pp. LCCN ????

Technical:1988:IPS

[Int88r] International Organization for Standardization. Technical and Information Processing Systems Committee ISO/TC 97. Information processing systems: computer graphics, graphical kernel system (GKS) language bindings. International standard; ISO 8651-1. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 116 pp. LCCN ????

IBM:1989:VFL

[Int89a] International Business Machines Corporation. VS FORTRAN language and library reference version 2. IBM Corporation, New York, NY, USA, fifth edition, 1989. xvii + 465 pp.

IBM:1989:VFP

[Int89b] International Business Machines Corporation. VS FORTRAN programming guide version 2. IBM Corporation, New York, NY, USA, fifth edition, 1989. xx + 460 pp.

IBM:1989:VFVb

[Int89c] International Business Machines Corporation. VS FORTRAN Version 2, directory of programming interfaces for customers, release 3. IBM Corporation, New York, NY, USA, second edition, 1989. 4 + [3] pp.

IBM:1989:VFVc

[Int89d] International Business Machines Corporation. VS FORTRAN version 2: language and library reference, release 4. IBM Corporation, New York, NY, USA, fifth edition, 1989. xvii + 465 pp.

IBM:1989:VFVd

[Int89e] International Business Machines Corporation. VS FORTRAN version 2: programming guide, release 4. IBM Corporation, New York, NY, USA, fifth edition, 1989. xx + 460 pp.

ICCC:1984:FPU

[Ion84] Iona College. Computing Center. FORTRAN programming under MUSIC. Iona College, New Rochelle, NY, USA, September 1984. 74 pp.

IBM:1984:PF

[IR84] International Business Machines Corporation and Ryan-McFarland Corporation. Professional FOR-TRAN, 1984.

Iman:1984:FPUa

[IS84a] Ronald L. Iman and Michael J.

Shortencarier. A FORTRAN 77 program and user’s guide for the generation of Latin hypercube and random samples for use with computer models. Technical Report NUREG/CR-3624, SAND83-2365, U.S. Nuclear Regulatory Commission. Division of Risk Analysis and Operations., Washington, DC, USA, March 1984. vi + 50 + 10 pp.

Iman:1984:FPUb

[IS84b] Ronald L. Iman and Michael J. Shortencarier. A FORTRAN 77 program and user’s guide for the generation of Latin hypercube and random samples for use with computer models. Technical Report NUREG/CR-3624, SAND83-2365, U.S. Nuclear Regulatory Commission. Division of Risk Analysis and Operations., Washington, DC, USA, March 1984. vi + 50 + 10 pp. Available from GPO Sales Program, Division of Technical Information and Document Control, U.S. Nuclear Regulatory Commission; National Technical Information Service.

Iman:1985:FPU

[ISJ85] Ronald L. Iman, Michael J. Shortencarier, and Jay D. Johnson. A FORTRAN 77 program and user’s guide for the calculation of partial correlation and standardized regression coefficients. Technical Report ????, Division of Risk Analysis and Operations, Office of Nuclear Regulatory Research, U.S. Nuclear Regulatory Commission, Washington, DC, USA, June 1985. viii + 36 + 9 pp. Available from GPO Sales Program, Division of Technical Information and Document Control, U.S. Nuclear Regulatory Commission; National Technical Information Service.

Isner:1982:FPM

[Isn82] J. F. Isner. A Fortran programming methodology based on data abstraction. Comm. ACM, 25 (10):686–698, October 1982. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ISO:1988:IIPa

[ISO88] ISO. ISO 8651-1:1988: Information processing systems — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 1: FORTRAN. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 116 pp. LCCN ???? CHF 188. URL http://

www.ansi.org/;   http://www.
iso.ch/cate/d16024.html.

Iwan:1984:NUM

[Iwa84] W. D. (Wilfred D.) Iwan. NAT-FREQ users manual — a FOR-TRAN IV program for computing natural frequencies, mode shapes, and drag coefficients for taut strumming cables with attached masses and spring-mass combinations. Technical report, Naval Facilities Engineering Laboratory, Port Hueneme, CA, USA, 1984. xiv + 185 pp.

Incerti:1981:AFS

[IZP81] S. Incerti, F. Zirilli, and V. Parisi.

Algorithm 111: a FORTRAN subroutine for solving systems of nonlinear simultaneous equations. The Computer Journal, 24 (1):87–91, February 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/87.
tif; http://www3.oup.co.uk/
computer_journal/hdb/Volume_
24/Issue_01/tiff/88.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
89.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/90.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/
91.tif.

Jacob:1982:ROS

[Jac82] H. G. Jacob. Rechnergestuetzte Optimierung Statischer und Dynamischer Systeme. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-11641-9.

Jackson:1985:FSMb

[Jac85a] C. Jackson. A Fortran system to maintain a program library, adapt retrieve for retrieval of decks containing free format data. Computer Physics Communications, 38 (1):115, August/September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465585900505.

Jackson:1985:FSMa

[Jac85b] C. Jackson. A Fortran system to maintain a program library, adapt update to store decks containing free format data. Computer Physics Communications, 38(1):113–114, August/September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465585900499.

Jain:1984:CMS

[Jai84] Hemant Jain. A comprehensive model for the storage structure design of CODASYL databases. Information systems, 9(3), December 1984. CODEN INSYD6. ISSN 0306-4379 (print), 1873-6076 (electronic). ACM CR 8602-0146.

James:1986:FPVa

[Jam86a] J. H. James. Fortran program for vibration and sound radiation of spherical shell. Technical Report ARE-TM (N1) 86501, Admiralty Research Establishment; [ National Technical Information Service, distributor], Teddington, England, 1986. 22 pp.

James:1986:FPVb

[Jam86b] J. H. James. Fortran program for vibration and sound radiation of spherical shell. Technical Report ARE TM (N1) 86501, ARE, Teddington, Middlesex, England, 1986. 22 pp.

Janssens:1984:SPF

[Jan84] Hans Janssens. The strange power of FORTRAN: a very short calendar printing program. ACM SIGSOFT Software Engineering Notes, 9(1):50–51, January 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Janssens:1988:QPG

[Jan88] G. K. Janssens. Qualitätsprüfung der Pseudozufallszahlen-Generatoren von GPSS-Fortran. Angewandte Informatik (Elektron. Datenverarbeitung), 9:405–408, 1988. CO-DEN AWIFA7. ISSN 0013-5704.

JIS:1982:PLF

[Jap82] Japanese Industrial Standard. Programming Language FOR-TRAN, C6201 — 1982. ????, ????, 1982.

Jones:1984:ARI

[JBJ84] B. Jones, M. Banerjee, and L. Jones. Algorithm 118: Root isolation for transcendental equations. The Computer Journal, 27(2):184–187, May 1984. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.
oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/
184.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_02/tiff/185.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
27/Issue_02/tiff/186.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
02/tiff/187.tif.

Jost:1983:APA

[JBT83] A. C. Jost, R. Bolz, and G. Topping. Adventures in PASCAL — the academy approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 15(1):143–147, February 1983. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 14th SIGCSE Technical Symposium on Computer Science Education.

Jarvis:1982:EFP

[JC82] R. G. (Roger George) Jarvis and R. J. Cranston. EDDY, a FOR-TRAN program to extract significant features from eddy-current test data, the basis of the CAN-SCAN system. Technical report, Chalk River Nuclear Laboratories, Chalk River, Ontario, Canada, 1982. 32 pp.

Jones:1988:FTV

[JC88] Russell K. Jones and Tracy Crabtree. FORTRAN tools for VAX/ VMS and MS-DOS. John Wiley and Sons, New York, London, Sydney, 1988. ISBN 0-471-61976-0. xii + 447 pp. LCCN QA76.73.F25 J66 1988. US$29.95. Hash tables are used in the implementation of a FORTRAN preprocessor for macro statements.

Jones:1989:FTV

[JC89] Russell K. Jones and Tracy Crabtree. Fortran Tools: For VAX/ VMS and MS-DOS. John Wiley and Sons, New York, London, Sydney, February 1, 1989. ISBN 0-471-61976-0. 447 pp. LCCN ???? US$53.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471619760.

Jesshope:1982:PHD

[Jes82] C. R. Jesshope. Programming with a high degree of parallelism in Fortran. Computer Physics Communications, 26(3–4):237–246, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465582901114.

Jamieson:1987:CPA

[JGD87] Leah H. Jamieson, Dennis B. Gannon, and Robert J. Douglass, editors. The Characteristics of Parallel Algorithms, volume 30 of MIT Press series in scientific computation. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-10036-3. LCCN QA76.6 .C42981 1987.

Jazayeri:1986:OCH

[JH86] M. Jazayeri and M. Haden. Optimizing compilers are here (mostly). ACM SIGPLAN Notices, 21(5):61–63, May 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:1986:IRM

[Jia86] Benhuang Oliver Jiang. The implementation of rule modifier and definition of hypothetical parent FORTRAN. Thesis (M.S.), Auburn University, Auburn, AL, USA, 1986. ix + 95 pp.

Johanson:1982:MTH

[JK82] Robert C. Johanson and David Kliewer. Maintenance and testing of hydrological simulation program — FORTRAN (HSPF). Technical report, Hydrocomp, Inc., Mountain View, CA, USA, 1982. 85 pp.

Jones:1983:XSE

[JK83] Rondall E. Jones and David K. Kahaner. XERROR, the SLATEC error-handling package. Software —Practice and Experience, 13(3): 251–257, March 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jones:1981:XAF

[JL81a] C. B. Jones and R. I. Lawson. XRDPLT: A Fortran IV program for the graphical representation of X-ray powder diffraction data. Computers and Geosciences, 7:115–122, 1981. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Jones:1981:XFI

[JL81b] C. B. Jones and R. I. Lawson. XRDPLT: a Fortran IV program for the graphical representation of X-ray powder diffraction data. Computers and Geosciences, 7:115–122, 1981. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Johnson:1981:TFT

[Joh81] Carroll K. Johnson. OR TEP-II: a FORTRAN thermal-ellipsoid plot program for crystal structure illustrations. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1981. iv + 125 pp.

Johnson:1983:FRG

[Joh83] Evelyn C. Johnson. FORTRAN 77 reference guide: DOC4029-192. Prime Computer, Inc., Framingham, MA, USA, 3rd rev., rev. 19.2 edition, 1983. 213 pp.

Johnson:1984:FIP

[Joh84] Dean D. Johnson. FORTRAN instruction, PLATO versus lecture: a research paper. Paper (M.S.), North Dakota State University, Fargo, ND, USA, 1984. iv + 49 pp.

Johnson:1985:FPC

[Joh85a] C. H. J. Johnson. Fortran programming on the Cyber 205. Introductory guide / CSIRONET; no. 3 Introductory guide (CSIRONET); no. 3. CSIRONET, Canberra, Australia, 1985. ISBN 0-643-03891-4. 41 pp.

Johnson:1985:WFM

[Joh85b] Douglas Aubert Johnson. Weirel: a Fortran Monte Carlo program for the availability of a system with time-dependent instantaneous rates for failure and repair. Thesis (M.S.E. (nuclear engineering)), University of Washington, Seattle, WA, USA, 1985. viii + 105 pp.

Johnson:1986:CFF

[Joh86] Eric Stewart Johnson. A comparison of fast Fourier transform algorithms implemented in Fortran and 8086 assembly language on an IBM XT computer. Thesis (M.S.), San Diego State University, San Diego, CA, USA, 1986. viii + 178 pp.

Johnson:1987:FP

[Joh87a] A. S. Johnson. Fortran preprocessors. Computer Physics Communications, 45(1–3):275–281, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465587901640.

Johnson:1987:FSVa

[Joh87b] Peter J. Johnson. FORTRAN subroutines for VAX/VMS block I/O. Open-file report 87-641, U.S. Dept. of the Interior, Geological Survey; Books and Open-File Reports Section, distributor, Menlo Park, CA, USA, 1987. various pp.

Johnson:1987:FSVb

[Joh87c] Peter J. Johnson. FORTRAN subroutines for VAX/VMS block I/O. Open-file report 87-641, U.S. Geological Survey, Menlo Park, CA, USA, 1987. various pp.

Jordan:1986:SPA

[Jor86] Harry F. Jordan. Structuring parallel algorithms in an MIMD, shared memory environment. Parallel Computing, 3(2):93–110, May 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jachens:1981:DFP

[JR81] Robert C. Jachens and Carter W.

Roberts. Documentation of a FORTRAN program, “isocomp”, for computing isostatic residual gravity. Open-file report 81-574, U.S. Geological Survey, Menlo Park, CA, USA, 1981. 26 pp.

Jesshope:1989:CME

[JR89] C. R. Jesshope and K. D. Reinartz, editors. CONPAR 88 (1988: Manchester, England), The British Computer Society workshop series. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-37177-5. LCCN QA76.5 .C6194 1988.

Jaensch:1988:MEA

[JRS88a] Christian R. Jaensch, Ulrich Ruede, and Klaus Schnepper. Macro expansion, A tool for the systematic development of scientific software. Technical report, Technische Universität München (??), München, Germany, 1988.

Jaensch:1988:MET

[JRS88b] Christian R. Jaensch, Ulrich Ruede, and Klaus Schnepper. Macro expansion, a tool for the systematic development of scientific software. Technical report, Technische Universität München (??), München, Germany, 1988.

Jain:1988:FAS

[JS88] R. K. Jain and R. P. Suri. Fortran 77: with applications to science and engineering. Tata McGraw-Hill, New Delhi, India, 1988. ISBN 0-07-460164-4. viii + 191 pp.

James:1985:ANM

[JSW85a] M. L. (Merlin L.) James, G. M. Smith, and J. C. Wolford. Applied Numerical Methods for Digital Computation. Harper & Row, New York, NY, USA, third edition, 1985. ISBN 0-06-043281-0. xiii + 753 pp. LCCN QA297 .J3 1985.

James:1985:SMA

[JSW85b] M. L. (Merlin L.) James, G. M.

(Gerald M.) Smith, and J. C. Wolford. Solutions manual to accompany applied numerical methods for digital computation with FORTRAN and CMSP. Harper & Row, New York, NY, USA, third edition, 1985. ISBN 0-06-363256-X. 426 pp.

Jain:1988:EAD

[JT88] S. Jain and C. Thompson. An efficient approach to data flow analysis in a multiple pass global optimizer. ACM SIGPLAN Notices, 23(7):154–163, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org:80/pubs/
citations/proceedings/pldi/
53990/p154-jain/.

Jurs:1986:CSA

[Jur86] Peter C. Jurs. Computer Software Applications in Chemistry. John Wiley and Sons, New York, London, Sydney, December 1986. ISBN 0-471-84735-6. xiv + 253 pp. LCCN QD39.3.E46 J8731 1986. US$37.50 (est.); US$79.95.

Justice:1988:EFS

[Jus88] Ernest David Justice. An environment for FORTRAN software development on the personal computer. Thesis (M.S.C.S.E.), University of Arkansas, Fayetteville, Fayetteville, AR, USA, 1988. viii + 123 pp.

Jansen:1986:HAA

[JW86] Paul Jansen and Peter Weidner. High-accuracy arithmetic software—some tests of the ACRITH problem-solving routines. ACM Transactions on Mathematical Software, 12(1):62–70, March 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Yi:1989:FCH

[jYsS89] Pyong jae Yi and Hye suk So. Flow chart haesol BASIC/FORTRAN programming. Toso Sanopsa, Soul, Korea, 4-pan edition, 1989. ISBN ???? 403 pp. LCCN ????

Kahaner:1980:AFI

[Kah80] D. K. Kahaner. Algorithm 561: FORTRAN implementation of heap programs for efficient table maintenance [Z]. ACM Transactions on Mathematical Software, 6 (3):444–449, September 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaller:1985:VBG

[Kal85a] L. Kaller. Voruebersetzer für die Benutzung von grossen Reihungen (¿32k) auf 16-Bit-Rechnern in Fortran. Studienarbeit, 1985.

Kallin:1985:F

[Kal85b] Sten Kallin. FORTRAN. Computacion 389. El Ateneo, Buenos Aires, Argentina, 2. edition, 1985. ISBN 950-02-5224-4. 157 pp.

Kantaris:1988:PF

[Kan88] Noel Kantaris. Programming in Fortran 77. Babani, London, UK, 1988. ISBN 0-85934-195-X (paperback). 140 pp.

Karp:1986:PP

[Kar86] Alan H. Karp. Programming for parallelism. Technical Report G320-3490, IBM Palo Alto Scientific Center, San Jose, CA, USA, June 1986.

Karjala:1987:ACP

[Kar87a] E. Karjala. APL and SAS in cooperation. In Leroy J. Dickey and Lynne C. Shaw, editors, APL 88 Conference proceedings: APL, its past, its present, its future: 1-5 February 1988, Sydney, Australia, Members US$21.00, pages 202–208. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-253-5. LCCN QA 76.73 A27 A195 v.18 no.2 1988. US$28.00.

Karjala:1987:ASC

[Kar87b] E. Karjala. APL and SAS in cooperation. In Leroy J. Dickey and Lynne C. Shaw, editors, APL 88 Conference proceedings: APL, its past, its present, its future: 1– 5 February 1988, Sydney, Australia, Members US$21.00, pages 202–208. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-253-5. LCCN QA 76.73 A27 A195 v.18 no.2 1988. US$28.00.

Karp:1987:PP

[Kar87c] Alan H. Karp. Programming for parallelism. Computer, 20(5):43– 57, May 1987. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Katzan:1982:F

[Kat82] Harry Katzan. Fortran 77. Van Nostrand Reinhold, New York, NY, USA, February 1982. ISBN 0-442-25428-8. ???? pp. LCCN ???? US$14.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0442254288.

Kawabata:1984:SFP

[Kaw84] Hugh M. Kawabata. Savage floating point benchmark in Fortran in 16BST. Dr. Dobb’s Journal of Software Tools, 9(1):83–??, January 1984. CODEN DDJOEB. ISSN 1044-789X.

Karp:1988:CPF

[KB88] Alan H. Karp and Robert G. Babb. A comparison of 12 parallel Fortran dialects. IEEE Software, 5(5): 52–67, September 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kawamura:1986:NAC

[KBRM^{+}86] K. Kawamura, G. Beale, J. Rodriguez-Moscoso, B-J. Hsieh, S. Padalkar, M. Johnson, F. Vinz, and K. R. Fernandez. NESS: a coupled simulation expert system. In International Symp. on Methodologies for Intelligent Systems; University Tennessee. ????, ????, October 1986.

Kreitzberg:1984:IF

[KC84a] Charles B. Kreitzberg and Norman T. Carpenter. Introduction to Fortran. Books for professionals Harcourt Brace Jovanovich college outline series. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, October 1984. ISBN 0-15-600036-9 (paperback). viii + 211 pp. LCCN ???? US$12.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0156000369.

Kreitzberg:1984:IFH

[KC84b] Charles B. Kreitzberg and Norman T. Carpenter. Introduction to Fortran (Harcourt Brace Jovanovich College Outline Series). Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, October 1984. ISBN 0-15-600036-9. ???? pp. LCCN ???? US$12.25.

Keith:1984:FPPa

[KD84a] Laura A. Keith and Paul T. Delaney. FORTRAN programs for properties of water and steam, and solubilities of solid silica phases. Open-file report 84-690, U.S. Dept. of the Interior, Geological Survey, Flagstaff, AZ, USA, 1984. ???? pp.

Keith:1984:FPPb

[KD84b] Laura A. Keith and Paul T. Delaney. FORTRAN programs for properties of water and steam, and solubilities of solid silica phases. Open-file report 84-690, U.S. Geological Survey, Menlo Park, CA, USA, 1984. iii + 58 pp.

Kee:1988:FCCb

[Kee88a] R. J. Kee. A FORTRAN computer code package for the evaluation of gas-phase, multicomponent transport properties. Sandia report; sand86-8246, Sandia National Laboratories, Albuquerque, NM, USA, 1988. 46 pp.

Kee:1988:FPM

[Kee88b] R. J. Kee. A FORTRAN program for modeling steady laminar one-dimensional premixed flames. Sandia report; sand85-8240, Sandia National Laboratories, Albuquerque, NM, USA, 1988. 116 pp.

Kempi:1985:FPR

[Kem85] V. Kempi. A FORTRAN program for ranking and for calculation of Spearman’s correlation coefficient. Computer Methods and Programs in Biomedicine, 21(2 123-):5–??, November 1985. CODEN CMP-BEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Kempi:1987:FPD

[Kem87] V. Kempi. A FORTRAN program for deconvolution analysis using the matrix algorithm method with special reference to renography. Computer Methods and Programs in Biomedicine, 24(2 107-): 16–??, April 1987. CODEN CMP-BEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Kerridge:1982:FIC

[Ker82] Jon M. Kerridge. A FOR-TRAN implementation of Concurrent PASCAL. Software—Practice and Experience, 12(1):45–55, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kettleborough:1982:NF

[Ket82] Ian Kettleborough. Nevada Fortran, 1982.

Kettleborough:1984:UF

[Ket84] Ian D. Kettleborough. Utah Fortran, 1984.

Kettleborough:1985:UF

[Ket85a] Ian D. Kettleborough. Utah Fortran, 1985.

Kettleborough:1985:UFP

[Ket85b] Ian D. Kettleborough. UTAH FORTRAN programmer’s reference manual. Ellis Computing, San Francisco, CA, USA, 1985. ISBN 0-923877-05-3. 222 pp.

Keys:1981:ICP

[Key81] William J. Keys. Introduction to computer programming basic Fortran IV: a practical approach. Anaheim Pub. Co., Fullerton, CA, USA, 1981. ISBN ???? various pp. LCCN ????

Koffman:1987:PSS

[KF87] Elliot B. Koffman and Frank L.

Friedman. Problem solving and structured programming in FOR-TRAN 77. Addison-Wesley, Reading, MA, USA, third edition, 1987. ISBN 0-201-11561-1 (paperback). xiv + 518 + 9 pp. LCCN QA76.73.F25 K64 1987 Sci-Eng. US$26.95. Rev. ed. of: Problem solving and structured programming in FORTRAN / Frank L. Friedman. Includes index.

Koffman:1988:PSS

[KF88] Elliot B. Koffman and Frank L. Friedman. Problem solving and structured programming in FOR-TRAN 77. Addison-Wesley, Reading, MA, USA, third edition, 1988. ISBN 0-201-11561-1. xiv + 518 + 9 pp.

Kincaid:1981:IFP

[KGRY81] David R. Kincaid, Roger G. Grimes, John R. Respess, and David M. Young. ITPACK 2B: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. Report CNA-173, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, September 1981. (Also, Report CCSN–44, Computation Center, University of Texas at Austin.).

Huang:1984:FCH

[kH84] Ping kang Huang. FORTRAN 77 cheng hsu she chi. Shang-hai ko hsueh chi shu wen hsien chu pan she: Hsin hua shu tien Shang-hai fa hsing so fa hsing, Shang-hai shih, ti 1 pan edition, 1984. ISBN ???? 2 + 2 + 335 pp. LCCN ????

Khailany:1981:BPF

[Kha81] Asad S. O. Khailany. Business programming in FORTRAN IV and ANSI FORTRAN 77: a structured approach. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1981. ISBN 0-13-107607-8 (paperback). xvii + 440 pp. LCCN HF5548.5.F2 K48. US$36.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0131076078.

Kiefhaber:1983:IES

[Kie83] Sarah Hildebrandt Kiefhaber. An implementation and evaluation of a screen debugger for Fortran programs. Thesis (M.S.), University of Colorado, Boulder, CO, USA, 1983. vi + 195 pp.

Kieffer:1986:FPC

[Kie86] Susan W. Kieffer. FORTRAN program for calculation of thermodynamic properties of minerals from vibrational spectra. Open-file report 86-475, U.S. Geological Survey, Menlo Park, CA, USA, 1986. 47 pp.

Kim:1986:ETF

[Kim86] Jung Sik Kim. Evaluation of the TOOLPACK Fortran programming environment. Thesis (M.S. in computer science), Naval Postgraduate School, Monterey, CA, USA, 1986. 120 pp. ADA173943.

Kipp:1982:CPE

[Kip82] G. W. Kipp. Combat performance evaluation of fighter aircraft: a suite of FORTRAN IV programs based on energy manoeuvrability theory. Mechanical engineering report; 160, Aeronautical Research Laboratories, Melbourne, Victoria, Australia, 1982. ISBN 0-642-86965-0. 101 + [43] pp.

Kirstein:1985:IFT

[Kir85] Dean Allen Kirstein. An interactive Fortran tutorial for the IBM personal computer. Project (M.S.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1985. v + 167 pp.

Kirk:1989:UFS

[Kir89] Bernadette L. Kirk. UP-DATE, a FORTRAN 77 source file manipulator: adapted for the Data General MV Series Eclipse computers under AOS/VS. Technical Report NUREG/CR-4478, ORNL/TDMC-4, NRC FIN A9100, Division of Information Support Services, Office of Administration and Resources Management, U.S. Nuclear Regulatory Commission, Washington, DC, USA, January 1989. iii + 89 pp.

Kee:1989:CGP

[KJM89] R. J. Kee, Thomas H. Jefferson, and James A. (James Angus) Miller. CHEMKIN: a generalpurpose, problem-independent, transportable, FORTRAN chemical kinetics code package. Sandia report; sand80-8003, Sandia National Laboratories, Albuquerque, NM, USA, 1989. 204 pp.

Kaneko:1989:PFS

[KK89a] Toshiaki Kaneko and Setsuya Kawabata. A preprocessor for Fortran source code produced by REDUCE. Computer Physics Communications, 55(2): 141–147, September 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Klappholz:1989:CCF

[KK89b] David Klappholz and Xiangyun Kong. CFTP (Cray Fortran/ANSI FORTRAN ’77 prefine): a tool to aid in hand-parallelizing sequential code. Digest of Papers — IEEE Computer Society International Conference, pages 92–97, February 1989. CODEN DCSIDU. ISBN 0-8186-1909-0. Available from IEEE Service Cent. Piscataway.

Kwok:1989:RPM

[kK89c] Yue kuen Kwok. A regular perturbation method for subcritical flow over a two-dimensional airfoil. IMA Journal of Applied Mathematics, 43(1):71–81, ???? 1989. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Klappholz:1989:RFU

[KKK89] D. Klappholz, X. Kong, and A. D.

Kallis. Refined Fortran: An update. In ACM [ACM89b], pages 607–615. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Kagiwada:1985:NDN

[KKRK85] Harriet Kagiwada, Robert Kalaba, Nima Rasakhoo, and Spingarn Karl. Numerical Derivatives and Nonlinear Analysis, volume 31 of Mathematical Concepts and Methods in Science and Engineering. Plenum Press, Inc., New York, NY, USA, 1985. ISBN 0-306-42178-X. x + 201 pp. LCCN QA297.N854 1986.

Kiesling:1983:PF

[KL83a] I. Kiesling and M. Lowes. Programmierung mit Fortran 77.

Teubner, Stuttgart, Leipzig, second edition, 1983. ISBN 3-519-10089-4. ???? pp.

Kiesling:1983:PMF

[KL83b] I. Kiesling and M. Lowes. Programmierung mit Fortran 77. Teubner, Stuttgart, Leipzig, second edition, 1983. ISBN 3-519-10089-4. ???? pp.

Kiessling:1985:PF

[KL85a] I. Kiessling and M. Lowes. Programmierung mit Fortran 77. Teubner, Stuttgart, Leipzig, 1985. ISBN 3-519-20089-9. ???? pp.

Kiessling:1985:PMF

[KL85b] I. Kiessling and M. Lowes. Programmierung mit Fortran 77. Teubner, Stuttgart, Leipzig, 1985. ISBN 3-519-20089-9. ???? pp.

Klein:1989:UGFb

[Kle89a] Fred W. Klein. User’s guide to five VAX FORTRAN programs for manipulating HYPOIN-VERSE summary and archive files: SELECT, EXTRACT, SUM-LIST, ARCPRINT and FORCON. Open-file report 89-313, U.S. Dept. of the Interior, Geological Survey: [Books and Open-File Reports Section, distributor], Denver, Colo.?, 1989. 1 microfiche pp.

Klein:1989:UGFa

[Kle89b] Fred W. Klein. User’s guide to five VAX FORTRAN programs for manipulating summary and archive files: SLECT, EXTRACT, SUM-LIST, ARCPRINT and FORCON. Open-file report 89-313, U.S. Dept. of the Interior, U.S. Geological Survey, Menlo Park, CA, USA, 1989. 11 pp.

Kliewer:1989:HPP

[Kli89] B. D. Kliewer. HOOPS: Powerful portable 3-D graphics. BYTE Magazine, 14(7):193–194, 196, 198, July 1989. CODEN BYTEDJ. ISSN 0360-5280.

Kulisch:1983:NAS

[KM83] U. Kulisch and W. Miranker, editors. A new approach to scientific computation: Proceedings of the Symposium on a New Approach to Scientific Computation (1982: IBM Thomas J. Watson Research Center). Academic Press, New York, NY, USA, 1983. ISBN 0-12-428660-7. LCCN QA297 .S847 1982.

Koonin:1989:CPF

[KM89] Steven E. Koonin and Dawn Meredith. Computational physics (FORTRAN version). Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-12779-2. xvi + 636 pp.

Kahaner:1989:NMS

[KMN89] David Kahaner, Cleve Moler, and Stephen Nash. Numerical Methods and Software. Prentice-Hall series in computational mathematics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1989. ISBN 0-13-627258-4. xii + 495 pp. LCCN TA345 .K341 1989. US$50.

Knaak:1984:FAH

[Kna84] U. Knaak. Fortran-Programmpaket zur Ansprache einer Hardware-Testschnittstelle. Studienarbeit, 1984.

Kneis:1981:DSI

[Kne81] Wilfried Kneis. Draft standard industrial real-time FORTRAN. ACM SIGPLAN Notices, 16(7): 45–60, July 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knuth:1984:FIL

[Knu84] Donald E. Knuth. FORTRAN implementations (letter). Annals of the History of Computing, 6(4): 402–403, October/December 1984. CODEN AHCOE5. ISSN 0164-1239.

Keller:1987:SSF

[KODG^{+}87] J. P. Keller, E. Omodeo, V. Donzeau-Gouge, E.-E. Doberkat, and P. Spirakis. The Sed Software Factory. Technical report, Universität Hildesheim (??), Hildesheim, Germany, 1987.

Kolbig:1982:CE

[Köl82] K. S. Kölbig. Closed expressions for ∫_{0}^{1} t^{−1} log^{n−1t} log^{p}(1 − t)dt. Mathematics of Computation, 39(160):647–654, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kowalik:1985:PMC

[Kow85] J. S. Kowalik, editor. Parallel MIMD Computation: the HEP Supercomputer and Its Applications. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-11101-2. 411 pp. LCCN QA76.8.D436 P37 1985.

Kaylen:1986:MFS

[KP86] Michael S. Kaylen and Paul V.

Preckel. MINTDF: a FORTRAN subroutine for computing parametric integrals. Aewp; 1986-22 agricultural economics working paper; 1986-22, University of Missouri-Columbia, Dept. of Agricultural Economics, Columbia, MO, USA, 1986. ii + [15] pp.

Kreutz:1986:PSO

[Kre86a] Gregg Kreutz. Problem Solving for Oil Painters. Watson-Guptill Pubns, ????, November 1, 1986. ISBN 0-8230-4408-4. 144 pp. LCCN ???? US$24.95.

Kreutz:1986:POP

[Kre86b] Gregg Kreutz. Problemsolving for oil painters. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1986. ISBN 0-8230-4408-4. 144 pp. LCCN ND1500 .K74 1986.

Krebs:1988:FPE

[Kre88] Charles J. Krebs. FORTRAN programs for Ecological methodology. Krebs, Vancouver, BC, Canada, 1988. ISBN ???? 156 pp. LCCN ????

Krishnan:1983:RTF

[Kri83] G. Krishnan. Real time Fortran compatible routines for Apple II plus computers. Thesis (M.S.), Department of Chemical and Environmental Engineering, University of Louisville, Louisville, KY, USA, 1983. vii + 84 pp.

Katz:1988:MSC

[KRW88] S. Katz, W. A. Ray, and G. Walder. Multiprocessor software for the CYBERPLUS high performance system. Parallel Computing, 8(1–3):231–244, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kincaid:1982:IAF

[KRYG82a] D. R. Kincaid, J. R. Respess, D. M. Young, and R. G. Grimes. ITPACK 2C: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8 (3):302–322, September 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kincaid:1982:AIA

[KRYG82b] David R. Kincaid, John R. Respess, David M. Young, and Roger G. Grimes. Algorithm 586: ITPACK 2C: A FORTRAN package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8 (3):302–322, September 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kincaid:1982:AIF

[KRYG82c] David R. Kincaid, John R. Respess, David M. Young, and Roger G. Grimes. Algorithm 586: ITPACK 2C: A FORTRAN package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8 (3):302–322, September 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kincaid:1982:IFP

[KRYG82d] David R. Kincaid, John R. Respess, David M. Young, and Roger G. Grimes. ITPACK 2C: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8(3):302–322, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). Algorithm 586.

Kanada:1981:LSS

[KS81a] Y. Kanada and T. Sasaki. LISPbased ‘big-float’ system is not slow. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (2):123–119, May 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kanada:1981:LBB

[KS81b] Yasumasa Kanada and Tateaki Sasaki. LISP-based “big-float” system is not slow. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(2):13–19, May 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Knodel:1981:PPR

[KS81c] Walter Knödel and Hans Jürgen Schneider, editors. Parallel processes and related automata = Parallele Prozesse und damit zusammenhängende Automaten, volume 3 of Computing. Supplementum. Springer, Wien / New York, 1981. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Kreitzberg:1982:FP

[KS82a] Charles B. Kreitzberg and Ben Shneiderman. Fortran Programming. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, second edition, April 1982. ISBN 0-15-528012-0. x + 437 pp. LCCN ???? US$14.63. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0155280120.

Kreitzberg:1982:FPS

[KS82b] Charles B. Kreitzberg and Ben Shneiderman. FORTRAN programming, a spiral approach: compatible with WATFOR/WATFIV and FORTRAN 77. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, second edition, April 1982. ISBN 0-15-528015-5 (paperback). x + 437 pp. LCCN ???? US$26.60. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0155280155.

Kulisch:1988:SCA

[KS88] Ulrich Kulisch and Hans J. Stetter, editors. Scientific computation with automatic result verification, volume 6 of Computing. Supplementum. Springer, Wien / New York, 1988. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Knuth:1984:CQD

[KTW84] Donald E. Knuth, Erwin Tomash, and M. R. Williams. Comments, queries, and debate: FORTRAN implementations; the Madrid promptuary; Edouard Lucas vindicated; photo essays. Annals of the History of Computing, 6(4):402–405, October/ December 1984. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a4402.pdf;
 http://www.computer.org/annals/
an1984/a4402abs.htm.

Kulisch:1983:FES

[Kul83] U. Kulisch. A FORTRAN extension for scientific computation. In Kulisch and Miranker [KM83], pages 199–223. ISBN 0-12-428660-7. LCCN QA297 .S847 1982.

Kumar:1986:PF

[Kum86] Ram Kumar. Programming with Fortran 77. Tata McGraw-Hill, New Delhi, India, 5eme reimpr. (1990) edition, 1986. ISBN 0-07-451859-3. xiv + 386 pp.

Kurbel:1985:PPC

[Kur85] Karl Kurbel. Programmierstil in Pascal, Cobol, Fortran, Basic, PL/I. Springer Compass. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-15138-9 (U.S.). xii + 328 pp.

Kuck:1984:DRF

[KW84] David J. Kuck and Michael Wolfe. A debate: Retire FORTRAN?: No. Physics Today, 37(5):67–75, May 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

                                        physicstoday.org/resource/1/
                                        phtoad/v37/i5/p67_s1.

Kaagstrom:1987:GRG

[KW87a] B. Kågström and L. Westin. GSYLV—Fortran routines for the generalized Schur method with dif^{−1}-estimators for solving the generalized Sylvester equation. Report UMINF-132.86, Institute of Information Processing, University of Umeå, Umeå, Sweden, 1987.

Kaagstrom:1987:GFR

[KW87b] B. Kågström and L. Westin. GSYLV- Fortran routines for the generalized Schur method with dif^{−1} estimators for solving the generalized Sylvester equation. Technical Report UMINF-132.86, Information Processing, University of Umeå, S-901 87 Umeå, Sweden, 1987.

Kuehme:1987:F

[KW87c] Thomas Kuehme and Peter Witschital. Die Fortran-fibel. R. Oldenbourg, München, Germany, 1987. ISBN 3-486-20362-2. ???? pp.

Kramer:1989:FSF

[KW89] W. Krämer and W. Walter.

FORTRAN-SC: a FORTRAN extension for engineering/scientific computation with access to ACRITH: General information notes and sample programs. Technical report, IBM Deutschland GmbH, ????, 1989.

Kee:1988:FCCa

[KWM88] R. J. Kee, J. Warnatz, and James A. (James Angus) Miller. A FOR-TRAN computer code package for the evaluation of gas-phase viscosities, conductivities, and diffusion coefficients. Sandia report; sand83-8209, Sandia National Laboratories, Albuquerque, NM, USA, 1988. 37 pp.

Kent:1986:TAM

[KWWK86] D. C. (Douglas C.) Kent, Jan Wagner, Fred E. Witz, and Robert S. Kerr. Two-dimensional analytical model (FORTRAN) for prediction of contaminant movement in ground water. Technical report, Robert S. Kerr Environmental Research Laboratory, Ada, OK, USA, 1986. various pp.

LaPalmaObservatory:1987:AFP

[La 87] La Palma Observatory. Adam Fortran programmers manual for the Isaac Newton telescope: May 1987. Royal Greenwich Observatory, Herstmonceux, ??, 1987. various pp.

Laguitton:1985:SMC

[Lag85] D. (Daniel) Laguitton. The SPOC manual: chapter 5.2, unit models (part C): unit models and FORTRAN simulators of ore and coal process equipment; miscellaneous. [CANMET special publication]; SP85-1/5.2E CANMET special publication; SP85-1/5.2E. Canada Centre for Mineral and Energy Technology, Ottawa, Ontario, Canada, 1985. ISBN 0-660-11866-1. ix + 63 pp.

Lahey:1986:F

[Lah86] Lahey Computer Systems, Inc. FORTRAN 77, 1986. ca. 1 data file + 18 program files on 1 computer disk + ca. 2 data files, 23 program files on 1 computer disk + advertising material + 1 disk mailer.

Lahey:1987:FS

[Lah87a] T. Lahey. The Fortran 8X standard. ACM SIGPLAN FOR-TRAN Forum, 6(3):27–30, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Lahey:1987:LPF

[Lah87b] Lahey Computer Systems, Inc. Lahey personal Fortran 77, 1987.

Lahey:1988:LFF

[Lah88a] Lahey Computer Systems, Inc. Lahey F77L Fortran 77, 1988.

Lahey:1988:LPFa

[Lah88b] Lahey Computer Systems, Inc. Lahey personal Fortran 77: full FOR-TRAN 77 language system for PCs. Lahey Computer Systems, Incline Village, NV, USA, b, June 1988 edition, 1988. various pp.

Lahey:1988:LPFb

[Lah88c] Lahey Computer Systems, Inc. Lahey personal Fortran 77 full FOR-TRAN 77 language system for PCs, 1988.

Lahey:1988:LPFc

[Lah88d] Lahey Computer Systems, Inc. Lahey personal Fortran 77 toolkit reference manual, 1988.

Lamprecht:1981:EPF

[Lam81] Guenther Lamprecht. Einführung in die Programmiersprache Fortran 77. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1981. ISBN 3-528-03345-2. ???? pp.

Lampton:1984:FB

[Lam84] Christopher Lampton. Fortran for Beginners. A Computer literacy skills book. Franklin Watts, New York, NY, USA, March 1984. ISBN 0-531-04747-4. 106 pp. LCCN ???? US$17.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0531047474.

Lamprecht:1986:IF

[Lam86] Gunther Lamprecht. Introduction to Fortran 77. Friedrich Vieweg und Sohn, Braunschweig, Germany, December 1986. ISBN 3-528-03360-6 (paperback). 150 pp. LCCN QA76.73.F25L35 1986. US$22.00; US$28.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

3528033606. Translated from the German.

Lam:1989:FPC

[Lam89] S. S. W. Lam. A FORTRAN program for the calculation of the calibration coefficients of a six-component strain gauge balance. Flight mechanics technical memorandum; 410, Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratory, Melbourne, Victoria, Australia, 1989. ISBN 0-642-15255-1 (not printed on item). ii + 46 pp.

Lang:1988:SCB

[Lan88] U. Lang. Simple CPU-benchmarks in Fortran 77 on different machines. Supercomputer, 5(6):26– 33, November 1988. CODEN SP-COEL. ISSN 0168-7875.

Larmouth:1981:FP

[Lar81] J. Larmouth. Fortran 77 portability. Software—Practice and Experience, 11(10), October 1981. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lau:1986:CHA

[Lau86] H. T. (Hang Tong) Lau. Combinatorial Heuristic Algorithms With Fortran, volume 280 of Lecture notes in economics and mathematical systems ; 280. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., November 1986. ISBN 0-387-17161-4 (paperback). viii + 126 pp. LCCN QA402.5.L37 1986. US$25.60; DM29.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0387171614.

Lavigne:1983:DFI

[Lav83] A. Lavigne. Du Fortran IV au Fortran V: A.N.S.I. 77. ABC des langages. Masson, Masson, France, 1983. ISBN 2-225-79904-0. 87 pp.

Law:1983:AFI

[Law83] Victor J. Law. ANSI FORTRAN 77: An Introduction to Structured Software Design. Wm. C. Brown Publishers, Dubuque, IA, USA, January 1983. ISBN 0-697-08167-2. ???? pp. LCCN ???? US$47.85. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0697081672.

Lawson:1988:SRR

[Law88] Charles L. Lawson. Series reversion as the reversed chain rule. ACM SIGNUM Newsletter, 23(1): 7–9, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lang:1986:CPN

[LB86] B. F. Lang and G. Burger. A collection of programs for nucleic acid and protein analysis, written in FORTRAN 77 for IBM-PC compatible microcomputers. Nucleic Acids Research, 14(1):455– 465, January 10, 1986. CODEN NARHAD. ISSN 0305-1048.

Leredde:1989:DDS

[LB89] Henri Leredde and Helene Bigot.

Définition d’un sous-ensemble de règles pour la réalisation d’un logiciel portable en Fortran 77: version 2.1: normes modulad. Document normalisation. Institut national de recherche en informatique et en automatique, Le Chesnay, France, 1989. ISBN 2-7261-0594-7. 50 pp.

Luecke:1988:CSK

[LCH^{+}88] G. Luecke, J. Coyle, W. Haque, J. Hoekstra, H. Jespersen, and R. Schmidt. A comparative study of KAP and VAST-2: two automatic vector preprocessors with Fortran 8x output. Supercomputer, 5(6):15–25, November 1988. CO-DEN SPCOEL. ISSN 0168-7875.

Lacoume:1988:SPI

[LCMM88] J. L. Lacoume, A. Chehikian, N. Martin, and J. Malbos, editors. Signal Processing IV: Theories and Applications Proceedings of EUSIPCO-88. Fourth European Signal Processing Conference (Grenoble, France, September 5–8, 1988). Elsevier, Amsterdam, The Netherlands, 1988. ISBN 0-444-70516-3. LCCN TK5102.5.E9 1988. Three volumes.

Li:1983:FCH

[LcY83] Mu Li and Chia chuan Yang. FOR-TRAN cheng hsu she chi tao yin. Liao-ning ko hsueh chi shu chu pan she: Liao-ning sheng hsin hua shu tien fa hsing, Shen-yang, ti 1 pan edition, 1983. ISBN ???? 2 + 2 + 354 pp. LCCN ????

LePera:1987:FVO

[LD87] T. A. LePera and H. E. Donley. FORTRAN: a viable option for the introductory computer science course. ACM SIGPLAN Notices, 22(3):102–110, March 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Le:1983:CAI

[Le83] Gia-Loi Thi Le. A computer assisted instruction system to teach the Fortran language. Thesis (M.S.), University of Houston — University Park, Houston, TX, USA, 1983. viii + 112 pp.

Lecot:1989:AGL

[Lec89] C. Lecot. An algorithm for generating low discrepancy sequences on vector computers. Parallel Computing, 11(1):113–116, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lee:1984:CFG

[Lee84a] Baiksun Lee. A comparison of FORTRAN, GPSS, and SLAM on discrete simulation mo[d]els. Thesis (M.S.), Dept. of Industrial Engineering, Cullen College of Engineering, University of Houston — University Park, Houston, TX, USA, 1984. 162 pp.

Lee:1984:ABF

[Lee84b] J. A. N. Lee. An annotated bibliography of FORTRAN. Annals of the History of Computing, 6(1):49– 58, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1049.pdf;
http://www.computer.org/annals/
an1984/a1049abs.htm.

Leeson:1984:IFE

[Lee84c] Daniel N. Leeson. IBM FOR-TRAN exhibit and film. Annals of the History of Computing, 6(1):41– 48, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1041.pdf;
 http://www.computer.org/annals/
an1984/a1041abs.htm.

Lee:1985:ELC

[Lee85] Tae Wuk Lee. The effects of library-based computer assisted instruction on student programming performance and an implementation of the grading support system in FORTRAN language courses at the university. Thesis (Ph.D.), Florida Institute of Technology, Melbourne, FL, USA, 1985. ix + 149 pp.

Lehmkuhl:1983:F

[Leh83a] Nonna Lehmkuhl. Fortran 77. MacMillan Publishing Company, New York, NY, USA, January 1983. ISBN 0-02-369390-8. ???? pp. LCCN ???? US$21.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0023693908.

Lehmkuhl:1983:FTA

[Leh83b] Nonna Kliss Lehmkuhl. Fortran 77, a top-down approach. Collier Books, Macmillan Publishing Company, New York, NY, USA, 1983. ISBN 0-02-369390-8 (paperback). xiv + 466 pp. LCCN QA76.73.F25L44 1983.

Lehman:1986:PSS

[Leh86] Richard S. Lehman, editor. Programming for the Social Sciences: Algorithms and Fortran 77 Coding. Computer Science for the Behavioral Sciences. Lawrence Erlbaum Associates, Hillsdale, N.J., November 1986. ISBN 0-89859-588-6. xvi + 575 pp. LCCN QA76.6 .L44 1986. US$99.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0898595886.

Leigh:1987:FES

[Lei87] William Leigh. Fortran: The Engineers and Science Language. MacMillan Publishing Company, New York, NY, USA, February 1987. ISBN 0-938188-67-4. ???? pp. LCCN ???? US$28.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0938188674.

Leppin:1986:UAS

[Lep86] R. Leppin. Ueberwachung und Auswertung eines Simulationslaufes. Studienarbeit, 1986.

Lerner:1983:FPW

[Ler83] Doug Lerner. FORTRAN programming: WATFOR on WITS, with Pascal. Kendall/Hunt Pub., Dubuque, IA, USA, second edition, January 1983. ISBN 0-8403-2932-6. 167 pp. LCCN ???? US$11.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0840329326.

Levy:1989:AFP

[Lev89] Stephen H. Levy. Ada, the Fortran Programmer’s Companion. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, October 1989. ISBN 0-929306-02-3. various pp. LCCN ???? US$27.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0929306023.

Lewis:1981:PPF

[Lew81a] William E. Lewis. Problemsolving principles for FORTRAN programmers: applied logic, psychology, and grit. Hayden Book Co., Rochelle Park, NJ, USA, 1981. ISBN 0-8104-5430-0. 177 pp. LCCN QA76.73.F25 L48x.

Lewis:1981:PSP

[Lew81b] William E. Lewis. Problem-Solving Principles for Fortran Programmers: Applied Logic, Psychology and Grit. Howard W. Sams, Indianapolis, IN 46268, USA, April 1981. ISBN 0-8104-5430-0. ???? pp. LCCN ???? US$15.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0810454300.

Lewis:1982:F

[Lew82a] Anthony Lewis. FORTRAN. Prime Computer, Inc., Framingham, MA, USA, revision 18 edition, 1982. 40 pp.

Lewis:1982:FRG

[Lew82b] Anthony Lewis. The FORTRAN reference guide. Prime Computer, Inc., Framingham, MA, USA, 1982. 97 pp.

LeGland:1986:SNE

[LG86] Francois LeGland and Antoine Gondel. Systematic numerical experiments in nonlinear filtering with automatic Fortran code generation. Proceedings of the IEEE Conference on Decision and Control Including The Symposium on Adaptive Pro, pages 638–642, 1986. CODEN PCDCDZ. ISSN 0191-2216. IEEE Service Cent. Piscataway, NJ, USA.

Lewis:1981:BFS

[LH81] Robert Jay Lewis and David G.

Hart. Business FORTRAN: a structured approach. Wadsworth, Pacific Grove, CA, USA, 1981. ISBN 0-534-00778-3. xiv + 457 pp.

Leigh:1987:SFb

[LH87] William Leigh and David Huffman. Structured Fortran ’77. Mitchell Publishing, ????, June 1987. ISBN 0-07-554168-8. ???? pp. LCCN ???? US$31.95.

URL http://www.cbooks.com/
sqlnut/SP/search/gtsumt?source=
&isbn=0075541688.

Larus:1988:RLP

[LH88] James R. Larus and Paul N. Hilfinger. Restructuring Lisp Programs for Concurrent Execution. In PPEALS 1988, Parallel Programming: Experience with Applications, Languages and Systems (New Haven, CT, USA), volume 23(9) of SIGPLAN Notices, pages 100–110. ACM Press, New York, NY 10036, USA, 1988.

Lawson:1979:ABL

[LHKK79a] C. L. Lawson, R. J. Hanson, D. R. Kincaid, and F. T. Krogh. Algorithm 539: Basic Linear Algebra Subprograms for Fortran usage [F1]. ACM Transactions on Mathematical Software, 5(3):324–325, September 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [DG82, Dod83, HK87a, LN88b].

Lawson:1979:BLA

[LHKK79b] C. L. Lawson, R. J. Hanson, D. R. Kincaid, and F. T. Krogh. Basic Linear Algebra Subprograms for Fortran usage. ACM Transactions on Mathematical Software, 5 (3):308–323, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Leigh:1987:SFa

[LHP87] William E. Leigh, G. David Huffman, and Noemi M. Paz. Structured Fortran ’77. Mitchell Publishing, ????, June 1987. ISBN 0-394-39067-9 (paperback). xii + 468 pp. LCCN QA76.73.F25L45 1987. US$14.96. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=

0394390679. Programs by Noemi Paz.

IMSL:1984:ILUa

[Lib84a] International Mathematical and Inc. Statistical Libraries. IMSL library: User’s Manual. Problemsolving software system for mathematical and statistical FORTRAN programming. IMSL, Houston, TX, USA, ed. 9.2 edition, 1984. 4 v. pp.

IMSL:1984:ILUb

[Lib84b] International Mathematical and Inc. Statistical Libraries. IMSL library: user’s manual: problemsolving software system for mathematical and statistical FORTRAN programming. IMSl, Houston, TX, USA, 9.2 edition, 1984. 4 v. pp.

IMSL:1987:IPS

[Lib87] International Mathematical and Inc. Statistical Libraries. IMSL problem-solving software systems: user’s manual, FORTRAN subroutines for evaluating special functions. IMSL, Houston, TX, USA, 1987. ???? pp.

IMSL:1989:MLFc

[Lib89a] International Mathematical and Inc. Statistical Libraries. Math/ library: FORTRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, version 1.1 edition, 1989. vi + 1152 + [79] pp.

IMSL:1989:SLFf

[Lib89b] International Mathematical and Inc. Statistical Libraries. SFUN/ library: FORTRAN subroutines for evaluating special functions. IMSL, Houston, TX, USA, version 2.1 edition, 1989. v + 60 + [24] pp.

IMSL:1989:SLFe

[Lib89c] International Mathematical and Inc. Statistical Libraries. Stat/ library: FORTRAN subroutines for statistical analysis. IMSL, Houston, TX, USA, version 1.1 edition, 1989. vi + 1232 + [75] pp.

Lignelet:1982:FLF

[Lig82a] Patrice Lignelet. Fortran 77: langage Fortran V. Masson, Masson, France, 1982. ISBN ???? xiii + 269 pp. LCCN ????

Lignelet:1982:PDF

[Lig82b] Patrice Lignelet. La Pratique du Fortran 77: 77 exercices résolus de Fortran V. Masson, Masson, France, 1982. ISBN 2-903607-11-7. 288 pp.

Lignelet:1984:FLF

[Lig84] Patrice Lignelet. Fortran 77: langage Fortran V. Masson, Masson, France, 2e edition, 1984. ISBN 2-903607-31-1. 189 pp.

Lignelet:1985:FLF

[Lig85a] Patrice Lignelet. Fortran 77: le langage Fortran V. Masson, Masson, France, 3e edition, 1985. ISBN 2-903115-20-6. 191 pp.

Lignelet:1985:PDF

[Lig85b] Patrice Lignelet. La Pratique du Fortran 77. Masson, Masson, France, 2e edition, 1985. ISBN 2-903607-46-X. 231 pp.

Lignelet:1988:FLF

[Lig88a] Patrice Lignelet. Fortran 77: le langage Fortran. Masson, Masson, France, 4e edition, 1988. ISBN 2-903607-69-9. 192 pp.

Lignelet:1988:PDF

[Lig88b] Patrice Lignelet. La Pratique du Fortran 77. Masson, Masson, France, 2 eme edition, 1988. ISBN 2-903607-46-X. 213 pp.

Lignelet:1988:FFA

[Lig88c] Patrice Lignelet. Les Fichiers en Fortran 77: avec exercices résolus. Masson, Masson, France, 1988. ISBN 2-903607-62-1. 234 pp.

Lin:1983:EFS

[Lin83] Frank C. Lin. Elementary Fortran With Scientific and Business Applications. Reston Publishing Co., Inc., Reston, VA, USA, August 1983. ISBN 0-8359-1697-9 (hardcover), 0-8359-1696-0 (paperback). xv + 336 pp. LCCN QA76.73.F25L56 1983. US$30.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835916979.

Lioen:1985:NFL

[Lio85] W. M. Lioen. NUMVEC FOR-TRAN library manual: chapter, elliptic PDEs: routine, MGZEB. Report. Centrum voor Wiskunde en Infomatica NM-8518, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, 1985. 17 pp.

Leis:1988:AOO

[LK88] Jorge R. Leis and Mark A.

Kramer. Algorithm 658: ODESSA: An ordinary differential equation solver with explicit simultaneous sensitivity analysis. ACM Transactions on Mathematical Software, 14(1):61–67, March 1988. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/citations/journals/
                                        toms/1988-14-1/p61-leis/.

Lutz:1988:SFP

[LKM88] Andrew E. Lutz, R. J. Kee, and James A. (James Angus) Miller. SENKIN: a FORTRAN program for predicting homogeneous gas phase chemical kinetics width sensitivity analysis. Sandia report; sand87-8248, Sandia National Laboratories, Albuquerque, NM, USA, 1988. 32 pp.

Louter-Nool:1987:TAB

[LN87] M. Louter-Nool. Translation of algorithm 539: basic linear algebra subprograms for FORTRAN usage in FORTRAN 200 for the Cyber 205 / F. Centrum voor Wiskunde en Informatica Report NM-R8702, CWI, Amsterdam, The Netherlands, 1987. 14 pp.

Louter-Nool:1988:TAB

[LN88a] M. Louter-Nool. Translation of Algorithm 539: Basic Linear Algebra Subprograms for Fortran usage in Fortran 200 for the Cyber 205. ACM Transactions on Mathematical Software, 14(2):177–195, June 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Louter-Nool:1988:ATA

[LN88b] Margreet Louter-Nool. Algorithm 663: Translation of Algorithm 539: Basic Linear Algebra Subprograms for FORTRAN usage in FORTRAN 200 for the Cyber 205. ACM Transactions on Mathematical Software, 14(2):177–195, June 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/
0098-3500/45058.html.  See

[LHKK79a, DG82, Dod83, HK87a].

Legler:1989:VFP

[LN89] David M. Legler and I. M. Navon.

VARIATM: a Fortran program for objective analysis of pseudo-stress with large-scale conjugate-gradient minimization. Technical report, Florida State University, Supercomputer Computations Research Institute, Tallahassee, FL, USA, 1989. 18 + [18] pp.

Lusk:1985:UMFa

[LO85a] E. L. Lusk and R. A. Overbeek. Use of monitors in Fortran: a tutorial on the barrier, self-scheduling DO-loop, and Askfor monitors. Technical Report ANL84-51, Rev.1, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June 1985.

Lusk:1985:UMFb

[LO85b] E. L. Lusk and R. A. Overbeek.

Use of monitors in FORTRAN: a tutorial on the barrier, selfscheduling DO-loop, and askfor monitors. In Kowalik [Kow85], pages 367–411. ISBN 0-262-11101-2. LCCN QA76.8.D436 P37 1985.

Lewis:1986:ASS

[LOU86] Peter A. W. Lewis, E. J. (Endel John) Orav, and Luis Uribe.

Advanced simulation and statistics package IBM professional FOR-TRAN version, 1986. ISBN 0-534-05304-1 (book).

Leff:1985:BCP

[LP85a] Lawrence S. Leff and Arlene Podos. Barron’s Computer programming in FORTRAN the easy way. Barron’s Educational Series, Inc., Woodbury, NY, USA, 1985. ISBN 0-8120-2800-7 (paperback). 326 pp. LCCN ???? US$11.95; US$10.75. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0812028007.

Leff:1985:CPF

[LP85b] Lawrence S. Leff and Arlene Podos. Computer Programming in Fortran the Easy Way. Barrons Educational Series, ????, January 1985. ISBN 0-8120-2800-7. 326 pp. LCCN ???? US$11.95.

Lipschutz:1987:PFT

[LP87] Seymour Lipschutz and Arthur Poe. Programmation Fortran: théorie et applications. Serie Schaum. McGraw-Hill, New York, NY, USA, 7e tirage edition, 1987. ISBN 2-7042-1013-6. ix + 313 pp.

Lipschutz:1983:PFT

[LPJ83] Seymour Lipschutz, Arthur Poe, and Sylvie Jacoud. Programmation Fortran: théorie et applications. Serie Schaum. McGraw-Hill, New York, NY, USA, 4e tirage edition, 1983. ISBN 2-7042-1013-6. ix + 313 pp.

Lozier:1989:SPC

[LR89] D. W. Lozier and R. G. Rehm.

Some performance comparisons for a fluid dynamics code. Parallel Computing, 11(3):305–320, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Li:1985:VCV

[LS85] Kuo-Cheng Li and Herb Schwetman. Vector C: a vector processing language. Journal of Parallel and Distributed Computing, 2(2):132– 169, May 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:1987:PVP

[LS87] B. Liu and N. Strother. Peak vector performance from VS Fortran. Research report RC 12849, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1987. 30 pp.

Liu:1988:PVF

[LS88] Bowen Liu and Nelson Strother.

Programming in VS Fortran on the IBM 3090 for maximum vector performance. IEEE Computer, 21(6):65–77, June 1988.

Lioen:1988:OMF

[LtW88] W. Lioen, H. te Riele, and D. Winter. Optimization of the MPQSfactoring algorithm on the Cyber 205 and the NEC SX-2. Supercomputer, 5(4):42–50, July 1988. CO-DEN SPCOEL. ISSN 0168-7875.

Ludwig:1981:UF

[Lud81] Herbert R. Ludwig. Understanding Fortran. An Alfred handy guide. Alfred, Sherman Oaks, CA, USA, January 1981. ISBN 0-88284-148-3. 62 pp. LCCN ???? US$3.50.

Levesque:1988:GFS

[LW88a] John M. Levesque and Joel W. Williamson. A Guidebook to Fortran on Supercomputers. Academic Press, New York, NY, USA, 1988. ISBN 0-12-444760-0. vii + 218 pp. LCCN QA76.73.F25 L4751 1988.

Lugosi:1988:SWP

[LW88b] Erzsebet Lugosi and Arthur T. Winfree. Simulation of wavepropagation in three dimensions using Fortran on the CYBER 205. Journal of Computational Chemistry, 9(6):689–701, September 1988. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Levesque:1989:GFS

[LW89] John M. Levesque and Joel W. Williamson. A Guidebook to Fortran on Supercomputers. Academic Press, New York, NY, USA, January 1989. ISBN 0-12-444760-0. vii + 218 pp. LCCN QA76.73.F25 L4751 1988. US$59.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124447600.

LaBonte:1982:CFC

[LZ82] Edward LaBonte and Richard J. Zinkl. CALFLT: a FORTRAN computer program to plot aerial radiometric data for the National Uranium Resource Evaluation. Technical Report 194-82 GJBX, U.S. Dept. of Energy, Grand Junction, CO, USA, 1982. 67 pp.

Mojena:1989:F

[MA89] Richard Mojena and Roy Ageloff.

Fortran 77. Wadsworth, Pacific Grove, CA, USA, September 1989. ISBN 0-534-11742-2. ???? pp. LCCN ???? US$45.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534117422.

Maany:1989:FAD

[Maa89] Z. A. Maany. FORTRAN automatic differentiation package for the optimization of functions of many variables. 1989. To appear.

MacDonald:1981:AGC

[Mac81] Alexander E. MacDonald. AFOS graphics creation from FOR-TRAN. NOAA Western Region computer programs and problems NWS WRCP 18, National Oceanic and Atmospheric Administration, National Weather Service, Western Region, Salt Lake City, UT, USA, 1981. 22 pp.

Maine:1981:PMMa

[Mai81a] Richard E. Maine. Programmer’s manual for MMLE3, a general FORTRAN program for maximum likelihood parameter estimation. NASA technical paper 1690, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1981. 113 pp.

For sale by the National Technical Information Service.

Maine:1981:PMMb

[Mai81b] Richard E. Maine. Programmer’s manual for MMLE3, a general FORTRAN program for maximum likelihood parameter estimation. NASA technical paper 1690, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1981. 113 pp. For sale by the National Technical Information Service.

Maine:1987:MGV

[Mai87] Richard E. Maine. Manual for GetData version 3.1 a FOR-TRAN utility program for time history. NASA technical memorandum 88288, National Aeronautics and Space Administration, Ames Research Center, Dryden Flight Research Facility, Edwards, CA, USA, 1987. ???? pp. For sale by the National Technical Information Service.

Malloch:1985:VFP

[Mal85] A. J. C. (Andrew John Cadoux) Malloch. Vespan: FORTRAN programs for handling and analysis of vegetation data and species distributions. Technical report, University of Lancaster, Lancaster, UK, 1985. ISBN 0-901272-18-3 (paperback). 131 pp.

Mann:1982:FIP

[Man82] Richard A. Mann. A FORTRAN IV primer. T. Y. Crowell: Harper and Row, New York, NY, USA, 1982. ISBN ???? vii + 207 pp. LCCN ????

Marquess:1981:CPC

[Mar81] Ken Marquess. Computer program for the computation of age and growth statistics of fish populations, (IBM 370, Fortran IV level G1): user’s guide. Management series 1, Department of Fisheries and Wildlife Sciences, Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1981. i + 92 pp.

Marra:1982:FRG

[Mar82a] Stephen R. Marra. FORTRAN 77 reference guide. Prime Computer Inc., Framingham, MA, USA, second edition, 1982. 194 pp.

Marshall:1982:LGP

[Mar82b] Howard Z. Marshall. The linear graph package, a compiler building environment. In Proceedings of the SIGPLAN ’82 Symposium on Compiler Construction [ACM82], pages 294–300. ISBN ???? LCCN ???? Available as SIGPLAN Notices 17(6) June 1982.

Marateck:1983:F

[Mar83a] Samuel L. Marateck. Fortran 77. Academic Press, New York, NY, USA, second edition, February 1983. ISBN 0-12-470463-8 (paperback). xviii + 584 pp. LCCN QA76.73.F25 M37 1983. US$29.26. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124704638.

Marateck:1983:IMF

[Mar83b] Samuel L. Marateck. Instructor’s Manual for Fortran 77. Academic Press, New York, NY, USA, second edition, February 1983. ISBN 0-12-470464-6. xiii + 64 pp. LCCN ???? US$13.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124704646.

Marshall:1983:FMa

[Mar83c] Garry Marshall. FORTRAN for micros. Newnes programming books. Newnes Technical, London, UK, 1983. ISBN 0-408-01336-2 (paperback). 80 pp.

Marshall:1983:FMb

[Mar83d] Garry Marshall. Fortran for Micros. Butterworth-Heinemann, Boston, MA, USA, December 1983. ISBN 99947-752-0-0. ???? pp. LCCN ???? US$11.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
9994775200.

Marca:1984:ASE

[Mar84a] David Marca. Applying Software Engineering Principles. Little, Brown computer systems series. Little, Brown and Co., Boston, MA, USA, February 1984. ISBN 0-316-54574-0. xviii + 270 pp. LCCN QA76.6.M358 1984. US$34.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0316545740. With FORTRAN examples.

Martin:1984:ASP

[Mar84b] Dave Martin. Avionics simulation package: a large systems application in Ratfor. In Software Tools Users Group [Sof84], pages 352–?? ISBN ???? LCCN QA76.8.U65 U83 1984. Abstract only.

Mashaw:1983:PBB

[Mas83] Bijan Mashaw. Programming Byte by Byte: Structured Fortran 77. Little, Brown computer systems series. Little, Brown and Co., Boston, MA, USA, August 1983. ISBN 0-316-54908-8 (paperback). xxv + 549 pp. LCCN QA76.73.F25 M39 1983 Sci-Eng. US$23.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0316549088.

Mashaw:1987:PBB

[Mas87] Bijan Mashaw. Programming Byte by Byte: Structured Fortran 77. American Computer Press, Livermore, CA, USA, second edition, 1987. ISBN 0-934433-02-X (paperback). xvii + 541 pp. LCCN QA76.73.F25 M39 1987. US$39.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
093443302X.

Mattis:1983:FPAa

[Mat83a] Richard L. Mattis. A FOR-TRAN program for analysis of data from microelectronic test structures. Semiconductor measurement technology National Bureau of Standards special publication 400-75, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1983. various pp. For sale by the Supt. of Docs., U.S. G.P.O.

Mattis:1983:FPAb

[Mat83b] Richard L. Mattis. A FOR-TRAN program for analysis of data from microelectronic test structures. NBS special publication 400-75 SeF, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1983. various pp. For sale by the Supt. of Docs., U.S. G.P.O.

MATH:1989:MRA

[MAT89a] MATH. MATH77, Release 3.0, A library of mathematical subprograms for FORTRAN 77. Internal Document D–134, Rev. B, Jet Propulsion Laboratory, Pasadena, CA, USA, May 1989. Also available as Program No. NPO–18120 from COSMIC (Computer Software Management and Information Center), The University of Georgia, Athens, GA.

MATH:1989:MRL

[MAT89b] MATH. MATH77, Release 3.0, A library of mathematical subprograms for FORTRAN 77. Internal Document D–134, Rev. B, Jet Propulsion Laboratory, Pasadena, CA 91105, May 1989. Also available as Program No. NPO–18120 from COSMIC (Computer Software Management and Information Center), The University of Georgia, Athens, GA.

Maurer:1984:UI

[Mau84] H. Maurer. Ueberblicke Informationsverarbeitung. Bibliographisches Institut, Mannheim, Germany, 1984. ISBN 3-411-01689-2.

May:1989:PPP

[May89] Edward N. May. Portable parallel programming in a Fortran environment. Computer Physics Communications, 57(1–3):278– 284, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465589902282.

Meinke:1981:ATF

[MB81] John G. Meinke and John A. Beidler. Alternatives to the traditional first course in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 13(1):57–60, February 1981. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th SIGCSE Symposium on Computer Science Education.

Mellichamp:1985:RCE

[MBP^{+}85a] Duncan Mellichamp, Dave Bedworth, Odd Pettersen, Peter Rony, Lew Bezanson, Walter Higgins, and Granino Korn. Real-time computing and the engineering support system. IEEE Micro, 5(5):27– 35, September/October 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mellichamp:1985:RTC

[MBP^{+}85b] Duncan Mellichamp, Dave Bedworth, Odd Pettersen, Peter Rony, Lew Bezanson, Walter Higgins, and Granino Korn. Real-time computing and the engineering support system. IEEE Micro, 5(5):27– 35, September/October 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

More:1980:ABF

[MC80] J. J. Moré and M. Y. Cosnard. Algorithm 554: BRENTM, A Fortran subroutine for the numerical solution of nonlinear equations [F5]. ACM Transactions on Mathematical Software, 6(2):240–251, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Mosher:1982:FP

[MC82] David A. Mosher and Robert P. Corbett. F77 performance. ;login: the USENIX Association newsletter, 7(3):9–14, June 1982. CODEN LOGNEM. ISSN 1044-6397.

McAndrew:1986:FPC

[McA86] C. C. McAndrew. FORTRAN program coding guidelines: Research and Development Department. Report / State Electricity Commission of Victoria, Research and Development Department; no. GO/86/96 Report (State Electricity Commission of Victoria. Research and Development Dept.); no. 86/96. State Electricity Commission of Victoria, Research and Development Dept., Engineering Research Division, Melbourne, Victoria, Australia, 1986. ISBN 0-7241-6286-0. 10 pp.

Chung:1984:TTC

[mCaLjH84] Ying ming Chung, Shih an Liang, and Hung jen Huang. Tien tzu chi suan chi cheng shih yu yen FOR-TRAN. Tao ming chu pan she: Tsung ching hsiao Sung kang tien nao tu shu tzu liao yu hsien kung ssu, Tai-pei shih, ti 12 pan edition, 1984. ISBN ???? 4 + 366 pp. LCCN ????

McCarty:1981:IFI

[McC81] Robert D. McCarty. Interactive Fortran IV computer programs for the thermodynamic and transport properties of selected cryogens (fluids pack). NBS technical note 1025, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1981. iv + 107 pp. For sale by the Supt. of Docs., U.S. G.P.O.

McCracken:1984:CES

[McC84a] Daniel D. McCracken. Computing for Engineers and Scientists With Fortran 77. John Wiley and Sons, New York, London, Sydney, 1984. ISBN 0-471-09701-2 (paperback). xi + 361 pp. LCCN TA345.M395 1984; QA76.73 .F25 M34 1984. US$14.95; US$22.95

(est.). URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0471097012.

McCracken:1984:EHF

[McC84b] Daniel D. McCracken. The early history of FORTRAN publications. Annals of the History of Computing, 6(1):33–34, January/March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1028.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1028abs.htm.

McCracken:1984:IMA

[McC84c] Daniel D. McCracken. Instructor’s manual to accompany Computing for engineers and scientists with Fortran 77. John Wiley and Sons, New York, London, Sydney, 1984. ISBN 0-471-80090-2 (paperback). 117 pp.

McCracken:1985:CFN

[McC85a] D. D. McCracken. Computerpraxis mit Fortran 77 in Naturwissenschaft und Technik. Carl Hanser, München, Germany, 1985. ISBN 3-446-14216-9.

McCracken:1985:CMF

[McC85b] D. D. McCracken. Computerpraxis mit Fortran 77 in Naturwissenschaft und Technik. Carl Hanser, München, Germany, 1985. ISBN 3-446-14216-9.

McCracken:1985:DAB

[McC85c] Daniel D. (Daniel Delbert) Mc-Cracken. Dalil al-barmajah bi lughat al-Fortran. John Wiley and Sons, New York, London, Sydney, 1985. ISBN 0-471-82382-1 (paperback). 409 pp.

McCarty:1986:IFP

[McC86] Robert D. McCarty. Interactive FORTRAN programs for micro computers to calculate the thermophysical properties of twelve fluids (MIPROPS). NBS technical note 1097, National Bureau of Standards, Gaithersburg, MD, USA, 1986. iii + 84 pp. For sale by the Supt. of Docs. U.S. G.P.O.

McDaniel:1983:IF

[McD83] Kenneth L. McDaniel. Introduction to FORTRAN. The Center, Pensacola, FL, USA, 1983. various pp.

McDaniel:1985:IF

[McD85] Kenneth L. McDaniel. Introduction to Fortran. The Center, Pensacola, FL, USA, 1985. various pp.

McDowell:1989:PAS

[McD89] Charles E. McDowell. A practical algorithm for static analysis of parallel programs. Journal of Parallel and Distributed Computing, 6 (3):515–536, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

McGraw:1984:DRF

[McG84] James R. McGraw. A debate: Retire FORTRAN?: Yes. Physics Today, 37(5):66–74, May 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic).  URL http://www.
physicstoday.org/resource/1/
phtoad/v37/i5/p66_s1.

McGirt:1987:MMCa

[McG87a] Frank McGirt. Mixing Microsoft C and Microsoft FORTRAN program modules on the IBM Personal Computer. Technical report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1987. 30 pp.

McGirt:1987:MMCb

[McG87b] Frank McGirt. Mixing Microsoft C and Microsoft FORTRAN program modules on the IBM personal computer. Technical report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1987. 22 pp. Available from Supt. of Docs. U.S. G.P.O; National Technical Information Service [distributor].

McKay:1983:FPC

[McK83] Steven F. McKay. A FORTRAN program for computing primary insurance amounts. Actuarial note (United States. Social Security Administration) 116, SSA 11-11500, U.S. Dept. of Health and Human Services, Social Security Administration, Baltimore, Md.?, 1983. 33 pp.

McKeown:1985:SPW

[McK85a] Patrick McKeown. Structured Programming Wsing Watfiv. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, January 1985. ISBN 0-15-584414-8. ???? pp. LCCN ???? US$29.26.

McKeown:1985:SPU

[McK85b] Patrick G. McKeown. Structured Programming Using Fortran 77. Harcourt, Brace, Jovanovich, San Diego, CA, USA, January 1985. ISBN 0-15-584411-3 (paperback). x + 482 pp. LCCN QA 76.73 F25 M395 1985.

US$38.57; US$20.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0155844113.

McMahon:1986:LFK

[McM86] Frank H. McMahon. The Livermore Fortran kernels: a computer test of the numerical performance range. Technical report, Lawrence Livermore National Laboratory, Livermore, CA, USA, 1986. iv + 203 pp.

McNitt:1983:IFT

[McN83] Lawrence L. McNitt. Invitation to Fortran for the TRS-80. Petrocelli Books, New York, NY, USA, August 1983. ISBN 0-89433-210-4 (paperback). viii + 234 pp. LCCN QA76.8.T18 M373 1983. US$15.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0894332104.

Mehta:1988:LFI

[MD88] Manish A. Mehta and N. De Leon. LUCY: a FORTRAN implementation of semiclassical spectral quantization. Computer Physics Communications, 51(1– 2):115–134, September/October 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465588900665.

Meissner:1984:FFSa

[Mei84] Loren P. Meissner. Fortran 77 Featuring Structured Programming. Addison-Wesley, Reading, MA, USA, June 1984. ISBN 0-07-582328-4. ???? pp. LCCN ???? US$13.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0075823284.

Meissner:1987:CFD

[Mei87] Lorer P. Meissner. Concerning the Fortran 8X draft. ACM SIGPLAN FORTRAN Forum, 6(3):1–31, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1988:MFP

[Mei88] Loren Meissner. More on Fortran 8X pointer proposals. ACM SIG-PLAN FORTRAN Forum, 7(1): 28–36, April 1988. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:FD

[Mei89a] Loren P. Meissner. Fortran 8X draft. ACM SIGPLAN FOR-TRAN Forum, 8(4):1–200, December 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:PIO

[Mei89b] Loren P. Meissner. Partialrecord input and output. ACM SIGPLAN FORTRAN Forum, 8 (3):20–23, August 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:SF

[Mei89c] Loren P. Meissner. Summary of Fortran 88. ACM SIGPLAN FOR-TRAN Forum, 8(2):15–31, May 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:XAD

[Mei89d] Loren P. Meissner. X3J3 avoids deadlock. ACM SIGPLAN FOR-TRAN Forum, 8(1):1–3, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Melhem:1988:PSL

[Mel88] Rami Melhem. Parallel solution of linear systems with striped sparse matrices. Parallel Computing, 6 (2):165–184, February 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Merchant:1981:FLS

[Mer81] Michael J. Merchant. FOR-TRAN 77: language and style: a structured guide to FOR-TRAN 77. Wadsworth, Pacific Grove, CA, USA, May 1981. ISBN 0-534-00920-4. xiii + 447 pp. LCCN QA76.73.F25 .M463. US$36.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534009204.

Merchant:1985:WLS

[Mer85] Michael J. Merchant. WATFIV S, language and style. Wadsworth series in FORTRAN and WAT-FIV. Wadsworth, Pacific Grove, CA, USA, February 1985. ISBN 0-534-04728-9 (paperback). xi + 449 pp. LCCN QA76.73.F25 M465 1985 Sci-Eng. US$31.95.

GarciaMerayo:1986:PF

[Mer86] Felix Garcia Merayo. Programacion en FORTRAN 77. Paraninfo, Madrid, Spain, 1986. ISBN 84-283-1470-5. 368 pp.

GarciaMerayo:1988:PF

[Mer88a] Felix Garcia Merayo. Programacion en FORTRAN 77. Paraninfo, Madrid, Spain, segunda edition, 1988. ISBN 84-283-1470-5. 368 pp.

Mercer:1988:CFC

[Mer88b] Randall Mercer. The Convex Fortran 5.0 compiler. In Proc. 1988 Intl Conf. on Supercomputing, volume 2, pages 164–175. ????, ????, 1988.

Metcalf:1982:FO

[Met82] Michael Metcalf. Fortran Optimization, volume 17 of A. P. I. C. Studies in Data Processing. Academic Press, New York, NY, USA, 1982. ISBN 0-12-492480-8. ISSN 0067-2483. xii + 242 pp. LCCN QA76.73.F25 M48 1982. US$35.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124924808.

Metcalf:1985:EF

[Met85a] Michael Metcalf. Effective Fortran 77. Oxford science publications. Oxford University Press and Clarendon Press, Walton Street, Oxford OX2 6DP, UK and Oxford, UK, June 1, 1985. ISBN 0-19-853709-3 (paperback), 0-19-853710-7. xii + 231 pp. LCCN QA76.73.F25 M478 1985. US$19.90, UK£9.95, US$40.00. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0198537093;
 http://www.cbooks.com/sqlnut/

SP/search/gtsumt?source=&isbn=
0198537107.

Metcalf:1985:FO

[Met85b] Michael Metcalf. Fortran Optimization, volume 25 of A.P.I.C. studies in data processing. Academic Press, New York, NY, USA, revised edition, August 1985. ISBN 0-12-492482-4. xii + 253 pp. LCCN QA76.73.F25 M48 1985. US$46.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124924824.

Metcalf:1985:FF

[Met85c] Michael Metcalf. Has Fortran a future? Computer Physics Communications, 38(2):199–210, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465585900864.

Metcalf:1986:EF

[Met86] Michael Metcalf. Effective FOR-TRAN 77. Oxford science publications. Clarendon Press and Oxford University Press, Oxford, UK and Walton Street, Oxford OX2 6DP, UK, repr. with corrections edition, 1986. ISBN 0-19-853709-3 (paperback). xii + 231 pp.

Metcalf:1987:FES

[Met87a] M. Metcalf. Fortran 8X–the emerging standard. ACM SIG-PLAN FORTRAN Forum, 6(1): 28–47, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1987:EF

[Met87b] Michael Metcalf. Effective FOR-TRAN 77. Oxford science publications. Clarendon Press; Oxford University Press, Oxford, Oxfordshire, UK, 1987. ISBN 0-19-853709-3 (U.S. paperback). xii + 231 pp.

MSC:1989:MRM

[Met89a] Metagraphics Software Corporation. MetaWINDOW reference manual: version 3: Fortran. Metagraphics, Scotts Valley, CA, USA, 1989. v + 246 pp.

Metcalf:1989:EF

[Met89b] Michael Metcalf. Effective FOR-TRAN 77. Oxford science publications. Clarendon Press; Oxford University Press, Oxford, Oxfordshire, UK, 1989. ISBN 0-19-853709-3 (paperback). xii + 231 pp.

Metcalf:1989:RPF

[Met89c] Michael Metcalf. Recent progress in Fortran standardization. Computer Physics Communications, 57 (1–3):78–83, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/001046558990194X.

Meyer:1984:EIP

[Mey84] S. Meyer. Entwurf und Implementierung eines Programms zur Strukturanalyse von Fortran 77-Programmen. Master’s thesis, Technische Universität Braunschweig (??), Braunschweig, Germany, 1984.

Maurer:1984:UIG

[MF84] Hermann A. Maurer and W.- D Fellner. Uberblicke Informationsverarbeitung 1984: Grafik auf Mikros, medizinische Informatik, Fehlertolerante Systeme, Mustererkennung, Layout Probleme, Robotertechnologie, Kryptographie, Neues zu FORTRAN, Bericht aus Silicon Valley = Surveys in computer science. Bibliographisches Institut, Mannheim, Germany, 1984. ISBN 3-411-01689-2. 369 pp.

More:1981:FST

[MGH81a] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. Fortran subroutines for testing unconstrained optimization software. ACM Transactions on Mathematical Software, 7(1):136–140, March 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1981:AFS

[MGH81b] Jorge J. Moré, Burton S. Garbow, and Kenneth E. Hillstrom. Algorithm 566: FORTRAN subroutines for testing unconstrained optimization software [C5 [E4]]. ACM Transactions on Mathematical Software, 7(1):136–140, March 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [AFS94].

Michel:1982:SFP

[MH82] W. L. Michel and J. Duncan Hepburn. SLACINPT: a FORTRAN program that generates boundary data for the SLAC gun code. Technical Report AECL-7614, Accelerator Physics Branch, Chalk River, ON, Canada, 1982. iii + 52 pp.

Maurer:1986:QAM

[MHK86] M. Maurer, A. Hayd, and H. J. Kaeppeler. Quasi-analytical method for solving nonlinear differential equations for turbulent self-confined magneto-plasma. Journal of Computational Physics, 66(1):151–172, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0021999186900586.

Mighell:1981:NFP

[MHS81] A. D. Mighell, C. R. Hubbard, and J. K. Stalick. NBS*AIDS80: a FORTRAN program for crystallographic data evaluation. NBS Technical note 1141, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1981. iii + 50 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Microsoft:1982:FCa

[MI82a] Microsoft and International Business Machines Corporation. Fortran compiler. IBM Corporation, Boca Raton, FL, USA, 1982. various pp.

Microsoft:1982:FCb

[MI82b] Microsoft and International Business Machines Corporation. FOR-TRAN compiler, 1982.

Microsoft:1984:FC

[MI84] Microsoft and International Business Machines Corporation. FOR-TRAN compiler. IBM Corporation, Boca Raton, FL, USA, version 2.00 edition, 1984. various pp.

Microsoft:1981:MFC

[Mic81] Microsoft. Microsoft FORTRAN compiler, 1981.

MTUACS:1983:PFL

[Mic83] Michigan Technological University.

Academic Computing Services. PDP-11 FORTRAN language reference manual. MTU Academic Computing Services, Houghton, MI, USA, 1983. various pp.

Microsoft:1984:MF

[Mic84a] Microsoft. Microsoft FORTRAN 77, 1984.

Microsoft:1984:MFCc

[Mic84b] Microsoft. Microsoft FORTRAN compiler, 1984.

Microsoft:1984:MFCa

[Mic84c] Microsoft. Microsoft FORTRAN compiler for the MS-DOS operating system: user’s guide. Microsoft Press, Bellevue, WA, USA, 1984. xiii + 165 pp.

Microsoft:1984:MFR

[Mic84d] Microsoft. Microsoft FORTRAN reference manual. Microsoft Press, Bellevue, WA, USA, 1984. ix + 233 pp.

Microsystems:1984:F

[Mic84e] Softech Microsystems. Fortran 77. IBM PC Apprentice Personal Computer Learning Series. Prentice-Hall Software, Englewood Cliffs, NJ 07632, USA, November 1984. ISBN 0-13-452889-1. ???? pp. LCCN ???? Beginning/Book and 128K Diskette/Self-Teach Edition.

Microsystems:1984:FBB

[Mic84f] Softech Microsystems. Fortran 77-Beginning/Book and 128K Diskette/Self-Teach Edition. Prentice-Hall Software, Englewood Cliffs, NJ 07632, USA, November 1984. ISBN 0-13-452889-1. ???? pp. LCCN ???? US$26.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0134528891.

Microsoft:1985:MFCc

[Mic85a] Microsoft. Microsoft FORTRAN compiler, 1985.

Microsoft:1985:MFCb

[Mic85b] Microsoft. Microsoft FORTRAN compiler for the MS-DOS operating system, 1985.

Microsoft:1985:MFCa

[Mic85c] Microsoft. Microsoft FORTRAN compiler for the MS-DOS operating system: user’s guide. Microsoft Press, Bellevue, WA, USA, 1985. xvii + 269 pp.

Microsoft:1985:MF

[Mic85d] Microsoft. Microsoft FOR-TRAN77, 1985.

Microsoft:1985:MFR

[Mic85e] Microsoft Corporation. Microsoft FORTRAN reference manual. Microsoft Press, Bellevue, WA, USA, 1985. ix + 217 pp.

Microsoft:1986:MFC

[Mic86] Microsoft. MS FORTRAN 2.1 compiler, 1986.

Microsoft:1987:MFOd

[Mic87a] Microsoft. Microsoft FORTRAN optimizing compiler, 1987.

Microsoft:1987:MFOe

[Mic87b] Microsoft. Microsoft FORTRAN optimizing compiler, 1987. Program files on 10 computer disks, Codeview manual + 1 language reference manual + 2 booklets + 1 reference guide + 2 keyboard templates.

Microsoft:1987:MFOf

[Mic87c] Microsoft. Microsoft FORTRAN optimizing compiler: for the MS-DOS operating system, 1987.

Microsoft:1987:MFOa

[Mic87d] Microsoft. Microsoft FORTRAN optimizing compiler: language reference: mixed-language programming guide. Microsoft Press, Bellevue, WA, USA, 1987. ???? pp.

Microsoft:1987:MFOc

[Mic87e] Microsoft. Microsoft FORTRAN optimizing compiler: Microsoft Codeview and utilities: Microsoft Editor. Microsoft Press, Bellevue, WA, USA, 1987. ???? pp.

Microsoft:1987:MFOb

[Mic87f] Microsoft. Microsoft FORTRAN optimizing compiler: user’s guide.

Microsoft Press, Bellevue, WA, USA, 1987. ???? pp.

Microsoft:1988:MFO

[Mic88] Microsoft. Microsoft FORTRAN optimizing compiler, 1988.

Microsoft:1989:MF

[Mic89a] Microsoft. Microsoft FORTRAN, 1989.

Microsoft:1989:MFA

[Mic89b] Microsoft. Microsoft FORTRAN advanced topics. Microsoft Press, Bellevue, WA, USA, 1989. xviii + 301 pp.

Microsoft:1989:MFO

[Mic89c] Microsoft. Microsoft FORTRAN optimizing compiler, 1989.

Microsoft:1989:MFR

[Mic89d] Microsoft. Microsoft FORTRAN reference. Microsoft Press, Bellevue, WA, USA, 1989. xxv + 523 pp.

Middlebrooks:1984:VF

[Mid84] C. H. (Charlotte H.) Middlebrooks. VAX Fortran. Reston Publishing Co., Inc., Reston, VA, USA, October 1984. ISBN 0-8359-8245-9 (hardcover), 0-8359-8243-2 (paperback). xii + 516 pp. LCCN QA 76.8 V37 M53 1984. US$21.95 (hardcover); US$15.95 (paperback); US$37.51.

Miller:1982:FPS

[Mil82a] Alan R. Miller. FORTRAN programs for scientists and engineers. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1982. ISBN 0-89588-082-2 (paperback). xvii + 280 pp. LCCN QA76.95.M534 1982.

Miller:1982:BRW

[Mil82b] C. D. F. Miller. Book review: E.

W. Martin and W. C. Perkins, Fortran for Business Students: a Programmed Instruction Approach. The Computer Journal, 25(4):492, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
25/4/492.full.pdf+html.

Miller:1987:PFA

[Mil87a] David Donald Miller. From Pascal to Fortran 77: Applications for Scientists and Engineers. Harcourt, Brace, Jovanovich, San Diego, CA, USA, January 1987. ISBN 0-15-529175-0 (paperback). xiii + 287 pp. LCCN QA 76.73 P2 M54 1987. US$21.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0155291750.

Miller:1987:SMP

[Mil87b] David Donald Miller. Solutions manual for From Pascal to FOR-TRAN 77: applications for scientists and engineers. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1987. ISBN 0-15-529176-9. 170 pp.

Miller:1988:F

[Mil88a] Alan R. Miller. Fortran. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, October 1988. ISBN 0-89588-082-2. xvii + 280 pp. LCCN QA76.95.M534 1982. US$12.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0895880822.

Miller:1988:FPS

[Mil88b] Alan R. Miller. Fortran: Programs for Scientists and Engineers. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, October 1988. ISBN 0-89588-571-9. xvii + 280 pp. LCCN QA76.95 .M534 1988. US$22.95; US$18.36. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0895885719.

Miller:1989:PGE

[Mil89] B. R. Miller. A program generator for efficient evaluation of Fourier series. In Gonnet [Gon89], pages 199–206. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Minoux:1988:LSL

[Min88] Michel Minoux. LTUR: a simplified linear-time unit resolution algorithm for Horn formulae and computer implementation. Information Processing Letters, 29(1): 1–12, September 15, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miyawaki:1987:FA

[Miy87] S. Miyawaki. Fortran 8X abridgement. ACM SIGPLAN FOR-TRAN Forum, 6(3):41–78, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Mizoguchi:1983:PBE

[Miz83] F. Mizoguchi. PROLOG based expert system. New Generation Computing, 1(1):99–104, 1983. CODEN NGCOE5. ISSN 0288-3635 (print), 1882-7055 (electronic). QA 76 N48.

Kim:1984:CKW

[mK84] Song mun Kim. Chonja kyesan wolli: computer, Basic, Fortran, Cobol. Kyohaksa, Soul Tukpyolsi, Korea, chopan edition, 1984. ISBN ???? xii + 324 pp. LCCN ????

Mamikonov:1986:SOM

[MK86] A. G. Mamikonov and V. V.

Kulba. Sintez Optimalnykh Modulnykh Sistem Obrabotki Dannykh. USSR Academy of Science, Izdatelstvo Nauka, Moscow, USSR, 1986. ISBN ???? LCCN ????

Marcotty:1987:WPL

[ML87] M. Marcotty and H. Ledgard.

The World of Programming Languages. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-96440-1.

Martin:1988:SPN

[ML88] Joanne L. Martin and Stephen F.

Lundstrom, editors. Supercomputing ’88: proceedings, November 14–18, 1988, Orlando, Florida, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Moore:1981:SFW

[MM81] John B. Moore and Leo J. Makela. Structured Fortran With Watfiv: Text and Reference. Reston Publishing Co., Inc., Reston, VA, USA, alternate edition, February 1981. ISBN 0-8359-7104-X, 0-8359-7103-1 (paperback). xvi + 567 pp. LCCN QA76.73.F25 .M65 1981. US$25.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
083597104X.

Mai:1984:CSC

[mM84] Sung mei Mai. Chi suan chi cheng shih: FORTRAN 77. Ta Chungkuo tu shu kung ssu, Tai-pei shih, chu pan edition, 1984. ISBN ???? v + 195 pp. LCCN ????

Matsuura:1985:SPT

[MMM85] Toshihiko Matsuura, Kenichi Miura, and Mitsuhiro Makino. Supervector performance without toil: FORTRAN implemented vector algorithms on the VP-100/ 200. Computer Physics Communications, 37(1–3):101–107, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.

sciencedirect.com/science/article/
pii/0010465585901419.

Miller:1988:ADA

[MMS88] Keith W. Miller, Larry J. Morell, and Fred Stevens. Adding data abstraction to FORTRAN software. IEEE Software, 5(6):50– 58, November 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Meissner:1982:FFS

[MO82] Loren P. Meissner and Elliott Irving Organick. Fortran 77: featuring structured programming. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, reprinted with corrections edition, 1982. ISBN 0-201-05499-X. xi + 500 pp. LCCN QA 76.73 F25 O68 1982.

Meissner:1984:FFSb

[MO84] Loren P. Meissner and Elliott Irving Organick. Fortran 77: featuring structured programming. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, repr. with corrections edition, 1984. ISBN 0-201-05499-X. xi + 500 pp.

Mong:1982:FV

[Mon82a] Sao Khai Mong. A Fortran version of . TUGboat, 3(2):25, October 1982. ISSN 0896-3207.

Mong:1982:FVM

[Mon82b] Sao Khai Mong. A Fortran version of . TUGboat, 3(2):25, October 1982. ISSN 0896-3207.

Monro:1982:F

[Mon82c] Donald M. Monro. Fortran 77. Edward Arnold, London, UK, August 1982. ISBN 0-7131-2794-5. vii + 360 pp. LCCN ???? US$29.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0713127945.

Monro:1983:F

[Mon83] Donald M. Monro. FORTRAN 77. E. Arnold, London, reprinted (with corrections) edition, 1983. ISBN 0-7131-2794-5 (paperback). vii + 360 pp. LCCN QA76.73.F25 M642 1983 Sci-Eng.

Monagan:1989:FTO

[Mon89a] Michael B. Monagan. Fortran translation and optimization. Maple Newsletter, 0(4): ??, January 1989. ISSN 1074-

3790. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/mtn4.
html.

Monro:1989:CCF

[Mon89b] Donald M. Monro. A Crash Course in Fortran 77. Edward Arnold, London, UK, February 1989. ISBN 0-7131-3582-4. vi + 218 pp. LCCN QA76.73.F25 M66 1989. US$15.90. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0713135824.

Mooney:1981:VOD

[Moo81] Jim Mooney. A Varian output driver in VAX/VMS Fortran.

TUGboat, 2(3):14–15, November 1981. ISSN 0896-3207.

Moore:1982:IMA

[Moo82a] Elmo Moore. Instructor’s manual to accompany Introduction to FORTRAN and its applications. Allyn and Bacon, Needham Heights, MA, USA, 1982. ISBN 0-205-07721-8 (paperback). 63 + [51] pp.

Moore:1982:IFA

[Moo82b] Elmo Moore. Introduction to Fortran and Its Applications. Allyn and Bacon computer science series. Allyn and Bacon, Needham Heights, MA, USA, June 1982. ISBN 0-205-07720-X (paperback). xiv + 315 pp. LCCN QA76.73 F25 M648 1982. US$27.84. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
020507720X.

Mooney:1983:RVO

[Moo83] Jim Mooney. Revised Varian output driver in VAX/VMS Fortran. TUGboat, 4(2):76, September 1983. ISSN 0896-3207.

Moon:1985:SPS

[Moo85a] T. T. Moon. Stellar parameters from Strömgren photometry: Fortran programs. Communications from the University of London Observatory 78 0458-2128, Department of Physics and Astronomy; University College London, London, UK, 1985. 23 pp.

Moore:1985:SFW

[Moo85b] John B. Moore. Structured Fortran with WATFIV: text and reference. Reston Publishing Co., Inc., Reston, VA, USA, third edition, February 1985. ISBN 0-8359-7118-X (paperback). xvii + 502 pp. LCCN QA 76.73 F25M65 1985. US$31.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
083597118X.

Moore:1986:CAF

[Moo86] Stacy Georgette Moore. Compositional applications of FORTRAN in music. Thesis (M.M.), University of South Florida, Tampa, FL, USA, 1986. xi + 115 pp.

Moore:1988:IFA

[Moo88a] Elmo Moore. Introduction to FOR-TRAN and its applications. Wm. C. Brown Publishers, Dubuque, IA, USA, 1988. ISBN 0-697-06842-0 (paperback). xiv + 315 pp.

Moore:1988:RCR

[Moo88b] Ramon E. Moore, editor. Reliability in Computing: the Role of Interval Methods in Scientific Computing, volume 19 of Perspectives in computing. Academic Press, New York, NY, USA, 1988. ISBN 0-12-505630-3. xv + 428 pp. LCCN QA76.9.E94 R45 1988.

Morris:1981:CAR

[Mor81a] Alfred H. Morris, Jr. Can Ada replace FORTRAN for numerical computation? ACM SIG-PLAN Notices, 16(12):10–13, December 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1981:NDL

[Mor81b] Alfred H. Morris, Jr. NSWC/ DL library of mathematics subroutines. Report NSWC/TR-79-338, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA; Silver Spring, MD 20903-5000, USA, 1981. 235 pp. URL

https://ntrl.ntis.gov/NTRL/
dashboard/searchResults/titleDetail/
ADA108106.xhtml. See also later

editions [Mor90, Mor93].

Morgan:1982:NCC

[Mor82] Howard Lee Morgan, editor. 1982 National Computer Conference: June 7–10, 1982, Houston, Texas. AFIPS Press, Arlington, VA, USA, 1982. ISBN 0-88283-035-X. LCCN TK7885.A1 J6 1982.

Moriguchi:1984:JFN

[Mor84] Shigeichi Moriguchi. JIS FOR-TRAN nyumon. Tokyodaigakushuppankai, Tokyo, Japan, 1984. ISBN 4-13-062030-4. 282 pp.

Morris:1990:NLM

[Mor90] Alfred H. Morris, Jr. NSWC library of mathematics subroutines. Report NSWC TR 90-21, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA; Silver Spring, MD 20903-5000, USA, January 1990. xii + 492 +

9 pp. URL https://apps.dtic.
mil/sti/citations/ADA476840;
https://apps.dtic.mil/sti/
pdfs/ADA476840.pdf;   https:
//people.math.sc.edu/Burkardt/
f_src/nswc/nswc.f90;   https:
//people.math.sc.edu/Burkardt/

f_src/nswc/nswc.html. See also

later edition [Mor93].

Morris:1993:NLM

[Mor93] Alfred H. Morris, Jr. NSWC library of mathematics subroutines. Report NSWCDD/TR-92/425, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA; Silver Spring, MD 20903-5000, USA, January 1993. 464 pp. URL

https://ntrl.ntis.gov/NTRL/
dashboard/searchResults/titleDetail/
ADA261511.xhtml. See also earlier

edition [Mor90].

Moses:1988:EAS

[Mos88] Gregory A. Moses. Engineering Applications Software Development: Using Fortran 77. John Wiley and Sons, New York, London, Sydney, 1988. ISBN 0-471-63851-X. xii + 320 pp. LCCN QA76.73.F25 M671 1988. US$74.95; US$79.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
047163851X.

Martin:1981:FBS

[MP81] E. Wainright (Edley Wainright) Martin and William C. Perkins. FORTRAN for business students: a programmed instruction approach. John Wiley and Sons, New York, London, Sydney, 1981. ISBN 0-471-04622-1. xiv + 811 pp. LCCN HF5548.5.F2 .M33. US$15.95 (est.).

Mehta:1986:AFF

[MP86a] Cyrus R. Mehta and Nitin R. Patel. Algorithm 643: FEX-ACT: A FORTRAN subroutine for Fisher’s exact test on unordered r × c contingency tables. ACM Transactions on Mathematical Software, 12(2):154–161, June 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/214326.html.

Menta:1986:FAF

[MP86b] C. R. Menta and N. R. Patel.

FEXACT: A Fortran subroutine for Fisher’s exact test on unordered r × c contingency tables. ACM Transactions on Mathematical Software, 12(2):154–161, June 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Menta:1986:AFF

[MP86c] Cyrus R. Menta and Nitin R. Patel. Algorithm 643: FEXACT: A FORTRAN subroutine for Fisher’s exact test on unordered r × c contingency tables. ACM Transactions on Mathematical Software, 12(2):154–161, June 1986. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Marse:1983:IPF

[MR83] K. Marse and S. D. Roberts. Implementing a portable FORTRAN uniform (0,1) generator. Simulation, 41, 4:135–139, 1983. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Modi:1986:SPE

[MR86] J. J. Modi and J. S. Rollett. Some problems of exploiting a pipeline processor. Parallel Computing, 3 (3):263–265, July 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Metcalf:1987:FE

[MR87] Michael Metcalf and John Ker Reid. Fortran 8x Explained. Oxford University Press and Clarendon Press, Walton Street, Oxford OX2 6DP, UK and Oxford, UK, 1987. ISBN 0-19-853751-4 (hardcover), 0-19-853731-X (paperback). xiv + 262 pp. LCCN QA76.73.F26 M48 1987. UK£12.50 (US$25.00 U.S.), 0198537514. See also [Ame87b].

Metcalf:1988:FE

[MR88] Michael Metcalf and John Ker Reid. Fortran 8x explained. Oxford science publications. Clarendon Press; Oxford University Press, Oxford, UK, reprinted (with corrections) edition, 1988. ISBN 0-19-853751-4, 0-19-853731-X (paperback). xiv + 262 pp.

Metcalf:1989:FE

[MR89] Michael Metcalf and John Ker Reid. Fortran 8X Explained. Oxford science publications. Oxford University Press and Clarendon Press, Walton Street, Oxford OX2 6DP, UK and Oxford, UK, revised edition, December 1989. ISBN 0-19-853824-3. xv + 315 pp. LCCN QA76.73.F26 M48 1989. US$24.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0198538243.

Majocchi:1984:TOE

[MRS84] Luca Majocchi, Luca Rodda, and Giacomo R. Sechi. A technique to obtain effective locality of computation in FORTRAN IV. ACM SIGPLAN Notices, 19(1):60–62, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malley:1981:EFP

[MS81] John C. Malley and Ralph M.

Stair, Jr. Essentials of Fortran Programming. The Irwin series in information and decision sciences. R. D. Irwin, Homewood, IL, USA, 1981. ISBN 0-256-02388-3 (paperback), 0-256-02390-5 (paperback). xii + 169 pp. LCCN QA76.73.F25 M34. US$29.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0256023905.

More:1983:CTR

[MS83] Jorge J. Moré and D. C. Sorensen.

Computing a trust region step. SIAM Journal on Scientific and Statistical Computing, 4(3):553– 572, September 1983. CODEN SI-JCD4. ISSN 0196-5204.

Malley:1984:EFP

[MS84a] John C. Malley and Ralph M.

Stair. Essentials of Fortran Programming. The Irwin series in information and decision sciences. R. D. Irwin, Homewood, IL, USA, revised edition, January 1984. ISBN 0-256-02994-6 (paperback). xii + 153 pp. LCCN QA 76.73 F25 M34 1984. US$6.50. URL http:

//www.cbooks.com/sqlnut/SP/

                                        search/gtsumt?source=&isbn=
                                        0256029946.

Moeller:1984:SGV

[MS84b] D. Moeller and Bernd Schmidt. Der Simulator GPSS-Fortran Version 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13782-3.

McComb:1988:ESS

[MS88a] J. McComb and S. Schmidt. Engineering and Scientific Subroutine Library for the IBM 3090 Vector Facility. IBM Systems Journal, 27 (4):404–415, November 1988. CO-DEN IBMSA7. ISSN 0018-8670.

McCracken:1988:IMA

[MS88b] Daniel D. McCracken and W. (William) Salmon. Instructor’s manual to accompany Computing for engineers and scientists with Fortran 77. John Wiley and Sons, New York, London, Sydney, second edition, 1988. ISBN 0-471-60018-0 (paperback). iv + 252 pp.

McCracken:1988:CES

[MS88c] Daniel D. McCracken and William I. Salmon. Computing for Engineers and Scientists With Fortran 77. John Wiley and Sons, New York, London, Sydney, second edition, 1988. ISBN 0-471-62552-3 (paperback). xiii + 730 pp. LCCN TA345 .M3951 1988. US$59.75; US$65.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0471625523.

Murthy:1986:MMI

[MSA86] T. S. Murthy, Y-K Shyy, and J. S.

Arora. MIDAS: Management of information for design and analysis of systems. Advances in engineering software, 8(3), July 1986. CO-DEN AESODT. ISSN 0141-1195, 0965-9978.

Morel-Seytoux:1986:UMM

[MSG86] Hubert J. Morel-Seytoux and Bruno Grawitz. User’s manual for MUNIT: a Fortran V to identify the discrete kernels of a timeinvariant linear system with several excitations. Interim report for 1986. Hydrology Days Publications, Fort Collins, CO, USA, 1986. ISBN ???? vi + 50 pp. LCCN ????

McHenry:1984:CEE

[MSM84] Albert L. McHenry, Alp T. Sabunis, and Lyle B. McCurdy. Cooperative engineering education program project evaluation report on PLATO LDEC course “Structured programming with FOR-TRAN 77”: Fall 1983 semester (August 22–December 16, 1983). Report CR-R; 84009, Dept. of Electronics and Computer (Engineering) Technology College of Engineering and Applied Sciences, Arizona State University, Tempe, AZ, USA, 1984. 12 pp.

McClelland:1987:IGA

[MSR87] Donna McClelland, J. Denbigh Starkey, and Rockford Ross. Instructor’s guide to accompany fundamental programming with Fortran 77: a science and engineering approach. West Publishing Company, St. Paul, MN, USA, 1987. ISBN 0-314-87243-4.

Magnenat-Thalmann:1982:CIL

[MT82a] Nadia Magnenat-Thalmann. Choosing an implementation language for automatic translation. Computer Languages, 7(3-4):161–170, ???? 1982. CODEN COLADA. ISSN 0096-0551.

Ting:1982:CKH

[mT82b] Ming ming Ting. Chieh kou hua fu chuan cheng shih yu yen: CDC FORTRAN 5 (ANSI fu chuan 77). Sung kang tien nao tu shu tzu liao yu hsien kung ssu: Tao ming chu pan she, min kuo 71 [1982], Taipei shih, 1982. ISBN ???? 340 pp. LCCN ????

Marsaglia:1984:FEI

[MT84a] George Marsaglia and Wai Wan Tsang. A fast, easily implemented method for sampling from decreasing or symmetric unimodal density functions. SIAM Journal on Scientific and Statistical Computing, 5(2):349–359, June 1984. CODEN SIJCD4. ISSN 0196-5204.

Myklebust:1984:FVQa

[MT84b] R. L. Myklebust and B. B. Thorne. A Fortran version of the quantitative energy-dispersive electron beam X-ray analysis program FRAME C. NBS technical note 1200, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1984. 39 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Mulders:1983:SOO

[Mul83] H. Mulders. Some observations on the in- and output in high-level languages. ACM SIGPLAN Notices, 18(9):55–58, September 1983. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mular:1985:SMC

[Mul85] Andrew L. Mular. The SPOC manual: chapter 5, unit models (part A): unit models and FORTRAN simulators of ore and coal process equipment: comminution and flotation. CAN-MET special publication SP85-1/5E, Canada Centre for Mineral and Energy Technology, Ottawa, Ontario, Canada, 1985. ISBN 0-660-11864-5. viii + 53 pp.

Mulder:1988:FPM

[Mul88] P. G. Mulder. A FORTRAN program for multivariate survival analysis on the personal computer. Computer Methods and Programs in Biomedicine, 27(2):175–188, September–October 1988. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Mares:1983:VCF

[MW83] Reinhard Mares and Rainer Wojcieszynski. Vektorisieren in Cyber 200, Fortran. Number 3 0723-2187 in Bochumer Schriften zur parallelen Datenverarbeitung. Rechenzentrum der Ruhr-Universität Bochum, Bochum, Germany, 1983. ISBN ???? 315 pp. LCCN ????

Marsr:1984:FPC

[MW84a] F. N. Marsr and I. R. Williams. I. A Fortran program for calculating degenerate Raman bands of symmetric tops with an adaptation for infrared bands. Computer Physics Communications, 35(1–3):C–96, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584823553.

Masri:1984:IFP

[MW84b] F. N. Masri and I. R. Williams. II. A Fortran program for calculating degenerate Raman bands of spherical tops with an adaptation for infrared bands. Computer Physics Communications, 35(1–3):C–97, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584823565.

Myers:1983:IFSa

[Mye83a] Arthur Scott Myers. An implementation of a FORTRAN source code rearranger and documentation generator program. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1983. 112 pp.

Myers:1983:IFSb

[Mye83b] Arthur Scott Myers. An implementation of a FORTRAN source code rearranger and documentation generator program. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1983. 112 pp.

Microsoft:1984:MFCb

[MZ84] Microsoft Corporation and Zenith Data Systems. Microsoft FOR-TRAN compiler (MS-DOS), 1984.

Nadrchal:1986:BRB

[Nad86] J. Nadrchal. Book review: Effective Fortran 77: M. Metcalf, Clarendon Press, Oxford, 1985, xii + 231 p., £9.95. Computer Physics Communications, 42(1): 149–150, September 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465586902390.

Nag:1981:GSS

[Nag81a] Nag. Graphical supplement: a simple plotting system for use with NAG FORTRAN Library. NAg, Downers Grove, IL, USA, 1981. various pp.

Nag:1981:NFM

[Nag81b] Nag. NAg FORTRAN mini manual. NAg [Downers Groves, Ill], Oxford, UK, mark 9 edition, 1981. various pp. Distributed by NAg (USA).

Nag:1985:NGSa

[Nag85] Nag. NAg graphical supplement: a suite of versatile plotting routines for use with the NAg FORTRAN Library. NAg, Oxford, UK, mark 2 edition, 1985. various pp.

Nair:1984:FPC

[Nai84] K. P. Rajappan Nair. A Fortran program for the calculation of hyperfine structure and stark effect in the rotational transition of a 2Σ diatomic molecule. Computer Physics Communications, 34(1– 2):163–174, November/December 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465584901681.

Nair:1986:FPC

[Nai86] K. P. R. Nair. A Fortran program for the calculation of hyperfine structure in the rotational transition of a ^{2}Σ diatomic molecule: II. Magnetic and electric quadrupole interaction from both nuclei. Computer Physics Communications, 41(1):59–73, July 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465586900214.

Nanney:1981:CPA

[Nan81a] T. Ray Nanney. Computing: a problem-solving approach with FORTRAN 77. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1981. ISBN 0-13-165209-5. xiv + 530 pp.

Nanney:1981:CPS

[Nan81b] T. Ray Nanney. Computing: a Problem-Solving Approach With Fortran 77. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, January 1981. ISBN 0-13-165209-5. ???? pp. LCCN ???? US$34.00. URL http:

//www.cbooks.com/sqlnut/SP/

search/gtsumt?source=&isbn=
0131652095.

Neukermans:1981:MCS

[NE81] A. P. Neukermans and S. G. Eaton.

Monte Carlo simulations for electron beam exposures. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12): 24–25, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Nobayashi:1989:CSA

[NE89] H. Nobayashi and C. Eoyang. A comparison study of automatically vectorizing Fortran compilers. In ACM [ACM89b], pages 820–825. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Niederdrenk:1984:EFW

[NEM84] Klaus Niederdrenk and Gisela Engeln-Mullges. Die endliche Fourier- und Walsh-Transformation mit einer Einfuhrung in die Bildverarbeitung: eine anwendungsorientierte Darstellung mit FOR-TRAN 77-Programmen. Friedr. Vieweg and Sohn, Braunschweig/Wiesbaden, 2., neubearb. und erweiterte Aufl. edition, 1984. ISBN 3-528-18535-X. xiii + 208 pp.

Nguyen:1981:NMD

[Ngu81] T. C. Nguyen. Numerical methods in data analysis. BYTE Magazine, 6(??):435–446, ?? 1981. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nickerson:1982:FFP

[Nic82] Robert C. Nickerson. Fundamentals of FORTRAN programming.

Little, Brown and Co., Boston, MA, USA, 1982. ISBN 0-316-60644-8 (paperback). xi + 450 pp.

Nickerson:1985:FFPa

[Nic85a] Robert C. Nickerson. Fundamentals of FORTRAN 77 programming: a structured approach. Little, Brown computer systems series. Little, Brown and Co., Boston, MA, USA, third edition, February 1985. ISBN 0-316-60653-7 (paperback). xiv + 537 pp. LCCN QA76.73.F25 N49 1985. US$16.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0316606537. Rev. ed. of: Fundamentals of FORTRAN programming. 2nd ed. c1980. Includes index.

Nickerson:1985:FFPb

[Nic85b] Robert C. Nickerson. Fundamentals of Fortran 77 Programming: a Structured Approach. Scott, Foresman and Company, Glenview, IL, USA, third edition, February 1985. ISBN 0-673-39039-X. ???? pp. LCCN ???? US$49.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
067339039X.

Nickerson:1985:IMT

[Nic85c] Robert C. Nickerson. Instructor’s manual with transparency masters to accompany Fundamentals of FORTRAN 77 programming: a structured approach. Little, Brown and Co., Boston, MA, USA, third edition, 1985. ISBN 0-316-60654-5. 140 pp.

Nguyen:1981:TAE

[NJLB81] Sang Nguyen, Linda James-Lefebvre, and Andre Babin. TRAFFIC ANSI: an equilibrium traffic assignment program: ANSI FORTRAN version. Publication / Université de Montréal, Centre de recherche sur les transports 220, Université de Montréal, Centre de recherche sur les transports, Montréal, PQ, Canada, 1981. iii + 51 pp.

Nyhoff:1983:PSF

[NL83] Larry R. Nyhoff and Sanford Leestma. Problem Solving With Fortran 77. MacMillan Publishing Company, New York, NY, USA, January 1983. ISBN 0-02-388720-6. xi + 352 pp. LCCN QA 76.73 F25 N93 1983. US$26.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0023887206.

Nakagawa:1985:RMFb

[NL85a] Tomoyasu Nakagawa and Hung-Chi Lai. Reference manual of FORTRAN program ILLOD- (NOR-B) for optimal NOR networks. Technical Report UILU-ENG-85-1727, Univ. of Illinois, Urbana, IL, USA, 1985. 68 pp.

Nakagawa:1985:RMFa

[NL85b] Tomoyasu-Taguti Nakagawa and Hung-Chi Lai. Reference manual of Fortran program ILLOD-(NOR-B) for optimal NOR networks. Technical Report UIUCDCS-R-85-1129, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, July 1985. 73 pp. Revised from Rept. 488.

Nyhoff:1985:FES

[NL85c] Larry R. Nyhoff and Sanford Leestma. FORTRAN 77 for engineers and scientists. Collier Books, Macmillan Publishing Company, New York, NY, USA, 1985. ISBN 0-02-388620-X (paperback). xii + 513 pp. LCCN QA 76.73 F25 N9 1985.

Nyhoff:1985:SMF

[NL85d] Larry R. Nyhoff and Sanford Leestma. Solutions manual, FOR-TRAN 77 for engineers and scientists. Collier Books, Macmillan Publishing Company, New York, NY, USA, 1985. ISBN 0-02-388630-7. v + 99 pp.

Nyhoff:1988:FES

[NL88] Larry R. Nyhoff and Sanford Leestma. Fortran 77 for Engineers and Scientists. Collier Books, Macmillan Publishing Company, New York, NY, USA, second edition, 1988. ISBN 0-02-388631-5 (paperback). xii + 590 pp. LCCN QA76.73.F25 N9 1988. US$40.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0023886315.

Nethercote:1985:FPF

[NM85] W. C. E. Nethercote and G. T. Millar. FORTRAN program for fast surface ship resistance and power estimation: version 3. Technical memorandum 85/212, Defence Research Establishment Atlantic, Dartmouth, NS, Canada, 1985. v + 40 pp.

Nohr:1984:MRF

[Noh84] Elliott C. Nohr. Meetings in retrospect; FORTRAN Celebration at IBM Santa Teresa Laboratory; FORTRAN Activities at SHARE 59 Meeting. Annals of the History of Computing, 6(1):65–69, January/March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1065.pdf;
 http://www.computer.org/annals/
an1984/a1065abs.htm.

Northeast:1983:UIU

[Nor83] E. Northeast. Usersin: an interactive user-interface for Fortran SIN. Report / Department of Defence, Materials Research Laboratories; MRL-R-877 Report (Materials Research Laboratories (Australia)) MRL-R-877, Materials Research Laboratories, Ascot Vale, Victoria, Australia, 1983. ISBN 0-642-07961-7. 3 + [16] + [1] pp.

Norred:1984:MPA

[Nor84] Michael Norred. Mine planning applications in Ratfor. In Software Tools Users Group [Sof84], pages 354–?? ISBN ???? LCCN QA76.8.U65 U83 1984. Abstract only.

Notkin:1989:ASP

[Not89] D. Notkin. Applying software process models to the full lifecycle is premature. ACM SIGSOFT Software Engineering Notes, 14(4): 116–117, June 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nordstrom:1982:LFF

[NSB2 ] Mats Nordstrom, Erik Sandewall, and Diz Breslaw. LISP F1: a FORTRAN implementation of LISP 1.5. Technical report, Uppsala University, Datalogilaboratoriet, Dept. of Computer Science, Uppsala, Sweden, 1982 (??). 73 pp.

Natesan:1981:FP

[NSV1 ] T. R. Natesan, N. Sabapathy, and K. Viswanathan. Fortran programming. Computer Centre, College of Engineering, Guindy, Madras, India, 1981 (??). various pp. Two volumes.

NAG:1981:NGS

[Num81] Numerical Algorithms Group. NAG graphical supplement mark 1. NAG, Oxford, UK, microfiche edition, 1981. various pp.

NAG:1983:FLM

[Num83a] Numerical Algorithms Group, Oxford Eng. Downer’s Grove, IL, USA. FORTRAN Library Manual, Mark 10, 1983. various pp.

NAG:1983:NFL

[Num83b] Numerical Algorithms Group. NAg FORTRAN library manual. Numerical Algorithms Group, Oxford, United Kingdom, MARK 10 edition, 1983. ???? pp.

NAG:1983:NFP

[Num83c] Numerical Algorithms Group. The NAG FORTRAN PC50 library handbook. Numerical Algorithms Group, Oxford, UK, release 1 edition, 1983. 26 + 324 + 8 pp.

NAG:1984:NFL

[Num84a] Numerical Algorithms Group. NAg FORTRAN library manual. Numerical Algorithms Group, Oxford, United Kingdom, MARK 11 edition, 1984. various pp.

NAG:1984:NFMa

[Num84b] Numerical Algorithms Group.

NAG Fortran manual mark II. Numerical Algorithms Group, Oxford, UK, 1984. 36 microfiche. pp.

NAG:1984:NFMb

[Num84c] Numerical Algorithms Group. NAg FORTRAN mini manual: an introductory guide to the NAg FOR-TRAN manual. Numerical Algorithms Group, Downers Grove, IL, USA, mark 11 edition, 1984. various pp.

NAG:1984:NFP

[Num84d] Numerical Algorithms Group. The NAG FORTRAN PC50 library handbook: mathematical problem solving for FORTRAN programmers using IBM personal computers. Numerical Algorithms Group Ltd., Oxford, Oxfordshire, second edition, 1984. various pp.

NAG:1985:NGSb

[Num85a] Numerical Algorithms Group.

NAG graphical supplement mark 2. NAG, Oxford, UK, microfiche edition, 1985. various pp.

NRS:1985:NR

[Num85b] Numerical Recipes Software. Numerical recipes, 1985. ISBN 0-521-30958-1 (v.1), 0-521-30957-3 (v.2).

NRS:1985:NRF

[Num85c] Numerical Recipes Software. Numerical recipes FORTRAN diskette VI.1, 1985. ISBN 0-521-30958-1.

NRS:1986:NRF

[Num86] Numerical Recipes Software. Numerical Recipes Fortran. Cambridge University Press, Cambridge, UK, January 1986. ISBN 0-521-30958-1. ???? pp. LCCN ???? US$34.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0521309581.

NAG:1987:NFL

[Num87] Numerical Algorithms Group. NAg FORTRAN library manual. Numerical Algorithms Group, Oxford, United Kingdom, MARK 12 edition, 1987. various pp.

NAG:1988:HSM

[Num88a] Numerical Algorithms Group. The Harwell sparse matrix library manual: a selection of 19 Fortran 77 packages for handling sparse linear and nonlinear problems. UKAEA, Oxfordshire, UK, release, 1st edition, 1988. various pp.

NAG:1988:NFLa

[Num88b] Numerical Algorithms Group. NAG FORTRAN library introductory guide: Mark 13. Numerical Algorithms Group, Oxford, UK, 1988. various pp.

NAG:1988:NFLb

[Num88c] Numerical Algorithms Group. NAg FORTRAN library manual, mark 13. Numerical Algorithms Group, Oxford, United Kingdom, 1988. various pp.

NRS:1988:NRF

[Num88d] Numerical Recipes Software. Numerical recipes F Macintosh FOR-TRAN subroutines and functions from Numerical Recipes, the Art of Scientific Computing, 1988. ISBN 0-521-35469-2.

NRS:1989:NRF

[Num89] Numerical Recipes Software. Numerical recipes FORTRAN diskette, 1989. ISBN 0-521-30958-1. Distributed by Numerical Recipes Software.

Oblow:1985:GGS

[Obl85] E. M. Oblow. GRESS: Gradientenhanced software system. Version D user’s guide. Tech. report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1985.

Offutt:1987:FIM

[OK87] A. Jefferson Offutt, VI and K. N.

King. A Fortran 77 interpreter for mutation analysis. In Wexelblat [Wex87], pages 177–188. ISBN 0-89791-235-7. LCCN QA76.7 .S54 v.22:7. US$23.00. URL

http://www.acm.org:80/pubs/
citations/proceedings/plan/
29650/p177-offutt/. SIGPLAN

Notices, v. 22, no. 7 (July 1987).

Oliarnyk:1981:ISF

[Oli81] M. Oliarnyk. An introduction to Spectre and Fortran. Media Algonquin, Nepean, Ontario, Canada, 1981. ISBN 0-88620-014-8. 95 + 16 pp.

Olsson:1983:MUL

[Ols83] O. Olsson. The memory usage of a LISP system: the Belady life-time function. ACM SIGPLAN Notices, 18(12):112–119, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OFlaherty:1982:SAA

[OM82] M. O’Flaherty and G. MacKenzie. Statistical algorithms: Algorithm AS 172: Direct simulation of nested Fortran DO-LOOPS. Applied Statistics, 31(1):71–74, March 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/172.

ONeal:1981:FIP

[O’N81] T. R. O’Neal. A FORTRAN IV program to scale and plot logarithmic and linear data. Nrl memorandum report; 4480, Naval Research Laboratory, Washington, DC, USA, 1981. iii + 43 pp.

Onibere:1985:WPF

[Oni85] E. A. Onibere. Writing portable Fortran programs for microcomputer. Software—Practice and Experience, 15(4):321–326, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Okada:1986:PFM

[OO86] M. Okada and M. Okada. Portable file management system in FOR-TRAN. II. the input/output routine for free-format text. Computer Methods and Programs in Biomedicine, 22(2 199-):207–??, April 1986. CODEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

OReilly:1981:FCP

[O’R81] Liam O’Reilly. A Fortran computer programme to analyse fluctuations of currencies in the EMS. Technical paper 2/RT/81, Research Dept., Central Bank of Ireland, Dublin, Ireland, February 1981. 11 + [5] pp.

Osipov:1982:IAE

[Osi82a] L. A. (Lev Aleksandrovich) Osipov. IAzyk analitik i ego sravnenie s iazykami ALGOL i FOR-TRAN. Bibliotechka programmista. “Nauka,” Glav. red. fizikomatematicheskoi lit-ry, Moskva, USSR, 1982. ISBN ???? 160 pp. LCCN ????

Osipov:1982:JAE

[Osi82b] L. A. (Lev Aleksandrovich) Osipov. Jazyk analitik i ego sravnenie s jazykami Algol i Fortran. Biblioteka programmista. Nauka, Moscow, Russia, 1982. ISBN ???? 160 pp. LCCN ????

Osyczka:1984:MOE

[Osy84] Andrzej Osyczka. Multicriterion optimization in engineering with FORTRAN programs. Ellis Horwood series in engineering science.

Ellis Horwood series in mechanical engineering. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, March 1984. ISBN 0-85312-481-7 (Ellis Horwood), 0-470-20019-7 (Halsted Press). viii + 178 pp. LCCN TA342 .O89 1984. UK£21.00. Translated from the Polish by B. J. Davies.

Ottenstein:1981:SDS

[Ott81] Linda M. Ottenstein. Software defects — a software science perspective. ACM SIGMETRICS Performance Evaluation Review, 10(1): 153–155, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Otter:1987:FEH

[Ott87] Martin Otter. FORTRAN 77 error handling of the RASP library. Mitteilung / Deutsche Forschungs — und Vesuchsanstalt fur Luftund Raumfahrt 87-14, Deutsche Forschungs- und Versuchsanstalt fur Luft- und Raumfahrt, Köln, 1987. 51 pp.

Owen:1986:APA

[Owe86] Susan M. Owen. Adding pipeline awareness to a FORTRAN compiler. Thesis (M.S.), University of Pittsburgh, Pittsburgh, PA, USA, 1986. v + 68 pp.

Owen:1987:

[Owe87] Albert Kimsey Owen. A FOR-TRAN code for the calculation of probe volume geometry changes in a laser anemometry system caused by window refraction. Nasa technical memorandum; 100210 avscom technical report; 87-c-27, National Aeronautics and Space Administration, Washington, DC, USA, 1987. ???? pp. For sale by the National Technical Information Service.

Perrott:1983:PFP

[PA83] Ronald H. Perrott and Donald Allison. Pascal for Fortran Programmers. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, December 1983. ISBN 0-914894-09-9. ???? pp. LCCN ???? US$26.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0914894099.

Perrott:1984:PFP

[PA84] Ronald H. Perrott and Donald C. S. Allison. PASCAL for FOR-TRAN programmers. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, 1984. ISBN 0-914894-09-9 (paperback). xi + 335 pp. LCCN QA76.73.P2 P47 1984.

Paddock:1985:SFB

[Pad85] Charles E. Paddock. Structured Fortran for Business. Prentice-Hall series in computer programming languages. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1985. ISBN 0-13-854233-3 (paperback). xv + 253 pp. LCCN ???? US$32.80. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0138542333.

Page:1983:FPG

[Pag83] Clive Page. FORTRAN 77: pocket guide. Programming pocket guides. Pitman Publishing Ltd., London, UK, 1983. ISBN 0-273-01973-2 (paperback). ii + 62 pp.

Page:1984:PGF

[Pag84] Ulive Page. Pocket Guide: Fortran 77. Pitman Publishing Ltd., London, UK, July 1984. ISBN 0-273-01973-2. ???? pp. LCCN ???? US$6.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0273019732.

Pagiola:1987:FVA

[Pag87] E. Pagiola. FORTRAN vs? answer; casual. Computer Physics Communications, 45(1– 3):485, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465587901949.

Page:1988:PPG

[Pag88] Clive Page. The professional programmers guide to Fortran 77. The Professional programmers guide to-. Pitman Publishing Ltd., London, UK, 1988. ISBN 0-273-02856-1 (paperback). ix + 180 pp. LCCN QA76.73.F25 P34 1988.

Palmer:1986:FPD

[Pal86] J. A. B. Palmer. A Fortran procedure for drawing some space-filling curves. Software— Practice and Experience, 16(6):

559–574, June 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Parlett:1984:SSE

[Par84] B. N. Parlett. The software scene in the extraction of eigenvalues from sparse matrices. SIAM Journal on Scientific and Statistical Computing, 5(3):590–604, September 1984. CODEN SIJCD4. ISSN 0196-5204.

Parker:1986:SFCa

[Par86a] J. R. Parker. A subset FORTRAN compiler for a modified Harvard architecture. Technical Report 86/230/4, University of Calgary, Calgary, Alberta, Canada, Febru-

ary 1986. URL mailto::parin@
cpsc.ucalgary.ca.

Parker:1986:SFCb

[Par86b] J. R. Parker. A subset FOR-TRAN compiler for a modified Harvard architecture. ACM SIGPLAN Notices, 21(9):57–62, September 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Patterson:1989:EFG

[Pat89] M. R. Patterson. EZVIDEO, FORTRAN graphics routines for the IBM AT. Technical Report ORNL/CSD/TM-265, Martin Marietta Energy Systems, Oak Ridge, TN, USA, October 1989. v + 29 pp.

Payne:1984:DAPa

[Pay84a] Kendall Robert Payne. The development and analysis of a portable runtime library accessible to all FORTRAN, COBOL and PAS-CAL compilers under the UNIX system 5 operating system. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1984. 70 pp.

Payne:1984:DAPb

[Pay84b] Kendall Robert Payne. The development and analysis of a portable runtime library accessible to all FORTRAN, COBOL and PAS-CAL compilers under the UNIX system 5 operating system. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1984. 70 pp.

Payne:1984:FPC

[PB84] G. L. Payne and P. L. Von Behren. Fortran program to calculate finite-range no-recoil DWBA transfer cross sections. Computer Physics Communications, 35 (1–3):C–223–C–224, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584824637.

Plamondon:1986:ORH

[PB86] R. Plamondon and R. Baron. On-line recognition of handprint schematic pseudocode for automatic FORTRAN code generation. In Association française de cybernetique économique et technique [Ass86], pages 741–744. ISBN 0-8186-0742-4 (paperback), 0-8186-4742-6 (microfiche), 0-8186-8742-8 (hardcover). LCCN Q 327 I615 1986. IEEE Computer Society order number 742. IEEE catalog number 86CH2342-4. On spine: 1986 International Conference on Pattern Recognition. Includes bibliographies and index.

Parsian:1988:ATT

[PBB^{+}88] Mahmoud Parsian, Brayan Basdell, Yusuf Bhayat, Ian Caldwell, Neva Garland, Bruce Jubanowsky, and Jeanne Robinette. Ada translation tools development: Automatic translation of FORTRAN to Ada. ACM SIGADA Ada Letters, 8(6):57–71, November/December 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pardalos:1989:PAQ

[PC89] P. M. Pardalos and J. V. Crouse.

A parallel algorithm for the quadratic assignment problem. In ACM [ACM89b], pages 351–360. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Petkov:1984:RDE

[PCK84] P. Hr. Petkov, N. D. Christov, and M. M. Konstantinov. Remark on “Algorithm 590: DSUBSP and EXCHQZ: FORTRAN subroutines for computing deflating subspaces with specified spectrum”. ACM Transactions on Mathematical Software, 10(2):207, June 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Van82].

Perrott:1981:EFP

[PD81] R. H. Perrott and P. S. Dhillon. An experiment with Fortran and Pascal. Software—Practice and Experience, 11(5):491–496, May 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Page:1983:FH

[PDA83a] Rex L. Page, Richard L. Didday, and Elizabeth Alpert. Fortran ’77 for Humans. West Publishing Company, St. Paul, MN, USA, second edition, 1983. ISBN 0-314-69672-5 (paperback). xii + 450 pp. LCCN QA76.73.F25 P3 1983. US$22.95; US$14.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0314696725.

Page:1983:IMA

[PDA83b] Rex L. Page, Richard L. Didday, and Elizabeth Alpert. Instructor’s manual to accompany FORTRAN 77 for humans. West Publishing Company, St. Paul, MN, USA, second edition, 1983. ISBN 0-314-71115-5 (paperback). iii + 94 pp.

Page:1986:FH

[PDA86] Rex L. Page, Richard L. Didday, and Elizabeth Alpert. Fortran ’77 for Humans. West Publishing Company, St. Paul, MN, USA, third edition, April 1986. ISBN 0-314-93404-9 (paperback). xiii + 462 pp. LCCN QA76.73.F25 P3 1986. US$43.25; US$24.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0314934049.

Peerless:1984:FSS

[Pee84a] Peerless. Fortran Scientific Subroutine Library: IBM PC. John Wiley and Sons Software, New York, NY, USA; London, UK; Sydney, Australia, October 1984. ISBN 0-471-81457-1. ???? pp. LCCN ???? US$250.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0471814571. Includes two 128K floppy disks.

PES:1984:FSSa

[Pee84b] Peerless Engineering Service.

FORTRAN: scientific subroutine library. John Wiley and Sons, New York, London, Sydney, 1984. ISBN 0-471-81457-1. x + 411 pp. LCCN QA76.73.F25F678 1984. US$150.00. System requirements for floppy disks: IBM PC or PC/XT or compatible computer; 128K; PC DOS version 1.1 or later or MS DOS compatible with FOR-TRAN compiler; copy of Microsoft FORTRAN 77 compiler, version 3.13 or later; 2 dual sided disk drives; Intel 8087 NDP optional.

PES:1984:FSSb

[Pee84c] Peerless Engineering Service.

FORTRAN; scientific subroutine library. John Wiley and Sons, New York, London, Sydney, 1984. ISBN 0-471-81457-1. various pp.

PES:1984:FSSc

[Pee84d] Peerless Engineering Service.

FORTRAN scientific subroutine library, 1984. ISBN 0-471-81457-1.

PES:1985:MFSa

[Pee85a] Peerless Engineering Service. 50 more: FORTRAN scientific subroutines and utilities. John Wiley and Sons, New York, London, Sydney, 1985. ISBN ???? 1 v. disks (5 1/4 in.) pp. LCCN ????

PES:1985:MFSb

[Pee85b] Peerless Engineering Service. 50 more: FORTRAN scientific subroutines and utilities. John Wiley and Sons, New York, London, Sydney, 1985. ISBN 0-471-00910-5. 400 pp.

PES:1985:MFSc

[Pee85c] Peerless Engineering Service. 50 more: FORTRAN scientific subroutines and utilities. John Wiley and Sons, New York, London, Sydney, 1985. ISBN 0-471-84449-7. 1 v. disks (5 1/4 in.) pp.

PES:1985:PFSa

[Pee85d] Peerless Engineering Service. Professional FORTRAN; scientific subroutine library. John Wiley and Sons, New York, London, Sydney, 1985. ISBN 0-471-82954-4. various pp.

PES:1985:PFSb

[Pee85e] Peerless Engineering Service. Professional FORTRAN scientific subroutine library, 1985.

PES:1986:PFS

[Pee86] Peerless Engineering Service. Professional FORTRAN scientific subroutine library, 1986. ISBN 0-471-82954-4.

PES:1989:FSS

[Pee89] Peerless Engineering Service.

FORTRAN: scientific subroutine library. John Wiley and Son, New York, NY, USA, version 2.0 edition, 1989. ISBN 0-471-51499-3. x + 414 pp.

Pemberton:1983:TCT

[Pem83] Steven Pemberton. Technical correspondence: On Tanenbaum, van Staveren, and Stevenson’s “Using Peephole Optimization on Intermediate Code”. ACM Transactions on Programming Languages and Systems, 5(3):499, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [TvSS82, Tan83b].

Perry:1981:FCF

[Per81] William J. Perry. FABSTAT: a card-imaged FORTRAN program to be used for axially distributed planar-structural data. Open-file report 81-158, U.S. Geological Survey, Denver, CO, USA, 1981. 29 pp.

Perez:1983:BWC

[Per83a] Aram Perez. Byte-wise CRC calculations. IEEE Micro, 3(3):40– 50, May/June 1983. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Perko:1983:RDS

[Per83b] A. Perko. A representation of disjoint sets with fast initialization. Information Processing Letters, 16 (1):21, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Perrott:1987:PP

[Per87] R. H. Perrott. Parallel Programming. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-14231-7. LCCN QA76.6.P463 1987.

Petersen:1983:VFN

[Pet83] W. P. Petersen. Vector Fortran for numerical problems on Cray-1. Comm. ACM, 26(11): 1008–1021, November 1983. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pettit:1987:FLA

[Pet87] F. R. (Francis Richard) Pettit. Fortran lectures at Oxford. A Chartwell-Bratt student text. Chartwell-Bratt, Bromley, Kent, UK, 1987. ISBN 0-86238-122-3 (paperback), 91-44-26521-2 (Sweden). 136 pp.

Petti:1988:RSM

[Pet88] R. Petti. Role of symbolic mathematics software in mathematical modeling. American Society of Mechanical Engineers, Heat Transfer Division, (Publication) HTD, 105:13–20, 1988. CODEN ASMHD8. ISSN 0272-5673.

Petersen:1989:PTS

[Pet89] Paul M. Petersen. PDE tutor: a system for the automatic solution of partial differential equations. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1989. v + 50 pp.

Plitt:1985:SMC

[PF85] L. R. Plitt and Brian C. Flintoff.

The SPOC manual: chapter 5.1, unit models (part B): unit models and FORTRAN simulators of oreand coal process equipment: classification and coal processing. CAN-MET special publication SP85-1/5.1E, Canada Centre for Mineral and Energy Technology, Ottawa, Ontario, Canada, 1985. ISBN 0-660-11865-3. x + 132 pp.

Pattnaik:1983:INS

[PFF83] P. C. Pattnaik, G. Fletcher, and J. L. Fry. Improved numerical stability for norm-conserving ion-Ure pseudopotentials. Physical Review B: Condensed Matter and Materials Physics, 28(6):3364–3365, 1983. CODEN PRBMDO. ISSN 1098-0121.

Press:1986:NRA

[PFTV86] William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes: The Art of Scientific Computing (FORTRAN Version). Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-38330-7. xx + 702 pp. LCCN QA297 .N866 1989.

Phillips:1986:NLB

[Phi86] Jen Phillips. The NAG library: a beginner’s guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-853263-6. viii + 245 pp. LCCN QA297.P53 1986. US$14.50 (U.S.).

Phillips:1987:NLB

[Phi87] Jen Phillips. The Nag Library: a Beginner’s Guide. Clarendon Press, Oxford, UK, June 1987. ISBN 0-19-853263-6. 245 pp. LCCN ???? US$35.00.

Pierchala:1985:IMU

[Pie85]  Carl D. Pierchala.  An im-

provement for the McGill University Random Number Package. Computational Statistics & Data Analysis, 2(4):317–322, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Patel:1984:PPR

[PJ84] N. R. Patel and H. F. Jordan. A parallelized point rowwise successive over-relaxation method on a multiprocessor. Parallel Computing, 1(3–4):207–222, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pignataro:1985:ELO

[PLR85] M. Pignataro, A. Luongo, and N. Rizzi. On the effect of the local overall interaction on the postbuckling of uniformly compressed channels. Thin-Walled Structures, 3:292–321, 1985. CO-DEN TWASDE. ISSN 0263-8231.

Pease:1987:FTC

[PM87] J. M. Pease and M. F. Morselli. A FORTRAN technique for correlating a circular environmental variable with a linear physiological variable in the sugar maple. Computers in Biology and Medicine, 17 (5):363–367, ???? 1987. CODEN CBMDAW. ISSN 0010-4825.

Pleasant:1982:RFIa

[PMBK82a] James C. Pleasant, L. M.

McDowell-Boyer, and G. G. Killough. RAGBEEF, a FORTRAN IV implementation of a timedependent model for radionuclide contamination of beef. Technical report, The Commission: Available from GPO Sales Program, Division of Technical Information and Document Control, U.S. Nuclear Regulatory Commission; Available from National Technical Information Service, Washington, DC, USA, 1982. v + 140 pp.

Pleasant:1982:RFIb

[PMBK82b] James C. Pleasant, L. M.

McDowell-Boyer, and G. G. Killough. RAGBEEF, a FORTRAN IV implementation of a timedependent model for radionuclide contamination of beef. Technical report, The Commission: Available from GPO Sales Program, Division of Technical Information and Document Control, U.S. Nuclear Regulatory Commission; Available from National Technical Information Service, Washington, DC, USA, 1982. v + 140 pp.

Pollack:1981:SFP

[Pol81] Seymour V. Pollack. Structured Fortran 77 Programming. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, February 1981. ISBN 0-87835-095-0. xvi + 496 pp. LCCN QA76.73.F25P64 1982. US$27.50. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0878350950.

Pollack:1982:SFP

[Pol82] Seymour V. Pollack. Structured FORTRAN 77 programming. Boyd and Fraser computer science series. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, 1982. ISBN 0-87835-095-0 (paperback). xvi + 496 pp. LCCN QA76.73.F25 P64 1982.

Pollack:1983:SFP

[Pol83] Seymour V. Pollack. Structured Fortran 77 Programming With Hewlett–Packard Computers. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, October 1983. ISBN 0-87835-130-2 (paperback). xix + 524 pp. LCCN QA76.73.F25 P64 1982. US$37.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0878351302.

Polychronopoulos:1987:ARF

[Pol87] Constantine D. Polychronopoulos. Automatic restructuring of Fortran programs for parallel execution. Report CSRD 665, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL, USA, 1987. 26 pp.

SIGPLAN:1982:CRN

[POP82] Conference Record of the Ninth Annual ACM Symposium on Principles of Programming Languages. ACM Press, New York, NY 10036, USA, January 1982. ISBN ???? LCCN ????

Pountain:1987:OOF

[Pou87] Dick Pountain. Object-Oriented Forth: Implementation of Data Structures. Academic Press, New York, NY, USA, September 1987. ISBN 0-12-563570-2. vii + 119 pp. LCCN QA76.73.F24 P681 1987. US$37.00.

Prothero:1982:TDR

[PP82a] J. Prothero and J. Prothero.

Three-dimensional reconstruction from serial sections. I. a portable microcomputer-based software package in Fortran. Computers and Biomedical Research, 15: 598–604, December 1982. CO-DEN CBMRB7. ISSN 0010-4809 (print), 1090-2368 (electronic).

Prothero:1982:TRS

[PP82b] J. Prothero and J. Prothero.

Three-dimensional reconstruction from serial sections. I. A portable microcomputer-based software package in Fortran. Computers and Biomedical Research, 15: 598–604, December 1982. CO-DEN CBMRB7. ISSN 0010-4809 (print), 1090-2368 (electronic).

Prothero:1985:TSP

[PP85] J. D. Prothero and J. W. Prothero.

TOPPER, a software package in FORTRAN for scaling studies. International Journal of Bio-Medical Computing, 17(3-4):185– 191, November 1985. CODEN IJBCBT. ISSN 0020-7101 (print), 1878-366X (electronic).

Pratt:1984:PL

[Pra84] T. W. Pratt. Programming Languages. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1984. ISBN 0-13-730580-X. xix + 604 pp. LCCN QA76.7 .P7 1984.

Pratt:1985:PEP

[Pra85] Terrence W. Pratt. Pisces: an environment for parallel scientific computation. IEEE Software, 2 (4):7–20, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pracht:1989:TCI

[Pra89] Bregitte R. Pracht. Translation of common image processing transforms into Image Algebra and Image Algebra FORTRAN. Thesis (M.S.), University of Florida, Gainesville, FL, USA, 1989. vii + 206 pp.

Press:1987:NRFb

[Pre87a] W. H. Press. Numerical Recipes Fortran Disk for MacIntosh V1.1. Cambridge University Press, Cambridge, UK, December 1987. ISBN 0-521-35469-2. ???? pp. LCCN ???? US$34.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0521354692.

Press:1987:NRFa

[Pre87b] William H. Press. Numerical recipes FORTRAN diskette, 1987.

Press:1988:NRE

[Pre88a] W. H. Press. Numerical Recipes Example Disk (Fortran for MacIntosh). Cambridge University Press, Cambridge, UK, September 1988. ISBN 0-521-35468-4. ???? pp. LCCN ???? US$34.50.

Press:1988:NRC

[Pre88b] William H. Press. Numerical recipes in C: the art of scientific computing. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35465-X, 0-521-35466-8 (diskette). xxii + 735 pp. LCCN QA76.73.C15 N865 1988.

Pressman:1988:PGF

[Pre88c] Michael H. Pressman. Programmer’s Guide to Fortran: a Second Course. Wm. C. Brown Publishers, Dubuque, IA, USA, June 1988. ISBN 0-697-07835-3 (paperback). xvi + 383 pp. LCCN QA76.73.F25P73 1988. US$43.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0697078353.

Press:1989:NRF

[Pre89] William H. Press. Numerical recipes FORTRAN format, Apple Macintosh, 1989. ISBN 0-521-35469-2.

Price:1989:BT

[Pri89] Walter J. Price. A benchmark tutorial. IEEE Micro, 9(5):28– 43, September/October 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Prothero:1985:TRS

[PRL^{+}85] J. S. Prothero, M. Riggins, A. Lindsay, R. Harris, and J. W.

Prothero. Three-dimensional reconstruction from serial sections III. AUTOSCAN, a software package in FORTRAN for semiautomated photomicrography. Computers in Biomedical Research, 18 (2):132–136, April 1985. CODEN CPBRAF. ISSN 0010-4809 (print), 1090-2368 (electronic).

Pruett:1987:FPF

[Pru87] James M. Pruett. Fundamentals of Programming With Fortran 77. An Independent learning module from the Instrument Society of America. Instrument Society of America, Research Triangle Park, NC, USA, January 1987. ISBN 0-87664-958-4. viii + 301 pp. LCCN QA76.73.F25P78 1987. US$50.00, US$39.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0876649584.

Platt:1989:RAP

[PRWB89] D. E. Platt, A. R. Rossi, J. W. Wells, and J. Becker. Realities associated with parallel processing. In ACM [ACM89b], pages 162–174. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Perrott:1981:CDA

[PS81] R. H. Perrott and D. K. Stevenson. Considerations for the design of array processing languages. Software—Practice and Experience, 11 (6):683–688, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perry:1982:PIF

[PS82] K. A. Perry and J. G. Szekely.

PSOD: an interactive FORTRAN program to simulate the radiation dose responses of membrane populations. Technical report, Whiteshell Nuclear Research Establishment, Pinawa, Manitoba, Canada, 1982. 91 pp.

Perry:1983:MIF

[PS83] K. A. Perry and J. G. Szekely.

MONTEC: an interactive FOR-TRAN program to simulate radiation dose and dose rate responses of populations. Technical Report AECL-7794, Whiteshell Nuclear Research Establishment, Pinawa, Manitoba, Canada, 1983. 67 pp.

Pomponiu:1984:FAS

[PS84] C. Pomponiu and M. Sararu.

Fourier analysis with splines, a Fortran program. Computer Physics Communications, 35(1– 3):C–523, ???? 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584827046.

Parkinson:1984:CAG

[PW84] D. Parkinson and M. Wunderlich.

A compact algorithm for Gaussian elimination over GF(2) implemented on highly parallel computers. Parallel Computing, 1(1):65– 73, August 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Padua:1986:ACO

[PW86] D. Padua and M. Wolfe. Advanced compiler optimizations for supercomputers. Comm. ACM, pages 1184–1201, December 1986. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pyramid:1984:FRD

[Pyr84] Pyramid Technology Corporation. Fortran 77 release description. Pyramid Technology Corporation, Mountain View, CA, USA, revision A edition, 1984. various pp.

Perrott:1986:SL

[PZA86] R. H. Perrott and A. Zarea-Aliabadi. Supercomputer languages. ACM Computing Surveys, 18(1):5–22, March 1986. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Radford:1981:CPF

[Rad81] Arthur S. Radford. Computer Programming Fortran (Teach Yourself Books). Teach Yourself, ????, June 1981. ISBN 0-679-10378-3. ???? pp. LCCN ???? US$8.95.

Radford:1983:CPFb

[Rad83a] Arthur S. Radford. Computer programming in FORTRAN. Teach Yourself, Sevenoaks, UK, 1983. ISBN 0-340-27587-1 (paperback). x + 238 pp.

Radford:1983:CPFa

[Rad83b] Arthur S. Radford. Computer Programming in Fortran: Teach Yourself. Teach yourself books. Hodder and Stoughton, London, UK, December 1983. ISBN 0-340-27587-1. 238 pp. LCCN ???? US$5.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0340275871.

Rager:1986:GPT

[Rag86] John E. Rager. Graphics packages for teaching graphics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(1): 225–231, February 1986. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/
pub/mirrors/ftp.ira.uka.de/
bibliography/Graphics/imager/
1986.bib;   ftp://ftp.math.
utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Graphics/
siggraph/1986.bib;   ftp://
ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/
Misc/DBLP/1986.bib. Proceed-

ings of the 17th SIGCSE Symposium on Computer Science Education.

Raicevic:1984:GLF

[Rai84] Nebojsa Raicevic. Govorite li Fortran. Pegaz ’84; 75. Knizevna omladina Srbije, Beograd, Yugoslavia, 1984. ISBN ???? 38 pp. LCCN ????

Rempfer:1988:ALD

[RAKK88] Dietmar Rempfer, Monika Auweter-Kurtz, and Hans J. Käppeler. Aufstellung und Lösung der Dispersionsgleichung zur Untersuchung von Instabilitäten in MPD-Triebwerken mit Hilfe eines MACSYMA-FORTRANHybrid-Codes. (German) [List and solution of dispersion equations for the investigation of instabilities in MPD-engines with the help of MACSYMA-FORTRAN hybrid codes]. Report 88-S18, Institut für Raumfahrtsysteme, Universität Stuttgart, Stuttgart, Germany, 1988. ii + 100 pp.

Rao:1981:DMD

[Rao81a] Natti S. Rao. Designing machines and dies for polymer processing with computer programs: FOR-TRAN and BASIC. Hanser International, München, Germany, 1981. ISBN 0-02-949630-6. 207 pp. LCCN TP1135 .R37. Distributed by Macmillan.

Rao:1981:CPF

[Rao81b] P. V. S. Rao. Computer programming, FORTRAN and other languages. Tata McGraw-Hill, New Delhi, India, 1981. ISBN 0-07-096569-2 (paperback). xx + 304 pp.

Rao:1982:CPF

[Rao82] P. V. Rao. Computer Programming in Fortran and Other Languages. McGraw-Hill, New York, NY, USA, June 1982. ISBN 0-07-096569-2. ???? pp. LCCN ???? US$3.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0070965692.

Rao:1983:DMD

[Rao83] Natti S. Rao. Designing machines and dies for polymer processing with computer programs: FOR-TRAN and BASIC. Hanser International, München, Germany, second edition, 1983. ISBN 3-446-13500-6, 0-02-949630-6 (NY). 207 pp. Distributed by Macmillan Publishing Co. Inc.

Rao:1986:CADa

[Rao86a] Natti S. Rao. Computer Aided Design of Plasticating Screws: Programs in Fortran and Basic. Carl Hanser, München, Germany, June 1986. ISBN 0-19-520754-8. ???? pp. LCCN ???? US$49.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0195207548.

Rao:1986:CADb

[Rao86b] Natti S. Rao. Computer Aided Design of Plasticating Screws: Programs in Fortran and Basic. Carl Hanser, München, Germany, June 1986. ISBN 0-02-947600-3 (NY), 3-446-14331-9 (Munich), 1-56990-082-5. 134 pp. LCCN ???? US$47.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

1569900825. Distributed in the U.S. by MacMillan Pub. Co.

Rao:1987:PF

[Rao87] P. V. S. Rao. Programming in Fortran. Tata McGraw-Hill, New Delhi, India, 2eme edition, 1987. ISBN 0-07-451857-7. xvi + 377 pp.

Rapp:1982:FPC

[Rap82a] Richard H. Rapp. A FORTRAN program for the computation of gravimetric quantities from high degree spherical harmonic expansions. Technical report, Ohio State University, Columbus, OH, USA, 1982. 23 pp.

Rapport:1982:FPCa

[Rap82b] Amy Rapport. FORTRAN program to compute chemical geothermometers for geothermal fluids. Open-file report 82-308, The Survey, Menlo Park, CA, USA, 1982. ???? pp.

Rapport:1982:FPCb

[Rap82c] Amy Rapport. FORTRAN program to compute chemical geothermometers for geothermal fluids. Open-file report 82-308, U.S. Geological Survey, Menlo Park, CA, USA, 1982. 25 pp.

Rasmussen:1984:FPS

[Ras84] J. L. Rasmussen. A Fortran program for statistical evaluation of pseudorandom number generators. Behavior Research Methods and Instrumentation, 16(??): 63–64, ???? 1984. CODEN BR-MIAC. ISSN 0005-7878.

Ratzer:1981:FC

[Rat81] Gerald F. G. Ratzer. A FOR-TRAN 77 course. Kendell, Hunt, Dubuque, IA, USA, 1981. ISBN 0-8403-2427-8. 145 pp.

Ratzer:1986:FC

[Rat86] Gerald F. G. Ratzer. A FOR-TRAN 77 course. Kendall/Hunt Pub., Dubuque, IA, USA, second edition, 1986. ISBN 0-8403-3947-X. various pp.

Ratzer:1987:FC

[Rat87] Gerald Ratzer. A Fortran 77 Course. Kendall/Hunt Pub., Dubuque, IA, USA, May 1987. ISBN 0-8403-2427-8. ???? pp. LCCN ???? US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0840324278.

RTI:1989:IECa

[Rat89] Rational Technology, Inc. IN-GRES/EQUEL companion guide for BASIC; INGRES/ embedded SQL companion guide for BA-SIC; INGRES/EQUEL companion guide for COBOL; INGRES/ embedded SQL companion guide for COBOL; INGRES/EQUEL companion guide for FORTRAN; IN-GRES/ embedded SQL companion guide for FORTRAN. Relational Technology Inc., Alameda, CA, USA, 1989. 6 v. in 1 pp.

Rees:1982:BRBb

[RB82] Michael J. Rees and D. W. Barron. Book review: Programming in standard FORTRAN 77, A. Balfour and D. H. Marwick, Heinemann Educational Books, London, 1979. No. of pages: 388. Price: £4.50 (softback). Software —Practice and Experience, 12(10): 983–984, October 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rouse:1983:PIPa

[RB83] Robert A. Rouse and Thomas L.

Bugnitz. Programming the IBM Personal Computer: FORTRAN 77. IBM personal computer series.

Programming. Holt, Reinhart, and Winston, New York, NY, USA, 1983. ISBN 0-03-062042-2 (paperback). xiv + 240 pp. LCCN QA76.8.I2594 R68 1983.

Redner:1982:FPC

[Red82] S. Redner. A FORTRAN program for cluster enumeration. Journal of Statistical Physics, 29(2): 309–315, October 1982. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/BF01020788.

Reddy:1986:SGE

[Red86] T. S. R. Reddy. Symbolic generation of elastic rotor blade equations using a FORTRAN processor and numerical study on dynamic inflow effects on the stability of helicopter rotors. NASA technical memorandum 86750, National Aeronautics and Space Administration, Ames Research Center, Moffett Field, CA, USA, 1986. ???? pp. For sale by the National Technical Information Service.

Reid:1984:PSM

[Rei84a] J. K. Reid. TREESOLVE: a Fortran package for solving large sets of linear finite-element equations. with a discussion. In PDE software: modules, interfaces and systems (Söderköping, 1983), pages 1–17. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984.

Reid:1984:TFP

[Rei84b] J. K. Reid. TREESOLVE, a Fortran package for solving large sets of linear finite element solutions. Technical Report CSS 155, 1984, AERE Harwell, Computer Science and Systems Division, Harwell, Berkshire, UK, 1984.

Reid:1987:EPU

[Rei87a] J. Reid. The exploitation of parallelism by using Fortran 8X features. Supercomputer, 19:8–18, 1987. CODEN SPCOEL. ISSN 0168-7875.

Reid:1987:JRRa

[Rei87b] John Reid. John Reid reports: X3J3 meeting, August 1986. ACM SIGPLAN FORTRAN Forum, 6(1):4–7, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRc

[Rei87c] John Reid. John Reid reports:

X3J3 meeting, February 1987. ACM SIGPLAN FORTRAN Forum, 6(1):12–14, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRb

[Rei87d] John Reid. John Reid reports:

X3J3 meeting, November 1986. ACM SIGPLAN FORTRAN Forum, 6(1):8–11, April 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:XMN

[Rei87e] John Reid. X3J3 meeting, November 1987. ACM SIGPLAN FOR-TRAN Forum, 6(3):4–6, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:ERC

[Rei89a] Edward Reid. Edward Reid: some comments on Fortran 8X. ACM SIGPLAN FORTRAN Forum, 8 (1):33–36, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:ERS

[Rei89b] Edward Reid. Edward Reid: some comments on Fortran 8X. ACM SIGPLAN FORTRAN Forum, 8 (1):33–36, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:JRR

[Rei89c] John Reid. John Reid reports: X3J3 Meeting 108, 9–13 May 1988, Urbana IL; X3J3 Meeting 109, 8–12 Aug 1988, Jackson WY; ISO/WG5 Meeting, 19– 23 Sep 1988, Paris; X3J3 Meeting 110, 14–18 Nov 1988, Boston MA. ACM SIGPLAN FORTRAN Forum, 8(1):6–17, January 1989. ISSN 1061-7264 (print), 1931-1311 (electronic).

RTI:1983:EFU

[Rel83] Relational Technology, Inc. EQUEL/ FORTRAN user’s guide. Relational Technology Inc., Berkeley, CA, USA, 1983. various pp.

RTI:1986:IEF

[Rel86] Relational Technology, Inc. IN-GRES/EQUEL/FORTRAN user’s guide. Relational Technology, Alameda, CA, USA, 1986. various pp.

RTI:1988:IEC

[Rel88] Relational Technology, Inc. IN-GRES/EQUEL/C user’s guide; INGRES/ embedded SQL companion guide for C; INGRES/ embedded SQL companion guide for FORTRAN. Relational Technology, Alameda, CA, USA, revised July, 1988 edition, 1988. 3 v. in 1 pp.

RTI:1989:IESa

[Rel89a] Relational Technology, Inc. IN-GRES/ embedded SQL companion guide for COBOL; INGRES/ embedded SQL companion guide for BASIC; INGRES/ embedded SQL companion guide for ADA; IN-GRES/ embedded SQL companion guide for FORTRAN; INGRES/ embedded SQL companion guide for PL/I. Relational Technology Inc., Alameda, CA, USA, 1989. 5 v. in 1 pp.

RTI:1989:IESb

[Rel89b] Relational Technology, Inc. IN-GRES/ embedded SQL user’s guide and reference manual; IN-GRES/ embedded SQL companion guide for C; INGRES/ EQUEL companion guide for FORTRAN. Relational Technology Inc., Alameda, CA, USA, 1989. 3 v. in 1 pp.

RTI:1989:IECb

[Rel89c] Relational Technology, Inc. IN-GRES/EQUEL companion guide for FORTRAN; INGRES/EQUEL companion guide for BASIC; IN-GRES/EQUEL companion guide for ADA; INGRES/EQUEL companion guide for PL/I; INGRES/ EQUEL companion guide for C; INGRES/EQUEL companion guide for COBOL; INGRES/ EQUEL companion guide for PAS-CAL. Relational Technology Inc., Alameda, CA, USA, 1989. 7 v. in 1 pp.

RTI:1989:ISR

[Rel89d] Relational Technology, Inc. IN-GRES/SQL reference manual; IN-GRES/ embedded SQL companion guide for FORTRAN; INGRES/ QUEL reference manual; IN-GRES/EQUEL companion guide for C. Relational Technology Inc., Alameda, CA, USA, 1989. 4 v. in 1 pp.

Renka:1984:TIA

[Ren84] R. J. Renka. Triangulation and interpolation at arbitrarily distributed points in the plane. ACM Transactions on Mathematical Software, 10(4):440–442, 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

REA:1986:HGC

[Res86] Research and Education Association. Handbook and guide for comparing and selecting the computer languages: BASIC, FORTRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, revised print. edition, 1986. ISBN ???? v + 122 pp. LCCN ????

Reynolds:1980:ECS

[Rey80] John H. Reynolds. Evaluation of contemporary software engineering techniques for a large Fortran simulation. The Journal of Systems and Software, 1(2):131–140, February 1980. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Roberts:1985:DCI

[RG85] M. L. Roberts and P. D. Griffiths.

Design considerations for IBM Personal Computer Professional FOR-TRAN, an optimizing compiler. IBM Systems Journal, 24(1):49– 60, 1985. CODEN IBMSA7. ISSN 0018-8670.

Rice:1984:ARK

[RH84a] John R. Rice and Richard J.

Hanson. Algorithm 620: References and keywords for Collected Algorithms of the ACM. ACM Transactions on Mathematical Software, 10(4):359–360, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Ham85, HM90].

Rossbach:1984:IWI

[RH84b] E. O. Rossbach and S. I. Hyman.

Interchangeability, what it means: a useful program in FORTRAN and its development, presented at A.G.A. Distribution Conference, Denver, Colorado, May 22, 1978. Technical Report ????, American Gas Association, Arlington, VA, USA, 1984. iii + 132 pp.

Richards:1982:CEM

[Ric82a] Thomas C. Richards. Cost effective methods for teaching introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 14(1):128–132, February 1982. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 13th SIGCSE Symposium on Computer Science Education.

Richards:1982:FPR

[Ric82b] W. F. Richards. A Fortran program for rectangular microstrip antennas. Technical Report RADC-TR-82-78, Univ. of Illinois, Urbana, IL, USA, 1982. 84 pp.

Rice:1984:FEP

[Ric84] John R. Rice. Fortran extensions for parallel and vector computation. Technical report TR-470, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1984. 8 pp. URL http:/

                                        /www.cs.purdue.edu/research/
                                        technical_reports/1984/TR%2084-
                                        470.pdf.

Rice:1986:LIP

[Ric86] J. R. Rice. Language independent PROTRAN. Technical Report SERC-TR-0-P, Software Engineering Research Centre, Utrecht, The Netherlands (??), October 6, 1986.

Ridler:1982:APG

[Rid82a] Philip Ridler. Addison-Wesley pocket guide to FORTRAN. Programming pocket guides Micro books. Addison-Wesley, Reading, MA, USA, 1982. ISBN 0-201-07746-9. iii + 57 pp. LCCN ???? US$6.95. Cover title. Spine title: Pocket guide to FORTRAN. Pages 29-57 printed on versos of pages 1-28.

Ridler:1982:PGFa

[Rid82b] Philip Ridler. Pocket guide to FORTRAN. Pitman programming pocket guides. Pitman Publishing Ltd., London, UK, 1982. ISBN 0-273-01683-0. iii + 57 pp.

Ridler:1982:PGFb

[Rid82c] Philip Ridler. Pocket Guide to Fortran. Addison-Wesley, Reading, MA, USA, May 1982. ISBN 0-201-07746-9. ???? pp. LCCN ???? US$6.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201077469.

Riley:1983:ACI

[Ril83] Robert Riley. Applications of a computer implementation of Poincaré’s theorem on fundamental polyhedra. Mathematics of Computation, 40(162):607–632, April 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ring:1983:FCP

[Rin83] D. R. Ring. A FORTRAN computer program for determining start date and base temperature for degree day models. Technical report, Texas Agricultural Experiment Station, the Texas A and M University System, College Station, TX, USA, 1983. 10 pp.

Ritz:1989:BRH

[Rit89] C. L. Ritz. Book review: Handbook for Matrix Computations, by Thomas F. Coleman and Charles Van Loan. Fortran Journal, 1(??): ??, May/June 1989. ISSN 1060-0221.

Rizzi:1985:VCF

[Riz85] Arthur Rizzi. Vector coding the finite-volume procedure for the CYBER 205. Parallel Computing, 2(4):295–312, December 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rule:1984:FSY

[RkC84] Wilfred P. Rule and Yun kung Chung. Fortran 77: Shih yung cheng shih she chi fa. Chu pi shu chu, Tai-pei shih, chu pan edition, 1984. ISBN ???? 7 + 573 pp. LCCN ????

Rogers:1981:CQD

[RL81] James L. Rogers and J. A. N. Lee. Comments, queries, and debate: The Sumador Chino; history of FORTRAN. Annals of the History of Computing, 3(4):408–410, October/December 1981. CODEN AHCOE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1981/pdf/a4408.pdf.

Riedel:1985:SPI

[RMFG85] Neal K. Riedel, David A. McAninch, Cameron Fisher, and Nahum B. Goldstein. A signal processing implementation for an IBM-PC-based workstation. IEEE Micro, 5(5):52–67, September/ October 1985. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic).

Ross:1982:UMS

[RMS82] D. Lauren Ross and Hubert J.

Morel-Seytoux. User’s manual for SOILMOP: a FORTRAN IV program for prediction of infiltration and water content profiles under variable rainfall conditions. Hydrowar interim reports, Hydrowar Program, Colorado State University, Fort Collins, CO, USA, 1982. 129 pp.

Reasenberg:1985:FFF

[RO85] Paul Reasenberg and D. Oppenheimer. FPFIT, FPPLOT, and FPPAGE FORTRAM computer programs for calculating and displaying earthquake fault-plane solutions. Open-file report 85-739, U.S. Dept. of the Interior, Geological Survey: Books and Open-File Reports Section, distributor, Denver, Colo.?, 1985. 2 microfiches pp.

Reasenberg:1986:FFF

[RO86] Paul Reasenberg and D. Oppenheimer. FPFIT, FPPLOT and FPPAGE: Fortran computer programs for calculating and displaying earthquake fault-plane solutions. Open-file report 85-739, U.S. Geological Survey, Menlo Park, CA, USA, 1986. 109 pp.

Roberts:1982:MBP

[Rob82] Tim N. Roberts. More on bit processing with FORTRAN. ACM SIGPLAN Notices, 17(2):49–52, February 1982. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1983:RFS

[Rob83] K. V. Roberts. The regeneration of Fortran software. Computer Physics Communications, 29(1):7–13, March 1983. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465583900243.

Roche:1986:CPF

[Roc86] Timothy Edward Roche. The conversion of the process flowsheet simulator, FLOWTRAN, for FOR-TRAN 66 based mainframe computers. Thesis (M.S.), New Jersey Institute of Technology, Newark, NJ, USA, 1986. vii + [111] pp.

Rodich:1984:SWAb

[Rod84] Grover Rodich. Student workbook to accompany First course in data processing with BASIC, COBOL, FORTRAN, and RPG, 3d ed., and First course in data processing with BASIC, 2nd ed. [by] Couger and McFadden. John Wiley and Sons, New York, London, Sydney, 1984. ISBN ???? v + 218 pp. LCCN ????

Rodman:1986:FIP

[Rod86] James P. Rodman. FORTRAN IV programming for microcomputers. J. P. Rodman, Alliance, OH, USA, 1986. ISBN ???? vi + 66 + [45] + [1] pp. LCCN QA76.73.F25 R59

MelendezRodriguez:1987:DCP

[Rod87] Marcos Melendez Rodriguez. Development of a computer program in FORTRAN IV using the Newton–Raphson method for fast load flow studies. Report (M.E.), University of Puerto Rico, Mayaguez Campus, Mayaguez, Puerto Rico, 1987. ix + 161 pp.

Romaya:1981:FSI

[Rom81] A. M. Romaya. FORTRAN subset interpreter for a microprocessor gas control system. Computer Physics Communications, 22 (2–3):307–309, April 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465581900655.

Rosenblatt:1984:SFW

[Ros84] Bruce Rosenblatt. The successors to FORTRAN—why does FOR-TRAN survive? Annals of the History of Computing, 6(1):39–40, January/March 1984. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1984/pdf/a1028.pdf;
 http://www.computer.org/annals/
an1984/a1028abs.htm.

Rosenblum:1987:UFS

[Ros87] Paul Rosenblum. Using the Fortran 8x style of programming. Technical report TMC-148, PL87-7, Thinking Machines Corporation, Cambridge, MA, USA, 1987. 12 pp. Presented at the Usenix C++ Conference, Sante Fe December 1987.

Rouse:1983:PIPb

[Rou83] Robert A. Rouse. Programming the IBM Personal Computer: Fortran 77. Henry Holt & Co, ????, January 1983. ISBN 0-03-063668-X. ???? pp. LCCN ???? US$17.50. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        003063668X.

Rouse:1986:IF

[Rou86] Robert A. Rouse. Introduction to Fortran 77. Holt, Reinhart, and Winston, New York, NY, USA, June 1986. ISBN 0-03-063634-5. ???? pp. LCCN ???? US$26.54. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0030636345.

Roy:1988:PPF

[Roy88] Joydeep Roy. A parallel preprocessor for the FORTRAN language. Thesis (M.S.), University of Alabama at Birmingham, Birmingham, AL, USA, 1988. v + 44 pp.

Rao:1989:NFP

[RRC89] K. Srinivasa Rao, V. Rajeswari, and Charles B. Chiu. A new Fortran program for the 9-j angular momentum coefficient. Computer Physics Communications, 56(2):231–248, December 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0010465589900210.

Radicati:1988:DLS

[RRS88] G. Radicati, Y. Robert, and P. Sguazzero. Dense linear systems FORTRAN solvers on the IBM 3090 vector multiprocessor. Parallel Computing, 8(1–3):377–384, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Routti:1981:FPS

[RS81] J. T. Routti and J. V. Sandberg.

Fortran program spall for computing spallation reaction cross sections. Computer Physics Communications, 23(4):411–426, August 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465581901818.

Roggenkamrp:1982:RMS

[RS82] B. Roggenkamrp and A. Schewe.

Rechnerprogramm zur Messung der Spezifischen Waerme von Metallen Bei Tiefen Temperaturen mit Hilfe der Pulsmethode. Studienarbeit, 1982.

Routti:1984:FPS

[RS84] J. T. Routti and J. V. Sandberg. Fortran program spall for computing spallation reaction cross sections. Computer Physics Communications, 35(1–3):C–739, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584828738.

Rhoades:1985:EME

[RS85] Clifford E. Rhoades, Jr. and K. G. Stevens, Jr. Early MIMD experience on the CRAY X-MP. Computer Physics Communications, 37(1–3):215–221, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465585901559.

Ronsch:1987:TRS

[RS87] W. Ronsch and H. Strauss. Timing results of some internal sorting algorithms on vector computers. Parallel Computing, 4(1):49– 61, February 1987. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ranai:1989:CAB

[RS89] K. Ranai and B. Srinivasan. Complexity analysis of benchmark software for supercomputers. The Journal of Supercomputing, 3(3): 235–242, September 1989. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.springerlink.
                                        com/openurl.asp?genre=article&
                                        issn=0920-8542&volume=3&issue=
                                        3&spage=235.

Rizzi:1985:SMB

[RT85] N. Rizzi and A. Tatone. Symbolic manipulation in buckling and postbuckling analysis. Computers and structures, 21:691–700, 1985. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Rubenstein:1983:FCP

[Rub83] Nancy Ordelheide Rubenstein. A FORTRAN computer program for transcribing Franconian rhythm. Thesis (Ph.D.), Washington University, St. Louis, MO, USA, 1983. v + 134 pp.

Rudolph:1983:ODA

[Rud83] Bruce L. Rudolph. An overview of the design of an Ada ballistics system. ACM SIGADA Ada Letters, 2 (5):60–61, March/April 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rule:1983:FPAa

[Rul83a] Wilfred P. Rule. Fortran 77: a Practical Approach. Van Nostrand Reinhold, New York, NY, USA, fourth edition, July 1983. ISBN 0-442-27810-1. x + 486 pp. LCCN QA76.73.F25 R84 1983. US$46.95; US$30.45. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0442278101. Rev. ed. of: FOR-TRAN, a practical approach with style and structure. 3rd ed. c1980. Includes index.

Rule:1983:FPAb

[Rul83b] Wilfred P. Rule. Fortran 77, a Practical Approach. PWS-Kent Pub. Co., Boston, MA, USA, fourth edition, June 1983. ISBN 0-87150-390-5 (paperback). x + 486 pp. LCCN QA 76.73 F25 R835 1983. US$27.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0871503905.

Rust:1987:FFP

[Rus87] Roland T. Rust. FLEX: a FOR-TRAN program for flexible regression. Working paper 87/88-5-2, Dept. of Marketing Administration, College of Business Administration and Graduate School of Business, University of Texas at Austin, Austin, TX, USA, 1987. 25 pp.

Rao:1984:NFP

[RV84] K. Srinivasa Rao and K. Venkatesh. New Fortran programs for angular momentum coefficients. Computer Physics Communications, 35 (1–3):C–498–C–499, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/S0010465584826855.

RadicatidiBrozolo:1989:CSI

[RV89] Giuseppe Radicati di Brozolo and Marcello Vitaletti. Conjugategradient subroutines for the IBM 3090 Vector Facility. IBM Journal of Research and Development, 33 (2):125–135, March 1989. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Robinson:1988:ERF

[RV8 ] K. Robinson and P. Van Santen. EMR report: Feasibility study into a parallel extracting Fortran compiler. Technical Report P3, Software Engineering Research Centre, Utrecht, The Netherlands (??), 1988 (??).

Reid:1985:AFF

[RW85] J. K. Reid and A. Wilson. The array features in FORTRAN 8X with examples of their use. Computer Physics Communications, 37 (1–3):125–132, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Robinet:1986:ESP

[RW86] B. (Bernard) Robinet and R. (Reinhard) Wilhelm, editors. European Symposium on Programming, Saarbrucken, Federal Republic of Germany, March 17–19, 1986: Proceedings, volume 213 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-16442-1. LCCN QA76.6.E976 1986.

Ruppelt:1989:ATH

[RW89] Th. Ruppelt and G. Wirtz. Automatic transformation of highlevel object-oriented specifications into parallel programs. Parallel Computing, 10(1):15–28, March 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rodich:1984:SWAa

[RWA84] Grover Rodich, Collin J. Watson, and John M. Anderson. Student workbook to accompany First course in data processing with BASIC, COBOL, FORTRAN, and RPG, 3d ed., and First course in data processing with BASIC, 2nd ed. [by] Couger and McFadden. John Wiley and Sons, New York, London, Sydney, 1984. ISBN 0-471-86951-1. v + 218 pp.

Ryan-McFarland:1986:RF

[Rya86] Ryan-McFarland Corporation. RM/Fortran, 1986.

Reddy:1989:FAS

[RZ89a] Rama N. Reddy and Carol A. Ziegler. Fortran 77 with applications for scientists and engineers. West Publishing Company, St. Paul, MN, USA, 1989. ISBN 0-314-48135-4. xix + 565 pp. LCCN QA76.73.F25 R421 1989.

Reddy:1989:IMA

[RZ89b] Rama N. Reddy and Carol A. Ziegler. Instructor’s manual to accompany Fortran 77 with applications for scientists and engineers. West Publishing Company, St. Paul, MN, USA, 1989. ISBN 0-314-52503-3. [6] + 77 pp. LCCN QA 76.73 F25R42 1989.

Rzytka:1984:PMC

[Rzy84] Jan Rzytka. Programowanie maszyn cyfrowych w problematyce inzynierskiej: podstawy programowania w jezyku FORTRAN. Prace Glownego Instytutu Gornictwa. Seria dodatkowa. Glowny Instytut Gornictwa, Katowice, Czechoslovakia, 1984. ISBN ???? vii + 226 pp. LCCN ????

Sobek:1988:ALI

[SAB88] S. Sobek, M. Azam, and J. C. Browne. Architectural and language independent parallel programming: a feasibility demonstration. In Proceedings of the 1988 International Conference on Parallel Processing, volume II, Software, pages 80–83. Penn State, University Park, PA, USA, August 1988. U. Texas.

Salem:1984:CFF

[Sal84] M. Salem. Converter of Fortran format and data statements to standard form. Computer Physics Communications, 35(1–3):C–371, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584825862.

Sammet:1981:HIT

[Sam81] Jean E. Sammet. History of IBM’s technical contributions to high level programming languages. IBM Journal of Research and Development, 25(5):520–534, September 1981. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Shaw:1981:CPL

[SAN^{+}81] Mary Shaw, Guy T. Almes, Joseph M. Newcomer, Brian K. Reid, and William W. Wulf. A comparison of programming languages for software engineering. Software—Practice and Experience, 11(1):1–52, January 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sanders:1982:VUA

[San82] U. Sanders. Vergleichende Untersuchung von Algorithmen zur Erkennung Spezifischer Muster im Elektroencephalogramm. Master’s thesis, Technische Universität Braunschweig (??), Braunschweig, Germany, 1982.

Sassa:1981:PMR

[Sas81] Masataka Sassa. A poor man’s realization of recursive structured FORTRAN. ACM SIGPLAN Notices, 16(5):43–53, May 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sass:1983:SAFa

[Sas83a] C. Joseph Sass. A structural approach to FORTRAN 77 programming: with WATFIV. Allyn and Bacon, Needham Heights, MA, USA, February 1983. ISBN 0-205-07918-0 (paperback). x + 427 pp. LCCN QA76.73.F25S28 1983. US$34.80. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0205079180.

Sass:1983:SAFb

[Sas83b] C. Joseph Sass. A structured approach to Fortran 77 programming with WATFIV. Allyn and Bacon, Needham Heights, MA, USA, 1983. ISBN 0-205-07918-0 (paperback). x + 427 pp.

SAS:1986:SDL

[SAS86] SAS Institute. System 2000 DBMS language specification manual for the Fortran Programming Language Extension (PLEX). SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN ???? various pp. LCCN ????

Sauer:1983:FPC

[Sau83a] F. W. Sauer. A Fortran program for the calculation of an extremal polynomial. ACM Transactions on Mathematical Software, 9 (3):381–383, September 1983. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sauer:1983:AAF

[Sau83b] Frederick W. Sauer. Algorithm 604: A FORTRAN program for the calculation of an extremal polynomial. ACM Transactions on Mathematical Software, 9(3): 381–383, September 1983. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sauer:1983:AFP

[Sau83c] Frederick W. Sauer. Algorithm 604: A FORTRAN program for the calculation of an extremal polynomial. ACM Transactions on Mathematical Software, 9(3): 381–383, September 1983. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Savage:1987:SGCb

[Sav87] William Z. Savage. In situ geomechanics of crystalline and sedimentary rocks: part VII, SLIP — a FORTRAN computer program for computing the potential for sliding on arbitrarily oriented weakness planes in triaxial stress states. Open-file report 87-82, U.S. Geological Survey, Denver, CO, USA, 1987. iv + 17 pp.

Schwar:1982:AFE

[SB82] James P. Schwar and Charles L. Best. Applied FORTRAN for engineering and science. Science Research Associates, Chicago, IL, USA, 1982. ISBN 0-574-21365-1. x + 226 pp. LCCN QA76.73.F25 S34 Sci-Eng. US$16.95.

Schwartz:1983:SPF

[SB83] Ronald D. Schwartz and David T. Basso. Statistical Programs in Fortran. Reston Publishing Co., Inc., Reston, VA, USA, March 1983. ISBN 0-8359-7096-5. xiv + 254 pp. LCCN QA276.4.S37 1983. US$21.95; US$18.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0835970965.

Schwar:1986:AFE

[SB86] James P. Schwar and Charles L. Best. Applied FORTRAN for engineering and science. Science Research Associates, Chicago, IL, USA, second edition, 1986. ISBN 0-574-21795-9 (paperback). x + 308 pp. LCCN QA76.73.F25 S34 1986 Sci-Eng. US$19.95.

Singh:1983:FSB

[SC83] Teja Singh and D. (David) Campbell. Fortran subroutines for biomass computation. Forest management note 22 0714-1181, Northern Forest Research Centre, Edmonton, Alberta, Canada, 1983. ISBN 0-662-12762-5. 7 pp.

Schrage:1979:MPF

[Sch79] Linus Schrage. A more portable Fortran random number generator.

ACM Transactions on Mathematical Software, 5(2):132–138, June 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Schryer:1981:TCF

[Sch81a] Norman L. Schryer. A test of a computer’s floating-point arithmetic unit. Computing Science Technical Report 89, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 4, 1981. ii + 64 pp.

Schwar:1981:AFE

[Sch81b] James P. Schwar. Applied FOR-TRAN for engineering and science. Science Research Associates, Chicago, IL, USA, 1981. ISBN 0-574-21365-1. ???? pp. LCCN ????

Scheer:1982:AFA

[Sch82a] Linda Sue Scheer. Ada, FOR-TRAN, ALGOL, JOVIAL, Pascal, PL/I, and LISP compared to Ada design requirements. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1982. x + 121 pp.

Schiller:1982:PSC

[Sch82b] Steven F. Schiller. A prototype system for certifying secure information flows in FORTRAN programs. Technical Report GWU-IIST-82-20, Institute for Information Science and Technology, Department of Electrical Engineering and Computer Science, School of Engineering and Applied Science, George Washington University, Washington, DC, USA, 1982. 14 + [13] pp.

Schmidt:1982:BKS

[Sch82c] B. Schmidt. Die Bestimmung der Konfidenzintervalle in der Simulation. Elektronische Rechenanlagen, 24, 3:118–124, 1982. CODEN EL-RAA4. ISSN 0013-5720.

Schmidt:1982:SDS

[Sch82d] B. Schmidt. Simulation of discrete systems using GPSS-Fortran. The Computer Journal, 25(1):84– 86, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Schmidt:1982:SM

[Sch82e] B. Schmidt. Systemanalyse und Modellbildung. Proc. 1. Symposium Simulationstechnik Erlangen, Informatik-Fachberichte, Springer-verlag, 56:106–115, 1982.

Schofield:1983:HBSb

[Sch83a] Charles G. Schofield. Homogenization/Blending Systems Design and Control for Minerals Processing. Trans Tech Publications, Ltd. and Gulf Publishing Company, Clausthal, Germany; Rockport, MA, USA and Houston, TX, USA, December 1983. ISBN 0-87201-360-X. xiii + 321 pp. LCCN TN500.S42 1983. US$60.00.

Schofield:1983:HBSa

[Sch83b] Charles G. Schofield. Homogenization/Blending Systems Design and Control for Minerals Processing: with FORTRAN programs, volume 2 of Series on bulk materials handling. Trans Tech Publications, Ltd., Clausthal, Germany; Rockport, MA, USA, 1983. ISBN 0-87849-030-2. xiii + 321 pp. LCCN TN500 .S42 1983.

Schmid:1984:PCB

[Sch84a] Hans Joachim Schmid. A DAP-FORTRAN code for the Hungarian method. In Parallel computing 83 (Berlin, 1983), pages 367–372. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984.

Schmidt:1984:SGV

[Sch84b] B. Schmidt. Der Simulator GPSS-Fortran Version 3, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13782-3. 336 pp.

Schmidt:1984:ESG

[Sch84c] B. Schmidt. Einführung in die Systemanalyse GPSS-FORTRAN Version 3, volume 1. Fachberichte Simulation, Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN ???? ???? pp.

Schmidt:1984:MGV

[Sch84d] B. Schmidt. Modellbildung mit GPSS-Fortran Version 3, volume 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13783-1. 307 pp.

Schmidt:1984:MMG

[Sch84e] B. Schmidt. Modellbildung mit GPSS-Fortran Version 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13783-1. ???? pp.

Schulz:1984:IDT

[Sch84f] B. Schulz. Implementierung, Dokumentation und Test eines Programms zur Berechnung von Antennenfeldern nach der Momentenmethode. Studienarbeit, 1984.

Schmidt:1985:SM

[Sch85a] B. Schmidt. Systemanalyse und Modellaufbau. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13784-X. ???? pp.

Schwar:1985:AFE

[Sch85b] James P. Schwar. Applied Fortran for Engineering and Science. Macmillan College Division, ????, second edition, November 1985. ISBN 0-02-408161-2. ???? pp. LCCN ???? US$47.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0024081612.

Schittkowski:1986:NFS

[Sch86] K. Schittkowski. NLPQL: a FOR-TRAN subroutine solving constrained nonlinear programming problems. Annals of Operations Research, 5:485–500, 1986. CO-DEN AOREEV. ISSN 0254-5330 (print), 1572-9338 (electronic).

Schmidt:1987:MCG

[Sch87a] Bernd Schmidt. Model Construction With GPSS-FORTRAN: Version 3. Advances in Simulation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., April 1987. ISBN 0-387-96503-3. ix + 293 pp. LCCN QA76.9.C65 S3617 1987. US$29.00 (est.). URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0387965033. Companion volume to: The simulator GPSS-FORTRAN version 3. Includes index. Bibliography: p. 191.

Schmidt:1987:SGV

[Sch87b] Bernd Schmidt. The simulator GPSS-FORTRAN version 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96504-1. ix + 336 pp. LCCN QA76.9.C65 S3618 1987. US$29.00 (est.). Companion volume to: Model construction with GPSS-FORTRAN version 3. Includes index. Bibliography: p. [323].

Schuster:1987:RPF

[Sch87c] Hildegard Schuster. Richtlinien zur Programmierung in FOR-TRAN 77. Mitteilung (Deutsche Forschungs- und Versuchsanstalt fur Luft- und Raumfahrt) 87-10, Deutsche Forschungs- und Versuchsanstalt fur Luft- und Raumfahrt, Köln, 1987. 41 pp.

Schrauf:1988:AGA

[Sch88a] Géza Schrauf. Algorithm 664:

A Gauss algorithm to solve systems with large banded matrices using random-access disk storage. ACM Transactions on Mathematical Software, 14(3):257– 260, September 1988. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/citations/journals/
                                        toms/1988-14-3/p257-schrauf/.

Schriber:1988:PSU

[Sch88b] T. J. Schriber. Perspectives on simulation using GPSS. Simulation news Europe, SCS, 3, 2:23–36, 1988. ISSN 0929-2268.

Schuster:1988:DPF

[Sch88c] Hildegard Schuster. Directions for programming in FORTRAN 77. NASA technical translation NASA TT-20317, National Aeronautics and Space Administration, Washington, DC, USA, 1988. ???? pp.

Schoen:1989:SSM

[Sch89a] Martin Schoen. Structure of a simple molecular dynamics FOR-TRAN program optimized for Cray vector processing computers. Computer Physics Communications, 52(2):175–185, January/February 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465589900027.

Schofield:1989:OFPb

[Sch89b] C. F. Schofield. Optimizing Fortran Programs. Ellis Horwood Series in Computers and Their Applications. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, July 1989. ISBN 0-13-638693-8. ???? pp.

LCCN ???? US$65.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0136386938.

Schofield:1989:OFPa

[Sch89c] Clive F. (Clive F.) Schofield. Optimising Fortran programs. Ellis Horwood series in computers and their applications. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, July 1989. ISBN 0-7458-0654-6, 0-470-21533-X (Halsted Press). x + 309 pp. LCCN QA76.73.F25 S334 1989 WITHDRAWN. US$40.95, US$74.95. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=047021533X.

Cover title: Optimizing Fortran programs. Includes index. Bibliography: p. [293]-296.

Schonfelder:1989:SEP

[Sch89d] J. L. Schonfelder. Semantic extension possibilities in the proposed new Fortran. Software— Practice and Experience, 19(6): 529–551, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schrader:1989:AFT

[Sch89e] Denis Schrader. Automated FORTRAN-C translation saves programmer effort. C Users Journal, 7(8):55–??, August 1989. ISSN 0898-9788.

Schulman:1989:DKC

[Sch89f] A. Schulman. A different kind of CASE tool. BYTE Magazine, 14 (5):217–220, May 1989. CODEN BYTEDJ. ISSN 0360-5280.

Simpson:1989:BFP

[SD89] D. G. Simpson and G. E. Dallal. BUMP: a FORTRAN program for identifying dose-response curves subject to downturns. Computers in Biomedical Research, 22 (1):36–43, February 1989. CODEN CPBRAF. ISSN 0010-4809 (print), 1090-2368 (electronic).

Strasser:1982:GIP

[SDC82] Andre Strasser, Eric Davaud, and Jean Charollais. GEOMAN: an interactive program for the management of geological thin section data, UNIVAC FORTRAN ASCII version. Publications du Département de geologie et de paleontologie de l’Université de Genève 1, l’Université de Genève, Genève, Switzerland, 1982. 78 pp.

Sheldon:1984:SFI

[SDH84] E. Sheldon, D. R. Donati, and H. R. Hiddleston. ’MIA’, a Fortran-IV program for making spin and parity assignments to high-lying single and coherent twin nuclear levels from (alpha, nucleon) angular distributions in on-resonance, compoundnuclear, channel-spin-1/2 reactions. Computer Physics Communications, 35(1–3):C–274–C–275, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0010465584825060.

Seeds:1981:SFB

[See81a] Harice L. Seeds. Structuring FOR-TRAN 77 for business and general applications. John Wiley and Sons, New York, London, Sydney, 1981. ISBN 0-471-07836-0 (paperback). x + 512 pp. LCCN QA76.73.F25 S39.

Sees:1981:SFB

[See81b] Harice L. Sees. Structuring Fortran 77 for Business and General Applications. John Wiley and Sons, New York, London, Sydney, July 1981. ISBN 0-471-07836-0. ???? pp. LCCN ???? US$41.92. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471078360.

SITFNTH:1983:HCP

[Sel83] Selskapet for industriell og teknisk forskning ved Norges tekniske høgskole. Handbook of computer programming: design, development and maintenance of engineering (Fortran) software. SIN-TEF, Trondheim, Norway, 1983. 344 pp.

Sequent:1989:AF

[Seq89] Sequent Computer Systems, Inc.

ATS FORTRAN. Sequent Computer Systems, Inc., Beaverton, OR, USA, 1989. various pp.

Service:1985:PFS

[Ser85] Peerless Engineering Service. Professional Fortran Scientific Subroutine Library. John Wiley and Sons, New York, London, Sydney, August 1985. ISBN 0-471-82954-4. ???? pp. LCCN ???? US$225.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471829544.

Service:1989:FSS

[Ser89] Peerless Engineering Service. Fortran Scientific Subroutine Library: IBM PC. John Wiley and Sons, New York, London, Sydney, August 1989. ISBN 0-471-51499-3. ???? pp. LCCN ???? US$610.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=

0471514993. Includes two 128K disks.

Solomon:1981:WPS

[SFKS81] Robert Solomon, Frank L. Friedman, Elliot B. Koffman, and Judith O’Shea Stebulis. Workbook for Problem solving and structured programming in FORTRAN. Addison-Wesley, Reading, MA, USA, second edition, 1981. ISBN 0-201-02465-9. 212 pp. LCCN QA76.73.F25 S6 1981.

Starr:1988:SBD

[SG88] G. Lynn Starr and J. Michael Geist. Soil bulk density and soil moisture calculated with a FOR-TRAN 77 program. General technical report PNW GTR-211, U.S. Dept. of Agriculture, Forest Service, Pacific Northwest Research Station, Portland, OR, USA, 1988. 6 + [2] pp.

Sackfield:1988:CMR

[SH88] A. Sackfield and D. A. Hills. Results in elasticity theory using complex FORTRAN. In Computers in mathematical research (Cardiff, 1986), volume 14 of Inst. Math. Appl. Conf. Ser. New Ser., pages 177–185. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988.

Shampine:1982:IRM

[Sha82] L. F. Shampine. Implementation of Rosenbrock methods. ACM Transactions on Mathematical Software, 8(2):93–113, June 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Shablygin:1987:IEH

[Sha87] E. Shablygin. Integral equation with hidden Eigenparameter Solver: REDUCE and FORTRAN in tandem. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 186– 191. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Shablygin:1989:EL

[Sha89] Eugene Shablygin. Integral equation with hidden eigenparameter solver: REDUCE + FORTRAN in tandem. In Davenport [Dav89], pages 186–191. CODEN LNCSD9. ISBN 0-387-51517-8 (New York), 3-540-51517-8 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E861 1987. US$39.40.

Sherman:1978:ANF

[She78] Andrew H. Sherman. Algorithm 533: NSPIV, A Fortran subroutine for sparse Gaussian elimination with partial pivoting [F4]. ACM Transactions on Mathematical Software, 4(4):391–398, December 1978. CODEN ACMSCU.

ISSN 0098-3500 (print), 1557-7295 (electronic).

Shen:1982:FF

[She82] Chen-Li Shen. Fortran 66 formalized. Project (M.S., Computer Science), California State University, Sacramento, CA, USA, 1982. 150 pp.

Shelley:1984:EF

[She84] John Shelley. Essentials of Fortran 77. John Wiley and Sons, New York, London, Sydney, October 1984. ISBN 0-471-90502-X (paperback). xiv + 204 pp. LCCN QA 76.73 F25 S447 1984. US$24.95; US$19.08. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
047190502X.

Shelley:1989:EF

[She89a] John Shelley. Essentials of Fortran 77. John Wiley and Sons, New York, London, Sydney, second edition, 1989. ISBN 0-471-92378-8. ix + 182 pp. LCCN QA76.73.F25 S447 1989. US$21.90; US$48.95; US$60.00. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0471923788.

Sherrod:1989:PDC

[She89b] Michael J. Sherrod. Part I. Docking calculations on the complexation of ferrocenes with cyclodextrins. Part II. SHADOW, a FORTRAN program for the calculation of molecular area: application monolayer packing. Part III. The synthesis and MM2 conformational analysis of a series of rigid and semi-rigid hydroxy olefins. Thesis (Ph.D.), Emory University, Atlanta, GA, USA, 1989. 228 pp. Director of this thesis, Dr. F. M. Menger, Dept. of Chemistry.

Shivers:1988:CFA

[Shi88] O. Shivers. Control flow analysis in Scheme. ACM SIGPLAN Notices, 23(7):164–174, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org:80/pubs/
citations/proceedings/pldi/
53990/p164-shivers/.

Shouse:1985:FCB

[Sho85] D. V. Shouse. ‘on the fly’ CRC-16 byte-wise calculation for 8088based computers. IEEE Micro, 5 (2):67–75, March/April 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

SIGPLAN:1984:FF

[SIG84] Fortran forum, page various, 1984.

ISSN 1061-7264 (print), 1931-1311 (electronic). ACM Press, New York, NY 10036, USA.

Simpson:1976:AFT

[Sim76] Joseph C. Simpson. Algorithm 501: Fortran translation of Algorithm 409, discrete Chebychev curve fit [E2]. ACM Transactions on Mathematical Software, 2(1):95–97, March 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Fut78].

Sims:1985:SCP

[Sim85] S. Eugene Sims. Selected computer programs in FORTRAN for fish stock assessment. FAO fisheries technical paper ; 259. Food and Agriculture Organization of the United Nations, Rome, Italy, 1985. ISBN 92-5-102272-0. v + 183 pp. LCCN SH331.5.C65 S441 1985.

Sims:1986:SCP

[Sim86] S. Eugene Sims, editor. Selected Computer Programs in Fortran for Fish Stock Assessment/F2812, volume 259 of Fisheries Technical Paper. Food & Agriculture Org, ????, April 1986. ISBN 92-5-102272-0. ???? pp. LCCN ???? US$6.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
9251022720.

Simeone:1988:FCN

[Sim88a] Bruno Simeone. Fortran codes for network optimization. Annals of operations research, v. 13:1-4 0254-5330. J. C. Baltzer, Basel, Switzerland, 1988. ISBN 3-905135-29-9. 432 pp. LCCN T57.85 .F67 1988.

Sims:1988:SCP

[Sim88b] S. Eugene Sims. Selected computer programs in FORTRAN for fish stock assessment: sample outputs. Food and Agriculture Organization of the United Nations, Rome, Italy, 1988. ISBN 92-5-102714-5. v + 96 pp. LCCN SH331.5.C65 S441 1988. FAO fisheries technical paper, 259 Suppl.

Singh:1981:PFF

[Sin81] Bhagat Singh. Programming in FORTRAN, FORTRAN 77 and WATFIV. Technical report, University of Wisconsin-Marshfield, Marshfield, WI, USA, 1981. iv + 403 pp.

Sir:1982:SHFa

[SIR82a] SIR, Inc. SIR/HOST: FORTRAN interface to SIR/DBMS. SIR, Inc., Skokie, IL, USA, version 2.1.1 edition, 1982. 134 pp.

Sir:1982:SHFb

[SIR82b] SIR, Inc. SIR/HOST FORTRAN interface to SIR/DBMS. Scientific Information Retrieval, Inc., Evanston, IL, USA, 1982. various pp.

Sissom:1985:ICA

[Sis85] Denny L. Sissom. Interactive creation and automatic conversion of Nassi-Schneiderman charts to FORTRAN 77 source code. Thesis (M.S.), Tennessee Technological University, Cookeville, TN, USA, 1985. viii + 98 pp.

Simpson:1983:AFPa

[SJB83a] Robert W. Simpson, Robert C.

Jachens, and Richard J. Blakely. AIRYROOT: a FORTRAN program for calculating the gravitational attraction of an Airy isostatic root out to 166.7 km. Openfile report 83-883, U.S. Dept. of the Interior, Geological Survey, Reston, VA, USA, 1983. ???? pp.

Simpson:1983:AFPb

[SJB83b] Robert W. Simpson, Robert C. Jachens, and Richard J. Blakely. AIRYROOT: a FORTRAN program for calculating the gravitational attraction of an Airy isostatic root out to 166.7 km. Openfile report 83-883, U.S. Geological SurveyF, Menlo Park, CA, USA, 1983. 66 pp.

Shepelenko:1983:FEB

[SK83] V. N. Shepelenko and V. I. (Valerii Ivanovich) Karnachuk. FOR-TRAN EVM BESM-6. Izdvo “Nauka,” Sibirskoe otd-nie, Novosibirsk, USSR, 1983. ISBN ???? 78 + [2] pp. LCCN ????

Scarborough:1986:VFC

[SK86] Randolph G. Scarborough and Harwood G. Kolsky. A vectorizing Fortran compiler. IBM Journal of Research and Development, 30 (2):163–171, March 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kuo:1981:CSY

[sKcH81] Te sheng Kuo and Shun chin Hsu. Cheng shih yu yen = FORTRAN 77. Tien tzu chi suan ko hsueh tsung shu Tien tzu chi suan chi ko hsueh tsung shu. Tien yeh chu pan she: Tsung ching hsiao Sung kang tien nao tu shu tzu liao yu hsien kung ssu, Tai-pei shih, 1981. ISBN ???? 8 + 196 pp. LCCN ????

Skovoroda:1988:RKZ

[Sko88] T. P. (Tatiana Petrovna) Skovoroda. Reshenie kraevykh zadach dlia uravneniia laplasa metodom granichnykh integralnykh uravnenii. Materialy po matematicheskomu obespecheniiu EVM. Seriia FOR-TRAN 0, Nauch. tsentr biologicheskikh issledovanii AN SSSR v Pushchine, Pushchino, USSR, 1988. 15 + [1] pp.

Schultz:1982:NSM

[SL82] Joel H. Schultz and Jonathan D.

Lettvin. Nonalgebraic symbol manipulators for use in scientific and engineering modeling: introducing the FORSE (FORtran symbol expander). Technical Report PFC/RR-82-4, U.S.D.O.E. DOE/ET-51013-31, Massachusetts Institute of Technology, Plasma Fusion Center, Cambridge, MA, USA, 1982. 32 pp.

Sloan:1988:FPP

[Slo88] S. W. Sloan. A FORTRAN program for profile and wavefront reduction. Research report 027.07.1988, Department of Civil Engineering and Surveying, University of Newcastle, Newcastle, NSW, Australia, 1988. ISBN 0-7259-0614-6. 46 pp.

Saltykov:1984:PIF

[SM84] A. I. (Albert Ivanovich) Saltykov and G. I. (Grigorii Ivanovich) Makarenko. Programmirovanie na iazyke Fortran dlia BESM-6. Bibliotechka programmista. Nauka, Glav. red. fiziko-matematicheskoi lit-ry, Moskva, USSR, izd. 2., perer. i dop. edition, 1984. ISBN ???? 270 pp. LCCN ????

Schonfelder:1987:PFS

[SM87] L. Schonfelder and S. Morgan. Pointer functionality sans datatype. ACM SIGPLAN FOR-TRAN Forum, 6(3):12–26, December 1987. ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:1988:IPFa

[SM88a] J. L. (J. Lawrie) Schonfelder and J. S. (J. Steve) Morgan. An Introduction to programming in Fortran 77. Computer science texts. Blackwell Scientific Publications, Oxford, UK, 1988. ISBN 0-632-01748-1. xii + 296 pp. LCCN QA76.73.F25 S335 1988.

Schonfelder:1988:IPFb

[SM88b] J. L. (J. Lawrie) Schonfelder and J. S. (J. Steve) Morgan. An Introduction to Programming in Fortran 77. Computer Science Texts. Blackwell Scientific Publications, Oxford, UK, December 1988. ISBN 0-632-01184-X (paperback). xii + 296 pp. LCCN ???? US$29.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        063201184X.

Sheldon:1984:CTDb

[SMD84] E. Sheldon, S. Mathur, and D. Donati. Computation of total, differential and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bγ) and (a, bγ − γ). (III) Fortran translations of the Algol programs ‘mandy’ and ‘barbara’. Computer Physics Communications, 35(1–3):C–91–C–93, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584823528.

Smetana:1981:FCC

[Sme81] Frederick O. Smetana. Fortran Codes for Classical Methods in Linear Dynamics. Mc-Graw-Hill, New York, NY, USA, November 1981. ISBN 0-07-058440-0. ???? pp. LCCN ???? US$12.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070584400.

Smillie:1981:PNA

[Smi81] K. W. Smillie. Programming notes: APL, FORTRAN, ALGOL, PASCAL. Technical report, Dept. of Computing Science, University of Alberta, Edmonton, Alberta, Canada, 1981. ii + 65 pp.

Smillie:1983:PNA

[Smi83a] K. W. Smillie. Programming notes: APL, FORTRAN, ALGOL, PASCAL. Technical report, Dept. of Computing Science, University of Alberta, Edmonton, Alberta, Canada, 1983. ii + 105 pp.

Smith:1983:SRF

[Smi83b] B. T. Smith. Status report on FORTRAN 8x as of August 1983. ACM SIGNUM Newsletter, 18(4): 19–21, October 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1983:FCL

[Smi83c] Brian T. Smith. Fortran 8x: a continuing language for numerical software? Technical Report CONF-830531-3., Argonne National Laboratory; National Technical Information Service, Argonne, IL, USA, 1983. 8 pp.

Smillie:1984:RFP

[Smi84] K. W. Smillie. Reviews: FOR-TRAN Papers from NCC ‘82 Proceedings; capsule reviews. Annals of the History of Computing, 6(1):74–80, January/ March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1074.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1074abs.htm.

Smith:1985:SPF

[Smi85a] B. T. Smith. Software parts in Fortran 8X. ACM SIGPLAN FOR-TRAN Forum, 4(1):33–38, March 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Smith:1985:IFI

[Smi85b] Graham Smith. INPROG, a FOR-TRAN input program generator. UNSW report 1985/OR/2, School of Mechanical and Industrial Engineering, University of New South Wales, Kensington, NSW, Australia, 1985. ISBN 0-909287-77-5. 48 pp.

Smith:1985:WPF

[Smi85c] Graham Smith. Writing portable FORTRAN 77 programs. UNSW report 1985/OR/5, School of Mechanical and Industrial Engineering, University of New South Wales, Kensington, NSW, Australia, 1985. ISBN 0-909287-80-5. ii + 65 pp.

Smith:1985:FPA

[Smi85d] Marilyn Z. Smith. FORTRAN 77: a problem-solving approach. Houghton Mifflin Co., Boston, MA, USA, 1985. ISBN 0-395-35041-7 (paperback). xii + 460 + [6] pp. LCCN QA 76.73 F25 S59 1985.

Smith:1985:FPS

[Smi85e] Marilyn Z. Smith. Fortran 77: a Problem-Solving Approach. Houghton Mifflin, Boston, MA, USA, January 1985. ISBN 0-395-35041-7. ???? pp. LCCN ???? US$39.16. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0395350417.

Smith:1987:RAFa

[Smi87a] Brian T. Smith. A review and analysis of Fortran 8x. Technical report, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1987. iii + 50 pp.

Smith:1987:RAFb

[Smi87b] Brian T. Smith. A review and analysis of Fortran 8x. Technical Report ANL-87-40, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1987. vii + 45 pp.

Smith:1988:FPR

[Smi88a] Brian T. Smith. Fortran 8x — its public review. ACM SIGNUM Newsletter, 23(1):2–6, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1988:RAF

[Smi88b] Brian T. Smith. A review and analysis of Fortran 8x. ACM SIGNUM Newsletter, 23(2):29–60, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1988:WGS

[Smi88c] Graham Smith. Writing Good Software in Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1988. ISBN 0-13-969601-6, 0-7248-1281-4. xv + 375 pp. LCCN QA76.73.F25S58 1988. US$38.00.

Smith:1988:PIF

[Smi88d] Kevin Scott Smith. PAT: an interactive Fortran parallelizing assistant tool. Thesis (Ph.D.), Georgia Institute of Technology, Atlanta, GA, USA, 1988. iv + 99 pp. Directed by William F. Appelbe (U.M. order no. 89-16,173).

Smolarski:1989:EF

[Smo89a] Dennis Chester Smolarski. The ESSENTIALS of FORTRAN. Research and Education Association, New York, NY, USA, 1989. ISBN 0-87891-663-6. viii + 120 pp. LCCN QA76.73.F25S593 1989.

Smorlarski:1989:EF

[Smo89b] D. Rev. Smorlarski. The Essentials of Fortran. Research & Education Association, ????, May 1989. ISBN 0-87891-663-6. ???? pp. LCCN ???? US$5.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0878916636.

Snelling:1988:SFP

[Sne88] D. F. Snelling. Standard FOR-TRAN 77 as a parallel language. Parallel Computing, 8(1–3):409– 414, October 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

STCUS:1983:VSR

[Sof83a] Software Testing Center (U. S.).

Validation summary report: IBM corporation IBM 4341 VS FOR-TRAN release 3.0. Technical Report ????, Federal Software Testing Center; National Technical Information Service, Falls Church, VA, USA, 1983. 6 pp.

STUG:1983:PUA

[Sof83b] Software Tools Users Group, editor. Proceedings: USENIX Association [and] Software Tools Users Group Summer Conference, Toronto 1983, July 1983, Toronto, Ontario, Canada. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1983. ISBN ???? LCCN QA76.8.U65 U74 1983. “Sponsored by USENIX Association in cooperation with Software Tools Users Group.”–p. iii.

STUG:1984:UAS

[Sof84] Software Tools Users Group, editor. USENIX Association [and] Software Tools Users Group Summer Conference, Salt Lake City 1984: proceedings, June 12–15, 1984, Salt Lake City, Utah, USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1984. ISBN ???? LCCN QA76.8.U65 U83 1984.

Centre:1987:MST

[Sof87] Software Engineering Research Centre. The mothra software testing environment user’s manual. Technical Report SERC-TR-4-P, Software Engineering Research Centre, Utrecht, The Netherlands (??), August 1987.

Solchenbach:1989:SFM

[Sol89]  K. Solchenbach.  Suprenum-

Fortran — an MIMD/SIMD language. Supercomputer, 6(2):25–30, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Sommeijer:1986:NFL

[Som86] B. P. Sommeijer. Numvec Fortran library manual: Chapter: parabolic PDEs routine: BDMG. Report NM. Centrum voor Wiskunde en Informatica. 8610, CWI, Amsterdam, The Netherlands, 1986. [2] + 16 pp.

Soni:1983:DAC

[Son83] Som R. Soni. A digital algorithm for composite laminate analysis — Fortran. Technical report, University of Dayton Research Institute; National Technical Information Service, Dayton, OH, USA, 1983. 138 pp.

Sorensen:1984:BVP

[Sor84] D. C. Sorensen. Buffering for vector performance on a pipelined MIMD machine. Parallel Computing, 1(2):143–164, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Skordalakis:1982:CSV

[SP82] E. Skordalakis and G. Papakonstantinou. Control structure for a variable number of nested loops. The Computer Journal, 25(1):48– 51, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Sinha:1984:RET

[SP84a] Mihir K. Sinha and Larry R.

Padgett. Reservoir Engineering Techniques Using Fortran. International Red Cross, ????, December 1984. ISBN 0-934634-50-5. ???? pp. LCCN ???? US$49.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0934634505.

Spielman:1984:HFPa

[SP84b] J. B. Spielman and David A.

Ponce. Handtc: a FORTRAN program to calculate inner-zone terrain corrections. Open-file report 84-777, U.S. Dept. of the Interior, Geological Survey; Open-File Services Section, Branch of Distribution, Menlo Park, CA, USA, 1984. ???? pp.

Spielman:1984:HFPb

[SP84c] J. B. Spielman and David A. Ponce. Handtc: a FORTRAN program to calculate inner-zone terrain corrections. Open-file report 84-777., U.S. Geological Survey, Menlo Park, CA, USA, 1984. 20 pp.

Sinha:1985:RETa

[SP85a] Mihir K. Sinha and Larry R. Padgett. Reservoir engineering techniques using FORTRAN. International Human Resources Development Corp., Boston, MA, USA, 1985. ISBN 0-934634-50-5. xi + 224 pp. LCCN TN871.S554 1985.

Sinha:1985:RETb

[SP85b] Mihir K. Sinha and Larry R. Padgett. Reservoir engineering techniques using FORTRAN. Reidel [for] International Human Resources Development Corporation, Dordrecht, The Netherlands, 1985. ISBN 90-277-1921-7. xi + 224 pp.

Stephens:1987:PFF

[SP87] Bruce R. Stephens and John D. Pryce. Passing functions in Fortran for automatic differentiation. Technical Report AM–87–13, Maths. Dept., Bristol University, Bristol, UK, 1987.

Spath:1985:CDAb

[Spa85a] Helmuth Spath. The Cluster Dissection and Analysis Theory: Fortran Program Examples. Ellis Horwood Series in Computers and Their Applications. Ellis Horwood, New York, NY, USA, March 1985.

ISBN 0-13-137985-2. ???? pp. LCCN ???? US$54.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0131379852.

Spath:1985:CDAa

[Spa85b] Helmuth Spath. The Cluster Dissection and Analysis Theory Fortran Programs Examples. Ellis Horwood series in Computers and their applications. Ellis Horwood, New York, NY, USA, March 1985. ISBN 0-470-20129-0 (Halsted Press), 0-85312-736-0. 226 pp. LCCN QA278 .S68213 1985. UK£22.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0470201290. Translation by Johannes Goldschmidt of Cluster Formation und Analyse.

Sperry:1982:FAL

[Spe82] Sperry Rand Corporation. Univac Division. FORTRAN (ASCII) level 10R1: programmer reference. Sperry Univac, St. Paul, MN, USA, 1982. various pp.

Sperry:1983:IFS

[Spe83] Sperry Corporation. Introduction to FORTRAN: self-study course. Sperry Corp., Princeton, NJ, USA, 1983. 176 pp.

Sperry:1985:FAL

[Spe85] Sperry Rand Corporation. Univac Division. Fortran (ASCII): level 11R1 reference. Sperry Univac, St. Paul, MN, USA, 1985. various pp.

Savage:1984:SGC

[SPS84] William Z. Savage, Philip S. Powers, and Henri S. Swolfs. In situ geomechanics of crystalline and sedimentary rocks: part V, RVT — a FORTRAN program for an exact elastic solution for tectonic and gravity stresses in isolated symmetric ridges and valleys. Openfile report 84-827, U.S. Geological Survey, Denver, CO, USA, 1984. 12 pp.

Sheldon:1984:CTDc

[SR84a] E. Sheldon and V. C. Rogers. Computation of total and differential cross section for compound nuclear reactions of the type (a, a), (a, a′), (a, b), (a, γ), (a, γ − γ), (a, bγ), and (a, bγ − γ) (IV) Fortran program ‘CINDY’. Computer Physics Communications, 35(1–3):C–212–C–213, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584824558.

Steinberg:1984:UVW

[SR84b] Stanley Steinberg and P. Roache. Using VAXIMA to write Fortran code. In V. Ellen Golden, editor, Proceedings of the 1984 MACSYMA User’s Conference, page 1. General Electric, Schenectady, NY, USA, 1984.

Steinberg:1986:UMW

[SR86] Stanly Steinberg and Patrick J. Roache. Using Macsyma to write FORTRAN subroutines. Journal of Symbolic Computation, 2 (2):213–216, June 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Starkey:1987:FPF

[SR87] J. Denbigh Starkey and Rockford Ross. Fundamental programming with Fortran 77: a science and engineering approach. West Publishing Company, St. Paul, MN, USA, January 1987. ISBN 0-314-77805-5 (paperback). xviii + 876 pp. LCCN QA 76.73 F25 S77 1987. US$54.75; US$19.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0314778055.

Steinberg:1988:AGF

[SR88] S. Steinberg and P. Roache. Automatic generation of finite difference code. American Society of Mechanical Engineers, Heat Transfer Division, (Publication) HTD, 105:81–86, 1988. CODEN ASMHD8. ISSN 0272-5673.

Swartztrauber:1979:AEF

[SS79] Paul N. Swartztrauber and Roland A.

Sweet. Algorithm 541: Efficient Fortran subprograms for the solution of separable elliptic partial differential equations [D3]. ACM Transactions on Mathematical Software, 5(3):352–364, September 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smetana:1982:FCC

[SS82] Frederick O. Smetana and Andrew O. Smetana. Fortran codes for classical methods in linear dynamics. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-058440-0. x + 394 pp. LCCN TA345 .S58 1982b. US$9.95 (est.).

Shinagawa:1984:FSB

[SS84] Yoshiya Shinagawa and Yasuko Shinagawa. FORTRAN sitisiti to BASIC puroguramu zitureisyu. Baifukan, Tokyo, Japan, 1984. ISBN 4-563-01305-6. 110 pp.

Savage:1987:SGCa

[SS87a] William Z. Savage and Henri S. Swolfs. In situ geomechanics of crystalline and sedimentary rocks. Open-file report 87-82, Dept. of the Interior, U.S. Geological Survey: [Books and Open-File Reports Section, distributor], Denver, Colo.?, 1987. 1 microfiche pp.

Spoletini:1987:LF

[SS87b] Alessandra Spoletini and Enrico Spoletini. Il linguaggio Fortran. Collana dei Quaderni di informatica; 7. F. Angeli, Milano, Italy, 3a edition, 1987. ISBN ???? 315 pp. LCCN ????

Sato:1988:BCM

[SS88] R. K. Sato and P. N. Swarztrauber. Benchmarking the Connection Machine 2. In IEEE [IEE88b], pages 304–309. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Sikorski:1984:AAF

[SSS84a] K. Sikorski, F. Stenger, and J. Schwing. Algorithm 614: A FORTRAN subroutine for numerical integration in H_{p}. ACM Transactions on Mathematical Software, 10(2):152–160, June 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sikorski:1984:AFS

[SSS84b] K. Sikorski, F. Stenger, and J. Schwing. Algorithm 614: A FORTRAN subroutine for numerical integration in H_{p}. ACM Transactions on Mathematical Software, 10(2):152–160, June 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sikorski:1984:FSI

[SSS84c] K. Sikorski, F. Stenger, and J. Schwing. A Fortran subroutine for integration in H_{p} spaces. ACM Transactions on Mathematical Software, 10(2):140–160, June 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tsai:1985:SFC

[sT85] Erl shin Tsai. Simulating FOR-TRAN for the Cyber-205 on a VAX-11/ 780. Thesis (M.S.), Dept. of Computer Science, University of Houston — University-Park, Houston, TX, USA, 1985. vii + 75 pp.

Stewart:1976:AHE

[Ste76] G. W. Stewart. Algorithm 506:

HQR3 and EXCHNG: Fortran subroutines for calculating and ordering the eigenvalues of a real upper Hessenberg matrix [F2]. ACM Transactions on Mathematical Software, 2(3):275–280, September 1976. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [FW82].

Steuerwalt:1979:CEF

[Ste79] Michael Steuerwalt. Certification of “Algorithm 541: Efficient Fortran subprograms for the solution of separable elliptic partial differential equations [D3]”. ACM Transactions on Mathematical Software, 5(3):365–371, September 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Steding:1987:EIP

[Ste87] Burkhard Steding. Entwicklung und Implementierung eines Programmsystems zur Fakturierung und Auftragsabwicklung. Studienarbeit, 1987.

Stoyan:1984:MCSa

[Sto84a] H. Stoyan. Maschinen-unabhaengige Codeerzeugung als semantikerhaltende beweisbare Programmtransformation. Technical report, Universität Erlangen (??), Erlangen, Germany, 1984.

Stoyan:1984:MCSb

[Sto84b] H. Stoyan. Maschinen-unabhaengige Codeerzeugung als semantikerhaltende beweisbare Programmtransformation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13895-1.

Stone:1985:IVE

[Sto85a] J. M. Stone. Introduction to the VM/EPEX FORTRAN preprocessor. Research report RC 11407, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1985. iv + 17 pp.

Stone:1985:VEF

[Sto85b] J. M. Stone. The VM/EPEX FOR-TRAN preprocessor reference. Research report RC 11408 (#51330), IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1985. 30 pp.

Stone:1985:NPP

[Sto85c] Janice M. Stone. Nested parallelism in a parallel FORTRAN environment. Research report RC 11506, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1985. 20 pp.

Strohmeier:1982:FAS

[Str82] Alfred Strohmeier. Fortran 77:

approche systématique illustrée d’exemples. Eyrolles, Paris, France, 1982. ISBN ???? xiii + 144 pp. LCCN ????

Strohmeier:1985:FAS

[Str85] Alfred Strohmeier. FORTRAN 77: approche systématique illustrée d’exemples. Eyrolles, Paris, France, 3. nouveau tirage edition, 1985. ISBN ???? [xiv] + 144 pp. LCCN ????

Sturt:1981:SAA

[Stu81a] E. Sturt. Statistical algorithms: Algorithm AS 165: An algorithm to construct a discriminant function in Fortran for categorical data. Applied Statistics, 30(3):313–325, September 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.
                                        cmu.edu/apstat/165.

Sturt:1981:CCF

[Stu81b] Elizabeth Sturt. Computerized construction in FORTRAN of a discriminant function for categorical data. Applied Statistics, 30(3): 213–222, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sullivan:1988:RHU

[Sul88] Kevin M. (Kevin Maurice) Sullivan. The redefinition of HP-FORTRAN and its use in the translation from VAX-11 FOR-TRAN to Microsoft FORTRAN 77. Thesis (M.S.), Auburn University, Auburn, AL, USA, 1988. vii + 67 pp.

Sullins:1982:IPU

[Sul2 ] Dee G. Sullins. An instructional package for using FORTRAN using CAI. Thesis (ed.s.), Georgia State University, Atlanta, GA, USA, 1982 (??). various pp.

Sun:1984:FPS

[Sun84] Sun Microsystems. FORTRAN and Pascal for the Sun workstation. Sun Microsystems, Inc., Mountain View, CA, USA, 1984. various pp.

Sun:1985:FPG

[Sun85] Sun Microsystems. FORTRAN programmer’s guide for the Sun workstation. Sun Microsystems, Inc., Mountain View, CA, USA, revision D of 15 May, 1985 edition, 1985. various pp.

Sun:1986:FPGa

[Sun86a] Sun Microsystems. FORTRAN programmer’s guide. Sun Microsystems, Inc., Mountain View, CA, USA, revision E of 17 February 1986 edition, 1986. various pp.

Sun:1986:FPGb

[Sun86b] Sun Microsystems. FORTRAN programmer’s guide. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, revised E edition, 1986. xvi + 197 pp.

Sun:1987:SMS

[Sun87] Sun Microsystems. SunINGRES manual set. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, revision A of January 16, 1987 edition, 1987. 4 v. pp.

Sun:1988:FSC

[Sun88a] Hong-Jie Sun. A Fortran subroutine for computing normal orthant probabilities of dimensions up to nine. Communications in Statistics: Simulation and Computation, 17(3):1097–1111, 1988. CODEN CSSCDB. ISSN 0361-0918.

Sun:1988:SFP

[Sun88b] Sun Microsystems. Sun FOR-TRAN programmer’s guide. Sun Microsystems, Inc., Mountain View, CA, USA, revision A of 6 May, 1988 edition, 1988. various pp.

Susman:1986:IPD

[Sus86] Brook Fereday Susman. Implementation of a pre-processor for data abstractions in a Fortran environment. Thesis (M.S.), University of Maryland, College Park, MD, USA, 1986. ix + 84 pp.

Slape:1983:CFA

[SW83] John K. Slape and Peter J. L. Wallis. Conversion of Fortran to Ada using an intermediate tree representation. The Computer Journal, 26(4):344–353, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Swarztrauber:1984:FAV

[Swa84] P. N. Swarztrauber. FFT algorithms for vector computers. Parallel Computing, 1(1):45–63, August 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Symbolics:1985:UGF

[Sym85] Symbolics, Inc., Cambridge, MA, USA. User’s Guide to the FOR-TRAN 77 Tool Kit, March 1985.

Symbolics:1986:UGF

[Sym86] Symbolics. User’s guide to the FORTRAN 77 tool kit. Technical Report 99 90 12, Symbolics, Inc., Cambridge, MA, USA, August 1986.

Symbolics:1988:UGS

[Sym88] Symbolics. User’s guide to Symbolics FORTRAN 77 tool kit (revised). Technical Report 99 90 62, Symbolics, Inc., Cambridge, MA, USA, March 1988.

Szymanski:1987:FPW

[Szy87] J. K. Szymanski. Fortran program for Weibull parameter estimation. Division report / Mineral and Earth Sciences Program, Coal Research Laboratories; CRL 87-27 (TR) Division report (Coal Research Laboratories (Canada)); CRL 87-27 (TR), Coal Research Laboratories, Ottawa, Ontario, Canada, 1987. v + 13 pp.

Tanaka:1981:ETF

[Tan81a] Akihiko Tanaka. Equivalence testing for Fortran mutation system using data flow analysis. Thesis (M.S. in I.C.S.), Georgia Institute of Technology, Atlanta, GA, USA, 1981. iv + 67 pp. Directed by Richard A. DeMillo.

Tanik:1981:PMC

[Tan81b] M. M. Tanik. Prediction models for cyclomatic complexity. ACM SIGPLAN Notices, 16(4):89–97, April 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanaka:1982:ETF

[Tan82] Akihiko Tanaka. Equivalence testing for Fortran mutation system using data flow analysis. Technical Report GIT-ICS; 82/10, School of Information and Computer Science, Georgia Institute of Technology, Atlanta, GA, USA, 1982. various pp.

Tandem:1983:FPG

[Tan83a] Tandem Computers. FORTRAN pocket guide. Tandem Computers Inc., Cupertino, CA, USA, 1983. various pp.

Tanenbaum:1983:TCT

[Tan83b] Andrew S. Tanenbaum. Technical correspondence: Tanenbaum’s reply. ACM Transactions on Programming Languages and Systems, 5(3):499–500, July 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [TvSS82, Pem83].

Tandem:1985:FPG

[Tan85a] Tandem Computers. FORTRAN pocket guide. Tandem Computers Inc., Cupertino, CA, USA, second edition, 1985. vii + 61 pp.

Tandem:1985:ISF

[Tan85b] Tandem Computers. Interim supplement to the FORTRAN 77 reference manual. Tandem Computers Inc., Cupertino, CA, USA, 1985. various pp.

Tan:1986:AFD

[Tan86] Hui-Qian Tan. Automatic formula derivation for Fortran code generation in finite element analysis. Thesis (Ph.D.), Kent State University, Kent, OH, USA, 1986. v + 149 pp.

Tan:1988:AED

[Tan88] H. Q. Tan. Automatic equation derivation and code generation in engineering analysis. In IEEE [IEE88a], pages 597–603.

Tate:1987:FPM

[Tat87] Peter M. (Peter Michael) Tate. A FORTRAN program for modelling bedload transport using current meter records. Technical report 14, Ocean Sciences Institute, University of Sydney, Sydney, NSW, Australia, 1987. ISBN 0-7316-3809-3 (not printed in book). 9 + [16] pp.

Taylor:1984:ACL

[Tay84] David Taylor. An alternative to current looping syntax. ACM SIG-PLAN Notices, 19(12):48–53, December 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1986:VIP

[Tay86] D. Taylor. Visionaries: an interview with Professor Michael O’Leary. ACM SIGPLAN Notices, 21(6):41–54, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thakur:1985:FPT

[TBM85] A. K. Thakur, K. J. Berry, and P. W. Mielke, Jr. A FORTRAN program for testing trend and homogeneity in proportions. Computer Programs in Biomedicine, 19 (2-3):229–233, ???? 1985. CO-DEN COPMBU. ISSN 0010-468X (print), 1878-3139 (electronic).

Teague:1981:POC

[Tea81] David B. Teague. A projectoriented course (Computer Programming II). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 13 (1):41–45, February 1981. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th SIGCSE Symposium on Computer Science Education.

Teichman:1986:IFL

[Tei86]  Bob Teichman.  IBM For-

tran language conversion program overview. IBM Corporation, New York, NY, USA, 1986. iv + 23 + [18] pp.

Tellier:1982:SCP

[Tel82] Gilles Tellier. Solutions a certains problèmes de FORTRAN structuré avec WATFIV. École polytechnique de Montréal, Montréal, PQ, Canada, 2e edition, 1982. ISBN 2-553-00150-9. 50 pp.

Temperton:1989:FMC

[Tem89a] C. Temperton. Further measurements of (r_{∞}, n_{1/2}) on the CRAY-1 and CRAY X-MP. Parallel Computing, 11(1):107–111, July 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Temperton:1989:FMR

[Tem89b] Clive Temperton. Further measurements of (r$-$INF$/, n$- $HLF$/) on the Cray-1 and Cray X-MP. Parallel Computing, 11(1):

107–111, July 1989. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Terry:1987:FP

[Ter87] Patrick D. Terry. Fortran from Pascal. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-17821-4 (paperback). ix + 224 pp. LCCN QA76.73.F25 T47 1987. US$12.95; US$21.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201178214.

Tew:1981:UGTb

[Tew81] Bernard V. Tew. User’s guide to taxes: a FORTRAN IV adaptation of depreciation schedules. Faculty series FS-81-2, Division of Agricultural Economics, College of Agriculture University of Georgia, Athens, GA, USA, 1981. 35 pp.

Tarter:1986:FIU

[TFH86] Michael E. Tarter, William Freeman, and Alan Hopkins. A Fortran implementation of univariate Fourier series density estimation. Communications in Statistics: Simulation and Computation, 15(3):855–870, 1986. CO-DEN CSSCDB. ISSN 0361-0918.

Triolet:1986:APF

[TFI86] Remi Triolet, Paul Feautrier, and Francois Irigoin. Automatic parallelizations of Fortran programs in the presence of procedure calls. In Robinet and Wilhelm [RW86], page ?? ISBN 0-387-16442-1. LCCN QA76.6.E976 1986.

Tharp:1982:SRP

[Tha82] Alan L. Tharp. Selecting the “right” programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 14(1):151–155, February 1982. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 13th SIGCSE Symposium on Computer Science Education.

Thames:1989:FCA

[Tha89a] Joseph M. Thames. FORTRAN CALCULUS: a New Implementation of Synthetic Calculus. Digital Calculus Corp., Torrance, CA 90505, 1989.

Thames:1989:FCN

[Tha89b] Joseph M. Thames. FORTRAN CALCULUS: A New Implementation of Synthetic Calculus. Digital Calculus Corp., Torrance, CA 90505, 1989.

Nguyen:1989:MOH

[Tha89c] Thi Phuong Thao. MS-FORTRAN 3.2 on-screen help book, 1989. ISBN 1-55064-059-3 (set), 1-55064-058-5 (disks) 1-55064-057-7 (printed version).

Theriault:1988:FAE

[The88] James A. Theriault. A FOR-TRAN algorithm for estimation of azimuthal and vertical arrival angles at a rotatable horizontal line array. Technical communication 88/304, Defence Research Establishment Atlantic, Dartmouth, NS, Canada, 1988. iv + 56 pp.

Thompkins:1981:FPC

[Tho81] William T. Thompkins. FOR-TRAN program for calculating three-dimensional, inviscid, rotational flows with shock waves in axial compressor blade rows I– user’s manual. GT and PDL report 162, Gas Turbine and Plasma Dynamics Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1981. 170 pp.

Thompkins:1982:FPC

[Tho82a] William T. Thompkins. A FOR-TRAN program for calculating three-dimensional, inviscid, rotational flows with shock waves in axial compressor blade rows user’s manual. NASA contractor report 3560, National Aeronautical and Space Administration, Scientific and Technical Information Office, Washington, DC, USA, 1982. viii + 170 pp. For sale by the National Technical Information Service.

Thomson:1982:DFP

[Tho82b] David Procter Thomson. DIET.FOR:

a FORTRAN program for dietary analysis. Thesis (M.S.), Ball State University, Muncie, IN, USA, 1982. 113 pp.

Thorkildsen:1984:EFP

[Tho84a] Gunnar Thorkildsen. EWALD: a FORTRAN program for determination of possible three-beam scattering situations in the vicinity of a specified three-beam point. Arbeidspapirer fra Rogaland distriktshøgskole 17, Rogaland distriktshøgskole, Stavanger, Norway, 1984. 34 pp.

Thorkildsen:1984:MFPa

[Tho84b] Gunnar Thorkildsen. M-D-test 1: a FORTRAN program for calculation of correction factor to the observed integrated intensity due to multiple diffraction. Arbeidspapirer fra Rogaland distriktshøgskole 18, Rogaland distriktshøgskole, Stavanger, Norway, 1984. 48 pp.

Thorkildsen:1984:MFPb

[Tho84c] Gunnar Thorkildsen. M-D-test 2: a FORTRAN program for calculation of the relative changes in peak intensity and integrated power for a primary reflection at a possible three-beam points. Arbeidspapirer fra Rogaland distriktshøgskole 19, Rogaland distriktshøgskole, Stavanger, Norway, 1984. 48 + [2] pp.

Thorkildsen:1984:MFPc

[Tho84d] Gunnar Thorkildsen. M-diff: a FORTRAN program for multiple diffraction correction of measured integrated intensities. Arbeidspapirer fra Rogaland distriktshøgskole 15, Rogaland distriktshøgskole, Stavanger, Norway, 1984. 47 pp.

Thorkildsen:1984:OFP

[Tho84e] Gunnar Thorkildsen. ORGT: a FORTRAN program for determination of crystal orientation giving three-beam scattering at a 4circle X-ray diffractometer. Arbeidspapirer fra Rogaland distriktshøgskole 16, Rogaland distriktshøgskole, Stavanger, Norway, 1984. 44 pp.

Thompson:1986:VVV

[Tho86a] W. K. (W. Kenneth) Thompson.

Validation and verification (V and V) of Fortran 77 software. Technical report, Georgia Tech Research Institute, Georgia Institute of Technology, Atlanta, GA, USA, 1986.

Thornburg:1986:KIC

[Tho86b] J. Thornburg. Keyword input for C. ACM SIGPLAN Notices, 21 (6):55–58, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thune:1986:AGS

[Thu86] Michael Thuné. Automatic GKS stability analysis. SIAM Journal on Scientific and Statistical Computing, 7(3):959–977, July 1986. CODEN SIJCD4. ISSN 0196-5204.

Tew:1981:UGTa

[TMjC81] Bernard V. Tew, Wesley N.

Musser, and Chia jen Chou. User’s guide to TAXES: a Fortran IV adaptation of depreciation schedules. Faculty series FS81-2, Division of Agricultural Economics, College of Agriculture University of Georgia, Athens, GA, USA, 1981. 35 pp.

Templeton:1988:FPSa

[TMS88a] Edna M. Templeton, John C. Mc-Connell, and Allan D. Stauffer. A FORTRAN primer with some useful algorithms. Captus Press, North York, Ontario, Canada, revised for use on the VAX computer edition, 1988. ISBN 0-921801-22-X. vii + 381 pp.

Templeton:1988:FPSb

[TMS88b] Edna M. Templeton, John C. Mc-Connell, and Allan D. Stauffer. A FORTRAN primer with some useful algorithms. Captus Press, North York, Ontario, Canada, second edition, 1988. ISBN 0-921801-07-6. vii + 382 pp.

Todd:1985:PFI

[Tod85] Michael J. Todd. PLALGO: a FORTRAN implementation of a piecewise-linear homotopy algorithm for solving systems of nonlinear equations. Technical report, School of Operations Research and Industrial Engineering, College of Engineering, Cornell University, Ithaca, NY, USA, 1985. 131 pp.

Tone:1982:FKO

[Ton82] Kaoru Tone. FORTRAN: kihon oyo. Baifukan, Tokyo, Japan, 1982. ISBN 4-563-00759-5. 319 pp.

Touzeau:1984:FCF

[Tou84] Roy F. Touzeau. A Fortran compiler for the FPS-164 scientific computer. ACM SIGPLAN Notices, 19(6):48–57, June 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Van84a].

Tao:1985:NFP

[TPR85] G. Y. Tao, P. A. Pella, and R. M. Rousseau. NBSGSC — a FOR-TRAN program for quantitative X-ray fluorescence analysis. NBS technical note 1213, U.S. Dept. of Commerce, National Bureau of Standards, Gaithersburg, MD, USA, 1985. iv + 117 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Toomey:1988:IPF

[TPS^{+}88] L. J. Toomey, E. C. Plachy, R. G.

Scarborough, R. J. Sahulka, and J. F. Shaw. IBM parallel FOR-TRAN. IBM Systems Journal, 27 (4):416–435, November 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Trouve:1984:XOG

[TR84] Jean-Bernard Trouvé and Pierre N.

Robillard. XREF, outil de gestion des sous-programmes FORTRAN: manuel de l’usager (version VMS). Rapport technique EPM/RT-84-6, École polytechnique de Montréal, Montréal, PQ, Canada, 1984. 33 pp.

Tricot:1984:MBFb

[Tri84] Laurence Tricot. Modulad: Bibliothèque Fortran pour l’analyse des données: document normalisation: F66 version 1.1. INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, 1984. ISBN 2-7261-0373-1. 63 pp.

Tricot:1989:MBF

[Tri89] Laurence Tricot. Modulad: bibliotheque Fortran 77 pour l’analyse des données: version 2.2, supplement a la brochure de documentation de la version 2.1. INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, 1989. ISBN 2-7261-0599-8. 247 pp.

Tropp:1984:FA

[Tro84] Henry S. Tropp. FORTRAN anecdotes. Annals of the History of Computing, 6(1):59–64, January/March 1984. CODEN AH-COE5. ISSN 0164-1239. URL

                                        http://dlib.computer.org/an/
                                        books/an1984/pdf/a1059.pdf;
                                         http://www.computer.org/annals/
                                        an1984/a1059abs.htm.

Tatsuta:1988:HNP

[TS88] Rumi Tatsuta and Katsuhiro Sasaki. Hajimete no puroguramingu: FORTRAN to BASIC nyumon. Yuhikaku, Tokyo, Japan, shohan edition, 1988. ISBN 4-641-07501-8. vi + 236 pp.

Tsuchiya:1985:AAD

[Tsu85] M. Tsuchiya. Availability analysis for the design of distributed processing networks. The Journal of Systems and Software, 5(3):221– 227, August 1985. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tanabe:1988:BFS

[TSU88] Kunio Tanabe, Masahiko Sagae, and Sumie Ueda. BNDE, FOR-TRAN subroutines for computing Bayesian nonparametric univariate and bivariate density estimator. Number 24 in Computer science monographs (Tokei Suri Kenkyujo (Tokyo, Japan)). Institute of Statistical Mathematics, Tokyo, Japan, 1988. ISBN ???? 52 pp. LCCN ????

Tanabe:1981:NFS

[TU81] K. (Kosai) Tanabe and Sumie Ueda. NOLLSI, a FORTRAN subroutine for nonlinear least squares by a quasi-Newton method. Number 17 in Computer science monographs (Tokei Suri Kenkyujo (Tokyo, Japan)). Institute of Statistical Mathematics, Tokyo, Japan, 1981. ISBN ???? 25 pp. LCCN ????

Tucker:1986:FI

[Tuc86] Allen B. Tucker. Fortran: une introduction. McGraw-Hill, New York, NY, USA, 1986. ISBN 2-7042-1146-9. 52 pp.

Turner:1986:SPUa

[Tur86] William M. Turner. Scientific programming: using Fortran 77. Hutchinson computer studies series. Hutchinson Education, London, UK, 1986. ISBN 0-09-161601-8 (paperback). 180 pp. LCCN Q183.9.T87 1986. UK£7.95.

Tanenbaum:1982:UPO

[TvSS82] Andrew S. Tanenbaum, Hans van Staveren, and Johan W. Stevenson. Using peephole optimization on intermediate code. ACM Transactions on Programming Languages and Systems, 4(1):21–36, January 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remarks [Pem83, Tan83b].

Toews:1987:IFI

[TW87] N. A. Toews and A. S. Wong.

Impresslib: Fortran impress interface library. Divisional report / Mining Research Laboratories (Canada) MRL 87-25(TR), Mining Research Laboratories, Ottawa, Ontario, Canada, 1987. iv + 17 pp.

Tsuji:1988:SFP

[TWI88] Tatsuo Tsuji, Katsumasa Watanabe, and Athushi Ikehata. Structured Fortran preprocessors generating optimized output. Software—Practice and Experience, 18 (5):427–442, May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ullrich:1984:RWF

[Ull84] C. Ullrich. Rechnerarithmetik und die Weiterentwicklung von FOR-TRAN. Elektronische Rechenanlagen, 26, 2:71–78, 1984. CODEN ELRAA4. ISSN 0013-5720.

Ullrich:1985:FAM

[Ull85] Ch. Ullrich. A Fortran 8x application module for optimal arithmetic. In Proc. 11th Imacs Congr. Syst. Simul. Sci. Comput. Oslo, Norway, Aug. 5–9, 1985, Vol. I, pages 175–178. ??, ??, 1985.

Ullrich:1986:FAM

[Ull86] Ch. Ullrich. A Fortran 8x application module for optimal arithmetic. In Vichnevetsky and Vignes [VV86], pages 175–178. ISBN 0-444-70067-6, 0-444-70082-X. LCCN QA 297 N857 1986.

USFHA:1981:UMX

[Uni81a] United States. Federal Highway Administration. User’s manual for XSRAIN: a FORTRAN IV program for calculation of flood hydrographs for ungaged watersheds. The Division; National Technical Information Service [distributor], Washington, DC, USA, 1981. v + 168 pp.

UMCC:1981:MFR

[Uni81b] University of Minnesota. Computer Center. M N F reference manual: Minnesota FOR-TRAN. Technical report, University of Minnesota, University Computer Center, Minneapolis, MN, USA, 1981. vi + 259 pp.

UCBCC:1982:FCG

[Uni82a] University of California, Berkeley.

Computing Services. FORTRAN conversion guide. CDC 14.2.2, Computing Services, University of California, Berkeley, CA, USA, February 1982. 61 pp.

UCBCC:1982:UFS

[Uni82b] University of California, Berkeley.

Computing Services. UNIX FOR-TRAN subroutines. Unx; 1.2.7 unx; 1.2.07., Computing Services, University of California, Berkeley, CA, USA, 1982. various pp.

UGCS:1982:FFO

[Uni82c] University of Glasgow. Computing Service. FORTRAN on the 2976: the F1 and optimising FORTRAN compilers. User guide G1/1, University of Glasgow Computing Service, Glasgow, Scotland, 1982. ii + 57 pp.

USNBS:1983:FFF

[Uni83a] United States. National Bureau of Standards. FIPS 69: FOR-TRAN; FIPS interpretation no. 1– nested parentheses in an expression. Technical report, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1983. various pp.

UMCC:1983:MRM

[Uni83b] University of Minnesota. Computer Center. M77 reference manual: Minnesota FORTRAN, 1977 standards version. The Center, Minneapolis, MN, USA, edition 1 edition, 1983. viii + 353 pp.

UWDCS:1983:VCF

[Uni83c] University of Waterloo. Dept. of Computing Services. VM/CMS, FORTRAN user’s guide. Technical report, Dept. of Computing Services, University of Waterloo, Waterloo, Ontario, Canada, 1983. iii + 36 pp.

USEPA:1984:HSPa

[Uni84a] United States. Environmental Protection Agency. Hydrological simulation program — FORTRAN (HSPF) user’s manual for release 8.0. Technical report, Environmental Research Laboratory, Office of Research and Development, U.S. Environmental Protection Agency, Athens, GA, USA, 1984. vi + 767 pp.

USEPA:1984:HSPb

[Uni84b] United States. Environmental Protection Agency. Hydrological simulation program — FORTRAN (HSPF): User’s manual for release 8.0. Technical report, Environmental Research Laboratory, Office of Research and Development, U.S. Environmental Protection Agency, Athens, GA, USA, 1984. vi + 767 pp.

UWDCS:1984:VCF

[Uni84c] University of Waterloo. Dept. of Computing Services. VM/CMS Fortran user’s guide. Technical report, Dept. of Computing Services, University of Waterloo, Waterloo, Ontario, Canada, 1984. iv + 50 pp.

UWMACC:1984:AIR

[Uni84d] University of Wisconsin — Madison. Academic Computing Center. Approximation and interpolation routines: subroutines for ASCII Fortran; user manual for the Sperry 1100. Mathematical routines series, University of Wisconsin, MACC Academic Computing Center, Madison, WI, USA, 1984. various pp.

UWMACC:1984:DER

[Uni84e] University of Wisconsin — Madison. Academic Computing Center. Differential equations routines: subroutines for ASCII Fortran; reference manual for the Sperry 1100. Mathematical routines series, University of Wisconsin — Madison, MACC Academic Computing Center, Madison, WI, USA, 1984. various pp.

USBR:1985:CSD

[Uni85a] United States. Bureau of Reclamation, Denver, CO, USA. Callable subroutines: documentation for FORTRAN subroutines, 1985. 57 pp.

USNBS:1985:Fa

[Uni85b] United States. National Bureau of Standards. Fortran. Federal information processing standards publication, FIPS PUB 69-1, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1985. 3 pp. For sale by the National Technical Information Service.

USNBS:1985:Fb

[Uni85c] United States. National Bureau of Standards. FORTRAN. Federal information processing standards publication, FIPS PUB 69-1, United States. National Bureau of Standards, Washington, DC, USA, 1985. 3 pp. For sale by the National Technical Information Service.

UACS:1986:MMI

[Uni86a] University of Alberta. Computing Services. MinIdent, mineral identification: a FORTRAN 77 program for mineral identification. Technical report, Computing Services, University of Alberta, Edmonton, Alberta, Canada, 1986. ii + 5 pp.

UWDCS:1986:VCF

[Uni86b] University of Waterloo. Dept. of Computing Services. VM/CMS, FORTRAN user’s guide. Technical report, Dept. of Computing Services, University of Waterloo, Waterloo, Ontario, Canada, 1986. iv + 52 pp.

UGCS:1987:FUV

[Uni87] University of Glasgow. Computing Service. Fortran 77 under VME. User guide G2/1, University of Glasgow Computing Service, Glasgow, Scotland, 1987. various pp.

UWMACC:1988:SSF

[Uni88] University of Wisconsin — Madison. Academic Computing Center. SPINC2/SPINS4 Fortran 77 subroutines for the shortest path tree problem: reference manual for the VMS VAX. Technical report, Madison Academic Computing Center, University of Wisconsin — Madison, Madison, WI, USA, 1988. various pp.

USENIX:1982:UAS

[Usr82] Usr Group, editor. /usr/group, USENIX Association, Software Tools Users Group Joint Conference Proceedings: Boston, July 1982. USENIX Association, Berkeley, CA, USA, 1982. ISBN ???? LCCN ????

Utter:1989:PD

[Utt89] Sue Utter. Parallel debugging.

Fortran Journal, 1(6):3–5, September/October 1989. ISSN 1060-0221.

Vagi:1989:AAF

[Vag89] Jon Curtis Vagi. Aftran: array Fortran programming language. Thesis (M.S.), Florida State University, Tallahassee, FL, USA, 1989. v + 24 pp.

Valentino:1985:FTE

[Val85] James Valentino. Fortran for Technologists and Engineers. Holt, Reinhart, and Winston, New York, NY, USA, September 1985. ISBN 0-03-060569-5 (paperback). xiii + 594 pp. LCCN QA 76.73 F25 V35 1985. US$36.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0030605695.

VanDooren:1982:ADE

[Van82] P. Van Dooren. Algorithm 590: DSUBSP and EXCHQZ: FOR-TRAN subroutines for computing deflating subspaces with specified spectrum. ACM Transactions on Mathematical Software, 8 (4):376–382, December 1982. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [PCK84].

VanDeusen:1984:CCP

[Van84a] M. Van Deusen, editor. Compiler construction: Proceedings of the ACM SIGPLAN ’84 symposium (Montréal, Canada, June 17–22, 1984), ACM SIGPLAN Notices, vol. 19, no. 6, June, 1984. ACM Press, New York, NY 10036, USA, 1984. ISBN 0-89791-139-3. LCCN QA 76.7 A1095 v.19 no.6. US$19 members; US$25 non-members.

VanTuyl:1984:EF

[Van84b] Robert R. Van Tuyl. On evolution of FORTRAN. ACM SIGPLAN Notices, 19(11):52–55, November 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanderplaats:1984:AFP

[Van84c] Garret N. Vanderplaats. ADS a FORTRAN program for automated design synthesis. NASA contractor report NASA CR-177985, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1984. ???? pp. For sale by the National Technical Information Service.

Vanderplaats:1985:AFP

[Van85] Garret N. Vanderplaats. ADS a FORTRAN program for automated design synthesis, version 1.10. NASA contractor report NASA CR-177985, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1985. ???? pp. For sale by the National Technical Information Service.

vanderVorst:1986:PFI

[van86] H. A. van der Vorst. The performance of FORTRAN implementations for preconditioned conjugate gradients on vector computers. Parallel Computing, 3(1):49– 58, March 1986. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vargas:1985:AVF

[Var85] Luis E. Vargas. Automated verification of FORTRAN 77 programs. Thesis (M.S.), Arizona State University, Tempe, AZ, USA, 1985. x + 92 + [1] pp.

Collani:1987:TRA

[vC87] Gernot von Collani. Testing and running algorithms for the detection of cliques (with a FOR-TRAN program). Technical report, Institute of Psychology, University of Technology, Brunswick, Brunswick, West Germany, 1987. 15 pp.

Vinette:1989:USC

[VC89] F. Vinette and J. Cizek. The use of symbolic computation in solving some nonrelativistic quantum mechanical problems. In Gianni [Gia89], pages 85–95. ISBN 3-540-51084-2. LCCN ????

Volpano:1984:EIC

[VD84] D. Volpano and H. Dunsmore. Empirical investigation of COBOL features. Information Processing and Management, 20(1/2): 277–291, 1984. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic). From Computing Reviews: “… (6) COBOL compilers can help solve some problems by:

(a) Coercing edited numeric data items in arithmetic expressions, as does FORTRAN. (b) Automatically correcting simple spelling errors with known techniques. (c) Better and more accurate diagnostics in compilers which are deficient in these areas.”.

vanderVorst:1985:CPT

[vdV85a] H. van der Vorst. Comparative performance tests of Fortran codes on the CRAY-1 and CYBER 205. In J. van Leeuwen and J. Lenstra, editors, Parallel Computers and Computations. CWI, Amsterdam, 1985. CWI Syllabus 9.

Vorst:1985:CPT

[vdV85b] H. van der Vorst. Comparative performance tests of Fortran codes on the CRAY-1 and CYBER 205. In J. van Leeuwen and J. Lenstra, editors, Parallel Computers and Computations. CWI, Amsterdam, The Netherlands, 1985. CWI Syllabus 9.

Vorst:1986:PFI

[vdV86] H. van der Vorst. The performance of Fortran implementations for preconditioned conjugate gradients on vector computers. Parallel Computing, 3:49–58, 1986. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Velasevic:1982:RLC

[Vel82] Dusan M. Velasević. Right-to-left code generation for arithmetic expressions. The Computer Journal, 25(3):316–326, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
25/3/316.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/
316.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_03/tiff/317.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
25/Issue_03/tiff/318.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/319.tif;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/
320.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_03/tiff/321.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
25/Issue_03/tiff/322.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_
03/tiff/323.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/
324.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_25/Issue_03/tiff/325.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
25/Issue_03/tiff/326.tif.

Vetterling:1985:NRE

[Vet85] William T. Vetterling. Numerical recipes example book (Fortran). Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-31330-9 (paperback). viii + 179 pp. LCCN QA76.73.F25 N84 1985.

Vogelsang:1987:MCF

[VH87] R. Vogelsang and C. Hoheisel. A Monte Carlo FORTRAN 200 programme for the determination of static properties of liquids vectorized to run on the CY-BER 205 vector processing computer. Computer Physics Communications, 46(2):209–216, August 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465587900026.

Vogel:1981:MRM

[VL81] John Vogel and Lawrence Liddiard. MNF reference manual: Minnesota FORTRAN. Technical report, The University, Minneapolis, MN, USA, 1981. vi + 259 pp.

VanRompuy:1986:IFR

[VLV^{+}86] L. L. Van Rompuy, C. Lesage, M. E. Vanderhaegen, M. P. Telemans, and M. F. and Zabeau. An improved FORTRAN 77 recombinant DNA database management system with graphic extensions in GKS. Computer applications in the biosciences: CABIOS, 2(4): 251–255, December 1986. CODEN COABER. ISSN 0266-7061.

vonMeerwall:1984:FCA

[vM84a] E. D. von Meerwall. A Fortran code for automatic spectrum analysis on medium-scale computers. Computer Physics Communications, 35(1–3):C–316, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584825412.

vonMeerwall:1984:FPRc

[vM84b] E. D. von Meerwall. A Fortran program for routine analysis of magnetic susceptibility data. Computer Physics Communications, 35(1–3):C–500, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584826867.

vonMeerwall:1984:FPC

[vM84c] E. D. von Meerwall. A Fortran program to collect histograms of microscopic scalar interactions. Computer Physics Communications, 35(1–3):C–444, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S001046558482651X.

vonMeerwall:1984:FPI

[vM84d] E. D. von Meerwall. A Fortran program to interpret pulsed field-gradient spin-echo diffusion data. Computer Physics Communications, 35(1–3):C–557, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/S0010465584827320.

vonMeerwall:1984:FPS

[vM84e] E. D. von Meerwall. A Fortran program to simulate quadrupoledistorted NMR powder patterns. Computer Physics Communications, 35(1–3):C–418, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/S0010465584826284.

vonMeerwall:1984:SFP

[vM84f] E. D. von Meerwall. A simple Fortran program to interpret cubic X-ray powder diffraction data. Computer Physics Communications, 35(1–3):C–383, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584825965.

vonMeerwall:1981:FPF

[vMF81] E. D. von Meerwall and R. D. Ferguson. A Fortran program to fit diffusion models to field-gradient spin echo data. Computer Physics Communications, 21(3):421–429, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465581900187.

vonMeerwall:1984:FPF

[vMF84] E. D. von Meerwall and R. D. Ferguson. A Fortran program to fit diffusion models to field-gradient spin echo data. Computer Physics Communications, 35(1–3):C–677, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0010465584828301.

Verdin:1981:UMA

[VMS81] L. Kristin Verdin and Hubert J.

Morel-Seytoux. User’s manual for AQUISM: FORTRAN IV PRO-GRAMS for DISCRETE KER-NELS GENERATION and for SIMULATION of an ISOLATED AQUIFER BEHAVIOR IN TWO DIMENSIONS. Hydrowar interim reports, Hydrowar Program, Colorado State University, Fort Collins, CO, USA, 1981. 199 pp.

vonMeerwall:1984:FPRa

[vMT84] E. von Meerwall and D. Thompson. A Fortran program for reduction of NMR relaxation data. Computer Physics Communications, 31(4):385–392, March 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465584900213.

Vetterling:1989:NREb

[VN89a] William T. Vetterling and Numerical Recipes Software. Numerical recipes example diskette (FOR-TRAN), 1989. ISBN 0-521-30957-3.

Vetterling:1989:NRF

[VN89b] William T. Vetterling and Numerical Recipes Software. Numerical recipes FORTRAN example diskette: format, Apple Macintosh, 1989. ISBN 0-521-35468-4.

Voelcker:1989:TS

[Voe89] J. Voelcker. Technology ’89: software. IEEE Spectrum, 26(1): 37–39, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vorontsov:1989:EAS

[Vor89] S. A. Vorontsov. An example of the application of symbolic computation in solid state physics. Journal of Symbolic Computation, 7 (2):179–182 (or 179–181??), February 1989. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Veran:1984:QAP

[VP84] M. Veran and D. Potier. QNAP2:

a portable environment for queueing systems modelling. Rapports de recherche 314, INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, 1984. 1–44 pp.

Vignes:1982:TPP

[VPH82] Jean Vignes, Michel La Porte, and Bernard Hallopeau. Theorie et pratique de la programmation Fortran. Editions Technip, Paris, France, 3eme edition, 1982. ISBN 2-7108-0115-9. 220 pp.

Vogelsang:1983:VMD

[VSH83] R. Vogelsang, M. Schoen, and C. Hoheisel. Vectorization of molecular dynamics Fortran programs using the Cyber 205 vector processing computer. Computer Physics Communications, 30(3):235–241, November 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465583900905.

Vetterling:1989:NREa

[VTP89] William T. Vetterling, Saul A.

(Saul Arno) Teukolsky, and William H. Press. Numerical recipes: example book (FOR-TRAN). Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-31330-9. viii + 179 pp. LCCN QA76.73.F25 N84 1985.

Vetterling:1987:NRE

[VTPF87] William T. Vetterling, Saul A. Teukolsky, William H. Press, and Brian P. Flannery. Numerical Recipes: Example Book (Fortran). Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-31330-9. viii + 179 pp. LCCN QA76.73.F25 N84 1987.

VuTienKhang:1989:FSC

[Vu 89] J. Vu Tien Khang. A FORTRAN subroutine to compute inbreeding and kinship coefficients according to the number of ancestral generations. Computer applications in the biosciences: CABIOS, 5 (3):199–204, July 1989. CODEN COABER. ISSN 0266-7061.

Vichnevetsky:1986:NMA

[VV86] Robert Vichnevetsky and Jean Vignes, editors. Numerical mathematics and applications: 11th IMACS World Congress on Scientific Computation, Oslo, Norway, 5–9 August 1985, volume 1 of IMACS transactions on scientific computation. Elsevier, Amsterdam, The Netherlands, 1986. ISBN 0-444-70067-6, 0-444-70082-X. LCCN QA 297 N857 1986.

vandenHeuvel:1987:AGFa

[vvHG87a] P. van den Heuvel, J. A. van Hulzen, and V. V. Goldman. Automatic generation of FORTRANcoded Jacobians and Hessians. Memorandum INF-87-4, Department of Computer Science, Twente University of Technology, Twente, The Netherlands, January 1987.

vandenHeuvel:1987:AGFb

[vvHG87b] P. van den Heuvel, J. A. van Hulzen, and Victor V. Goldman. Automatic generation of FORTRAN-coded Jacobians and Hessians. In J. Davenport, editor, Proceedings of EUROCAL 87, volume 378 of Lecture Notes in Computer Science, pages 120– 131. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Vanbegin:1989:FSC

[VVV89a] M. Vanbegin, P. Van Dooren, and M. Verhaegen. Fortran subroutines for computing the square root covariance filter and square root information filter in dense or Hessenberg forms. ACM Transactions on Mathematical Software, 15(3): 243–256, September 1989. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Vanbegin:1989:AFS

[VVV89b] Michel Vanbegin, Paul Van Dooren, and Michel Verhaegen. Algorithm 675: FORTRAN subroutines for computing the square root covariance filter and square root information filter in dense or Hessenberg forms. ACM Transactions on Mathematical Software, 15(3):243–256, September 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0098-
3500/69647.html.

Watson:1981:SWA

[WAD81] Collin J. Watson, John M. Anderson, and J. Daniel. Student workbook to accompany First course in data processing with BASIC, COBOL, FORTRAN, and RPG, second edition, and First course in data processing with BASIC, [by] Couger and McFadden. John Wiley and Sons, New York, London, Sydney, 1981. ISBN 0-471-09128-6. 248 pp.

Wagener:1984:SWT

[Wag84] Jerrold L. Wagener. Status of work toward revision of programming language Fortran. ACM SIGNUM Newsletter, 19(3):1–42, July 1984. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). Also in FORTEC forum, v. 3, no. 2 (June 1984).

Wagener:1985:IFG

[Wag85] J. L. Wagener. Is Fortran getting too big? ACM SIGPLAN FOR-TRAN Forum, 4(2):16–??, June 1985. ISSN 1061-7264 (print), 1931-1311 (electronic).

Walters:1981:CPS

[Wal81] Donald Richard Walters. A command processor and source card generator for a FORTRAN 77 preprocessor — TIDY 81. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1981. vi + 126 pp.

Waldman:1985:FPD

[Wal85] W. (Witold) Waldman. A Fortran program for the determination of unsteady air forces on general combinations of interfering lifting surfaces oscillating in subsonic flow. Structures report 412, Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratories, Melbourne, Victoria, Australia, 1985. ISBN 0-642-11263-0. 18 + [58] pp.

Waller:1983:TFI

[Wal3 ] Thomas R. Waller. Two FOR-TRAN II programs for the univariate and bivariate analysis of morphometric data. United States. National Museum. Bulletin 285, Smithsonian Institution Press, Washington, DC. USA, 1983 (??). v + 55 pp. For sale by the Supt. of Docs. U.S. Govt. Print. Off.

Wang:1984:VFL

[Wan84] Wang Laboratories, Inc. VS FORTRAN 77 language reference. Technical report, Wang Laboratories, Lowell, MA, USA, 1984. various pp.

Wang:1985:CSN

[Wan85] Paul S. Wang. Combining symbolic and numerical computational techniques on modern workstations. Proceedings of the Hawaii International Conference on System Science, pages 248–??, 1985. CODEN PHISD7. ISSN 0073-1129.

Wang:1986:FSS

[Wan86] Paul S. Wang. FINGER: a symbolic system for automatic generation of numerical programs in finite element analysis. Journal of Symbolic Computation, 2(3):305– 316, September 1986. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Ward:1986:CDE

[War86] R. C. (Roy C.) Ward. The catchwater drain experimental catchment: FORTRAN listing of a hydrological model. Miscellaneous series; no.29, Department of Geography, University of Hull, Hull, UK, 1986. ISBN 0-85958-115-2 (paperback). ii + 35 pp.

Watts:1982:FPCa

[Wat82a] Raymond D. Watts. FORTRAN ’77 programs for computing data fitting functions based on a principle of minimum integrated squared curvature. Open-file series 82-831, U.S. Geological Survey, Reston, VA, USA, 1982. ???? pp.

Watts:1982:FPCb

[Wat82b] Raymond D. Watts. FORTRAN ’77 programs for computing data fitting functions based on a principle of minimum integrated squared curvature. Open-file series 82-831, U.S. Geological Survey, Denver, CO, USA, 1982. 41 pp.

WatcomSystems:1985:WF

[Wat85] Watcom Systems. WATCOM Fortran, 1985.

Watson:1986:IDS

[Wat86] Clayton A. Watson. An implementation of dynamic structures using VAX FORTRAN and VMS. Technical report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1986. 8 pp.

Watanabe:1987:APN

[Wat87] T. Watanabe. Architecture and performance of NEC supercomputer SX system. Parallel Computing, 5(1–2):247–255, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Whitney:1985:IGA

[WB85] David C. Whitney and Michel H.

Boillot. Instructor’s guide to accompany Understanding FOR-TRAN. West Publishing, St. Paul, MN, USA, third edition, 1985. ISBN 0-314-87094-6. 247 pp.

Ward:1989:FAP

[WB89] Tim Ward and Eddie Bromhead.

FORTRAN and the Art of PC Programming. John Wiley and Sons, New York, London, Sydney, December 1, 1989. ISBN 0-471-92253-6. xi + 337 pp. LCCN QA76.73.F25 W37 1989. US$39.90; US$72.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471922536.

Weston:1981:EFPa

[WD81a] Suzanne I. Weston and T. S. Dyman. EXTRACT a FORTRAN program to retrieve selected Well History Control System (WHCS) data. Open-file report 81-29, U.S. Dept. of the Interior, Geological Survey, Denver, CO, USA, 1981. i + 48 pp.

Weston:1981:EFPb

[WD81b] Suzanne I. Weston and T. S. Dyman. EXTRACT, a FORTRAN program to retrieve selected Well History Control System (WHCS) data. Open-file report 81-29, U.S. Geological Survey, Denver, CO, USA, 1981. i + 48 pp.

Webring:1985:SFPa

[Web85a] Michael Webring. SAKI: a Fortran program for generalized linear inversion of gravity and magnetic profiles. Open-file series 85-122, U.S. Dept. of the Interior, Geological Survey, Reston, VA, USA, 1985. ???? pp.

Webring:1985:SFPb

[Web85b] Michael Webring. SAKI: a Fortran program for generalized linear inversion of gravity and magnetic profiles. Open-file report 85-122, U.S. Geological Survey, Denver, CO, USA, 1985. 108 pp.

Webber:1988:SF

[Web88] Robert P. Webber. Short FOR-TRAN. Kendall/Hunt Pub., Dubuque, IA, USA, 1988. ISBN 0-8403-4670-0. 137 + [1] pp.

Weeks:1986:CEF

[Wee86] Cindy Lou Weeks. Concurrent extensions to the FORTRAN language for parallel programming of computational fluid dynamics algorithms. NASA technical memorandum 88363, National Aeronautics and Space Administration, Ames Research Center, Moffett Field, CA, USA, 1986. ???? pp.

For sale by the National Technical Information Service.

Weerawarana:1989:GAC

[Wee89] Sanjiva Weerawarana. GEN-CRAY: an automatic code generator for Cray Fortran. Thesis (M.S.), Kent State University, Kent, OH, USA, 1989. xi + 195 pp.

Wehnes:1985:FSP

[Weh85] H. Wehnes. Fortran 77 Strukturierte Programmierung mit Fortran 77. Carl Hanser, München, Germany, 1985. ISBN 3-446-14259-2. ???? pp.

Weinberger:1984:CDI

[Wei84] Peter J. Weinberger. Cheap dynamic instruction counting. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1815–1826, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Weinman:1986:IMA

[Wei86a] David G. Weinman. Instructor’s manual to accompany VAX FOR-TRAN. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, 1986. ISBN 0-87835-173-6. various pp.

Weinman:1986:VF

[Wei86b] David G. Weinman. VAX FOR-TRAN. The Boyd and Fraser programming language series. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, 1986. ISBN 0-87835-172-8 (paperback). xiii + 450 pp. LCCN QA76.8.V37 W45 1986. US$22.00.

Weinman:1989:FSE

[Wei89a] David G. Weinman. Fortran for Scientists and Engineers. PWS-Kent Pub. Co., Boston, MA, USA, April 1989. ISBN 0-87835-337-2. xv + 527 + 4 pp. LCCN QA76.73.F25 W45 1989. US$47.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0878353372.

Weinman:1989:VF

[Wei89b] David G. Weinman. VAX Fortran. PWS-Kent series in computer science. PWS-Kent Pub. Co., Boston, MA, USA, 1989. ISBN 0-87835-172-8 (paperback). xvi + 450 pp. LCCN QA76.8.V37W45 1986.

Wexelblat:1981:HPL

[Wex81] R. L. Wexelblat, editor. History of programming languages. Proceedings of the ACM SIGPLAN conference (Los Angeles, Calif., June 1–3, 1978). Academic Press, New York, NY, USA, June 1981. ISBN 0-12-745040-8. LCCN QA76.7 .H56 1978. US$45.

Wexelblat:1987:IIT

[Wex87] Richard L. Wexelblat, editor. Proceedings of the SIGPLAN ’87 Symposium on Interpreters and Interpretive Techniques, St. Paul, Minnesota, June 24–26, 1987. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-235-7. LCCN QA76.7 .S54 v.22:7. US$23.00. SIGPLAN Notices, v. 22, no. 7 (July 1987).

Wurgler:1985:FPC

[WF85] F. E. Wurgler and H. Frei. A FOR-TRAN program for the construction of Selby-Olson tables. Basic Life Sciences, 34(????):373–381, ???? 1985. CODEN BLFSBY. ISSN 0090-5542.

Wainwright:1984:BFP

[WG84] S. J. Wainwright and A. Grant.

BASIC and FORTRAN in parallel. Babani, London, UK, 1984. ISBN 0-85934-112-7 (paperback). 79 pp.

Wade:1987:NAB

[WH87] S. Wade and J. Henstridge. Note on Algorithm 117: Buffered output in FORTRAN. The Computer Journal, 30(3):276, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_30/Issue_03/tiff/276.

tif. See [HPB84].

Wherry:1984:EDFa

[Whe84a] Kenneth Lynn Wherry. Error detection in Fortran programs as a function of error type and location. Thesis (M.A.), University of Missouri–Columbia, Columbia, MO, USA, 1984. 29 pp.

Wherry:1984:EDFb

[Whe84b] Kenneth Lynn Wherry. Error detection in Fortran programs as a function of error type and location. Thesis (M.A.), University of Missouri–Columbia, Columbia, MO, USA, 1984. 29. pp.

Whitaker:1981:FFO

[Whi81a] Lt Col William A. Whitaker. FORTRAN-like formatted output with Ada. ACM SIGADA Ada Letters, 1(1):26–28, July/August 1981. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Whitaker:1981:FLF

[Whi81b] Lt Col William A. Whitaker. FORTRAN-like formatted output with Ada. ACM SIGADA Ada Letters, 1(1):26–28, July/August 1981. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

White:1989:NFW

[Whi89] Kenneth J. White. NAG Fortran Workstation Library: A review. Journal of Applied Econometrics, 4(1):93–96, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wichmann:1989:SPI

[Wic89] B. A. Wichmann. Scientific processing in ISO-pascal: a proposal to get the benefits of mixed precision floating-point. ACM SIGPLAN Notices, 24(6):20–22, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Widemann:1988:IUF

[Wid88] Thomas Widemann. Interprozedurale Ueberpruefung von Fortran-Programmen. Master’s thesis, Technische Universität Braunschweig (??), Braunschweig, Germany, 1988.

Wiecek:1982:CSV

[Wie82] Cheryl A. Wiecek. A case study of VAX-11 instruction set usage for compiler execution. ACM SIGPLAN Notices, 17(4):177–184, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wier:1983:ETS

[Wie83] Robert Reid Wier. An experiment in teaching structured FOR-TRAN. Thesis (Ph.D.), University of Texas at Arlington, Arlington, TX, USA, 1983. ix + 99 pp.

Wienecke:1985:UFE

[Wie85] U. Wienecke. Ueberarbeitung von Fortran-Programmen zur Erstellung und zur Pruefung von elektrischen Verbindungslisten. Studienarbeit, 1985.

Wiese:1986:UVDa

[Wie86a] Michael R. Wiese. USSAERO version D computer program development using ANSI standard FORTRAN 77 and DI-3000 graphics. NASA contractor report NASA CR-3980, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1986. viii + 97 pp. For sale by the National Technical Information Service.

Wiese:1986:UVDb

[Wie86b] Michael R. Wiese. USSAERO version D computer program development using ANSI standard FOR-TRAN 77 and DI-3000 graphics. NASA contractor report NASA CR-3980., National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1986. viii + 97 pp.

Wilson:1983:AIP

[Wil83] J. M. Wilson. Algorithm 114: Interrupted permutations in lexicographic order. The Computer Journal, 26(1):92, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Williams:1984:AR

[Wil84] David O. Williams. Alternate RE-TURNs. ACM SIGPLAN Notices, 19(10):52–57, October 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1987:ITI

[Wil87a] David Thomas Williams. In-line translation of input/output statements of Fortran to ADA. Thesis (M.S.), Dept. of Computer Science, Wichita State University, Wichita, KS, USA, 1987. 43 pp.

Wilson:1987:BRB

[Wil87b] J. D. Wilson. Book review: Fortran Optimization (revised edition), Michael Metcalf, Academic Press, 1985. No. of pages: 253. Price: £20.00. Software—Practice and Experience, 17(2):166– 167, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Winston:1985:BFC

[Win85] Elliot Winston. BTREE: a FOR-TRAN code for a B+ tree. Technical Report NSWC TR 85-54, Naval Surface Weapons Center, Silver Spring, MD, USA, 1985. iv + 6 + 22 + 30 pp.

Wist:1981:FTM

[Wis81] Peter Wist. Fortran-Routinen für den Test eines mikroprogrammierten Spezialprozessors. Studienarbeit, 1981.

Wittram:1981:HAS

[Wit81] M. Wittram. Hidden-line algorithm for scenes of high complexity. Computer Aided Design, 13, 4: 187–192, 1981. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

Wyatt:1976:PEP

[WLO76] W. T. Wyatt, Jr., D. W. Lozier, and D. J. Orser. A portable extended precision arithmetic package and library with Fortran precompiler. ACM Transactions on Mathematical Software, 2(3): 209–231, September 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Whitehead:1985:FPD

[WM85a] J. Whitehead and P. Marek. A FORTRAN program for the design and analysis of sequential clinical trials. Computers in Biomedical Research, 18(2):176–183, April 1985. CODEN CPBRAF. ISSN 0010-4809 (print), 1090-2368 (electronic).

Wolfe:1985:WOC

[WM85b] M. Wolfe and T. Macke. Where are the optimizing compilers? ACM SIGPLAN Notices, 20(11):64–68, November 1985. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whang:1987:EDN

[WN87] K. Y. Whang and S. B. Navathe. An extended disjunctive normal form approach for processing recursive logic queries in loosely coupled environments. Technical Report Res.R. No.R, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, March 1987.

Wolfe:1985:LPB

[Wol85] Carvel S. Wolfe. Linear Programming With Basic and Fortran. Reston Publishing Co., Inc., Reston, VA, USA, January 1985. ISBN 0-8359-4082-9 (paperback). x + 293 pp. LCCN T 57.74 W64 1985. US$33.00; US$19.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835940829.

Woo:1989:MFL

[Woo89] Richard Woo. Mathematical Function Library for Microsoft Fortran. John Wiley and Sons, New York, London, Sydney, January 1989. ISBN 0-471-63831-5. ???? pp. LCCN ???? US$595.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0471638315. Includes disk.

Inc:1984:IPA

[Wor84a] Wordware Pub., Inc. IBM PC Apprentice: Fortran 77 Beginning. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1984. ISBN 0-13-452244-3. ???? pp. LCCN ???? US$19.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0134522443.

WPI:1984:IPA

[Wor84b] Wordware Pub., Inc. IBM PC Apprentice: Fortran 77 Beginning (Tutorial Workbook). Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1984. ISBN 0-13-452244-3. ???? pp. LCCN ???? US$19.00.

Worland:1988:MFS

[Wor88] Peter B. Worland. Modern Fortran 77 for Scientists and Engineers. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, December 1988. ISBN 0-15-561156-9. ???? pp. LCCN ???? US$39.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0155611569.

Worland:1989:MF

[Wor89a] Peter B. Worland. Modern Fortran 77. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1989. ISBN 0-15-561156-9. ???? pp.

Worland:1989:MFS

[Wor89b] Peter B. Worland. Modern Fortran 77 for scientists and engineers. Harcourt Brace Jovanovich and its subsidiary, Academic Press, San Diego, CA, USA, 1989. ISBN 0-15-561156-9. xv + 410 pp. LCCN QA76.73.F25W66 1989.

Wossner:1982:PPM

[Wös82] Hans Wössner, editor. Programmiersprachen und Programmentwicklung München, 9./10. März 1982, 7. Fachtagung. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-11204-9. LCCN QA76.7 .P74 1982.

Wise:1984:FCP

[WP84] Robert A. Wise and Fletcher C. Parsons. Fortran 77 computer program for test procedure calculations of vented heaters. Nbsir; 84-2918, U.S. Department of Commerce, National Bureau of Standards National Engineering Laboratory, Center for Building Technology, Building Equipment Division, Washington, DC, USA, 1984. iii + 47 pp.

Wright:1989:ACA

[Wri89] M. Wright, editor. Aspects of Computation on Asynchronous Parallel Processors. Proceedings of the IFIP WG 2.5 Working Conference, Stanford, CA, USA, 22– 26 August, 1988. North-Holland Publishing Co., Amsterdam, The Netherlands, 1989. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Weiss:1984:PFP

[WS84] Robert Weiss and Charles Seiter.

Pascal for Fortran Programmers. Micro computer books. Addison-Wesley, Reading, MA, USA, July 1984. ISBN 0-201-08296-9 (paperback). xii + 237 pp. LCCN QA76.73.P2 W43 1984. US$16.95

(est.). URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0201082969.

Wasserman:1988:PMA

[WSL88] H. J. Wasserman, M. L. Simmons, and O. M. Lubeck. The performance of minisupercomputers: Alliant FX/8, Convex C-1, and SCS-40. Parallel Computing, 8(1–3): 285–293, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wu:1982:AFIa

[Wu82a] Nesa L’abbe Wu. ANSI FOR-TRAN IV and FORTRAN 77 programming with business applications. W. C. Brown Co. Pub., Dubuque, IA, USA, third edition, 1982. ISBN 0-697-08153-2. x + 321 pp. LCCN HF5548.2.W8 1982.

Wu:1982:AFIb

[Wu82b] Nesa L’abbe Wu. ANSI FOR-TRAN IV and FORTRAN 77: programming with business applications. Wm. C. Brown Publishers, Dubuque, IA, USA, third edition, 1982. ISBN 0-697-08153-2 (paperback). x + 321 pp. LCCN HF5548.2.W8 1982.

Wu:1983:AFI

[Wu83] Nesa L’Abbe Wu. ANSI Fortran IV and Fortran 77: Programming With Business Applications. Wm. C. Brown Publishers, Dubuque, IA, USA, third edition, February 1983. ISBN 0-697-08153-2. ???? pp. LCCN ???? US$38.28. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0697081532.

Weerawarana:1989:GPC

[WW89] S. Weerawarana and P. S. Wang. GENCRAY: a portable code generator for Cray Fortran. In ACM [ACM89a], pages 186–191. ISBN 0-89791-325-6. LCCN ????

Wylie:1986:RSS

[Wyl86] Ian Wakefield Wylie. A Raman spectroscopic study of @-carotene the development and application of a Fortran program to simulate a non steady state monochromatic transmission experiment. Thesis (M.Sc.), Carleton University, Ottawa, Ontario, Canada, 1986. 2 microfiches (138 fr.) pp.

ANSI:1989:FD

[X3J89] ANSI X3J3. Fortran 8X draft. ACM SIGPLAN Notices, 8(4):??, December 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Also published by Global Engineering.

Ying:1985:AAF

[YF85] Dao-Ning Ying and Xing Feng. Arbitrary area filling in a fast procedure. Computer Graphics Forum, 4(4):363–370, December 1985. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yamana:1989:PAS

[YHKM89] H. Yamana, T. Hagiwara, J. Kohdate, and Y. Muraoka. A preceding activation scheme with graph unfolding for the parallel processing system-harray. In ACM [ACM89b], pages 675–684. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Yang:1988:MFP

[YK88] M. H. Yang and S. J. Knapp. Microcomputer FORTRAN programs for calculating percentage points from t or F distributions. The Journal of Heredity, 79(2):122–??, March–April 1988. CODEN JO-HEA8. ISSN 0022-1503.

Young:1985:MCP

[YM85] Ming Huei Young and Saburo Muroga. Minimal covering problem and PLA minimization. International Journal of Computer and Information Sciences, 14(6): 337–364, December 1985. CODEN IJCIAH. ISSN 0091-7036.

Younglove:1982:IFP

[You82] Ben Younglove. Interactive Fortran program to calculate thermophysical properties of six fluids. NBS technical note 1048, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1982. iii + 49 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Young:1984:CAG

[YS84a] Jon I. Young and Paul L. Schlieve. Course administrator’s guide: FORTRAN 77 — beginning. IBM PC apprentice. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-452251-6. viii + 216 + 43 pp.

Young:1984:FB

[YS84b] Jon I. Young and Paul L. Schlieve. FORTRAN 77 — beginning. IBM PC apprentice IBM PC apprentice personal computer learning series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-452244-3. viii + 216 pp.

Young:1984:FCA

[YS84c] Jon I. Young and Paul L. Schlieve. FORTRAN 77-beginning course administrator’s guide, 1984.

Young:1984:FSE

[YS84d] Jon I. Young and Paul L. Schlieve. FORTRAN 77-beginning: selfteach edition, 1984. ISBN 0-13-452889-1 (Self-teach edition).

Wang:1985:FCH

[yW85] Ya yung Wang. FORTRAN 77 cheng hsu she chi yu yen. Yeh chin kung yeh chu pan she, Pei-ching, ti 1 pan edition, 1985. ISBN ???? 198 pp. LCCN ????

Zaki:1984:ASS

[Zak84] Ahmed S. Zaki. Algorithm 119: a space saving insertion algorithm for 2-3 trees. The Computer Journal, 27(4):368–372, November 1984. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/
368.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_27/Issue_04/tiff/369.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
27/Issue_04/tiff/370.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_
04/tiff/371.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/
372.tif.

Zima:1988:STS

[ZBG88] H. P. Zima, H.-J. Bast, and M. Gerndt. SUPERB: a tool for semi-automatic MIMD/SIMD parallelization. Parallel Computing, 6 (1):1–18, January 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zinkl:1981:FCP

[ZDS81a] Richard J. Zinkl, Ralph F.

D’Andrea, and Don L. Shettel. FORTRAN computer programs to read, sort, and edit Los Alamos hydrogeochemical data. Technical Report GJBX 11(81), U.S. Dept. of Energy, Grand Junction Office, Grand Junction, CO, USA, 1981. 53 pp.

Zinkl:1981:OFC

[ZDS81b] Richard J. Zinkl, Ralph F.

D’Andrea, and Don L. Shettel. OREDIT, a FORTRAN computer program to process Oak Ridge HSSR data. Technical Report GJBX 10(81), U.S. Dept. of Energy, Grand Junction Office, Grand Junction, CO, USA, 1981. 39 pp.

Zoicas:1988:PBG

[ZGK88] A. Zoicas, K. Grohe, and C. Kellerhoff. PC based general-purpose floating-point DSP mu PD77230 board with various analog front end options and application software packages. In Lacoume et al. [LCMM88], pages 1233–1236 (vol. 3). ISBN 0-444-70516-3. LCCN TK5102.5.E9 1988. Three volumes.

Zuo:1989:FSL

[ZGS89] J. M. Zuo, K. Gjonnes, and J. C. Spence. FORTRAN source listing for simulating three-dimensional convergent beam patterns with absorption by the Bloch wave method. Journal of Electron Microscopy Technique, 12(1):29–55, May 1989. CODEN JEMTEC. ISSN 0741-0581.

Zhou:1989:UGF

[Zho89] Jian Zhou. User’s guide for FSQP version 1.0B: a Fortran software for solving optimization problems with general inequality constraints and linear equality constraints, generating feasible iterates. Technical research report SRC TR 89-61r1, University of Maryland, Systems Research Center, College Park, MD, USA, 1989. 19 pp.

Zimmer:1986:DFI

[Zim86] Susan Jane Zimmer. Development of a FORTRAN 77 interpreter. Thesis (M.S.), Oklahoma State University, Stillwater, OK, USA, 1986. viii + 195 pp.

Zenios:1986:NNP

[ZM86] Stavros A. Zenios and John M.

Mulvey. Nonlinear network programming on vector supercomputers: a study on the Cray X-MP. Operations Research, 34(5):667– 682, September–October 1986. CODEN OPREAI. ISSN 0030-364X (print), 1526-5463 (electronic).

Zohni:1984:FPC

[Zoh84] O. Zohni. A Fortran program for the computation of the generalized Talmi coefficients. Computer Physics Communications, 35 (1–3):C–123–C–124, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/S0010465584823814.

Zinkl:1982:LFC

[ZSD82a] Richard J. Zinkl, Don L. Shettel, and Ralph F. D’Andrea. LAB-PLT: A FORTRAN computer system to plot and label sample location maps for hydrogeochemical and stream sediment reconnaissance data. Technical Report GJBX 192-82, U.S. Dept. of Energy, Grand Junction, CO, USA, 1982. 95 pp.

Zinkl:1982:PFC

[ZSD82b] Richard J. Zinkl, Don L. Shettel, and Ralph F. D’Andrea. PLT-SYM: a FORTRAN computer system to plot Canadian symbol location maps for hydrogeochemical and stream-sediment reconnaissance data. Technical Report GJBX 193-82, U.S. Dept. of Energy, Grand Junction, CO, USA, September 1982. 141 pp.

Zwass:1981:PFS

[Zwa81] Vladimir Zwass. Programming in FORTRAN: structured programming with FORTRAN IV and FORTRAN 77. The Barnes and Noble outline series; COS 194. Barnes and Noble, New York, NY, USA, 1981. ISBN 0-06-460194-3 (paperback). x + 213 pp. LCCN QA76.73.F25Z86 1981.

Zwass:1985:PFP

[Zwa85] Vladimir Zwass. Programando en FORTRAN: programacion estructurada con FORTRAN IV y FOR-TRAN 77. El Tutor del estudiante. CECSA, Mexico, DF, Mexico, 1a en espanol de la 1a en ingles edition, 1985. ISBN 968-26-0551-2. 229 pp.